### v1

Abstract
We tackle the hard problem of consciousness taking the naturally-selected, self-organising, embodied organism as our starting point. We provide a mathematical formalism describing how biological systems self-organise to hierarchically interpret unlabelled sensory information according to valence and specific needs. Such interpretations imply behavioural policies which can only be differentiated from each other by the qualitative aspect of information processing. Selection pressures favour systems that can intervene in the world to achieve homeostatic and reproductive goals. Quality is a property arising in such systems to link cause to affect to motivate real world interventions. This produces a range of qualitative classifiers (interoceptive and exteroceptive) that motivate specific actions and determine priorities and preferences. Building upon the seminal distinction between access and phenomenal consciousness, our radical claim here is that phenomenal consciousness without access consciousness is likely very common, but the reverse is implausible. To put it provocatively: Nature does not like zombies. We formally describe the multilayered architecture of self-organisation from rocks to Einstein, illustrating how our argument applies in the real world. We claim that access consciousness at the human level is impossible without the ability to hierarchically model i) the self, ii) the world/others and iii) the self as modelled by others. Phenomenal consciousness is therefore required for human-level functionality. Our proposal lays the foundations of a formal science of consciousness, deeply connected with natural selection rather than abstract thinking, closer to human fact than zombie fiction.
keywords:
hard problem of consciousness
1 Introduction
Why is anything conscious? [^1] This work was supported by a Fundaçao para a Ciencia e a Tecnologia (FCT) grant PTDC/FER-FIL/4802/2020 ; and 2020-02773 CEECIND FCT to A.C. Both biological and other physical [^2] Physical here just means non-biological. We are not suggesting biological systems are non-physical. systems process information, yet it seems that humans consciously experience in addition to merely process information. Why? Living organisms are constantly processing self- and world-related information to secure survival in an ever-changing world. Human bodies share with all other physical systems the property of being instantiated in time and space, (e.g., our body occupies a given position and volume in space at a given time). Yet, unlike physical systems, biological, living systems are dissipative systems using energy to self-organise in the face of entropic decay and environmental perturbation [1, 2].
Originally formalised in the field of cybernetics [3, 4] the notion of self-organisation has been subsequently applied to various disciplines including physics [5], biology [6, 7] and neuroscience [8, 9, 10]. Self-organization is typically defined as the spontaneous emergence of spatiotemporal order or pattern-formation processes in physical and biological systems resulting from interactions of its components with the environment [11, 12, 13].
Interestingly, much self- and world related information processing goes on behind the scenes, or “in the dark” so to speak, that is, without being constantly present to our conscious minds. But why doesn’t all information processing go in the dark?
This question is the subject of long-standing debates across disciplines [14]. For example, one highly influential view is that consciousness has two aspects [14]. The first is functional, by which we mean the ability to access and communicate information [15]. How exactly information processing is linked to consciousness is the “easy problem” of consciousness [2]. The second aspect is “what it is like” to consciously experience information processing, or phenomenal consciousness [2, 15, 1, 16, 17]. This doesn’t mean just global states like being awake, but more specific local states like smelling a cup of coffee. These local contents or “qualia” are characterised by what it is like to be in them [14]. It is unclear the extent to which functional and phenomenal aspects are independent. David Chalmers has influentially suggested that it may be possible to construct a “zombie” which acts in every way like a person but has no qualia [15]. For example, a thermostat certainly detects heat and so processes information, but there presumably is not anything it is like to be a thermostat. Hence the question “why is anything conscious” may be understood as “why is there sometimes a qualitative aspect to information processing?”. This is the “hard problem” of consciousness.
The hard problem has sparked a substantial body of work and a detailed discussion of these debates [14, 18] lies beyond the scope of our paper. Rather, in what follows we build upon the useful distinction between lower and higher order theories of information processing in relation to conscious experiences.
For example, higher order thought theory (HOT) [19, 20] holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states. Lower order states may include emotions and perceptions, while higher order meta-representations reflect upon those. The link between the two may explain something of the phenomenal character of states. Sense data is processed by the body resulting in lower order mental states, and then meta representations of those is where we might find more abstract conceptual or thought-like contents of consciousness.
The division between lower and higher orders is a good starting point to understand why some information processing goes on “in the dark”. Yet, the interesting question in our view is why do these lower order states arise? Can they occur in the absence of subjective, qualitative experience? It is widely agreed that higher order mental states such as desire to drink a coffee is accompanied by a qualitative aspect. But is there something it is like to “be in the dark”? That is: to process information at the lower bodily levels?
In this paper we suggest that a useful way to dissolve the “hard problem” of consciousness is to reverse the order and start with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states.
If an explanation of how consciousness functions is to reveal why information processing sometimes has a qualitative aspect, then it must explain how we get local states [21], not begin by assuming local states. We need to go a level down and begin at the level of the embodied organism. Longtime considered a fringe approach, the embodied cognition paradigm [22] has recently gained substantial influence in cognitive science and philosophy [23, 24, 25]. The key idea is that instead of considering the body as a mere device designed to fuel and contain the mind (a device that can be replaced with a vat or a robot, for example), one must consider the mind as serving the self-sustaining needs of a surviving body.
If this is so, then understanding consciousness must start with understanding the ‘humble’ lower bodily levels of information processing, and not the higher order levels of information processing only. The key idea is that conscious experiences do not merely depend on bodily experiences as an external factor that can be replaced with a vat or an artificial system.
One promising approach to this question is to try to rigorously define consciousness from first principles and show that some aspects of functional consciousness depend on phenomenal consciousness in a manner that makes zombies impossible, “dissolving” the hard problem by showing the phenomenal to be functional [^3] Note that we are far from the first to claim that the phenomenal is functional. For example, the proposed Conscious Turing Machine [26] based on Global Workspace Theory [27], and constructivist approaches to artificial intelligence [28] take similar position. It is our explanation of how and why that is novel.. Unifying them. We would need to establish axioms that hold in every possible environment and show that it is impossible to have the function of consciousness without the subjective experience of it. Such an answer must explain how consciousness functions, and why some information processing goes on “in the dark”.
This paper lays the grounds of a computational model formalizing the link between lower and higher levels of information processing in relation to conscious experiences in a way that is compatible with the basic principles of embodied cognition and enactivism. Enactivism is roughly the view that information processing arises through a dynamic interaction between an acting organism and its environment [29].
Now, the notion of computation is widely debated, and a detailed review of these discussions would lead us to a major digression [30, 31]. Here we define computation not in terms of symbol shuffling and representation, but in mechanistic terms. ‘Information processing’ in this sense is the causal relations dictating the transition of a system from one state to another, which concerns biological systems such as human bodies.
Pancomputationalism is the idea that all dynamic systems are constituted by computation. This paper builds upon and extends the formalism of pancomputational enactivism developed by one of us [32]. To answer “why is anything conscious” we proceed in two steps. First we develop a mathematical formalism in which lower and higher order theories of consciousness, and phenomenal and access consciousness, are all derived from first principles. Each follows because of scaling natural selection pressures and the ability to adapt [^4] This is loosely inspired by a scale-based framing of machine learning [33].. Second, we provide an argument building on that to dissolve the hard problem of consciousness, answering why there must a qualitative aspect to information processing and in what circumstances.
This paper is organised as follows. Section 2 is a list of formal mathematical definitions for pancomputational enactivism, which the reader may refer to as needed. For readers without formal background, the Section 2 can be skipped, without impacting the argumentative streamline of the paper. In section 3 we lay the foundations. We extend separately published work unifying pancomputationalism and enactivism [32], developing a model of self-organising systems that holds across all conceivable environments. We explain how this model formalises relevance realisation and unifies lower and higher order theories, citing separately published mathematical and experimental results [34]. We extend previous work [35] on causal learning and the development of self, to formalise analogues of subjective experience [36, 37, 38], access consciousness [15] and meta self-awareness [39]. We call these analogues first (1ST), second (2ND) and third (3RD) order selves respectively, and explain how they are constructed as a direct consequence of scaling the ability of a self-organising system to adapt with natural selection pressures. Section 4 has two main parts. First we explain how subjective experience requires a 1ST order self, and conversely why a 1ST order self implies there is “something” it is like to be an organism that has a 1ST order self. We then argue qualia are information processing without representation, and representationless information becomes qualia when there is a common 1ST order self be subject to these experiences. Phenomenal contents must precede representational contents, because representational contents are just interpretations from phenomenal experience. Quality precedes quantity. We call this the psychophysical principle of causality.
In the second part we describe the developmental stages of consciousness this implies as we scale up the capacity to adapt with natural selection pressures. We identify 6 stages including unconscious, hard-coded behaviour, learning, and then 1ST, 2ND and 3RD order selves. We provide examples of each from computers to jellyfish [40], to houseflies [38], to birds [41] and then humans. Finally, in section 5 we provide concluding remarks and outlook discussion for future research.
2 Pancomputational Enactivism Definitions
For convenience of reference, we have placed all definitions here. We have aimed to make the gist of the paper understandable without the math, hence the reader can either skip this section or refer back to it if needed. Many of the definitions have been adapted from a variety of preceding work [32, 34, 35, 42, 43, 44]. They are referred to in the body of the paper when they become relevant, in the order in which they appear here.
Definition 1 (environment).
  • We assume a set $\Phi$ whose elements we call states.
  • A declarative program is $f\subseteq\Phi$, and we write $P$ for the set of all declarative programs (the powerset of $\Phi$).
  • By a truth or fact about a state $\phi$, we mean $f\in P$ such that $\phi\in f$.
  • By an aspect of a state $\phi$ we mean a set $l$ of facts about $\phi$ s.t. $\phi\in\bigcap l$. By an aspect of the environment we mean an aspect $l$ of any state, s.t. $\bigcap l\neq\emptyset$. We say an aspect of the environment is realised [^5] Realised meaning it is made real, or brought into existence. by state $\phi$ if it is an aspect of $\phi$.
Definition 2 (abstraction layer).
  • We single out a subset $\mathfrak{v}\subseteq P$ which we call the vocabulary of an abstraction layer. The vocabulary is finite.
  • $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap l\neq\emptyset\}$ is a set of aspects in $\mathfrak{v}$. We call $L_{\mathfrak{v}}$ a formal language, and $l\in L_{\mathfrak{v}}$ a statement.
  • We say a statement is true given a state iff it is an aspect realised by that state.
  • A completion of a statement $x$ is a statement $y$ which is a superset of $x$. If $y$ is true, then $x$ is true.
  • The extension of a statement $x\in{L_{\mathfrak{v}}}$ is $E_{x}=\{y\in{L_{\mathfrak{v}}}:x\subseteq y\}$. $E_{x}$ is the set of all completions of $x$.
  • The extension of a set of statements $X\subseteq{L_{\mathfrak{v}}}$ is $E_{X}=\bigcup\limits_{x\in X}E_{x}$.
  • We say $x$ and $y$ are equivalent iff $E_{x}=E_{y}$.
(notation) $E$ with a subscript is the extension of the subscript666e.g. $E_{l}$ is the extension of $l$..
(intuitive summary) $L_{\mathfrak{v}}$ is everything which can be realised in this abstraction layer. The extension $E_{x}$ of a statement $x$ is the set of all statements whose existence implies $x$, and so it is like a truth table. Intuitively a sensorimotor system is an abstraction layer. Likewise, a computer (each statement asserting a state of the computer, or a part thereof).
Definition 3 ($\mathfrak{v}$-task).
For a chosen $\mathfrak{v}$, a task $\alpha$ is a pair $\langle{I}_{\alpha},{O}_{\alpha}\rangle$ where:
  • ${I}_{\alpha}\subset L_{\mathfrak{v}}$ is a set whose elements we call inputs of $\alpha$.
  • ${O_{\alpha}}\subset E_{I_{\alpha}}$ is a set whose elements we call correct outputs of $\alpha$.
${I_{\alpha}}$ has the extension $E_{I_{\alpha}}$ we call outputs, and ${O_{\alpha}}$ are outputs deemed correct. $\Gamma_{\mathfrak{v}}$ is the set of all tasks given $\mathfrak{v}$.
[…]
(notation) If $\omega\in\Gamma_{\mathfrak{v}}$, then we will use subscript $\omega$ to signify parts of $\omega$, meaning one should assume $\omega=\langle{I}_{\omega},{O}_{\omega}\rangle$ even if that isn’t written.
(intuitive summary) To reiterate and summarise the above:
  • An input is a possibly incomplete description of a world.
  • An output is a completion of an input [def. 2].
  • A correct output is a correct completion of an input.
A $\mathfrak{v}$-task is a formal, behavioural description of goal directed behaviour. For example, an organism could be described by all behaviour in which it remains alive. Likewise, a $\mathfrak{v}$-task could describe a Turing machine.
Definition 4 (inference).
  • A $\mathfrak{v}$-task policy is a statement $\pi\in L_{\mathfrak{v}}$. It constrains how we complete inputs.
  • $\pi$ is a correct policy iff the correct outputs $O_{\alpha}$ of $\alpha$ are exactly the completions $\pi^{\prime}$ of $\pi$ such that $\pi^{\prime}$ is also a completion of an input.
  • The set of all correct policies for a task $\alpha$ is denoted $\Pi_{\alpha}$. [^7] To repeat the above definition in set builder notation: ${\Pi_{\alpha}}=\{\pi\in L_{\mathfrak{v}}:{E}_{I_{\alpha}}\cap E_{\pi}={O_{\alpha}}\}$
Assume $\mathfrak{v}$-task $\omega$ and a policy $\pi\in L_{\mathfrak{v}}$. Inference proceeds as follows:
  1. 1.
    we are presented with an input ${i}\in{I}_{\omega}$, and
  2. 2.
    we must select an output $e\in E_{i}\cap E_{\pi}$.
  3. 3.
    If $e\in{O}_{\omega}$, then $e$ is correct and the task “complete”. $\pi\in{\Pi}_{\omega}$ implies $e\in{O}_{\omega}$, but $e\in{O}_{\omega}$ doesn’t imply $\pi\in{\Pi}_{\omega}$ (an incorrect policy can imply a correct output).
(intuitive summary) To reiterate and summarise the above:
  • A policy constrains how we complete inputs.
  • A correct policy is one that constrains us to correct outputs.
In functionalist terms, a policy is a “causal intermediary” between inputs and outputs.
Definition 5 (learning).
. A proxy $<$ is a binary relation on statements. In this paper we use only one proxy, called the weakness proxy, which compares the cardinality of a statement’s extension. For statements $l_{1},l_{2}$ we have $l_{1}<l_{2}$ iff $\lvert Z_{l_{1}}\rvert<\lvert Z_{l_{2}}\rvert$. Whenever we use $<$ to compare statements, we are referring to the aforementioned weakness proxy.
(generalisation) A statement $l$ generalises to a $\mathfrak{v}$-task $\alpha$ iff $l\in\Pi_{\alpha}$. We speak of learning $\omega$ from $\alpha$ iff, given a proxy $<$, $\pi\in{\Pi}_{\alpha}$ maximises $<$ relative to all other policies in ${\Pi}_{\alpha}$, and $\pi\in{\Pi}_{\omega}$.
(probability of generalisation) We assume a uniform distribution over $\Gamma_{\mathfrak{v}}$. If $l_{1}$ and $l_{2}$ are policies, we say it is less probable that $l_{1}$ generalizes than that $l_{2}$ generalizes, written $l_{1}<_{g}l_{2}$, iff, when a task $\alpha$ is chosen at random from $\Gamma_{\mathfrak{v}}$ (using a uniform distribution) then the probability that $l_{1}$ generalizes to $\alpha$ is less than the probability that $l_{2}$ generalizes to $\alpha$.
(sample efficiency) Suppose $\mathfrak{app}$ is the set of all pairs of policies. Assume a proxy $<$ returns $1$ iff true, else 00. Proxy $<_{a}$ is more sample efficient than $<_{b}$ iff
$\left(\sum_{(l_{1},l_{2})\in\mathfrak{app}}\lvert(l_{1}<_{g}l_{2})-(l_{1}<_{a}l_{2})\rvert-\lvert(l_{1}<_{g}l_{2})-(l_{1}<_{b}l_{2})\rvert\right)<0$
(optimal proxy) There is no proxy more sample efficient than weakness. The weakness proxy formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [34]).
(intuitive summary) Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy. Humans typically learn from “examples”. An example of a task is a correct output and input. A collection of examples is a child task, so “learning” is an attempt to generalise from a child to one of its parents. The lower level the child from which an agent generalises to parent, the “faster” it learns, the more sample efficient the proxy. The most sample efficient proxy is weakness [34, prop. 1, 2], which is why we’re using it here.
Definition 6 (organism).
We can describe the circumstances of an organism $\mathfrak{o}$ as $\langle\mathfrak{v}_{\mathfrak{o}},\mu_{\mathfrak{o}},\mathfrak{p}_{\mathfrak{o}},<_{\mathfrak{o}}\rangle$ where:
  • ${O}_{\mu_{\mathfrak{o}}}$ contains every output which qualifies as “fit” according to natural selection.
  • $\mathfrak{p}_{\mathfrak{o}}$ is the set of policies an organism knows, s.t. $\mathfrak{p}_{\mathfrak{o}}\subset\mathfrak{p}_{n.s.}\cup\mathfrak{p}_{\mathfrak{h}_{{<t}_{\mathfrak{o}}}}$ and:
    • $\mathfrak{p}_{n.s.}\subset L_{\mathfrak{v}_{\mathfrak{o}}}$ is reflexes hard coded from birth by natural selection.
    • $\mathfrak{p}_{\mathfrak{h}_{{<t}_{\mathfrak{o}}}}=\underset{\zeta\in\mathfrak{h}_{{<t}_{\mathfrak{o}}}}{\bigcup}\Pi_{\zeta}$ is the set of policies it is possible to learn from a history of past interactions represented by a task $\mathfrak{h}_{{<t}_{\mathfrak{o}}}$.
    • If $\Pi_{{\mathfrak{h}_{<t_{\mathfrak{o}}}}}\not\subset(\mathfrak{p}_{\mathfrak{o}}-\mathfrak{p}_{n.s.})$ then the organism has selective memory. It can “forget” outputs, possibly to productive ends if they contradict otherwise good policies.
  • $<_{\mathfrak{o}}$ is a binary relation over $\Gamma_{\mathfrak{v}_{\mathfrak{o}}}$ we call preferences.
(intuitive summary) Strictly speaking an organism $\mathfrak{o}$ would be a policy, but we can describe the circumstances of its existence as a task $\mu$ that describes all “fit” behaviour for that organism. We can also identify policies the organism “knows”, because these are implied by the policy that is the organism. Likewise, we can represent lossy memory by having the organism “know” fewer policies than are implied by its history of interactions. Finally, preferences are the particular “protosymbol” the organism will use to “interpret” an input in later definitions.
Definition 7 (protosymbol system).
Assume an organism $\mathfrak{o}$. For each policy $p\in\mathfrak{p}_{\mathfrak{o}}$ there exists a set $\mathfrak{s}_{p}=\{\alpha\in\Gamma_{\mathfrak{v}_{\mathfrak{o}}}:p\in\Pi_{\alpha}\}$ of all tasks for which $p$ is a correct policy. The union of all such sets is
$\mathfrak{s}_{\mathfrak{o}}=\underset{p\in\mathfrak{p}_{\mathfrak{o}}}{\bigcup}\{\alpha\in\Gamma_{\mathfrak{v}_{\mathfrak{o}}}:p\in\Pi_{\alpha}\}$
We call $\mathfrak{s}_{\mathfrak{o}}$ a “protosymbol system”. A $\mathfrak{v}$-task $\alpha\in\mathfrak{s}_{\mathfrak{o}}$ is called a “protosymbol”, and is “more abstract” if it is higher in the generational hierarchy.
Definition 8 (interpretation).
Interpretation is an activity undertaken by an organism $\mathfrak{o}=\langle\mathfrak{v}_{\mathfrak{o}},\mu_{\mathfrak{o}},\mathfrak{p}_{\mathfrak{o}},<_{\mathfrak{o}}\rangle$, as follows:
  1. 1.
    Assume an input $i\in L_{\mathfrak{v}_{\mathfrak{o}}}$.
  2. 2.
    We say that $i$ signifies a protosymbol $\alpha\in\mathfrak{s}_{\mathfrak{o}}$ if $i\in I_{\alpha}$.
  3. 3.
    $\mathfrak{s}_{\mathfrak{o}}^{i}=\{\alpha\in\mathfrak{s}_{\mathfrak{o}}:i\in I_{\alpha}\}$ is the set of all protosymbols which $i$ signifies.
  4. 4.
    If $\mathfrak{s}_{\mathfrak{o}}^{i}\neq\emptyset$ then $i$ means something to the organism in the intuitive sense that there is “affect” or “value” compelling the organism to act.
  5. 5.
    If $i$ means something, then $\mathfrak{o}$ chooses $\alpha\in\mathfrak{s}_{\mathfrak{o}}^{i}$ that maximises its preferences $<_{\mathfrak{o}}$.
  6. 6.
    The organism then infers an output $o\in E_{i}\cap E_{\Pi_{\alpha}}$.
(intuitive summary) Interpretation is inference, with the additional step of choosing policies according to preference. This allows for irrational and instinctive choices, as well as rational ones. Intuitively, $i$ is every aspect of the context in which the organism finds itself; everything that can influence its interpretation.
Definition 9 (to affect).
Suppose we have two organisms, $\mathfrak{a}$ (Alice) and $\mathfrak{b}$ (Bob). Suppose $\mathfrak{a}$ interprets $i\in L_{\mathfrak{v}_{\mathfrak{o}}}$ as an output $o$, then:
  • a statement $v\subset i$ affects $\mathfrak{a}$ if $\mathfrak{a}$ would have interpreted $e=i-v$ as a different output $g\neq o$.
  • an organism $\mathfrak{b}$ has affected $\mathfrak{a}$ by making an output $k$ if, because of $k$, there exists $v\subset s$ which affects $\mathfrak{a}$.
Definition 10 (intervention).
By event we mean a statement in $L_{\mathfrak{v}}$, and an event happens or is observed iff it is a true statement given a state $\phi$. If $obs\in L_{\mathfrak{v}}$ is sensorimotor activity we interpret as an “observed event”, and $int\in L_{\mathfrak{v}}$ is an intervention to cause that event, then $obs\subset int$ (because $int$ could not be said to cause $obs$ unless $obs\subset int$).
(intuitive summary) An intervention is action undertake an organism or other agency, in the sense described by Pearl [45]. Intuitively, if “$int$” and “$obs$” are events which have happened, then we say that $int$ has caused $obs$ if $obs$ would not have happened in the absence of $int$ (counterfactual).
Definition 11 (causal identity).
If $obs\in L_{\mathfrak{v}}$ is an observed event, and $int\in L_{\mathfrak{v}}$ is in intervention causing $obs$, then $c\subseteq int-obs$ “identifies” or “names” the intervening agency. If $c=\emptyset$ then we have no way of knowing the intervening agency, if there is one. We call $c$ a causal identity corresponding to $int$ and $obs$. Suppose $INT$ and $OBS$ are sets of statements, and we assume $OBS$ contains observed events and $INT$ interventions, then a causal identity corresponding to $INT$ and $OBS$ is $c\neq\emptyset$ s.t. $\forall i\in INT(c\subset int)$ and $\forall obs\in OBS(c\cap obs=\emptyset)$ (we can attempt to construct a causal identity for any $INT$ and $OBS$). If a policy is a causal identity, then the associated task is to classify interventions.
Definition 12 (purpose, goal or intent).
We consider a policy $c$ which is a causal identity corresponding to $INT$ and $OBS$ to be the intent, purpose or goal ascribed to the interventions. $c$ is what the interventions share in common, meaning the “name” or “identity” of behaviour is the “intent”, “goal” or “purpose” of behaviour. Just as an intervention caused an observation, the intent which motivated the agency undertaking the intervention is what caused it (to correctly infer intent, one must infer a causal identity that implies subsequent interventions).
Definition 13 (1ST order self).
If $c$ is the lowest level causal identity corresponding to $INT$ and $OBS$, and $INT$ is every intervention an organism could make (not just past interventions, but all potential future interventions), then we consider $c$ to be the system’s 1ST order self. If $c\in\mathfrak{p}_{\mathfrak{o}}$ then an organism has constructed a 1ST order self. A 1ST order self for an organism $\mathfrak{o}$ is denoted $\mathfrak{o^{1}}$. An organism has at most one 1ST order self.
(intuitive summary) Intuitively, $\mathfrak{o^{1}}$ is where we draw the line between what the organism can intend and what it cannot. It is conceivable we might have two “organisms” in the same body by this definition, each with its own 1ST order causal identity. Ultimately, where an organism begins or ends remains malleable.
Definition 14 (preconditions).
If $\mathfrak{o}$ is an organism, and $c$ is a causal identity, the $\mathfrak{o}$ will construct $c$ only if the representation and incentive preconditions below are met:
  • the scale precondition is met iff $c\in L_{\mathfrak{v_{o}}}$, and
  • the incentive precondition is met if $\mathfrak{o}$ must learn $c$ to remain “fit”.
(intuitive summary) If $c$ is a 1ST order self, then these are the preconditions that must be met for an organism to construct $c$. Likewise, any other sort of causal identity.
Definition 15 (chain notation).
Suppose we have two organisms, $\mathfrak{a}$ (Alice) and $\mathfrak{b}$ (Bob). $c^{\mathfrak{b}}_{\mathfrak{a}}$ denotes a causal identity for $\mathfrak{b}$ constructed by $\mathfrak{a}$ (what Alice thinks Bob intends). Subscript denotes the organism who constructs the causal identity, while superscript denotes the object. The superscript can be extended to denote chains of predicted causal identity. For example, $c^{\mathfrak{ba}}_{\mathfrak{a}}\subset c^{\mathfrak{b}}_{\mathfrak{a}}$ denotes $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a^{1}}$ (what Alice thinks Bob thinks Alice intends). The superscript of $c^{\mathfrak{*}}_{\mathfrak{a}}$ can be extended indefinitely to indicate recursive predictions; however the extent recursion is possible is determined by $\mathfrak{a}$’s vocabulary $\mathfrak{v}_{\mathfrak{a}}$. Finally, Bob need not be an organism. Bob can be anything for which Alice constructs a causal identity.
[…]
An $n^{th}$ order self for $\mathfrak{a}$ is $\mathfrak{a^{n}}=c^{\mathfrak{*a}}_{\mathfrak{a}}$ where * is replaced by a chain, and $n$ denotes the number of reflections. For example, a 2ND order self $\mathfrak{a^{2}}=c^{\mathfrak{ba}}_{\mathfrak{a}}$, and a 3RD order self $\mathfrak{a^{3}}=c^{\mathfrak{baba}}_{\mathfrak{a}}$. We use $\mathfrak{a^{2}}$ to refer to any 2ND order self, and chain notation to refer to a specific 2ND order self, for example $c^{\mathfrak{ba}}_{\mathfrak{a}}$. The union of two $n^{th}$ order selves is also considered to be an $n^{th}$ order self, for example $\mathfrak{a^{3}}=c^{\mathfrak{baba}}_{\mathfrak{a}}\cup c^{\mathfrak{dada}}_{\mathfrak{a}}$, and the weaker or higher level a self is in the generational hierarchy, the more selves there are of which it is part.
Definition 17 (stages of consciousness).
We argue the following stages by scaling the ability to learn weak policies:
  1. 1.
    Hard Coded: organism that acts but does not learn, meaning $\mathfrak{p}_{\mathfrak{o}}$ is fixed from birth.
  2. 2.
    Learning: an organism that learns, but $\mathfrak{o^{1}}\not\in\mathfrak{p}_{\mathfrak{o}}$ either because $\mathfrak{o^{1}}\not\in L_{\mathfrak{v_{o}}}$ (failing the “scale precondition”) or because the organism is not incentivised to construct $\mathfrak{o^{1}}$ (failing the “incentive precondition”).
  3. 3.
    1ST order self: reafference and phenomenal or core consciousness are achieved when $\mathfrak{o^{1}}\in\mathfrak{p}_{\mathfrak{o}}$ is learned by an organism because of attraction to and repulsion from statements in $L_{\mathfrak{v}_{\mathfrak{o}}}$.
  4. 4.
    Second order selves:
    1. (a)
      access or self-reflexive consciousness is achieved when $\mathfrak{o^{2}}\in\mathfrak{p}_{\mathfrak{o}}$.
    2. (b)
      hard consciousness is achieved when a phenomenally conscious organism learns a 2ND order self (an organism is consciously aware of the contents of 2ND order selves, which must have quality if learned through phenomenal conscious).
  5. 5.
    Third and higher order selves: meta self-reflexive consciousness (human level hard consciousness) is achieved when $\mathfrak{o^{3}}\in\mathfrak{p}_{\mathfrak{o}}$.
Proposition
An organism that uses weakness as its proxy will learn an $n^{th}$ order self if the incentive and scale preconditions are met for that order of self.
Proof sketch. Assume we have an organism $\mathfrak{a}$ that learns using “weakness” as a proxy. A $\mathfrak{v}_{\mathfrak{a}}$-task $\mathfrak{h}_{<t_{\mathfrak{a}}}$ represents the history of $\mathfrak{a}$ (meaning $\mathfrak{h}_{<t_{\mathfrak{a}}}\sqsubset\mu_{\mathfrak{a}}$ and $\mathfrak{h}_{<t_{\mathfrak{a}}}$ is an ostensive definition of $\mu_{\mathfrak{a}}$, because $\mathfrak{a}$ remains alive). The organism explores the environment, intervening to maintain homeostasis. As it does so, more and more inputs and outputs are included in $\mathfrak{h}_{<t_{\mathfrak{a}}}$. It follows that:
  1. 1.
    From the scale precondition we have that there exists a $n^{th}$ order self $\mathfrak{a^{n}}\in L_{\mathfrak{v}_{\mathfrak{a}}}$.
  2. 2.
    To remain fit, $\mathfrak{a}$ must “generalise” to $\mu_{\mathfrak{a}}$ from $\mathfrak{h}_{<t_{\mathfrak{a}}}$. According to the incentive precondition, generalisation to $\mu_{\mathfrak{a}}$ requires $\mathfrak{a}$ learn the $n^{th}$ order self, which is when $\mathfrak{a^{n}}\in\mathfrak{p_{\mathfrak{a}}}$.
  3. 3.
    From [34, prop. 3] we have proof that weakness is the optimal choice of proxy to maximise the probability of generalisation from child to parent is the weakest policy. It follows that $\mathfrak{a}$ will generalise from $\mathfrak{h}_{<t_{\mathfrak{a}}}$ to $\mu_{\mathfrak{a}}$ given the smallest history of interventions with which it is possible to do so (meaning the smallest possible ostensive definition, or cardinality $\lvert D_{\alpha}\rvert$).
Were we to assume learning under the above conditions does not construct an $n^{th}$ order self for $\mathfrak{a}$, then one of the three statements above would be false and we would have a contradiction. It follows that the proposition must be true. $\square$
3 Back to Foundations
Rather than presupposing local states or an abstraction layer, we must start from very basic first principles to formalise all conceivable environments [32, 43] in definition 1. Where there are things, we call those things an environment. Where things differ, we have different states of that environment. Hence, we begin by formalising the environment as a set of contentless global states $\Phi$. We don’t assume there is any internal structure to the states $\Phi$ contains, but rather define declarative “programs” in terms of relations between these irreducible, contentless states. The set $P$ contains all such programs. The powerset of $P$ is every aspect of the environment, because every aspect of any conceivable environment must be set of such declarative programs, which is a subset of $P$.
Axiom 1: When there are things, we call these things the environment.
Axiom 2: Where things differ, we have different states of the environment.
Universality Claim: Axioms 1 and 2 hold for every conceivable environment.
We don’t need to concern ourselves with the internal structure of environment states beyond this. Declarative programs return true or false, so a declarative program is a “fact” if it is true, and the current state of the environment is the set of all “facts”. Truth is determined with respect to states. If time is one way in which things differ, then there is only one state at a time (dimensions are implicit). The only meaning these programs have is how their truth values relate in different environmental states. Hence this is a representationless form of pancomputationalism [30].
3.1 Natural Selection and Embodiment
We then assume a process of natural selection. This produces embodied organisms.
A set of facts can represent anything [^8] Some may object, pointing out that this ignores composition. However, the application of a function is a fact regardless of whether it takes another function as input., so the body of an organism (like everything else in every conceivable environment) must be a set of declarative programs. The vocabulary can be seen as hardware and software operating together, avoiding issues related to solipsism [46]. No two organisms have the same vocabulary, because that would mean they are the same body.
However, it is not enough to just say a body is a set of declarative programs. Bodies tend to imply finite resource constraints, so we adopt one more axiom from [43].
Axiom 3: All aspects of the environment are spatially extended [43].
This means that a body can occupy only a finite number of states, which is formalised in definition 2. A body is a vocabulary $\mathfrak{v}$ with finitely many elements.
The vocabulary of an abstraction layer is a subset of the aforementioned $P$ in definition 1. $\mathfrak{v}$ implies a formal language $L_{\mathfrak{v}}$ of interaction between body and environment. A statement in this formal language is just a set of “programs”, which is an interaction, allowing us to simultaneously discuss cognition in enactivist and computational terms. Statements have truth conditions with respect to environmental states, and every statement has an extension. The extension of a statement is the set of all statements which are supersets of the first statement (the set of all other statements by which the first statement is implied). Extension is important we can relate statements by their truth conditions, forming a lattice. Most importantly, we avoid the distinction between software and hardware and so avoid computational dualism [^9] Computational dualism [32] is the distinction between software mind and hardware body commonly used in artificial intelligence. Computational dualism is a simplification that has led to erroneous claims regarding the behaviour of artificial intelligence [46]. Software is a state of hardware, not a distinct object. What software does depends entirely upon the hardware that interprets it..
3.1.1 Self-Organising Systems as Self and World Constraints
Both snowflakes and human bodies are self-organising systems, yet only the latter are regarded to display conscious experiences. What exactly in the self-organisation model of a body make it radically different from the snowflake? To tackle this question, we need to introduce the notion of self- and world constraints.
Given that we have an environment, and an abstraction layer that implies an embodied formal language, we can talk about computation with inputs and outputs by just treating everything as embodied statements embedded and enacted within the pancomputational environment. Given an input $i$, the set of all possible outputs is the extension $E_{i}$ of that input. This is because if $i$ is realised by the environment, then the environment is constrained to only those states that realise $i$, which constrains what other statements can be realised. We can use this fact to talk about “policies” as embodied [^10] As a state of hardware rather than software, which is how the term policy is normally used in reinforcement learning. constraints on behaviour. A policy is a statement whose extension constrains outputs, like a causal intermediary in machine functionalism [47]. If behaviour is “motivated”, then statements have valence determined by natural selection and only a subset of the statements a body could make are “fit”. As a result, an organism will self-organise to express some statements, but not others. This is formalised by the $\mathfrak{v}$-task as in definition 3, so called because a task exists in the context of a vocabulary $\mathfrak{v}$. If $\mathfrak{v}$ is the vocabulary of a body, and $\mathfrak{v}$-task $\mu=\langle I_{\mu},O_{\mu}\rangle$ is fit behaviour, then $I_{\mu}$ is all the statements that body can express in which it is possible to remain fit, and $O_{\mu}$ is all the statements that body can express in which it remains fit. The extension $E_{I_{\mu}}$ of $I_{\mu}$ would be every output that it is possible to choose given the inputs $I_{\mu}$, but only a subset of those $O_{\mu}\subset E_{I_{\mu}}$ are correct outputs. In other words, this defines a self-organising behaviour as a constraint on outputs, given inputs. Conversely, a system which considered all outputs to be correct would not be self-organising. When we say a system is self-organising, mean it is goal directed and can adapt to serve that goal, which for the purposes of this paper is the fundamental goal of self-survival and self-reproduction.
3.1.2 Inference
Every statement in $L_{\mathfrak{v}}$ implies a constraint, because there are only so many outputs that can be expressed by a body at the same time as any given statement. That is what the extension of a statement represents. If $i$ is a statement which is true, then the only possible statements a body can express is $E_{i}$. As a result, a body might be express a statement $\pi$ (meaning $\pi$ is true), and $\pi$ would then constrain the body to only correct outputs $O_{\mu}\subset E_{I_{\mu}}$ if $O_{\mu}=E_{I_{\mu}}\cap E_{\pi}$. We call a constraining statement a policy. A policy constrains outputs given inputs. A correct policy is one that constrains outputs to only correct outputs. For the sake of intuition, think of “correct” as “fit” according to natural selection (although this need not always be the case). This is more formally expressed in definition 4.
“The best model of the world is the world itself” - Rodney Brooks [48]
Importantly, to reiterate, one does not need a model of the world; one must embody a policy that constrains behaviour to only fit behaviour.
The history of an organism’s interactions is a subset of the statements that can be made in the abstraction layer; hence the history of an organism is also a $\mathfrak{v}$-task. The history of the organism that remains alive is a subset of fit behaviour for that organism. Examples of “fit” self-organising behaviour. Hence, we could denote an organism’s history to be a $\mathfrak{v}$-task $\mathfrak{h}$, where each input and output in that task is behaviour the organism has exhibited and everything that was involved in that behaviour (the interaction between organism and environment). Fit behaviour would be another task $\mu$, and if the organism remains fit then $\mathfrak{h}\sqsubset\mu$ [^11] One task is a child of the other, meaning the inputs and outputs are subsets of the parents’ inputs and outputs..
3.1.3 Learning
Learning, in computational terms, tends to be understood as the process of modelling the program which caused data. Constructing a “world” model. Here, we just need to explain how an organism gets $\mu$ from $\mathfrak{h}$. We don’t need a world model, but a way to constrain outputs to just those that are fit. Like everything else, that constraint is once again a statement in the formal language $L_{\mathfrak{v}}$.
Those examples in $\mathfrak{h}$ could have been generated by any one of a set of policies $\Pi_{\mathfrak{h}}$. These are not representations of possible causal intermediaries in the machine functionalist sense, but just parts of the outputs from which the outputs could be derived given the inputs. Every $\mathfrak{v}$-task (and thus every self-organising system) implies such a set of policies, and not all of those policies will constrain outputs the same way given new inputs. Some policies will “generalise” to imply fit behaviour in unfamiliar circumstances, meaning correct outputs given a new set of inputs (a parent task of the organism’s history $\mathfrak{h}$). If those policies imply fit behaviour given new inputs, then the organism will remain fit. If an organism learns, then it is adapting to embody a policy that generalises in accord with its motives (which, set by natural selection, will tend to favor fit behaviour).
“I survive therefore my model is viable.” - Mark Solms [49]
According to previous experimental and mathematical results, the optimal strategy to learn and adapt as fast as possible [^12] Meaning to converge on a fit policy given the smallest history possible, to realise fit behaviour. is to prefer “weaker” policies, meaning those with larger extensions [34]. Formally, the policy which generated outputs $O$ given inputs $I$ is most efficiently identified by constructing a policy $\pi$ such that $\pi$ generates $O$ from $I$, and $\pi$ implies the weakest constraint that can be implied while still generating $O$ from $I$. This maximises adaptability, because it allows a self-organising system to construct fit policies from a shorter history [35]. This is described in formal terms in definition 5.
4 Relevance Realisation Through Causal Learning
The enactive process of relevance realisation can be framed as policy learning. Natural selection prefers more adaptable organisms, so we assume it optimises for organisms that optimise for weaker policies. We call this weak policy optimisation (WPO). Policies determine how inputs are mapped to outputs. Interpretation. Fit policies must correctly predict causes of valence. Hence by constructing fit policies, an organism must realise what is relevant. A weaker policy implies all the more specific versions of itself, meaning those that more tightly constrain outputs by having a smaller extension. Hence policy learning implies a lattice of policies that vary in weakness. Every fit policy is a “causal identity” for something, for example an object like a “food” or a more abstract concept like “pain”.
For convenience, we’ll now expand upon the concept of organism given in definition 6. All organisms must have preferences (will make some decisions and not others), regardless of how those arise. All organisms must have policies that reflect those preferences, and every policy implies tasks, so we can define preferences as an binary relation over tasks. Correct or incorrect choice of policy affects the organism’s existence and survival.
We also need to introduce the notion of interaction between organisms. Organisms interact when they “affect” one another (in the physical rather than psychological sense of the word). This is formally defined in 9.
Now, artificial intelligence and machine learning [50, 51, 52, 53] are concerned with engineering adaptive agents. In that context, causality has now become a mainstream topic of research. Causal learning is demonstrably necessary to thrive in an interactive setting [45, 54]. Where the causal graph is known in advance (for example if we were to be measuring the efficacy of medical interventions), this issue could be resolved by the use of causal language (such as “Do Calculus” [55]) to represent an intervention from outside the system the graph describes. Such an intervention represent agency, in that an organism observing the environment draws conclusion about that environment, and then intervenes to change the environment. To illustrate this point, consider the following example [35].
Suppose an organism named Bob is attempting to learn and predict the environment. Now assume Bob has observed Alice wearing a raincoat only when it rains, and that Bob has observed rain only at those times when Alice has been observed wearing a raincoat. If we represent the raincoat observation with a binary variable $O\in\{true,false\}$ and the advent of rain with a similar variable $R\in\{true,false\}$, and if Bob draws conclusions according to Bayesian probability, then Bob’s observations will lead to the conclusion that $p(R=true\mid O=true)=1$. This means Bob believes that if Alice wears a raincoat, then it must be raining. Now lets permit Bob to interact with its environment. Assume Bob wants it to rain. Based on the belief $p(R=true\mid O=true)=1$, Bob may conclude that forcing $O=true$ by holding a gun to Alice’s head and demanding Alice wear a raincoat will cause it to rain. This is obviously absurd. $O=true$ does not represent the event $q=$“I coerced Alice into wearing a raincoat”, but an entirely different event $v=$“Alice decided to put on a raincoat for the same reason I have observed Alice wearing a raincoat in the past”. To accurately represent the environment, we need a way of representing that $p(R=true\mid q)=p(R=true)\neq p(R=true\mid v)=1$, meaning we need a way to represent $q$ and $v$. To illustrate the problem visually, we started with the acyclic graph
$O$$R$
and our intervention disconnected rain from the choice of clothing:
$O$$R$
This can be resolved by introducing a “do” operator that we apply to a variable $O$ to obtain $do[O=o]$, to represent the fact that an agency from outside the system has intervened to assign a value to $O$, so that we can represent $p(R=true\mid do[O=true])=p(R=true)\neq p(R=true\mid O=true)=1$. Thus, the aforementioned $q$ is equivalent to $do[O=true]$, while $v$ is equivalent to $O=true$.
This works very well for the purpose of evaluating treatments and interventions by humans. However, this merely establishes that we need to account for causality in cognition. It does not explain how one would come to know all the objects involved (to represent them as variables), or how they relate to one another causally.
The inference of causal relations can be understood in two steps. First, we must show that the “do” operator is equivalent to including additional variables in a causal graph. Second, we must show how it is possible to learn the objects that best represent cause and effect relations. The matter of which cause and effect relations are learned is determined by valence, and so the objects learned are statements classifying causes and valence.
Substituting The “Do” Operator
The very notion of an “intervention” from outside a system echoes mind-body dualism, in that it treats an organism performing an intervention as something apart from the system in which it intervenes (its environment). To illustrate what this means, consider another example involving the variables $O$ and $R$ from before.
Assume we again have Bob who constructs a causal graph of the environment. Assume Alice exists in that environment. From Bob’s perspective, Alice is just a part of the environment represented by a variable $A$ in Bob’s causal graph.
$O$$A$$R$
Now assume Bob observes Alice taking an action that changes an aspect of the environment, represented by the variable $O$ (for example, Bob observes Alice putting on the raincoat). From the perspective of Bob, Alice’s action is the assignation of a value to a variable $A$, and to the variable $O$. There is no need to involve a do operator in this scenario because we can already represent that $p(R=true\mid A=x,O=true)=p(R=true)\neq p(R=true\mid A=y,O=true)=1$ (because Alice is part of the causal graph). This raises the question; if we do not require a $do$ operator to represent the actions of Alice, then why would we need it to represent an intervention by Bob? Couldn’t we represent the same information by introducing a new variable? The answer is yes [35, 56].
$O$$B$$A$$R$
Of course, this does not solve the problem of how the causal graph is learned in the first place, and this is where relevance realisation comes in.
5 Relevant Causal Identities
The extension of fit behaviour is an extensional definition of what an organism is compelled to want by natural selection. Likewise, an intension of fit behaviour would be any policy that results in fit behaviour, delivering the organism what it is compelled to want. Indeed, this was how the formalism of tasks we employ originated, to formalise the inference of norms [57] and explain why some abstractions are formed, and not others.
As mentioned earlier in the context of heuristics or “proxies” for learning, the optimal choice of policy for generalisation is the weakest. Such a policy must, by definition, isolate those things which cause valence. Put another way, “fitness” in an interactive setting is likely to require an organism be able to distinguish passive observation of an event from having intervened to cause that event (reafference [36, 37]). A preference for weak hypotheses facilitates this via construction of causal identities. This divides the environment up into objects, events and anything else which is sufficiently relevant to the organism’s motivations.
To explain how, we must first explain how intervention can be differentiated from observation without presupposing variables, and then the circumstances under which it will be differentiated [35].
5.1 Causal Learning
In accord with [42, def. 6] we assume a vocabulary $\mathfrak{v}_{\mathfrak{a}}$ belonging to an organism we’ll denote $\mathfrak{a}$. A “cause” in the context of this formalisation is not a variable but a statement $l\in L_{\mathfrak{v}_{\mathfrak{a}}}$ in that formal language. The raincoat example would involve $obs,rain\in L_{\mathfrak{v}_{\mathfrak{a}}}$ such that:
[…]
$int$$obs$$rain$
In other words, so long as $c\neq int$, the intervention can be differentiated from the passive observation. We formally define this in definition 10.
This being the case, any set $c\subseteq int-obs$ could be used to identify the party undertaking the intervention, which is why $c$ is referred to as a “causal identity”. It distinguishes the intervention $int$ from the passively observed effect $obs$, like reafference in living organisms. However, the above only considers one intervention. A weaker or more general causal identity would be one that is shared by more intervention. For example, we might have two different interventions $a_{1}$ and $a_{2}$, with the observed effects $c_{1}\subset a_{1}$ and $c_{2}\subset a_{2}$, and causal identities $a_{1}-c_{1}=i_{1}$ and $a_{2}-c_{2}=i_{2}$. If $i_{3}=i_{1}\cap i_{2}\neq\emptyset$, then $i_{3}$ is a causal identity that is present in two interventions.
Being a question of classification, we can express a causal identity in relation to a $\mathfrak{v}$-task, for which the causal identity is a policy (expanding upon [35, def. 10]), giving us definition 11.
Example 1.
Suppose we have organisms $\mathfrak{a}$ (Alice) and $\mathfrak{b}$ (Bob), and that the inputs Alice has experienced so far are $I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}$. These can be divided into those in which Bob affected Alice $S^{\mathfrak{b}}_{\mathfrak{a}}$ and those in which Bob did not $S^{\lnot\mathfrak{b}}_{\mathfrak{a}}=I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}-S^{\mathfrak{b}}_{\mathfrak{a}}$. By affecting Alice, Bob has intervened in Alice’s experience. Alice can construct a causal identity $b$ for Bob corresponding to interventions $INT=S^{\mathfrak{b}}_{\mathfrak{a}}$ and observations $OBS=S^{\lnot\mathfrak{b}}_{\mathfrak{a}}$.
5.2 Ascribing Intent To Other Objects
The “do” operator is necessary to discern the difference between an event one has caused, and that same event passively observed. However, what is passive observation if not the result of an intervention by something other than oneself? The distinction between “intervention” and not is misleading. Everything is an intervention. The question is not “is this an intervention” but “by whom was this intervention made?”.
To illustrate this point we return to Alice, Bob and her raincoat. Earlier, we arrived at the following graph in which Bob’s intervention was given by $int$.
$int$$obs$$rain$
However, what if a third person Larry puts the coat on Alice? Surely Bob can observe this, and so Bob’s observation of Larry’s intervention is $v\in L_{\mathfrak{v}_{\mathfrak{a}}}$ such that $obs\subset v$. To account for this, Bob can construct a causal graph as below (with $b.$ representing Bob and $l.$ representing Larry).
$b.\ int$$l.\ int$$obs$$rain$
Bob’s causal identity for himself $c_{b}\subset b.\ int-obs$ only represents the intervention by himself. However, now we can see that Bob must also construct a causal identity $c_{l}$ for Larry, where the $c_{l}\subset l.\ int-obs$. More generally, for an organism $\mathfrak{a}$ with sensorimotor language $L_{\mathfrak{v}_{\mathfrak{a}}}$ to construct a causal identity for an object $\mathfrak{b}$, it must first be the case that $\mathfrak{a}$ is affected by $\mathfrak{b}$ [42], to satisfy the incentive precondition for causal identity. It depends on motive, or valence. Recall that to be affected is formally defined in definition 9.
Assume an organism $\mathfrak{a}$ is affected by $\mathfrak{b}$ given inputs $INT$, and not affected given inputs $OBS$. To then attribute the contents of $INT$ to one specific entity, there must be something in common between the members of $INT$ caused by $\mathfrak{b}$ that is not shared by any member of $OBS$ caused by something else (in other words it must be at least possible for $\mathfrak{a}$ to discern the existence of $\mathfrak{b}$). The contents of $INT$ are “interventions” by $\mathfrak{b}$ and by learning $c$, a corresponding causal identity, $\mathfrak{a}$ can discern the existence of $\mathfrak{b}$. This is not to say that $\mathfrak{b}$ decides (chooses an output) or is even alive. What is important is that $\mathfrak{b}$ affects $\mathfrak{a}$, making it possible to discern when these interventions are a consequence of $\mathfrak{b}$’s existence.
There are certain preconditions for the existence of a causal identity corresponding to $INT$ and $OBS$.
5.3 Preconditions
First, the vocabulary $\mathfrak{v}_{\mathfrak{a}}$ of an organism $\mathfrak{a}$ must be large and expressive enough to ensure that observations are distinguishable from interventions (in other words, the causal identity must be a subset of the vocabulary). We must have sufficient scale. Second, it must be in the organism’s interest to make this distinction.
Inference is only possible if some states are preferable to others. It is a value judgement. As Hume pointed out, one cannot derive what “ought” to be from a statement of what “is”. Natural selection provides a notion of what ought to be, by eliminating anything which ought not.
  1. 1.
    The scale precondition requires $\mathfrak{v}$ contain the causal identity.
  2. 2.
    The incentive precondition is that fitness demands the causal identity.
5.4 Realising Lower Order States And Higher Order Meta Representations
Causal identities are not constrained to other organisms, but we use other organisms for intuition. This is because the causal identity for an object is a policy predicting its behaviour. This makes intuitive sense when that object is a self-organising system with goals. However, it also applies to inanimate objects like rocks. A rock still has behaviour, but it does not have intent. Yet by constructing a policy representing the rock’s intent, one may predict what the rock will do (e.g. damage a tree when thrown). That ability to predict is what matters. The weaker the causal identity, the more pervasive the “identity” in the sense of being part of more interventions. Learning is not just about constructing policies but joining them together into weaker, more abstract policies. By learning policies that correctly identify the causes of valence at different levels of abstraction, the organism engages in relevance realisation.
Importantly, embodiment imposes severe limitations in the form of a finite vocabulary. In definition 6 we formalise this for the sake of explanation, formalising an organism $\mathfrak{o}$ with a set of policies the organism has realised. Each and every policy implies a $\mathfrak{v}$-task. A $\mathfrak{v}$-task is a triadic relation between inputs, outputs and policies which resembles Peircean semiosis [58, 42] of sign, referent and interpretant. Hence, in definition 7 we formalise this as a “protosymbol” [^13] Proto because it is something more primitive than a symbol as conceived of by Peirce. system $\mathfrak{s}_{\mathfrak{o}}$ for the organism $\mathfrak{o}$. Interaction is defined in terms of choosing an output, which means there is inevitably a policy and a policy implies a $\mathfrak{v}$-task, so definition 6 includes a preference order over tasks, which is used in definition 8 to formalise the interpretation of inputs (in terms of constraints, rather than an algorithm). In particular, the organism must always act according to an interpretation, and some interpretations imply others. Related as they are in a lattice, protosymbols are analogous lower order states and higher order meta representations. Tasks exist in a “generational hierarchy”. They are not mutually exclusive. Higher level tasks are more general, and have fewer policies because only very weak policies could complete them. Some have framed consciousness as a problem of moving from unary, to dyadic, to triadic relations [59]. Similarly, we have gone from unary states, to dyadic declarative programs, to triadic tasks and protosymbols.
The question now is when does an organism realise a causal identity for when it causes valence? When does it construct a causal identity for itself?
6 Multi-Layered Self-Organisation
In an interactive setting, it is necessary to correctly predict the effect of one’s actions [45] to satisfy even very basic goals. As capacity for weak policy optimisation (WPO) scales, a greater variety of concepts can be learned. Hence as WPO scales, progressively higher orders of ‘causal-identity’ for one’s self-related information processing are constructed [35]. This lets us frame the construction of embodied selves in developmental [60] and evolutionary terms. We turn to this discussion now.
6.1 The First Order Self
A first order self (1ST henceforth, see figure 1) is functionally equivalent to reafference, which has been observed in the mammalian mid-brain and insect central cortex [37, 38]. Reafference lets an organism discern the consequences of its actions, and so a 1ST order self is necessary for accurate inference in an interactive setting in the same sense as Pearl’s ‘do’ operator [45].
A 1ST order self formally defined in definition 13. It serves as the locus of self-related information processing and experience [35], allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction. But also, to anticipate future planned interactions (for example, an insect navigating its environment [38]), and recollections of past interactions (it is a subset of all those relevant “statements” in the formal language).
Figure 1: Visual intuition for a 1ST order self. The organism constructs a causal identity for itself and can relate that to observed events.
Natural selection prefers efficiency. The absence of a single, centralised causal identity can create inefficiencies [^14] The effect of insufficiently weak representations can be observed in large language models like GPT-4 [42, 61], where they fail to make consistent statements about an object because represent it as multiple objects.. Decentralised or asynchronous control might be advantageous in some circumstances (e.g. an ant colony), but not others (e.g. in an individual human body). An organism that has a distributed control system might have a “self” for each part of that system, and perhaps the co-ordination of parts might seem to suggest the existence of a centralised “self”. Redundancy might be useful in some circumstances, but there is also a cost - more data may be required to learn and adapt, because the optimal choice of policy for accurate generalisation is the “weakest” [34].
If the scale and incentive preconditions are met for a causal identity for one’s self, then the organism must realise one. This might be partly hard-coded, but our focus is on inference by the organism itself. Preconditions are formally defined in 14.
6.2 The Second Order Selves
Survival may demand organism $\mathfrak{a}$ infer $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s interventions (to see one’s self as if through another’s eyes [42]). This is called a second order self (2ND henceforth). We argue that if access conscious contents are available for communication in the human sense, then they must be communicable in the Gricean sense [62, 63]. Grice argued that communication is about the inference of intent. If person $\mathfrak{a}$ and $\mathfrak{b}$ are talking, then the meaning $m_{\mathfrak{a}}$ of what $\mathfrak{a}$ says is whatever $\mathfrak{a}$ intends $\mathfrak{b}$ understand. The meaning $m_{\mathfrak{b}}$ that $\mathfrak{b}$ understands is whatever $\mathfrak{b}$ thinks $\mathfrak{a}$ wants $\mathfrak{b}$ to think. $\mathfrak{b}$ has understood what $\mathfrak{a}$ means if $m_{\mathfrak{b}}$ approximates $m_{\mathfrak{a}}$. This can happen only if $\mathfrak{a}$ can predict with reasonable accuracy what $\mathfrak{b}$ thinks $\mathfrak{a}$ thinks, and $\mathfrak{b}$ can predict what $\mathfrak{a}$ thinks $\mathfrak{b}$ will think upon hearing an utterance. In other words, both $\mathfrak{a}$ and $\mathfrak{b}$ must have 2ND order selves that are good approximations. Yes, there are other aspects to communication.
However, here we are talking about consciousness. Access conscious contents are those available for reasoning and report. It follows [^15] If the definition of access consciousness is to be consistent with reasoning and report as exhibited by conscious humans. that access conscious contents must in principle be communicable in the sense Grice described.
As such, we argue contents available for communication can only be the contents of 2ND order selves, which means only an organism with 2ND order selves can be considered to have access consciousness. 2ND order selves also explain attention and self-awareness. An organism can have many 2ND order selves because they depend upon who or what the organism is interacting with, just as the availability of information depends on context.
Intuitively, where a 1ST order self might allow one to observe a cat and form plans regarding causal interactions with the cat, a 2ND order self would allow one to be consciously aware of the cat for the purpose of reasoning and report. One can know of the cat, and that another organism knows of the cat, but a 2ND order self is insufficient to be aware that one is aware of the cat [^16] This is ‘meta-self-reflexive consciousness’ as some have described it [39]..
Figure 2: Visual intuition for a 2ND order self. The organism constructs a causal identity for itself, and for another object, and the causal identity for the other object includes a prediction of the organism itself from that object’s perspective. This would occur, for example, if one was a predator trying to predict the movements of prey in response to one’s own actions. It is an extension of the 1ST order self.
More formally using the notation given in definition 15, assume $\mathfrak{a}$ and $\mathfrak{b}$ are organisms that evolved to accurately predict one another’s behaviour. Assume $\mathfrak{a}$ constructs a causal identity $c^{\mathfrak{b}}_{\mathfrak{a}}$ to predict $\mathfrak{b}$ given input $i_{\mathfrak{a}}\in I_{\mu_{\mathfrak{a}}}$, of which a second order self $c^{\mathfrak{ba}}_{\mathfrak{a}}$ is part. Likewise, $\mathfrak{b}$ constructs $c^{\mathfrak{a}}_{\mathfrak{b}}$ to predict $\mathfrak{a}$ given input $i_{\mathfrak{b}}\in I_{\mu_{\mathfrak{b}}}$, of which $c^{\mathfrak{ab}}_{\mathfrak{b}}$ is part. What is important here is that each organism’s intent is to some extent inferred by the other, and that fact inevitably changes the sorts of policies that will be “fit”. For example, second order self means each knows the other can anticipate manipulation, which means the optimal policy will often be to have rather than feign intent that aligns to some extent with the other party’s desires, to co-operate [64] [^17] Depending upon circumstances, for example organisms may co-operate in some circumstances but not others [42], and transient relations, information asymmetry and other factors can make deceit a more attractive option.. Repeated interaction creates an iterated prisoner’s dilemma, incentivising co-operation and signals that both parties interpret similarly (the beginnings of language) [42]. To communicate in Gricean terms, $\mathfrak{a}$ must intend to convey meaning $m_{\mathfrak{a}}$, and $\mathfrak{b}$ must recognise this intent. The incentive precondition explains why $\mathfrak{a}$ would form such intent (co-operation is often advantageous), while how may be understood as follows:
  • $c^{\mathfrak{ba}}_{\mathfrak{a}}$ lets $\mathfrak{a}$ predict what $\mathfrak{b}$ will come to believe when it observes $\mathfrak{a}$’s behaviour.
  • $c^{\mathfrak{ab}}_{\mathfrak{b}}$ then lets $\mathfrak{b}$ predict what $\mathfrak{a}$ intends that $\mathfrak{b}$ believe.
$\mathfrak{a}$ can use $c^{\mathfrak{ba}}_{\mathfrak{a}}$ to infer behaviour to which $\mathfrak{b}$ will ascribe the intent to communicate $m_{\mathfrak{a}}$, and $c^{\mathfrak{ab}}_{\mathfrak{b}}$ lets $\mathfrak{b}$ infer that this is what $\mathfrak{a}$ intends. The “utterance” Grice refers to is how $\mathfrak{a}$ affects $\mathfrak{b}$ in accord with earlier definitions. Put another way, $\mathfrak{a}$ encodes $m_{\mathfrak{a}}$ into its behaviour in a manner that $\mathfrak{b}$ can decode (their respective second order selves act as encoders and decoders). By encode and decode, we mean a loose approximation of $m_{\mathfrak{a}}$ is communicated. There are of course shortcuts, for example of $\mathfrak{a}$ and $\mathfrak{b}$ are of the same species then they likely have similar motives and experiences, and so the efficient thing for each to do would be to use its own intent as an approximation of what the other might think. However, that does not obviate the need for second order selves, it just makes such things easier to realise.
6.2.1 The Third Order Selves
We can continue scaling WPO indefinitely. 3RD and higher order selves can explain function. For example, meta self-awareness [39] is the awareness that one is self-aware. If self-awareness stems from 2ND order selves, then it follows that meta self-awareness requires 3RD order selves. Formally a third order self for $\mathfrak{a}$ reflecting off $c^{\mathfrak{ba}}_{\mathfrak{a}}$ lets $\mathfrak{b}$ is $c^{\mathfrak{baba}}_{\mathfrak{a}}$. It is $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$.
Figure 3: Visual intuition for a 3RD order self. The organism constructs a 2ND order self for its 2ND order self, and so becomes aware that it is aware.
7 The What and Why of Consciousness
Up to now we have developed the conceptual toolbox that one can use to dissolve the hard problem. We started with the basic observation that self-organizing systems such as the human bodies constantly process information to maintain oneself in the face of constant change both inside and outside one’s body. An organism does not spring into existence understanding number systems or objects, but constructs them through constant interactions with the environment. One can regard them as policies according to which the environment is interpreted. To learn how to interpret the environment an organism must differentiate between states, must react to change, and learn policies according to the valence associated with that change.
As we saw earlier, learned policy has valence, and it is a classifier of inputs. Those inputs are “information” in the mechanistic sense of the environment being in one state and not another, in the sense of axioms 1 and 2 (see section 3). Hence, a policy is a classifier of information, but that information is not in a language, and it is not yet something labelled or quantified (until there is a policy that labels or quantifies).
Clearly, prelinguistic self-organising system must classify and attach value and disvalue to states and anticipated states to prioritise and make decisions. To learn to label and classify information an organism must sense that “something” has changed, and motivated by valence construct a policy classifying that “something”. In the absence of language, the difference between two states can only be qualitative.
The bold claim here is that information processing at 1ST level biological self-organising systems such as the human body is necessarily qualitative by the very virtue of existing and experiencing (i.e. exploring one’s body and environment). This fundamental, basic way of engaging with the environment is what we call experience (see Ciaunica in prep.). Note that this is different from experience in the sense of experiential (i.e. phenomenal) as usually defined in the literature, which we take to be the second-order self-perception.
Importantly, experience in the sense of experiential content cannot exist without experience in the sense of embodied exploration of the body and world as we defined here. To put it provocatively quality precedes quantity, and quantity is nothing more than the interpretation of quality. Quality comes first and experiences should be regarded on a continuum rather than a switch on/ switch off phenomena. All living systems experience the world through their bodies and as such, there is something what it is like to experience the world in that basic way (even when one is asleep or a baby). One can access those phenomenal, experiential aspects at a higher level, true, but by accessing them, it doesn’t mean that one ‘constructs’ consciousness or one becomes a conscious being. One is already consciously experiencing the world before one can explicitly access one’s own experiences as a 2ND order selves.
It follows that every policy learned in this way must classify a quality. Hence every such policy is a local state. A causal identity imbued with meaning by valence. There would be a policy for one’s act of smelling coffee. For perceiving one’s friend. There would be something different it is like to interact with a hostile version of that very same friend. The 1ST order self-accompanies everything an organism does. It has a quality, so the 1ST order self is “what it is like” to be that organism. Put another way, Nagel’s question of “what it is like to be” a particular organism could be answered if one could somehow have that organism’s first order self [1].
One’s 2ND order selves would also have a certain qualitative character, and one’s 3RD order too.
There would, however, be a very clear delineation between conscious and not. The absence of a 1ST order self would mean there would be no policy linking all other policies. There could be no “self” to experience all of this. Hence a 1ST order self must precede all others.
Furthermore, there could be no 2ND order self without a 1ST. Our framework thus makes a zombie impossible (there can be no perfect unconscious replica of a conscious being). The hard problem has it backwards. The question is not why qualia exist, but why anyone thinks representational contents [^18] Policies for interpreting the environment. can exist without first being learned through qualitative experience and discrimination. This ‘dissolves’ the hard problem by going a level down to the fundamental drive: stay alive! The imperative ‘stay alive’ involves sensing, classifying, evaluating, prioritising and acting.
Our arguments align with those of Merker, Barron and Klein, who have linked subjective experience to reafference [36, 37, 38]. We agree that reafference is key, but provide a very different explanation of why and how. Their work presents biological evidence for subjective experience in organisms with reafference. In contrast, we derive the 1ST order self from first principals and explain why and how it is a classifies “what it is like” to be a particular organism. The 1ST order self also happens to be equivalent to reafference, so we arrive at the same conclusion as Merker from very different, mathematical premises. Hence these are very complementary positions.
We speculate an organism approaches the full richness of human subjective experience when different orders of self-interact [^19] As argued by Boltuc the full richness of human consciousness, or “hard” consciousness as he calls it, might be distinguished from phenomenal and access consciousness. It is this to which we refer now.. Different orders of learned self might interact through hierarchical planning in a distributed system, perhaps something like a connective core [65]. There is no possible way phenomenal consciousness can exist as we experience it without these different orders of self. There is also no way the adaptability we see in living organisms can be possible without the construction of such selves. We turn to this discuss now.
8 From Rocks to Einstein: The Hierarchy of Being
To illustrate how our argument applies in the real world we describe stages of conscious organism. Each stage follows from scaling up supply and demand for WPO, through natural selection (see figure 4):
$0:$ Unconscious (e.g. a rock)
$1:$ Hard Coded (e.g. protozoan)
$2:$ Learning (e.g. nematode)
$3:$ 1ST Order Self (e.g. housefly)
$4:$ 2ND Order Selves (e.g. raven)
$5:$ 3RD Order Selves (e.g. human)
Stage 0: Unconsciousness
Stage zero may be understood as the consciousness we are willing to assign to a rock: more exactly, the lack of consciousness so assigned. While some assert a thesis of panpsychism and claim everything, even a solitary hydrogen atom is conscious, we take rocks as expressing a baseline example of things in the universe that are not conscious at all. They do not sense. They do not think. They do not act. They have no 1ST order self.
  • Example: A rock.
Stage 1: Hard Coded
Stage one refers to adaptations “hard-coded” or hard-wired by natural selection. This is the sort of preset behaviour that allows complexity to persist [66] in a reasonably consistent environment.
  • What: Hard-coded adaptations. Habituation and sensitization.
  • How: The extension of fit behaviour is learned by natural selection and hard-coded into the organism as a policy (in DNA, form, the local environment etc).
  • Why: If the environment never changes, it makes more sense to just hard-code survivable behaviour.
  • Example: Single-celled protozoan.
Stage 2: Learning
Stage two introduces learning. To learn an organism must store, classify and order historical examples by valence. There is not something it is like to be stage two, because there is no locus of “self”. A biological example of such a decentralised nervous system is the cubozoan box jellyfish Tripedalia cystophora. Even Tripedalia cystophora was recently shown to be capable of associative learning [40]. An entirely distributed control system can “learn”. Likewise, stage two is exemplified by the nematode C. elegans [67, 68], which has a centralised nervous system and exhibits some ability to adapt with experience. However, the absence of a “self” prohibits cause and effect reasoning, which as others have already pointed out must limit spatial, navigational abilities [38]. When starved C. elegans exhibit “increased locomotion and dispersal in a random, rather than directed, search” [38, 69, 70], whereas something like a bee or an ant can recall and navigate to previously discovered food [71, 72, 73]. A nematode can learn, but the absence of a self for the purpose of causal reasoning limits adaptability.
  • What: Mindless learning, with no centralised intent or locus “self”, to feel.
  • How: Any system which learns will do. Search, approximation or biology. Learning is impossible without affect, reward or some other notion of value.
  • Why: An organism that can learn can survive more than one that cannot.
  • Examples: Jellyfish, nematode.
Stage 3: 1ST Order Self
This is where phenomenal (i.e. experiential) consciousness begins, with a 1ST order self. In biological terms this also implies reafference, which others have argued is the key to subjective experience, albeit for different reasons than what we have argued here [36, 37, 38]. They identified a housefly as a good example of where subjective experience may begin, and we concur. We also hold this is where an organism might be said to have intent. Intuitively, the policy that motivated behaviour is the intent of that behaviour [^20] In the same way declarative and imperative programs are equivalent [74].. The “weaker” a policy is, the more behaviours it motivates. Conversely, if one’s actions share anything in common it is the intent that motivated them. The more diverse the actions, the more “general” or “high level” the intent they share. For example, the action of eating tends to involve the intent of satisfying hunger. If a policy is implied by all of an organism’s behaviour, then it might have motivated that behaviour. While this is an unusual form of words, we contend that an organism is such a policy.
A stage three there is a self to feel simple hunger and pain. However, communication as described by Grice would be impossible [63], because a stage three organism has no self-awareness, to represent another’s perception of their intent [42]. Nor would a stage three organism be able to conceive of its own death, or shame, because it cannot conceive of itself.
  • What: A 1ST order self. Reafference. Phenomenal consciousness.
  • How: Embodiment in which intervention is not identical to observation.
  • Why: Accurate prediction of consequences of interventions. For example, a fly must distinguish between having moved, and the environment having moved, to navigate.
  • Example: Housefly.
Stage 4: 2ND Order Selves
Stage four is the 2ND order self, and importantly this is where we hold access consciousness begins because it is where information is available for report in the Gricean sense. In other words, access consciousness follows phenomenal consciousness. The ability of ravens to intentionally deceive [75] suggests they are at least stage four. Raven $\mathfrak{a}$, aware that it is being observed by raven $\mathfrak{b}$, will act as if it is hiding food in one location to mislead $\mathfrak{b}$, but will then move the food in another location unobserved by $\mathfrak{b}$. $\mathfrak{a}$ seems to predict not just the intent of $\mathfrak{b}$ (to steal the food), but $\mathfrak{b}$’s perception of $\mathfrak{a}$. It seems intuitively likely that dogs and cats have second order selves, as they must hunt reasonably intelligent animals and must anticipate how their actions are perceived. For example, a cat anticipates its prey will flee when it is observed, and hides.
  • What: Access consciousness. Theory of mind. Self-awareness. Inner narrative.
  • How: Selection pressures that demand theory of mind.
  • Why: A 2ND order self is necessary to anticipate, manipulate and communicate intent. It allows one to anticipate how others will perceive one’s behaviour, which may be necessary to survive in a social hierarchy.
  • Example: Raven.
Stage 5: 3RD Order Selves
A 2ND order self for one’s 2ND order self. Humans appear to possess this, because we are aware that we are aware.
  • What: Meta self-awareness. Inner narrative in which actors have inner narratives.
  • How: More accurate prediction and planning.
  • Why: Because a social organism must predict complex social dynamics.
  • Example: Human.
Figure 4: Overview of stages and orders of self.
Here we suggested that one single formalism can explain so many different theories of consciousness. This unifies lower and higher order theories of consciousness, by scaling simple axioms from first principles. Importantly, this formalism is compatible both with the idea the environment is software running on a single Turing machine (a strong interpretation of pancomputationalism), and the idea of an environment as co-created by the interaction of independent entities (enactivism or a distributed computing system). This ensures there is simultaneous compatibility with enactivism and pancomputationalism.
9 Unifying Lower and Higher Order Theories of Consciousness
Our earlier research argues that assuming lower order states is like assuming an abstraction layer in a computer [35, 44]. The behaviour of software is determined by the abstraction layer that “interprets” it. At the most foundational level software is machine code [^21] A “word” of machine code triggers a mechanistic process hardwired in the CPU by the human design. For example a word may copy the 32 bit value stored at memory address $X$ into the 32 bit register $Y$, which the next line of machine code “adds” to register $Z$ by looping over each bit in an “adder” circuit.. Machine code is interpreted by hardware, and that hardware determines every aspect of what that machine code does. A word of machine code is a mechanical trigger that only “means” whatever we have designed the hardware to do when we input that word. That hardware is an abstraction layer in which the software exists (including “higher level” abstraction layers like the Python interpreter). In other words, what we call software is nothing more than the state of hardware [32]. This has undermined all but the most subjective of claims regarding the behaviour of theorised, software superintelligence [76, 46, 77]. It is a flaw in the very idea of intelligent software. The distinction between software mind and hardware embodiment is subsequently called computational dualism [32], because it is reminiscent of how Cartesian dualism conceives of a mental substance distinct from physical substance. If lower order mental states function as an abstraction layer, then HOT is a form of computational dualism.
If we are to avoid computational dualism, then we should formalise mental activity as embedded, extending into the environment and enacted by interactions between bodies [29]. From the enactive perspective we cannot assume lower orders as pregiven and fixed. Instead, contents are co-created through interaction between an organism and its environment [22, 78]. The organism learns a world, or an “abstraction layer” in computing terms, that is relevant to its motivations. This process is sometimes called “relevance realisation” [79, 80, 81, 82].
Some have argued enactivism is incompatible with computationalism because the set of possible interpretations that must be searched for relevance realisation is intractable [82]. A computationalist who thinks of the mind as software must assume there is a hardware abstraction layer, and that abstraction layer hinders relevance realisation because it constrains the set of concepts that can be entertained [35]. This objection to cognitive computationalism [47] may be understood as an alternative framing of the objection to computational dualism raised the context of artificial intelligence [32]. However, just because relevance realisation is at odds with that sort of computational cognitivism does not mean enactivism is incompatible with the idea of computation [^22] There have also been efforts to combine enactivism with the Free Energy Principle (FEP) [83, 84]. While our explicandum and formalization are substantially different the overall approach, explaining behaviour in terms of optimisation and embodiment, is similar. Hence, our approach and the FEP should be understood as complementary. in general [85].
Importantly, our claim is not that everything is computation understood as minimal information processing (Williams & Ciaunica in prep.). Rather our claim is weaker, namely that one can provide one single unifying formalism in terms of computations linking lower to higher level information processing. We develop a formalisation [32] of representationless pancomputationalism. Computation as mechanics. Not representational, but physical in the sense described by Piccinini [30]. Intuitively, if this is information processing, then it is in the sense that an apple falling to the ground is information interpreted by physics. Whatever causes the environment to act as it does might then be thought of as an interpreter in this mechanistic sense.
Our radical and provocative claim is that phenomenal consciousness without access consciousness is likely very common, but the reverse is implausible. A zombie impossible. A data warehouse might have “access” to information, but that is not the same thing as access consciousness. We have classified this sort of mindless “access” to information as functional but not access consciousness.
For example, Block holds that phenomenally conscious content is phenomenal, whereas access conscious content is representational. In pancomputational enactivism there are only states and the programs they form. This suggests the abstract “representations” we construct are just organised phenomenal content, clustered according to what causes valence. We don’t discard these two sorts of consciousness, but unify them by showing how phenomenal consciousness gives rise to access.
Far from suggesting there is no such thing as qualia, this suggests instead that there is no such thing as purely representational content in anything but the tools we construct and our interpretations of their behaviour. Cells are a material with agency [86]. A human is a “multiscale competency architecture” [87, 88, 24]. Our intelligence is a swarm intelligence; the high level goal directed behaviour of a swarm of cells. When we embody human abstractions (e.g. arithmetic) in silico (e.g. the x86 instruction set), we disconnect the high level goal directed behaviour from the low level behaviour (and the motivating affect) that gave rise to it [44]. We call the information embodied in the computer “representational” because it means something to us. However, if we view the computer as an organism, then it amounts a set of stage one adaptations none of which are the sort of general purpose solution we observe in the “agentic material” we call cells. When we embody our abstractions in silico we disconnect them from the affect that motivated their construction [43]. This suggests there is no such thing as representational contents. They are a fiction we have invented because we struggle to reduce our own abstractions to their basic nature; the causes of valence.
10 Conclusions and Outlook: Why Nature Does Not Like Zombies
In this paper we provided a mathematical formalism uniting lower order states and higher order meta-representations, dissolving the hard problem of consciousness. Specifically, we described a multilayered formalism illustrating how biological self-organising systems become phenomenally conscious when they construct a 1ST order self. A human lacking a 1ST order self could not perform causal reasoning needed to adapt as humans evidently can [45].
As previous research pointed out [32], the computer metaphor with the seminal distinction between hardware and software is a simplification. Rather, software is a state of hardware. Nature privileges efficiency over abstract simplification. Compare the vast quantities of both training data and energy required by a language model, to the small quantities humans need to solve a problem. Biological systems are more efficient, because they are adaptive at a every level [24, 44]. Hence, rather than trying to explain the mind in the abstract like software, we started at the level of the embodied organism.
One direct consequence of our approach is that it places the phenomenal quality of conscious experience before access consciousness. We have shown a consistent definition of access consciousness implies 2ND order selves, which imply a 1ST order self. Phenomenal consciousness arises first, access comes later. Unlike panpsychism, we don’t believe rocks are conscious but solely those self-organising systems that need to adapt, motivated by valence, while keeping track of the self. Consciousness is a necessary adaptation, and a zombie is impossible because some behaviour cannot be achieved without consciousness.
There remain unanswered questions regarding whether the mere presence of 1ST, 2ND and 3RD order selves is sufficient for consciousness, but we hold they are at least necessary. Their absence should guarantee a system is not conscious. This serves to resolve questions about the consciousness of artificially intelligent systems, such as large language models. Such models are neither optimised to construct weak representations, nor have any incentive to construct selves, being passive mimics [42] of human behaviour. Future research should attempt to identify 2ND and 3RD order selves in biological systems. This will help establish levels of consciousness in different organisms. Merker, Barron and Klein have already nicely demonstrated the existence of a 1ST order self in humans and insects. Future research may also consider training organoids and artificial systems in a manner that should cause them to construct 1ST, 2ND and 3RD order selves. This may help whether what we have described is sufficient to achieve the outward behaviour of consciousness and what else, if anything, may be required to engineer it. Any difference between equivalent synthetic biological organoid and non-biological artificial intelligence will help resolve questions [44] about the difference between biological and non-biological substrates.
Another important future avenue for research in this area revolves around synchronicity and centralisation. For example, our formalism suggests a moment in which an output is chosen. This means that there exists a state of the environment in which all the declarative programs that make up a policy are facts. These declarative programs are physically manifest in the part of the environment at that moment, and a “decision” made by an agent is either determined by or determines the state at that moment. This centralises control, in that parts synchronise to affect a decision and a coherent whole (relative to a state). Importantly, this approach is analogous to how cells form an organ in a multiscale competency architecture [88]. When a cell is isolated from the informational structure of which it is part, it becomes cancer.
Now, one important question for future research is what happens if some part is isolated from the informational structure of the whole in the case of consciousness as we have portrayed it? Is this informational structure confined to a point in time? For example, can a policy be spread over time as a sequence much like how a CPU in a computer might process information over time, as sequence? What exactly is a moment? It is arguable that everything is centralised, and nothing is, depending upon the perspective we adopt with respect to space and time. There does not seem to be anything like an “event horizon” for consciousness; no cut-off point where something cannot affect something else and so prevent the formation of selves. This raises interesting questions about consciousness at different scales of space-time. Given our formalism, it is conceivable that what we perceive as continuous time is in fact disjointed, and there are large gaps in our apparently continuous perception we are unable to perceive. Perhaps we are “prompted” to process information by some states of the environment and not others, because some states are indistinguishable from one another given the abstraction layer in which we exist. Future research might explore such questions.
Looking further at the social aspect: if consciousness does not require some specific form of synchronicity, then what does that say about the action of populations? Where is the line between feeling and being? Feeling is being. But feeling comes first. We hold that valence informs learning at Stage Two. Being, at the human level, the Being of which the phenomenologists speak, does not emerge until Stage Three. Arguably, a group of humans is stage two, and yet an individual human is stage three. A group of humans cannot have a second order or first order self as we perceive it in time, but what of the effect of their actions in the environment across time? From the multiscale competency architecture perspective, the human mind is the mind of a collective of cells. It is at least conceivable, however implausible, that collectives of collectives of cells might be conscious, albeit at a very different timescale.
Another area future research might explore is how subjective experience depends on the ability to simultaneously learn and decide (simultaneous “learning and inference”, or “induction and abduction”). If an organism ceases to learn, it has only the memory of affect. Information retrieval must trigger affect to some extent in order to retrieve the quality of a protosymbol (otherwise, the quality would not be remembered). In other words, the memory of a painful experience must invoke pain, or it is not really remembered. Hence, an organism that has learned must still have subjective experience even if all it can do is remember (the absence of any subjective experience would mean it could not remember subjective experience). Recall of valence still compels one to act, meaning one must be subject to it. A human does not operate in two distinct phases of learning and inference as a machine learning system does. We appear to learn and infer at all times. Were we to have separate phases we would be unable to remember anything that happens in the inference phase, or else we would be motivated by the new memories and so would still be learning.
To put it provocatively, this suggests a “pure thinker” [89] cannot be conscious. Can we really consider the memory of consciousness to be the same thing as consciousness, if there is no awareness of the present which those memories are used for inference? If there are no new subjective experiences? Likewise, does “dropping Hume’s Guillotine” [44] to embody purely representative content in-silico eliminate the possibility of consciousness, by separating abstractions from the valence that motivated their construction?
Our paper sparks perhaps more questions that it hasn’t answered. But our proposal lays the foundations of a formal science of consciousness, deeply connected with natural selection rather than abstract thinking, closer to human fact than zombie fiction.

### v2

@@
-
+hard problem of consciousness
 Abstract
-We tackle the hard problem of consciousness taking the naturally-selected, self-organising, embodied organism as our starting point.
+We tackle the hard problem of consciousness taking the naturally selected, self-organising, embodied organism as our starting point.
 We provide a mathematical formalism describing how biological systems self-organise to hierarchically interpret unlabelled sensory information according to valence and specific needs.
@@
 Building upon the seminal distinction between access and phenomenal consciousness, our radical claim here is that phenomenal consciousness without access consciousness is likely very common, but the reverse is implausible.
-To put it provocatively: Nature does not like zombies.
+To put it provocatively: death grounds meaning, and Nature does not like zombies.
 We formally describe the multilayered architecture of self-organisation from rocks to Einstein, illustrating how our argument applies in the real world.
@@
 Why is anything conscious?
-[^1] This work was supported by a Fundaçao para a Ciencia e a Tecnologia (FCT) grant PTDC/FER-FIL/4802/2020 ; and 2020-02773 CEECIND FCT to A.C.
 Both biological and other physical [^1] Physical here just means non-biological.
 We are not suggesting biological systems are non-physical.
-systems process information, yet it seems that humans consciously experience in addition to merely process information.
+systems process information, yet it seems that humans consciously experience in addition to merely processing information.
 Why?
-Living organisms are constantly processing self- and world-related information to secure survival in an ever-changing world.
+Living organisms are constantly processing self and world related information to survive in an ever-changing world.
 Human bodies share with all other physical systems the property of being instantiated in time and space, (e.g., our body occupies a given position and volume in space at a given time).
-Yet, unlike physical systems, biological, living systems are dissipative systems using energy to self-organise in the face of entropic decay and environmental perturbation [1, 2].
-Originally formalised in the field of cybernetics [3, 4] the notion of self-organisation has been subsequently applied to various disciplines including physics [5], biology [6, 7] and neuroscience [8, 9, 10].
-Self-organization is typically defined as the spontaneous emergence of spatiotemporal order or pattern-formation processes in physical and biological systems resulting from interactions of its components with the environment [11, 12, 13].
-Interestingly, much self- and world related information processing goes on behind the scenes, or “in the dark” so to speak, that is, without being constantly present to our conscious minds.
+Yet, unlike physical systems, biological, living systems are dissipative systems using energy to self-organise in the face of entropic decay and environmental perturbation [nagel1974, chalmers1995].
+Initially formally introduced in the field of cybernetics [ashby1947, foerster1960] the notion of self-organisation has been applied to various disciplines including physics [haken1983], biology [camazine2003, bell2007] and neuroscience [kelso1997, Friston2010, tognoli2014].
+Self-organization is typically defined as the spontaneous emergence of spatiotemporal order or pattern-formation processes in physical and biological systems resulting from interactions of its components with the environment [camazine2001, seeley2002, rosas2018].
+Interestingly, much self and world related information processing goes on behind the scenes, or “in the dark” so to speak, that is, without being constantly present to our conscious minds.
 But why doesn’t all information processing go in the dark?
-This question is the subject of long-standing debates across disciplines [14].
-For example, one highly influential view is that consciousness has two aspects [14].
-The first is functional, by which we mean the ability to access and communicate information [15].
-How exactly information processing is linked to consciousness is the “easy problem” of consciousness [2].
-The second aspect is “what it is like” to consciously experience information processing, or phenomenal consciousness [2, 15, 1, 16, 17].
+This question is the subject of long-standing debates across disciplines [seth2022].
+One highly influential view is that consciousness has two aspects [seth2022].
+The first is functional, by which we mean the ability to access and communicate information [block1995].
+How information processing is linked to consciousness is the “easy problem” of consciousness [chalmers1995].
+The second aspect is “what it is like” to consciously experience information processing, or phenomenal consciousness [chalmers1995, block1995, nagel1974, gallagher2021, fuchs2017].
 This doesn’t mean just global states like being awake, but more specific local states like smelling a cup of coffee.
-These local contents or “qualia” are characterised by what it is like to be in them [14].
+These local contents or “qualia” are characterised by what it is like to be in them [seth2022].
 It is unclear the extent to which functional and phenomenal aspects are independent.
-David Chalmers has influentially suggested that it may be possible to construct a “zombie” which acts in every way like a person but has no qualia [15].
+David Chalmers has influentially suggested that it may be possible to construct a “philosophical zombie” which acts in every way like a person but has no qualia [block1995].
 For example, a thermostat certainly detects heat and so processes information, but there presumably is not anything it is like to be a thermostat.
@@
 This is the “hard problem” of consciousness.
-The hard problem has sparked a substantial body of work and a detailed discussion of these debates [14, 18] lies beyond the scope of our paper.
+The hard problem has sparked a substantial body of work, detailed discussion of which [seth2022, northoff2014] lies beyond the scope of our paper.
 Rather, in what follows we build upon the useful distinction between lower and higher order theories of information processing in relation to conscious experiences.
-For example, higher order thought theory (HOT) [19, 20] holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states.
+For example, higher order thought theory (HOT) [rosenthal2005, brown2019] holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states.
 Lower order states may include emotions and perceptions, while higher order meta-representations reflect upon those.
@@
 Sense data is processed by the body resulting in lower order mental states, and then meta representations of those is where we might find more abstract conceptual or thought-like contents of consciousness.
-The division between lower and higher orders is a good starting point to understand why some information processing goes on “in the dark”.
+The distinction between lower and higher orders is a good starting point to understand why most biological information processing goes on “in the dark”.
 Yet, the interesting question in our view is why do these lower order states arise?
@@
 That is: to process information at the lower bodily levels?
-In this paper we suggest that a useful way to dissolve the “hard problem” of consciousness is to reverse the order and start with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states.
-If an explanation of how consciousness functions is to reveal why information processing sometimes has a qualitative aspect, then it must explain how we get local states [21], not begin by assuming local states.
+In this paper we suggest a way to dissolve the “hard problem” of consciousness is to reverse the order and start with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states.
+If an explanation of how consciousness functions is to reveal why information processing sometimes has a qualitative aspect, then it must explain how we get local states [thompson2016], not begin by assuming local states.
 We need to go a level down and begin at the level of the embodied organism.
-Longtime considered a fringe approach, the embodied cognition paradigm [22] has recently gained substantial influence in cognitive science and philosophy [23, 24, 25].
-The key idea is that instead of considering the body as a mere device designed to fuel and contain the mind (a device that can be replaced with a vat or a robot, for example), one must consider the mind as serving the self-sustaining needs of a surviving body.
-If this is so, then understanding consciousness must start with understanding the ‘humble’ lower bodily levels of information processing, and not the higher order levels of information processing only.
-The key idea is that conscious experiences do not merely depend on bodily experiences as an external factor that can be replaced with a vat or an artificial system.
-One promising approach to this question is to try to rigorously define consciousness from first principles and show that some aspects of functional consciousness depend on phenomenal consciousness in a manner that makes zombies impossible, “dissolving” the hard problem by showing the phenomenal to be functional [^3] Note that we are far from the first to claim that the phenomenal is functional.
-For example, the proposed Conscious Turing Machine [26] based on Global Workspace Theory [27], and constructivist approaches to artificial intelligence [28] take similar position.
+Longtime considered a fringe approach, the embodied cognition paradigm [varela2016] has recently gained substantial influence in cognitive science and philosophy [ciaunica2023, levin2024, seth2018].
+The idea is that instead of considering the body as a mere device designed to fuel and contain the mind (a device that can be replaced with a vat or a robot, for example), one must consider the mind as serving the self-sustaining needs of a surviving body.
+If this is so then understanding consciousness must start with understanding the ‘humble’ lower bodily levels of information processing, not the higher order levels of information processing only.
+In other words, conscious experiences do not merely depend on bodily experiences as an external factor that can be replaced with a vat or an artificial system.
+One promising approach to this question is to try to rigorously define consciousness from first principles and show that some aspects of functional consciousness depend on phenomenal consciousness in a manner that makes zombies impossible, “dissolving” the hard problem by showing the phenomenal to be intrinsically functional [^2] Note that we are far from the first to claim that the phenomenal is functional.
+For example, the proposed Conscious Turing Machine [blum2020] based on Global Workspace Theory [baars1997], and constructivist approaches to artificial intelligence [wang2020] take similar position.
 It is our explanation of how and why that is novel..
 Unifying them.
-We would need to establish axioms that hold in every possible environment and show that it is impossible to have the function of consciousness without the subjective experience of it.
-Such an answer must explain how consciousness functions, and why some information processing goes on “in the dark”.
+We would need to establish axioms that hold in every conceivable environment and show that it is impossible to have the function of consciousness without the subjective experience of it.
+Such an answer must explain how consciousness functions, and why most information processing goes on “in the dark”.
 This paper lays the grounds of a computational model formalizing the link between lower and higher levels of information processing in relation to conscious experiences in a way that is compatible with the basic principles of embodied cognition and enactivism.
-Enactivism is roughly the view that information processing arises through a dynamic interaction between an acting organism and its environment [29].
-Now, the notion of computation is widely debated, and a detailed review of these discussions would lead us to a major digression [30, 31].
-Here we define computation not in terms of symbol shuffling and representation, but in mechanistic terms.
-‘Information processing’ in this sense is the causal relations dictating the transition of a system from one state to another, which concerns biological systems such as human bodies.
-Pancomputationalism is the idea that all dynamic systems are constituted by computation.
-This paper builds upon and extends the formalism of pancomputational enactivism developed by one of us [32].
+Enactivism is roughly the view that information processing arises through a dynamic interaction between an acting organism and its environment [thompson2007].
+Now, the notion of computation is widely debated, and a detailed review of these discussions would lead us to a major digression [piccinini2015, piccinini2021].
+Here we define computation not in terms of abstract symbol shuffling and representation, but in concrete mechanistic terms where ‘information processing’ refers to the actual physical transition of a system from one state to another; a system such as a human body.
+Pancomputationalism is the idea that all dynamic systems are constituted by computation, which for our purposes is defined as above.
+This paper builds upon the formalism of pancomputational enactivism [bennett2024a].
 To answer “why is anything conscious” we proceed in two steps.
-First we develop a mathematical formalism in which lower and higher order theories of consciousness, and phenomenal and access consciousness, are all derived from first principles.
-Each follows because of scaling natural selection pressures and the ability to adapt [^4] This is loosely inspired by a scale-based framing of machine learning [33]..
+First, we develop a formalism in which lower and higher order theories of consciousness, and phenomenal and access consciousness, are all derived from first principles.
+Each follows because of scaling natural selection pressures and the ability to adapt [^3] This is loosely inspired by a scale-based framing of machine learning [sutton2019] - we are applying Sutton’s “bitter lesson” to biology..
 Second, we provide an argument building on that to dissolve the hard problem of consciousness, answering why there must a qualitative aspect to information processing and in what circumstances.
 This paper is organised as follows.
-Section 2 is a list of formal mathematical definitions for pancomputational enactivism, which the reader may refer to as needed.
-For readers without formal background, the Section 2 can be skipped, without impacting the argumentative streamline of the paper.
-In section 3 we lay the foundations.
-We extend separately published work unifying pancomputationalism and enactivism [32], developing a model of self-organising systems that holds across all conceivable environments.
-We explain how this model formalises relevance realisation and unifies lower and higher order theories, citing separately published mathematical and experimental results [34].
-We extend previous work [35] on causal learning and the development of self, to formalise analogues of subjective experience [36, 37, 38], access consciousness [15] and meta self-awareness [39].
-We call these analogues first (1ST), second (2ND) and third (3RD) order selves respectively, and explain how they are constructed as a direct consequence of scaling the ability of a self-organising system to adapt with natural selection pressures.
-Section 4 has two main parts.
-First we explain how subjective experience requires a 1ST order self, and conversely why a 1ST order self implies there is “something” it is like to be an organism that has a 1ST order self.
-We then argue qualia are information processing without representation, and representationless information becomes qualia when there is a common 1ST order self be subject to these experiences.
-Phenomenal contents must precede representational contents, because representational contents are just interpretations from phenomenal experience.
-Quality precedes quantity.
+It should be noted first that the appendix at the end of this paper is a list of formal mathematical definitions, which the reader may refer to as needed.
+For readers without formal background, the appendix can be skipped.
+In Section 2 we extend separately published arguments, experiments and proof results [bennett2024a, bennett2023b, bennett2023c], developing a model of self-organising systems that holds across all conceivable environments.
+In Sections 3 and 4 we explain how this model formalises relevance realisation and unifies lower and higher order theories.
+Rather than assuming abstract objects are primary and trying to learn causal relations between them, this assumes valence is primary and that abstract objects are constructed to classify cause and anticipate valence.
+This takes us from attraction and repulsion to contentless states, through lower order thoughts to higher order meta representations.
 We call this the psychophysical principle of causality.
-In the second part we describe the developmental stages of consciousness this implies as we scale up the capacity to adapt with natural selection pressures.
-We identify 6 stages including unconscious, hard-coded behaviour, learning, and then 1ST, 2ND and 3RD order selves.
-We provide examples of each from computers to jellyfish [40], to houseflies [38], to birds [41] and then humans.
-Finally, in section 5 we provide concluding remarks and outlook discussion for future research.
-2 Pancomputational Enactivism Definitions
+In Section 5 we extend previous work [bennett2023c] on causal learning and the development of first (1ST), second (2ND) and third (3RD) order selves, and explain how they are constructed as a direct consequence of scaling the ability of a self-organising system to adapt with natural selection pressures.
+In Sections 6 and 7 we explain how subjective experience requires a 1ST order self, and conversely why a 1ST order self implies there is something it is like to be an organism that has a 1ST order self.
+We then argue that access consciousness requires the ability to communicate meaning, which requires both 1ST and 2ND order selves, and show how a philosophical zombie is impossible.
+In Section 8 we describe the development of consciousness as we scale up the capacity to adapt with natural selection pressures.
+We discuss examples of each stage from nematodes to humans.
+Section 9 is concluding remarks.
+2 Back to Foundations
+Rather than presupposing local states or an abstraction layer, we start from first principles to formalise all conceivable environments [bennett2024a, bennett2024b], as in definition 1.
+This is based on two assumptions:
+Axiom 1: When there are things, we call these things the environment.
+Axiom 2: Where things differ, we have different states of the environment.
+Hence, the environment is a set $\Phi$ of contentless global states $\Phi$.
+We define a declarative “program” as a set of states and treat the states themselves as irreducible [^4] We do not concern ourselves with imperative programs, because there is an equivalence between declarative and imperative [howard1980]..
+The set $P$ contains all such programs.
+The powerset of $P$ is every aspect of the environment, because every aspect of any conceivable environment must be set of such programs, which is a subset of $P$.
+Programs return true or false, so a program is a “fact” if it is true, and the current state of the environment is the set of all “facts”.
+Truth is determined with respect to states, and states are assumed to be mutually exclusive.
+For example, if time is one way in which things differ, then there is one state at a time.
+It follows that every aspect of the environment is a set of programs, and those programs are facts relative to any physical state they contain.
+An aspect of the environment is realized by a state if the programs it contains are facts.
+When we say a state is contentless, we mean we assume nothing about what states contain or are.
+Their only significance is in how they relate to one another, and each such relation is a program.
+This is intentionally like to post-structuralism [eacho2020, hayles2005], in which the only meaning words have is given via their relations to each other.
+This is why previous work claims it applies to all conceivable environments [bennett2024a].
+Universality Claim: Axioms 1 and 2 hold for every conceivable environment.
+Intuitively, one might understand the set of all programs (the powerset of all states) as the set of all relations between these otherwise meaningless states.
+2.1 Natural Selection and Embodiment
+We then assume a process of natural selection. This produces embodied organisms.
+A set of facts can represent anything [^5] Some may object, pointing out that this ignores composition.
+However, the application of a function is a fact regardless of whether it takes another function as input., hence the body of an organism must be a set of programs.
+We call this set a vocabulary, and it is formalised in definition 2 (see Appendix).
+The vocabulary can be seen as hardware and software operating together.
+Piccinini [piccinini2015, piccinini2021] identifies two sorts of computation: abstract and concrete.
+Abstract computations mean or represent whatever we say they do.
+In contrast concrete computations are physical.
+They are the action of physical systems over time.
+The programs we talk about here are concrete, physical computations in that they are the relations between physical states.
+It follows that the computations a body performs are also concrete, and so no two organisms have the same vocabulary, because that would mean they are the same body.
+This might seem inconvenient, as intuitively we might prefer to think about bodies as an abstract ‘type’.
+However, doing so would undermine any claim we might make, as earlier results in artificial intelligence (AI) show [bennett2023c, bennett2024c].
+The meaning of abstractions depends upon interpretation, much like the behavior of software is determined by the abstraction layer that “interprets” it.
+At its least abstract software is machine code [^6] A “word” of machine code triggers a mechanistic process hardwired in the CPU by the human design.
+For example a word may copy the 32 bit value stored at memory address $X$ into the 32 bit register $Y$, which the next line of machine code “adds” to register $Z$ by looping over each bit in an “adder” circuit..
+Machine code is interpreted by hardware, and that hardware determines every aspect of what that machine code does.
+A word of machine code is a mechanical trigger that only “means” whatever we have designed the hardware to do when we input that word.
+That hardware is an abstraction layer in which the software exists (including “higher level” abstraction layers like the Python interpreter).
+In other words, what we call software is nothing more than the state of hardware [bennett2024a].
+This has undermined all but the most subjective of claims regarding the behavior of theorised, software superintelligence [hutter2010, leike2015, leike2018].
+It is a flaw in the very idea of intelligent software.
+The distinction between software mind and hardware embodiment is subsequently called computational dualism [bennett2024a], because it is reminiscent of how traditional cartesianism conceived of a mental substance distinct from physical.
+Here we suggest that must avoid computational dualism.
+If we are to dissolve the hard problem by starting with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states, then we cannot presuppose organisms use particular abstractions.
+Instead we must explain why particular abstractions are formed, in terms of contentless states, as the product of interaction between an organism and its environment [thompson2007, varela2016, rolla2021].
+This process is sometimes called “relevance realisation” [vervaeke2012, vervaeke2013a, vervaeke2013b].
+Some have argued enactivism is incompatible with computation because the set of possible abstractions that must be searched for relevance realisation is intractable [jaeger2024].
+For tractability, one must first isolate a “small world” of relevant information from the “big world” of all information.
+Part of the reason we can address this is because we formalise embodiment, in concrete terms from the level of irreducible physical states.
+Embodiment implies finite resource constraints, so we adopt one more axiom from [bennett2024b].
+Axiom 3: All aspects of the environment are spatially extended [bennett2024b].
+This means that a body can occupy only a finite number of states, which is formalised in definition 2.
+A body is a vocabulary $\mathfrak{v}\subset P$ with finitely many elements.
+$\mathfrak{v}$ implies a formal language $L_{\mathfrak{v}}$ of interaction between body and environment.
+A statement in this formal language is an aspect of the environment.
+A set of “programs” from our vocabulary.
+Statements have truth conditions with respect to environmental states, and every statement has an extension.
+The extension of a statement is the set of all statements which are supersets of the first statement (the set of all other statements by which the first statement is implied).
+Extension is important we can relate statements by their truth conditions, forming a lattice.
+Most importantly, we avoid the distinction between software and hardware and so avoid computational dualism.
+Embodiment is dispositional.
+As natural selection evolves a vocabulary, it begins to isolate a “small world” relevant to survival [jaeger2024, bennett2024b, friston2023], from the intractable “big world” $P$ of all information.
+2.2 Self-Organising Systems as Self and World Constraints
+Both snowflakes and human bodies are self-organising systems, yet we regard only the latter as conscious.
+What separates the two?
+To tackle this question, we need to introduce the notion of self and world constraints.
+Given that we have an embodied formal language, we can talk about computation with inputs and outputs by just treating everything as embodied statements embedded and enacted within the environment.
+Given an input $i$, the set of all possible outputs is the extension $E_{i}$ of that input.
+This is because if $i$ is realised by the environment, then the environment is constrained to only those states that realise $i$, which constrains what other statements can be realised.
+We can use this fact to talk about “policies” as embodied [^7] As a state of hardware rather than software, which is how the term policy is normally used in reinforcement learning.
+constraints on behaviour.
+A policy is a statement whose extension constrains outputs, like a causal intermediary in functionalism [putnam1967].
+If behaviour is “motivated”, then statements have valence determined by natural selection and only a subset of the statements a body could make are “fit”.
+As a result, an organism will self-organise to express some statements, but not others.
+Though our environmental states are contentless and their only significance is in how they relate to one another, natural selection must make them attractive or repulsive to a self-organising system.
+This is formalised by the $\mathfrak{v}$-task as in definition 3888A $\mathfrak{v}$-task is called because a $\mathfrak{v}$-task because it exists in the context of a vocabulary $\mathfrak{v}$..
+If $\mathfrak{v}$ is the vocabulary of a body, and $\mathfrak{v}$-task $\mu=\langle I_{\mu},O_{\mu}\rangle$ is fit behaviour, then $I_{\mu}$ is all the statements that body can express in which it is possible to remain fit, and $O_{\mu}$ is all the statements that body can express in which it remains fit.
+The extension $E_{I_{\mu}}$ of $I_{\mu}$ would be every output that it is possible to choose given the inputs $I_{\mu}$, but only a subset of those $O_{\mu}\subset E_{I_{\mu}}$ are correct outputs.
+In other words, this defines a self-organising behaviour as a constraint on outputs, given inputs [^9] Conversely, a system which considered all outputs to be correct would not be self-organising..
+2.3 Inference
+Every statement in $L_{\mathfrak{v}}$ implies a constraint, because there are only so many outputs that can be expressed by a body at the same time as any given statement.
+Assume we want to constrain the body to $O_{\mu}$ given $I_{\mu}$.
+A body could express a statement $\pi$ (meaning $\pi$ is true), and $\pi$ could then constrain the body to only correct outputs $O_{\mu}\subset E_{I_{\mu}}$ if $O_{\mu}=E_{I_{\mu}}\cap E_{\pi}$.
+We call a constraining statement a policy.
+A policy constrains outputs given inputs.
+A correct policy is one that constrains outputs to only correct outputs.
+For the sake of intuition, think of “correct” as “fit” according to natural selection [^10] This need not always be the case..
+This is formally expressed in definition 4.
+“I survive therefore my model is viable.” - Mark Solms [solms2021]
+If an organism remains alive, then its history is an ostensive definition of “fit” self-organising behaviour.
+That history of interactions is just a set of statements in the organism’s embodied language, which can be expressed as a $\mathfrak{v}$-task $\mathfrak{h}$.
+Each input and output in $\mathfrak{h}$ is behaviour the organism has exhibited and everything that was involved in that behaviour (the interaction between organism and environment).
+The set of all fit behaviour would then be another task $\mu$, and if the organism remains fit then $\mathfrak{h}\sqsubset\mu$ [^11] One task is a child of the other, meaning the inputs and outputs are subsets of the parents’ inputs and outputs..
+2.4 Learning
+Learning can be understood as the process of approximating the program which caused data.
+Constructing a “world” model.
+Here, we just need to explain how an organism gets $\mu$ from $\mathfrak{h}$.
+We don’t need a world model, but a way to constrain outputs to just those that are fit.
+Like everything else, that constraint is once again a statement in the formal language $L_{\mathfrak{v}}$.
+“The best model of the world is the world itself” - Rodney Brooks [dreyfus2007]
+Those examples in $\mathfrak{h}$ could have been generated by any one of a set of policies $\Pi_{\mathfrak{h}}$.
+These are not representations of possible causal intermediaries in the machine functionalist sense, but just parts of the outputs from which the outputs could be derived given the inputs.
+Every $\mathfrak{v}$-task (and thus every self-organising system) implies such a set of policies, and not all of those policies will constrain outputs the same way given new inputs.
+Some policies will “generalise” to imply fit behaviour in unfamiliar circumstances, meaning correct outputs given a new set of inputs (a parent task of the organism’s history $\mathfrak{h}$).
+If those policies imply fit behaviour given new inputs, then the organism will remain fit.
+If an organism learns, then it is adapting to embody a policy that it hopes will generalise to fit behaviour $\mu$.
+Earlier work [bennett2023b] showed that the optimal strategy to learn and adapt as fast as possible [^12] Meaning to converge on a fit policy given the smallest history possible, to realise fit behaviour.
+is to prefer “weaker” policies, meaning those with larger extensions.
+The policy which generated outputs $O$ given inputs $I$ is most efficiently identified by constructing a policy $\pi$ such that $\pi$ generates $O$ from $I$, and $\pi$ implies the weakest constraint that can be implied while still generating $O$ from $I$.
+This maximises adaptability, because it allows a self-organising system to construct fit policies from a shorter history [bennett2023c].
+This formalised in definition 5 and illustrated by figure 1.
+Figure 1: Illustration of how child tasks are examples of parent tasks.
+3 Relevance Realisation Through Causal Learning
+When an embodied organism learns a policy, it learns a relevant interpretation of that information, labelling it through behaviour.
+To efficiently adapt, an organism must correctly anticipate valence, which means correctly identifying what causes valence.
+For convenience, we will now expand upon the concept of organism given in definition 6.
+All organisms must have preferences, because they will make some decisions and not others.
+Behavior implies policies, and every policy implies tasks, so we can define preferences as an binary relation over tasks.
+Natural selection favours adaptability, hence we assume it prefers organisims that prefer weaker policies, generalising to remain fit.
+When a system optimises for weak policies, we call it weak policy optimisation (WPO).
+If a policy is fit, then it must correctly predict causes of valence.
+Hence by constructing fit policies, an organism divides the world into relevant objects and properties.
+A weaker policy implies all more specific versions of itself, meaning those that more tightly constrain outputs by having a smaller extension.
+Hence policy learning implies a lattice of policies that vary in weakness.
+Every fit policy would be a “causal identity” for something relevant, for example a specific object like “this coffee” or a weaker and more general concept like “all coffee”.
+We also need to introduce the notion of interaction between organisms.
+Organisms interact when they “affect” one another (in the physical rather than psychological sense of the word).
+This is formally defined in 9.
+Now, Artificial Intelligence (AI) and machine learning [russell2005ai, bishop2006, sutton2018, goertzel2014] are concerned with engineering adaptive agents.
+In that context, causality has become a mainstream topic of research.
+Causal learning is demonstrably necessary to thrive in an interactive setting [pearl2018, richens2024robust].
+Where the causal graph is known in advance (for example if we were to be measuring the efficacy of medical interventions), this issue could be resolved using causal language (such as “Do Calculus” [pearl2009]) to represent an intervention from outside the system the graph describes.
+Such an intervention represent agency, in that an organism observing the environment draws conclusion about that environment, and then intervenes to change the environment.
+To illustrate this point, consider the following example from [bennett2023c].
+Suppose an organism named Bob is attempting to learn and predict the environment.
+Now assume Bob has observed Alice wearing a raincoat only when it rains, and that Bob has observed rain only at those times when Alice has been observed wearing a raincoat.
+If we represent the raincoat observation with a binary variable $O\in\{true,false\}$ and the advent of rain with a variable $R\in\{true,false\}$, and if Bob is a Bayesian, then Bob’s observations will lead to the conclusion that $p(R=true\mid O=true)=1$.
+This means Bob believes that if Alice wears a raincoat, then it must be raining.
+Now lets permit Bob to interact with its environment.
+Assume Bob wants it to rain.
+Based on the belief $p(R=true\mid O=true)=1$, Bob may conclude that forcing $O=true$ by holding a gun to Alice’s head and demanding Alice wear a raincoat will cause it to rain.
+This is absurd.
+$O=true$ does not represent the event $q=$“I coerced Alice into wearing a raincoat”, but an entirely different event $v=$“Alice decided to put on a raincoat for the same reason I have observed Alice wearing a raincoat in the past”.
+To accurately represent the environment, we need to represent that $p(R=true\mid q)=p(R=true)\neq p(R=true\mid v)=1$, meaning both $q$ and $v$.
+To illustrate the problem visually, we started with the acyclic graph
+$O$$R$
+and our intervention disconnected rain from the choice of clothing:
+$O$$R$
+This can be resolved by introducing a “do” operator that we apply to a variable $O$ to obtain $do[O=o]$, to represent the fact that an agency from outside the system has intervened to assign a value to $O$, so that we can represent $p(R=true\mid do[O=true])=p(R=true)\neq p(R=true\mid O=true)=1$.
+Thus, the aforementioned $q$ is equivalent to $do[O=true]$, while $v$ is equivalent to $O=true$.
+The Psychophysical Principle of Causality
+We have explained why we need to account for causality.
+However, we have not explained how one might come to know all the objects involved (where did the variables come from?), or how they relate to one another causally (to form a graph).
+There are two ways to learn causal graphs.
+One can either assume a set of variables and learn relations between them, or assume relations and learn the objects that fit them [bennett2023c].
+Beginning as we are with contentless states, we do the latter.
+Rather than assuming objects are primary and trying to learn causal relations between them, this assumes valence is primary and that abstract objects are constructed according to what causes valence.
+We call this the psychophysical principle of causality.
+This is what will take us from attraction to and repulsion from contentless physical states, through lower order thoughts to higher order meta representations.
+We start with the complete absence of content [thompson2016], and once a system learns policies there are distinct “contents”.
+Policies serve to classify objects and properties.
+Death grounds meaning.
+To develop this idea, we first show adding additional nodes to a causal graph is an adequate substitute for the “do” operator [dawid2002].
+Assume we again have Bob who constructs a causal graph of the environment.
+Assume Alice exists in that environment.
+From Bob’s perspective, Alice is just a part of the environment represented by a variable $A$ in Bob’s causal graph.
+$O$$A$$R$
+Now assume Bob observes Alice taking an action that changes an aspect of the environment, represented by the variable $O$ (for example, Bob observes Alice putting on the raincoat).
+From the perspective of Bob, Alice’s action is the assignation of a value to a variable $A$, and to the variable $O$.
+There is no need to involve a do operator in this scenario because we can already represent that $p(R=true\mid A=x,O=true)=p(R=true)\neq p(R=true\mid A=y,O=true)=1$ (because Alice is part of the causal graph).
+This raises the question; if we do not require a $do$ operator to represent the actions of Alice, then why would we need it to represent an intervention by Bob?
+Couldn’t we represent the same information by introducing a new variable?
+The answer is yes [bennett2023c, dawid2002].
+$O$$B$$A$$R$
+Now we will show how it is possible to learn the relevant nodes to add, in the form of statements.
+The matter of which cause-and-effect relations are learned is determined by valence, and so the objects learned are statements classifying causes and valence.
+4 Relevant Causal Identities
+The formalism we employ began as an attempt to explain why some abstractions are formed, and not others [bennettmaruyama2022a].
+The optimal choice of policy for efficient adaptation is the weakest because it is implied by my statements in an embodied language.
+That language is dispositional, shaped by natural selection, and so a weaker policy is likely to be a fitter policy.
+As we have already shown, correct identification of cause and effect is often necessary to make accurate predictions and survive.
+WPO will thus correctly identify cause and effect where relevant.
+This is not just a case of substituting the “do” operator, but of simplifying the entire environment into objects and properties relevant to the organism’s motivations.
+We will now explain how relevant objects are classified, as “causal identities” [bennett2023c] and protosymbols [bennett2023d].
+4.1 Causal Learning
+In accord with [bennett2023d, def.
+6] we assume a vocabulary $\mathfrak{v}_{\mathfrak{b}}$ belonging to an organism $\mathfrak{b}$ (Bob).
+A “cause” in the context of this formalisation is not a variable but a statement $l\in L_{\mathfrak{v}_{\mathfrak{b}}}$ in that formal language.
+The raincoat example would involve $obs,rain\in L_{\mathfrak{v}_{\mathfrak{b}}}$ such that:
+[…]
+[…]
+[…]
+$int$$obs$$rain$
+In other words, so long as $obs\neq int$, the intervention can be differentiated from the passive observation. We formally define this in definition 10.
+This being the case, any set $c\subseteq int-obs$ could be used to identify the party undertaking the intervention, which is why $c$ is referred to as a “causal identity”.
+It distinguishes the intervention $int$ from the passively observed effect $obs$, like reafference in living organisms.
+However, the above only considers one intervention.
+A weaker or more general causal identity would be one that is shared by more interventions.
+For example, we might have two different interventions $a_{1}$ and $a_{2}$, with the observed effects $c_{1}\subset a_{1}$ and $c_{2}\subset a_{2}$, and causal identities $a_{1}-c_{1}=i_{1}$ and $a_{2}-c_{2}=i_{2}$.
+If $i_{3}=i_{1}\cap i_{2}\neq\emptyset$, then $i_{3}$ is a causal identity that is present in two interventions.
+Being a question of classification, we can express a causal identity in relation to a $\mathfrak{v}$-task, for which the causal identity is a policy (expanding upon [bennett2023c, def.
+10]), giving us definition 11.
+Example 1.
+Suppose we have organisms $\mathfrak{a}$ (Alice) and $\mathfrak{b}$ (Bob), and that the inputs Alice has experienced so far are $I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}$.
+These can be divided into those in which Bob affected Alice $S^{\mathfrak{b}}_{\mathfrak{a}}$ and those in which Bob did not $S^{\lnot\mathfrak{b}}_{\mathfrak{a}}=I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}-S^{\mathfrak{b}}_{\mathfrak{a}}$.
+By affecting Alice, Bob has intervened in Alice’s experience.
+Alice can construct a causal identity $b$ for Bob corresponding to interventions $INT=S^{\mathfrak{b}}_{\mathfrak{a}}$ and observations $OBS=S^{\lnot\mathfrak{b}}_{\mathfrak{a}}$.
+4.2 Ascribing Intent to Other Objects
+The “do” operator is used to discern the difference between an event one has caused, and that same event passively observed.
+However, what is passive observation if not the result of an intervention by something other than one’s self?
+The distinction between “intervention” and not is misleading.
+Everything is an intervention.
+The question is not “is this an intervention” but “by whom was this intervention made?”.
+Earlier, we arrived at the following graph in which Bob’s intervention was given by $int$.
+$int$$obs$$rain$
+However, what if a third person Larry puts the coat on Alice?
+Surely Bob can observe this, and so Bob’s observation of Larry’s intervention is $v\in L_{\mathfrak{v}_{\mathfrak{a}}}$ such that $obs\subset v$.
+To account for this, Bob can construct a causal graph as below (with $b.$ representing Bob and $l.$ representing Larry).
+$b.\ int$$l.\ int$$obs$$rain$
+Bob’s causal identity for himself $c_{b}\subset b.\ int-obs$ only represents the intervention by himself.
+However, now we can see that Bob must also construct a causal identity $c_{l}$ for Larry, where the $c_{l}\subset l.\ int-obs$.
+More generally, for an organism $\mathfrak{a}$ with an embodied language $L_{\mathfrak{v}_{\mathfrak{a}}}$ to construct a causal identity for an object $\mathfrak{b}$, it must first be the case that $\mathfrak{a}$ is affected by $\mathfrak{b}$ [bennett2023d], to satisfy the incentive precondition for causal identity.
+It depends on motive, or valence.
+Recall that to be affected is formally defined in definition 9.
+Assume an organism $\mathfrak{a}$ is affected by $\mathfrak{b}$ given inputs $INT$, and not affected given inputs $OBS$.
+To then attribute the contents of $INT$ to one specific entity, there must be something in common between the members of $INT$ caused by $\mathfrak{b}$ that is not shared by any member of $OBS$ caused by something else (in other words it must be at least possible for $\mathfrak{a}$ to discern the existence of $\mathfrak{b}$).
+The contents of $INT$ are “interventions” by $\mathfrak{b}$ and by learning $c$, a corresponding causal identity, $\mathfrak{a}$ can discern the existence of $\mathfrak{b}$.
+This is not to say that $\mathfrak{b}$ decides (chooses an output) or is even alive.
+What is important is that $\mathfrak{b}$ affects $\mathfrak{a}$, making it possible to discern when these interventions are a consequence of $\mathfrak{b}$’s existence.
+4.3 Preconditions
+There are preconditions for the existence of a causal identity.
+First, the vocabulary $\mathfrak{v}_{\mathfrak{a}}$ of an organism $\mathfrak{a}$ must be large and expressive enough to ensure that observations are distinguishable from interventions, meaning we have a vocabulary of sufficient scale.
+Second, there must be an incentive to construct it.
+Inference is only possible if some states are preferable to others.
+It is a value judgement.
+As Hume pointed out, one cannot derive what “ought” to be from a statement of what “is”.
+Natural selection provides a notion of what ought to be, by eliminating anything which ought not.
+  1. 1.
+    The scale precondition requires $\mathfrak{v}$ contain the causal identity.
+  2. 2.
+    The incentive precondition is that fitness demands the causal identity.
+4.4 Realising Lower Order States And Higher Order Meta Representations
+The weaker the causal identity, the more pervasive the “identity” in the sense of being part of more interventions.
+Learning is not just about constructing policies but joining them together into weaker, more abstract policies.
+By learning policies that correctly identify the causes of valence at different levels of abstraction, the organism engages in relevance realisation.
+This is formalised in definition 6.
+An organism $\mathfrak{o}$ has a set of policies it has learned.
+Each and every policy implies a $\mathfrak{v}$-task.
+A $\mathfrak{v}$-task is a triadic relation between inputs, outputs and policies which resembles Peircean semiosis [sep-peirce-semiotics, bennett2023d] of sign, referent and interpretant.
+Hence, in definition 7 we formalise this as a “protosymbol” [^13] Proto because it is something more primitive than a symbol as conceived of by Peirce.
+system $\mathfrak{s}_{\mathfrak{o}}$ for the organism $\mathfrak{o}$.
+Interaction is defined in terms of choosing an output, which means there is inevitably a policy and a policy implies a $\mathfrak{v}$-task, so definition 6 includes a preference order over tasks, which is used in definition 8 to formalise the interpretation of inputs (in terms of constraints, rather than an algorithm).
+In particular, the organism must always act according to an interpretation, and some interpretations imply others.
+Related as they are in a lattice, protosymbols are analogous lower order states and higher order meta representations.
+Tasks exist in a “generational hierarchy”.
+They are not mutually exclusive.
+Higher level tasks are more general, and have fewer policies because only very weak policies could complete them.
+Some have framed consciousness as a problem of moving from unary, to dyadic, to triadic relations [goertzel2006].
+Similarly, we have gone from unary states to dyadic programs, to triadic tasks and protosymbols.
+Figure 2: Illustration of how the generational hierarchy of tasks is a lattice of protosymbols.
+The question now is when does an organism realise a causal identity for when it causes valence? When does it construct a causal identity for itself?
+5 Multi-Layered Self-Organisation
+As the vocabulary and capacity for WPO scale, a greater variety of concepts can be learned.
+Progressively higher orders of ‘causal-identity’ for one’s self related information processing are constructed [bennett2023c].
+This lets us frame the construction of embodied selves in developmental [ciaunica2019] and evolutionary terms.
+5.1 The First Order Self
+A first order self (1ST henceforth, see figure 3) is functionally equivalent to reafference, which has been observed in the mammalian mid-brain and insect central cortex [merker2007, barron2016].
+Reafference lets an organism discern the consequences of its actions, and so a 1ST order self is necessary for accurate inference in an interactive setting in the same sense as Pearl’s ‘do’ operator [pearl2018].
+A 1ST order self formally defined in definition 13.
+It serves as the locus of self related information processing and experience [bennett2023c], allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction.
+But also, to anticipate future planned interactions (for example, an insect navigating its environment [barron2016]), and recollect past interactions (it is a subset of all those relevant “statements” in the formal language).
+Figure 3: Visual intuition for a 1ST order self. The organism constructs a causal identity for itself and can relate that to observed events.
+Natural selection prefers efficiency.
+The absence of a single, centralised causal identity can create inefficiencies [^14] The effect of insufficiently weak representations can be observed in large language models like GPT-4 [bennett2023d, gpt], where they fail to make consistent statements about an object because represent it as multiple objects..
+Decentralised or asynchronous control might be advantageous in some circumstances (e.g.
+an ant colony), but not others (e.g.
+in an individual human body).
+An organism that has a distributed control system might have a “self” for each part of that system, and perhaps the co-ordination of parts might seem to suggest the existence of a centralised “self”.
+Redundancy might be useful in some circumstances, but there is also a cost - more data may be required to learn and adapt, because the optimal choice of policy for accurate generalisation is the “weakest” [bennett2023b].
+If the scale and incentive preconditions are met for a causal identity for one’s self, then the organism must learn one.
+This might be partly hard-coded, but our focus is on inference by the organism itself.
+5.2 The Second Order Selves
+Survival may demand organism $\mathfrak{a}$ infer $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s interventions (to see one’s self as if through another’s eyes [bennett2023d]).
+This is called a second order self (2ND henceforth).
+We argue that if access conscious contents are available for communication in the human sense, then they must be communicable in the Gricean sense [grice1957, grice1969].
+Grice argued that communication is about the inference of intent.
+If person $\mathfrak{a}$ and $\mathfrak{b}$ are talking, then the meaning $m_{\mathfrak{a}}$ of what $\mathfrak{a}$ says is whatever $\mathfrak{a}$ intends $\mathfrak{b}$ understand.
+The meaning $m_{\mathfrak{b}}$ that $\mathfrak{b}$ understands is whatever $\mathfrak{b}$ thinks $\mathfrak{a}$ wants $\mathfrak{b}$ to think.
+$\mathfrak{b}$ has understood what $\mathfrak{a}$ means if $m_{\mathfrak{b}}$ approximates $m_{\mathfrak{a}}$.
+This can happen only if $\mathfrak{a}$ can predict with reasonable accuracy what $\mathfrak{b}$ thinks $\mathfrak{a}$ thinks, and $\mathfrak{b}$ can predict what $\mathfrak{a}$ thinks $\mathfrak{b}$ will think upon hearing an utterance.
+In other words, both $\mathfrak{a}$ and $\mathfrak{b}$ must have 2ND order selves that are good approximations.
+Yes, there are other aspects to communication.
+However, here we are talking about consciousness.
+Access conscious contents are those available for reasoning and report.
+It follows [^15] If the definition of access consciousness is to be consistent with reasoning and report as exhibited by conscious humans.
+that access conscious contents must in principle be communicable in the sense Grice described.
+As such, we argue contents available for communication can only be the contents of 2ND order selves, which means only an organism with 2ND order selves can be considered to have access consciousness.
+2ND order selves also explain attention and self-awareness.
+An organism can have many 2ND order selves because they depend upon who or what the organism is interacting with, just as the availability of information depends on context.
+Where a 1ST order self might allow one to observe a cat and form plans regarding causal interactions with the cat, a 2ND order self would allow one to be consciously aware of the cat for the purpose of reasoning and report.
+One can know of the cat, and that another organism knows of the cat, but a 2ND order self is insufficient to be aware that one is aware of the cat [^16] This is ‘meta-self-reflexive consciousness’ as some have described it [morin2006]..
+Figure 4: Visual intuition for a 2ND order self.
+The organism constructs a causal identity for itself, and for another object, and the causal identity for the other object includes a prediction of the organism itself from that object’s perspective.
+This would occur, for example, if one was a predator trying to predict the movements of prey in response to one’s own actions.
+It is an extension of the 1ST order self.
+More formally using the notation given in definition 15, assume $\mathfrak{a}$ and $\mathfrak{b}$ are organisms that evolved to accurately predict one another’s behaviour.
+Assume $\mathfrak{a}$ constructs a causal identity $c^{\mathfrak{b}}_{\mathfrak{a}}$ to predict $\mathfrak{b}$ given input $i_{\mathfrak{a}}\in I_{\mu_{\mathfrak{a}}}$, of which a second order self $c^{\mathfrak{ba}}_{\mathfrak{a}}$ is part.
+Likewise, $\mathfrak{b}$ constructs $c^{\mathfrak{a}}_{\mathfrak{b}}$ to predict $\mathfrak{a}$ given input $i_{\mathfrak{b}}\in I_{\mu_{\mathfrak{b}}}$, of which $c^{\mathfrak{ab}}_{\mathfrak{b}}$ is part.
+What is important here is that each organism’s intent is to some extent inferred by the other, and that fact inevitably changes the sorts of policies that will be “fit”.
+For example, second order self means each knows the other can anticipate manipulation, which means the optimal policy will often be to have rather than feign intent that aligns to some extent with the other party’s desires, to co-operate [alexander2022] [^17] Depending upon circumstances, for example organisms may co-operate in some circumstances but not others [bennett2023d], and transient relations, information asymmetry and other factors can make deceit a more attractive option..
+Repeated interaction creates an iterated prisoner’s dilemma, incentivising co-operation and signals that both parties interpret similarly (the beginnings of language) [bennett2023d].
+To communicate in Gricean terms, $\mathfrak{a}$ must intend to convey meaning $m_{\mathfrak{a}}$, and $\mathfrak{b}$ must recognise this intent.
+The incentive precondition explains why $\mathfrak{a}$ would form such intent (co-operation is often advantageous), while how may be understood as follows:
+  • $c^{\mathfrak{ba}}_{\mathfrak{a}}$ lets $\mathfrak{a}$ predict what $\mathfrak{b}$ will come to believe when it observes $\mathfrak{a}$’s behaviour.
+  • $c^{\mathfrak{ab}}_{\mathfrak{b}}$ then lets $\mathfrak{b}$ predict what $\mathfrak{a}$ intends that $\mathfrak{b}$ believe.
+$\mathfrak{a}$ can use $c^{\mathfrak{ba}}_{\mathfrak{a}}$ to infer behaviour to which $\mathfrak{b}$ will ascribe the intent to communicate $m_{\mathfrak{a}}$, and $c^{\mathfrak{ab}}_{\mathfrak{b}}$ lets $\mathfrak{b}$ infer that this is what $\mathfrak{a}$ intends.
+The “utterance” Grice refers to is how $\mathfrak{a}$ affects $\mathfrak{b}$ in accord with earlier definitions.
+Put another way, $\mathfrak{a}$ encodes $m_{\mathfrak{a}}$ into its behaviour in a manner that $\mathfrak{b}$ can decode (their respective second order selves act as encoders and decoders).
+By encode and decode, we mean a loose approximation of $m_{\mathfrak{a}}$ is communicated.
+There are of course shortcuts, for example of $\mathfrak{a}$ and $\mathfrak{b}$ are of the same species then they likely have similar motives and experiences, and so the efficient thing for each to do would be to use its own intent as an approximation of what the other might think.
+However, that does not obviate the need for second order selves, it just makes them easier to realise.
+5.3 The Third Order Selves
+We can continue scaling the vocabulary and capacity for WPO indefinitely, however for the purposes of this paper we will stop at 3RD order selves, because this appears to be the highest level of human consciousness, known as Meta self-awareness [morin2006].
+This is the awareness that one is self-aware.
+If self-awareness stems from 2ND order selves, then it follows that meta self-awareness requires 3RD.
+Formally a 3RD order self for $\mathfrak{a}$ reflecting off $c^{\mathfrak{ba}}_{\mathfrak{a}}$ lets $\mathfrak{b}$ is $c^{\mathfrak{baba}}_{\mathfrak{a}}$.
+It is $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$.
+Figure 5: Visual intuition for a 3RD order self. The organism constructs a 2ND order self for its 2ND order self, and so becomes aware that it is aware.
+6 The What and Why of Consciousness
+Up to now we have developed the conceptual toolbox that we use to dissolve the hard problem.
+We started with the observation that self-organizing systems such as a human body processes information to maintain itself in the face of change both within and without.
+An organism does not spring into existence understanding number systems or objects, but constructs them through interaction with the environment.
+One can regard them as policies according to which the environment is interpreted.
+To learn how to interpret the environment an organism must differentiate between states, must react to change, and learn policies according to the valence associated with that change.
+We will now argue that: i) there is something it is like to be an organism with a 1ST order self, ii) that subjective experience is the process of learning and enacting a hierarchy or causal identities, and iii) that access consciousness requires both 1ST and 2ND order selves.
+Because we do not presuppose representations, we explain how they emerge from attraction to and repulsion from contentless physical states.
+This means the foundation of information processing is valence.
+As we saw earlier, learned policy has valence, and it is a classifier of inputs.
+Those inputs are “information” in the mechanistic sense of the environment being in one state and not another.
+Hence, a policy is a classifier of information, but that information is not in a language, and it is not yet something labelled or quantified (until there is a policy that labels or quantifies).
+Clearly, prelinguistic self-organising systems must classify and attach value and disvalue to states and anticipated states to prioritise and make decisions.
+To learn is to sense “something” and, motivated by valence, construct a policy classifying that “something”.
+Physical states are contentless.
+In the absence of language, the difference between any two states can only be qualitative.
+In other words we hold that quality is fundamentally valenced, i.e.
+the product of actively sensing and being physically attracted and repelled within a given environment.
+The bold claim here is that information processing (i.e.
+exploring one’s body and environment) in self-organising systems such as the human body is necessarily qualitative.
+Note that this is different from experience in the sense usually defined in the literature.
+Experience in the sense of experiential phenomenological content cannot exist without experience in the sense of embodied exploration of the body and world as we defined here (see also Ciaunica in prep).
+To put it provocatively quality precedes quantity, and quantity is nothing more than the interpretation of quality.
+Quality comes first and experiences should be regarded on a continuum rather than a switch on/ switch off phenomena.
+All living systems experience the world through their bodies and as such, there is something what it is like to experience the world in that basic way (even when one is asleep or a baby).
+One can access those phenomenal, experiential aspects at a higher level, true, but by accessing them, it doesn’t mean that one ‘constructs’ consciousness or one becomes a conscious being.
+One is already consciously experiencing the world before one can explicitly access one’s own experiences in 2ND order selves.
+It follows that every policy learned in this way must classify a quality.
+Hence every such policy is a local state.
+A causal identity imbued with disposition by valence.
+There would be a policy for one’s act of smelling coffee.
+For perceiving one’s friend.
+There would be something different it is like to interact with a hostile version of that very same friend.
+The 1ST order self accompanies everything an organism does.
+It has a quality, so the 1ST order self is “what it is like” to be that organism.
+Put another way, Nagel’s question of “what it is like to be” a particular organism could be answered if one could somehow have that organism’s first order self [nagel1974].
+An organism does not make a decision to interpret information, it just reacts.
+What we call subjective experience is the process of learning and enacting a lattice of causal identities, and there is something that “has” all of those subjective experiences once there is a 1ST order self that is part of them all.
+This is where phenomenal consciousness begins, with experience as learning and exploration.
+One’s 2ND order selves would also have a certain qualitative character, and one’s 3RD order too.
+There would, however, be a very clear delineation between conscious and not.
+The absence of a 1ST order self would mean there would be no one policy we might point at to say “this is what it is like” to be an organism.
+No policy linking all interventions together, and so no “self” that experiences making them.
+Hence a 1ST order self must precede all others.
+Our arguments align with those of Merker, Barron and Klein, who have linked subjective experience to reafference [merker2005, merker2007, barron2016].
+We agree that reafference is key, but we provide a very different explanation of why and how.
+Their work presents biological evidence for subjective experience in organisms with reafference.
+In contrast, we derive the 1ST order self from first principles and explain why and how it is a classifies “what it is like” to be a particular organism.
+The 1ST order self also happens to be equivalent to reafference, so we arrive at the same conclusion as Merker from very different, mathematical premises.
+Hence these are complementary positions.
+7 Unifying Lower and Higher Order Theories of Consciousness
+The psychophysical principle of causality takes us from attraction to and repulsion from contentless physical states, up to higher order meta representations.
+This produces different orders of self.
+We suggest that the full richness of human subjective experience depends upon these different orders of self, as does human behavior [^18] We additionally speculate that different orders of learned self may interact through hierarchical planning in a distributed system, perhaps something like a connective core [shanahan2012].
+Arguably [boltuc2012], the full richness of human consciousness may be distinguished from phenomenal and access consciousness alone..
+Conscious “access” as a human has it involves meaningful report in the sense described by Grice, which as we’ve established requires 2ND order selves.
+This is very different from the mere “access” to information.
+Hence, we hold that contents of which one is “access conscious”, in the sense that they are available for reasoning and report, can only be the contents of 2ND order selves.
+However, supposing we could somehow contrive a system with 2ND order selves but no 1ST, then 2ND order selves alone would still not allow one to reason about how one’s interventions might affect the contents of 2ND order selves to communicate meaning.
+For that, one requires a 1ST order self.
+Hence access consciousness requires both 1ST and 2ND order selves.
+Hence our radical and provocative claim is that phenomenal consciousness without access consciousness is likely very common, but the reverse is impossible.
+For example, Block holds that phenomenally conscious content is phenomenal, whereas access conscious content is representational.
+In pancomputational enactivism there are only contentless states and the programs they imply.
+Any abstract “representational” content is just organised phenomenal content, clustered according to what causes valence.
+We don’t discard these two sorts of consciousness, but unify them by showing how phenomenal consciousness gives rise to access.
+Our framework thus makes a philosophical zombie impossible because there can be no perfect unconscious replica of a conscious human.
+The hard problem has it backwards.
+The question is not why qualia exist, but why anyone thinks representational contents can exist without first being learned through qualitative experience and discrimination.
+Far from suggesting there is no such thing as qualia, this suggests instead that there is no such thing as purely representational content in anything but the tools we construct and our interpretations of their behaviour.
+Cells are a material with agency [ball2023].
+A human is a “multiscale competency architecture” [fields2020, levin2021, levin2024].
+Our intelligence is a swarm intelligence; the high level goal directed behaviour of a swarm of cells.
+When we embody human abstractions (e.g.
+arithmetic) in silico (e.g.
+the x86 instruction set), we disconnect the high level goal directed behaviour from the low level behaviour (and the motivating valence) that gave rise to it [bennett2024c].
+We call the information embodied in the computer “representational” because it means something to us.
+However when we embody our abstractions in silico we disconnect them from the valence that motivated their construction [bennett2024b].
+This suggests there is no such thing as representational contents.
+They are a fiction we have invented because we struggle to reduce our own abstractions to their basic nature: the causes of valence.
+This ‘dissolves’ the hard problem by going a level down to the fundamental drive: stay alive!
+8 From Rocks to Einstein: The Hierarchy of Being
+To illustrate how our argument applies in the real world we describe stages of conscious organism.
+Each stage follows from scaling up supply and demand for WPO, through natural selection (see figure 6).
+For each stage we point out animals which are likely to be at least so conscious:
+$0:$ Unconscious (e.g. a rock)
+$1:$ Hard Coded (e.g. protozoan)
+$2:$ Learning (e.g. nematode)
+$3:$ 1ST Order Self (e.g. housefly)
+$4:$ 2ND Order Selves (e.g. cat)
+$5:$ 3RD Order Selves (e.g. human)
+Stage 0: Unconsciousness
+Stage zero may be understood as the consciousness we are willing to assign to a rock: more exactly, the lack of consciousness so assigned.
+  • Example: A rock.
+Stage 1: Hard Coded
+Stage one refers to adaptations hard-wired by natural selection. This is the sort of preset behaviour that allows complexity to persist [heylighen2023b] in a reasonably consistent environment.
+  • What: Hard-coded adaptations. Habituation and sensitization.
+  • How: The extension of fit behaviour is learned by natural selection and hard-coded into the organism as a policy (in DNA, form, the local environment etc).
+  • Why: If the environment is very predictable, it may be more efficient to hard-code fit behavior.
+  • Example: Single-celled protozoan.
+Stage 2: Learning
+Stage two introduces learning.
+To learn an organism must store, classify and order historical examples by valence.
+However there is not something it is like to be stage two, because there is no locus of “self”.
+A biological example of such a decentralised nervous system is the cubozoan box jellyfish Tripedalia cystophora.
+Even Tripedalia cystophora was recently shown to be capable of associative learning [bielecki2023].
+An entirely distributed control system can “learn”.
+Likewise, stage two is exemplified by the nematode C.
+elegans [rankin2022, willett2018], which has a centralised nervous system and exhibits some ability to adapt with experience.
+However, the absence of a “self” prohibits cause and effect reasoning, which as others have already pointed out must limit spatial, navigational abilities [barron2016].
+When starved C.
+elegans exhibit “increased locomotion and dispersal in a random, rather than directed, search” [barron2016, luersen2014, artyukhin2015], whereas something like a bee or an ant can recall and navigate to previously discovered food [wehner2013, seeley1995page295, oades1978].
+A nematode can learn, but the absence of a self for the purpose of causal reasoning limits adaptability.
+  • What: Learning.
+  • How: Valence.
+  • Why: An organism that can learn can survive in more circumstances than one that cannot.
+  • Examples: Jellyfish, nematode.
+Stage 3: 1ST Order Self
+This is where phenomenal consciousness begins, with a 1ST order self.
+In biological terms this also implies reafference, which others have argued is the key to subjective experience, albeit for different reasons than what we have argued here [merker2005, merker2007, barron2016].
+They identified a housefly as a good example of where subjective experience may begin, and we concur.
+We also hold this is where an organism might be said to have intent.
+Intuitively, the policy that motivated behaviour is the intent of that behaviour [^19] In the same way declarative and imperative programs are equivalent [howard1980]..
+The “weaker” a policy is, the more behaviours it motivates.
+Conversely, if one’s actions share anything in common it is the intent that motivated them.
+The more diverse the actions, the more “general” or “high level” the intent they share.
+For example, the action of eating tends to involve the intent of satisfying hunger.
+If a policy is implied by all of an organism’s behaviour, then it might have motivated that behaviour.
+While this is an unusual form of words, we contend that an organism is such a policy.
+A stage three there is a self to feel simple hunger and pain.
+However, communication as described by Grice would be impossible [grice1969], because a stage three organism has no self-awareness, to represent another’s perception of their intent [bennett2023d].
+Nor would a stage three organism be able to conceive of its own death, or shame, because it cannot conceive of itself.
+  • What: A 1ST order self. Reafference. Phenomenal consciousness.
+  • How: Embodiment in which intervention is not identical to observation.
+  • Why: Accurate prediction of consequences of interventions. For example, a fly must distinguish between having moved, and the environment having moved, to navigate.
+  • Example: Housefly.
+Stage 4: 2ND Order Selves
+Stage four is the 2ND order self, and importantly this is where we hold access consciousness begins because it is where information is available for report in the Gricean sense.
+In other words, access consciousness follows phenomenal consciousness.
+The ability of ravens to intentionally deceive [bugnyar2002] suggests they are at least stage four.
+Raven $\mathfrak{a}$, aware that it is being observed by raven $\mathfrak{b}$, will act as if it is hiding food in one location to mislead $\mathfrak{b}$, but will then move the food in another location unobserved by $\mathfrak{b}$.
+$\mathfrak{a}$ seems to predict not just the intent of $\mathfrak{b}$ (to steal the food), but $\mathfrak{b}$’s perception of $\mathfrak{a}$.
+It seems intuitively likely that dogs and cats have second order selves, as they must hunt reasonably intelligent animals and must anticipate how their actions are perceived.
+For example, a cat anticipates its prey will flee when it is observed, and hides.
+  • What: Access consciousness. Theory of mind. Self-awareness. Inner narrative.
+  • How: Selection pressures that demand theory of mind.
+  • Why: A 2ND order self is necessary to anticipate, manipulate and communicate intent.
+It allows one to anticipate how others will perceive one’s behaviour.
+This could be anticipating the reactions of predator or prey, or navigating a social hierarchy.
+  • Example: Cats, dogs, ravens.
+Stage 5: 3RD Order Selves
+A 3RD order self facilitates complex planning, particularly of a social nature.
+A 3RD order self is a 2ND order self for one’s 2ND order self.
+Humans appear to possess this, because we are aware that we are aware.
+The seemingly altruistic behavior observed in Australian magpies [potvin2022] suggests they may also have 3RD order selves.
+  • What: Meta self-awareness. Inner narrative in which actors have inner narratives.
+  • How: More accurate prediction and planning.
+  • Why: Because a social organism must predict complex social dynamics.
+  • Example: Human. Highly intelligent animals such as Australian magpies may also be this conscious.
+Figure 6: Overview of stages and orders of self.
+9 Conclusions: Why Nature Does Not Like Zombies
+In this paper we formally unified lower order states and higher order meta-representations, dissolving the hard problem of consciousness.
+We described a multilayered formalism illustrating how biological self-organising systems become phenomenally conscious when they construct a 1ST order self.
+A human lacking a 1ST order self could not perform causal reasoning needed to adapt as humans evidently can [pearl2018].
+As previous research pointed out [bennett2024a], the computer metaphor with the seminal distinction between hardware and software is a simplification.
+Rather, software is a state of hardware.
+Nature privileges efficiency over abstract simplification.
+Compare the vast quantities of both training data and energy required by a large language model, to the small quantities humans need to solve a problem.
+Biological systems are more efficient, because they are adaptive at a every level [levin2024, bennett2024c].
+Hence, rather than trying to explain the mind in the abstract like software, we started at the level of the embodied organism.
+One consequence of our approach is that it places the phenomenal quality of conscious experience before access consciousness.
+We have shown a consistent definition of access consciousness requires 1ST and 2ND order selves.
+Phenomenal consciousness arises first, access comes later.
+Unlike panpsychism, we don’t believe rocks are conscious.
+Only self-organising systems that need to adapt, motivated by valence, while keeping track of the self are conscious.
+Consciousness is a necessary adaptation, and a philosophical zombie is impossible.
+Furthermore, the conceivability of zombies depends on the ability to duplicate disposition but not phenomenal experience.
+Carruth argued that if qualia are dispositional, then that is impossible [carruth2016].
+Our arguments align with Carruth’s.
+A causal identity is inherently dispositional, as is embodiment.
+In summary, we have argued that access consciousness at the human level is impossible without the ability to hierarchically model i) the self, ii) the world and others and iii) the self as modelled by others.
+These follow from learning and enacting a lattice of causal identities, if the scale and incentive preconditions are met.
+There is something it is like to be an organism with a self, and that is where phenomenal consciousness begins.
+Our proposal lays the foundation of a formal science of consciousness, deeply connected with natural selection rather than abstract thinking, closer to the organic facts of humanity than the conceptual fiction of zombies.
+Compliance with Ethical Standards
+Funding
+This work was supported by a Fundaçao para a Ciencia e a Tecnologia (FCT) grant PTDC/FER-FIL/4802/2020 ; and 2020-02773 CEECIND FCT to A.C.
+Ethics Approval
+This article does not contain any studies with human participants or animals performed by any of the authors.
+Appendix
 For convenience of reference, we have placed all definitions here.
 We have aimed to make the gist of the paper understandable without the math, hence the reader can either skip this section or refer back to it if needed.
-Many of the definitions have been adapted from a variety of preceding work [32, 34, 35, 42, 43, 44].
+Many of the definitions have been adapted from a variety of preceding work [bennett2024a, bennett2023b, bennett2023c, bennett2023d, bennett2024b, bennett2024c].
 They are referred to in the body of the paper when they become relevant, in the order in which they appear here.
@@
   • We assume a set $\Phi$ whose elements we call states.
-  • A declarative program is $f\subseteq\Phi$, and we write $P$ for the set of all declarative programs (the powerset of $\Phi$).
+  • A declarative program is $f\subseteq\Phi$, and we write $P$ for the set of all programs (the powerset of $\Phi$).
   • By a truth or fact about a state $\phi$, we mean $f\in P$ such that $\phi\in f$.
@@
   • We say $x$ and $y$ are equivalent iff $E_{x}=E_{y}$.
-(notation) $E$ with a subscript is the extension of the subscript666e.g. $E_{l}$ is the extension of $l$..
+(notation) $E$ with a subscript is the extension of the subscript212121e.g. $E_{l}$ is the extension of $l$..
 (intuitive summary) $L_{\mathfrak{v}}$ is everything which can be realised in this abstraction layer.
@@
 (optimal proxy) There is no proxy more sample efficient than weakness.
-The weakness proxy formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [34]).
+The weakness proxy formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [bennett2023b]).
 (intuitive summary) Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy.
@@
 The lower level the child from which an agent generalises to parent, the “faster” it learns, the more sample efficient the proxy.
-The most sample efficient proxy is weakness [34, prop.
+The most sample efficient proxy is weakness [bennett2023b, prop.
 1, 2], which is why we’re using it here.
@@
 If $obs\in L_{\mathfrak{v}}$ is sensorimotor activity we interpret as an “observed event”, and $int\in L_{\mathfrak{v}}$ is an intervention to cause that event, then $obs\subset int$ (because $int$ could not be said to cause $obs$ unless $obs\subset int$).
-(intuitive summary) An intervention is action undertake an organism or other agency, in the sense described by Pearl [45].
+(intuitive summary) An intervention is action undertake an organism or other agency, in the sense described by Pearl [pearl2018].
 Intuitively, if “$int$” and “$obs$” are events which have happened, then we say that $int$ has caused $obs$ if $obs$ would not have happened in the absence of $int$ (counterfactual).
@@
   4. 4.
-    Second order selves:
+    2ND order selves:
     1. (a)
@@
   5. 5.
-    Third and higher order selves: meta self-reflexive consciousness (human level hard consciousness) is achieved when $\mathfrak{o^{3}}\in\mathfrak{p}_{\mathfrak{o}}$.
+    3RD and higher order selves: meta self-reflexive consciousness (human level hard consciousness) is achieved when $\mathfrak{o^{3}}\in\mathfrak{p}_{\mathfrak{o}}$.
 Proposition
@@
   3. 3.
-    From [34, prop.
+    From [bennett2023b, prop.
 3] we have proof that weakness is the optimal choice of proxy to maximise the probability of generalisation from child to parent is the weakest policy.
@@
 $\square$
-3 Back to Foundations
-Rather than presupposing local states or an abstraction layer, we must start from very basic first principles to formalise all conceivable environments [32, 43] in definition 1.
-Where there are things, we call those things an environment.
-Where things differ, we have different states of that environment.
-Hence, we begin by formalising the environment as a set of contentless global states $\Phi$.
-We don’t assume there is any internal structure to the states $\Phi$ contains, but rather define declarative “programs” in terms of relations between these irreducible, contentless states.
-The set $P$ contains all such programs.
-The powerset of $P$ is every aspect of the environment, because every aspect of any conceivable environment must be set of such declarative programs, which is a subset of $P$.
-Axiom 1: When there are things, we call these things the environment.
-Axiom 2: Where things differ, we have different states of the environment.
-Universality Claim: Axioms 1 and 2 hold for every conceivable environment.
-We don’t need to concern ourselves with the internal structure of environment states beyond this.
-Declarative programs return true or false, so a declarative program is a “fact” if it is true, and the current state of the environment is the set of all “facts”.
-Truth is determined with respect to states.
-If time is one way in which things differ, then there is only one state at a time (dimensions are implicit).
-The only meaning these programs have is how their truth values relate in different environmental states.
-Hence this is a representationless form of pancomputationalism [30].
-3.1 Natural Selection and Embodiment
-We then assume a process of natural selection. This produces embodied organisms.
-A set of facts can represent anything [^8] Some may object, pointing out that this ignores composition.
-However, the application of a function is a fact regardless of whether it takes another function as input., so the body of an organism (like everything else in every conceivable environment) must be a set of declarative programs.
-The vocabulary can be seen as hardware and software operating together, avoiding issues related to solipsism [46].
-No two organisms have the same vocabulary, because that would mean they are the same body.
-However, it is not enough to just say a body is a set of declarative programs. Bodies tend to imply finite resource constraints, so we adopt one more axiom from [43].
-Axiom 3: All aspects of the environment are spatially extended [43].
-This means that a body can occupy only a finite number of states, which is formalised in definition 2. A body is a vocabulary $\mathfrak{v}$ with finitely many elements.
-The vocabulary of an abstraction layer is a subset of the aforementioned $P$ in definition 1.
-$\mathfrak{v}$ implies a formal language $L_{\mathfrak{v}}$ of interaction between body and environment.
-A statement in this formal language is just a set of “programs”, which is an interaction, allowing us to simultaneously discuss cognition in enactivist and computational terms.
-Statements have truth conditions with respect to environmental states, and every statement has an extension.
-The extension of a statement is the set of all statements which are supersets of the first statement (the set of all other statements by which the first statement is implied).
-Extension is important we can relate statements by their truth conditions, forming a lattice.
-Most importantly, we avoid the distinction between software and hardware and so avoid computational dualism [^9] Computational dualism [32] is the distinction between software mind and hardware body commonly used in artificial intelligence.
-Computational dualism is a simplification that has led to erroneous claims regarding the behaviour of artificial intelligence [46].
-Software is a state of hardware, not a distinct object.
-What software does depends entirely upon the hardware that interprets it..
-3.1.1 Self-Organising Systems as Self and World Constraints
-Both snowflakes and human bodies are self-organising systems, yet only the latter are regarded to display conscious experiences.
-What exactly in the self-organisation model of a body make it radically different from the snowflake?
-To tackle this question, we need to introduce the notion of self- and world constraints.
-Given that we have an environment, and an abstraction layer that implies an embodied formal language, we can talk about computation with inputs and outputs by just treating everything as embodied statements embedded and enacted within the pancomputational environment.
-Given an input $i$, the set of all possible outputs is the extension $E_{i}$ of that input.
-This is because if $i$ is realised by the environment, then the environment is constrained to only those states that realise $i$, which constrains what other statements can be realised.
-We can use this fact to talk about “policies” as embodied [^10] As a state of hardware rather than software, which is how the term policy is normally used in reinforcement learning.
-constraints on behaviour.
-A policy is a statement whose extension constrains outputs, like a causal intermediary in machine functionalism [47].
-If behaviour is “motivated”, then statements have valence determined by natural selection and only a subset of the statements a body could make are “fit”.
-As a result, an organism will self-organise to express some statements, but not others.
-This is formalised by the $\mathfrak{v}$-task as in definition 3, so called because a task exists in the context of a vocabulary $\mathfrak{v}$.
-If $\mathfrak{v}$ is the vocabulary of a body, and $\mathfrak{v}$-task $\mu=\langle I_{\mu},O_{\mu}\rangle$ is fit behaviour, then $I_{\mu}$ is all the statements that body can express in which it is possible to remain fit, and $O_{\mu}$ is all the statements that body can express in which it remains fit.
-The extension $E_{I_{\mu}}$ of $I_{\mu}$ would be every output that it is possible to choose given the inputs $I_{\mu}$, but only a subset of those $O_{\mu}\subset E_{I_{\mu}}$ are correct outputs.
-In other words, this defines a self-organising behaviour as a constraint on outputs, given inputs.
-Conversely, a system which considered all outputs to be correct would not be self-organising.
-When we say a system is self-organising, mean it is goal directed and can adapt to serve that goal, which for the purposes of this paper is the fundamental goal of self-survival and self-reproduction.
-3.1.2 Inference
-Every statement in $L_{\mathfrak{v}}$ implies a constraint, because there are only so many outputs that can be expressed by a body at the same time as any given statement.
-That is what the extension of a statement represents.
-If $i$ is a statement which is true, then the only possible statements a body can express is $E_{i}$.
-As a result, a body might be express a statement $\pi$ (meaning $\pi$ is true), and $\pi$ would then constrain the body to only correct outputs $O_{\mu}\subset E_{I_{\mu}}$ if $O_{\mu}=E_{I_{\mu}}\cap E_{\pi}$.
-We call a constraining statement a policy.
-A policy constrains outputs given inputs.
-A correct policy is one that constrains outputs to only correct outputs.
-For the sake of intuition, think of “correct” as “fit” according to natural selection (although this need not always be the case).
-This is more formally expressed in definition 4.
-“The best model of the world is the world itself” - Rodney Brooks [48]
-Importantly, to reiterate, one does not need a model of the world; one must embody a policy that constrains behaviour to only fit behaviour.
-The history of an organism’s interactions is a subset of the statements that can be made in the abstraction layer; hence the history of an organism is also a $\mathfrak{v}$-task.
-The history of the organism that remains alive is a subset of fit behaviour for that organism.
-Examples of “fit” self-organising behaviour.
-Hence, we could denote an organism’s history to be a $\mathfrak{v}$-task $\mathfrak{h}$, where each input and output in that task is behaviour the organism has exhibited and everything that was involved in that behaviour (the interaction between organism and environment).
-Fit behaviour would be another task $\mu$, and if the organism remains fit then $\mathfrak{h}\sqsubset\mu$ [^11] One task is a child of the other, meaning the inputs and outputs are subsets of the parents’ inputs and outputs..
-3.1.3 Learning
-Learning, in computational terms, tends to be understood as the process of modelling the program which caused data.
-Constructing a “world” model.
-Here, we just need to explain how an organism gets $\mu$ from $\mathfrak{h}$.
-We don’t need a world model, but a way to constrain outputs to just those that are fit.
-Like everything else, that constraint is once again a statement in the formal language $L_{\mathfrak{v}}$.
-Those examples in $\mathfrak{h}$ could have been generated by any one of a set of policies $\Pi_{\mathfrak{h}}$.
-These are not representations of possible causal intermediaries in the machine functionalist sense, but just parts of the outputs from which the outputs could be derived given the inputs.
-Every $\mathfrak{v}$-task (and thus every self-organising system) implies such a set of policies, and not all of those policies will constrain outputs the same way given new inputs.
-Some policies will “generalise” to imply fit behaviour in unfamiliar circumstances, meaning correct outputs given a new set of inputs (a parent task of the organism’s history $\mathfrak{h}$).
-If those policies imply fit behaviour given new inputs, then the organism will remain fit.
-If an organism learns, then it is adapting to embody a policy that generalises in accord with its motives (which, set by natural selection, will tend to favor fit behaviour).
-“I survive therefore my model is viable.” - Mark Solms [49]
-According to previous experimental and mathematical results, the optimal strategy to learn and adapt as fast as possible [^12] Meaning to converge on a fit policy given the smallest history possible, to realise fit behaviour.
-is to prefer “weaker” policies, meaning those with larger extensions [34].
-Formally, the policy which generated outputs $O$ given inputs $I$ is most efficiently identified by constructing a policy $\pi$ such that $\pi$ generates $O$ from $I$, and $\pi$ implies the weakest constraint that can be implied while still generating $O$ from $I$.
-This maximises adaptability, because it allows a self-organising system to construct fit policies from a shorter history [35].
-This is described in formal terms in definition 5.
-4 Relevance Realisation Through Causal Learning
-The enactive process of relevance realisation can be framed as policy learning.
-Natural selection prefers more adaptable organisms, so we assume it optimises for organisms that optimise for weaker policies.
-We call this weak policy optimisation (WPO).
-Policies determine how inputs are mapped to outputs.
-Interpretation.
-Fit policies must correctly predict causes of valence.
-Hence by constructing fit policies, an organism must realise what is relevant.
-A weaker policy implies all the more specific versions of itself, meaning those that more tightly constrain outputs by having a smaller extension.
-Hence policy learning implies a lattice of policies that vary in weakness.
-Every fit policy is a “causal identity” for something, for example an object like a “food” or a more abstract concept like “pain”.
-For convenience, we’ll now expand upon the concept of organism given in definition 6.
-All organisms must have preferences (will make some decisions and not others), regardless of how those arise.
-All organisms must have policies that reflect those preferences, and every policy implies tasks, so we can define preferences as an binary relation over tasks.
-Correct or incorrect choice of policy affects the organism’s existence and survival.
-We also need to introduce the notion of interaction between organisms.
-Organisms interact when they “affect” one another (in the physical rather than psychological sense of the word).
-This is formally defined in 9.
-Now, artificial intelligence and machine learning [50, 51, 52, 53] are concerned with engineering adaptive agents.
-In that context, causality has now become a mainstream topic of research.
-Causal learning is demonstrably necessary to thrive in an interactive setting [45, 54].
-Where the causal graph is known in advance (for example if we were to be measuring the efficacy of medical interventions), this issue could be resolved by the use of causal language (such as “Do Calculus” [55]) to represent an intervention from outside the system the graph describes.
-Such an intervention represent agency, in that an organism observing the environment draws conclusion about that environment, and then intervenes to change the environment.
-To illustrate this point, consider the following example [35].
-Suppose an organism named Bob is attempting to learn and predict the environment.
-Now assume Bob has observed Alice wearing a raincoat only when it rains, and that Bob has observed rain only at those times when Alice has been observed wearing a raincoat.
-If we represent the raincoat observation with a binary variable $O\in\{true,false\}$ and the advent of rain with a similar variable $R\in\{true,false\}$, and if Bob draws conclusions according to Bayesian probability, then Bob’s observations will lead to the conclusion that $p(R=true\mid O=true)=1$.
-This means Bob believes that if Alice wears a raincoat, then it must be raining.
-Now lets permit Bob to interact with its environment.
-Assume Bob wants it to rain.
-Based on the belief $p(R=true\mid O=true)=1$, Bob may conclude that forcing $O=true$ by holding a gun to Alice’s head and demanding Alice wear a raincoat will cause it to rain.
-This is obviously absurd.
-$O=true$ does not represent the event $q=$“I coerced Alice into wearing a raincoat”, but an entirely different event $v=$“Alice decided to put on a raincoat for the same reason I have observed Alice wearing a raincoat in the past”.
-To accurately represent the environment, we need a way of representing that $p(R=true\mid q)=p(R=true)\neq p(R=true\mid v)=1$, meaning we need a way to represent $q$ and $v$.
-To illustrate the problem visually, we started with the acyclic graph
-$O$$R$
-and our intervention disconnected rain from the choice of clothing:
-$O$$R$
-This can be resolved by introducing a “do” operator that we apply to a variable $O$ to obtain $do[O=o]$, to represent the fact that an agency from outside the system has intervened to assign a value to $O$, so that we can represent $p(R=true\mid do[O=true])=p(R=true)\neq p(R=true\mid O=true)=1$.
-Thus, the aforementioned $q$ is equivalent to $do[O=true]$, while $v$ is equivalent to $O=true$.
-This works very well for the purpose of evaluating treatments and interventions by humans.
-However, this merely establishes that we need to account for causality in cognition.
-It does not explain how one would come to know all the objects involved (to represent them as variables), or how they relate to one another causally.
-The inference of causal relations can be understood in two steps.
-First, we must show that the “do” operator is equivalent to including additional variables in a causal graph.
-Second, we must show how it is possible to learn the objects that best represent cause and effect relations.
-The matter of which cause and effect relations are learned is determined by valence, and so the objects learned are statements classifying causes and valence.
-Substituting The “Do” Operator
-The very notion of an “intervention” from outside a system echoes mind-body dualism, in that it treats an organism performing an intervention as something apart from the system in which it intervenes (its environment).
-To illustrate what this means, consider another example involving the variables $O$ and $R$ from before.
-Assume we again have Bob who constructs a causal graph of the environment.
-Assume Alice exists in that environment.
-From Bob’s perspective, Alice is just a part of the environment represented by a variable $A$ in Bob’s causal graph.
-$O$$A$$R$
-Now assume Bob observes Alice taking an action that changes an aspect of the environment, represented by the variable $O$ (for example, Bob observes Alice putting on the raincoat).
-From the perspective of Bob, Alice’s action is the assignation of a value to a variable $A$, and to the variable $O$.
-There is no need to involve a do operator in this scenario because we can already represent that $p(R=true\mid A=x,O=true)=p(R=true)\neq p(R=true\mid A=y,O=true)=1$ (because Alice is part of the causal graph).
-This raises the question; if we do not require a $do$ operator to represent the actions of Alice, then why would we need it to represent an intervention by Bob?
-Couldn’t we represent the same information by introducing a new variable?
-The answer is yes [35, 56].
-$O$$B$$A$$R$
-Of course, this does not solve the problem of how the causal graph is learned in the first place, and this is where relevance realisation comes in.
-5 Relevant Causal Identities
-The extension of fit behaviour is an extensional definition of what an organism is compelled to want by natural selection.
-Likewise, an intension of fit behaviour would be any policy that results in fit behaviour, delivering the organism what it is compelled to want.
-Indeed, this was how the formalism of tasks we employ originated, to formalise the inference of norms [57] and explain why some abstractions are formed, and not others.
-As mentioned earlier in the context of heuristics or “proxies” for learning, the optimal choice of policy for generalisation is the weakest.
-Such a policy must, by definition, isolate those things which cause valence.
-Put another way, “fitness” in an interactive setting is likely to require an organism be able to distinguish passive observation of an event from having intervened to cause that event (reafference [36, 37]).
-A preference for weak hypotheses facilitates this via construction of causal identities.
-This divides the environment up into objects, events and anything else which is sufficiently relevant to the organism’s motivations.
-To explain how, we must first explain how intervention can be differentiated from observation without presupposing variables, and then the circumstances under which it will be differentiated [35].
-5.1 Causal Learning
-In accord with [42, def.
-6] we assume a vocabulary $\mathfrak{v}_{\mathfrak{a}}$ belonging to an organism we’ll denote $\mathfrak{a}$.
-A “cause” in the context of this formalisation is not a variable but a statement $l\in L_{\mathfrak{v}_{\mathfrak{a}}}$ in that formal language.
-The raincoat example would involve $obs,rain\in L_{\mathfrak{v}_{\mathfrak{a}}}$ such that:
-[…]
-[…]
-[…]
-$int$$obs$$rain$
-In other words, so long as $c\neq int$, the intervention can be differentiated from the passive observation. We formally define this in definition 10.
-This being the case, any set $c\subseteq int-obs$ could be used to identify the party undertaking the intervention, which is why $c$ is referred to as a “causal identity”.
-It distinguishes the intervention $int$ from the passively observed effect $obs$, like reafference in living organisms.
-However, the above only considers one intervention.
-A weaker or more general causal identity would be one that is shared by more intervention.
-For example, we might have two different interventions $a_{1}$ and $a_{2}$, with the observed effects $c_{1}\subset a_{1}$ and $c_{2}\subset a_{2}$, and causal identities $a_{1}-c_{1}=i_{1}$ and $a_{2}-c_{2}=i_{2}$.
-If $i_{3}=i_{1}\cap i_{2}\neq\emptyset$, then $i_{3}$ is a causal identity that is present in two interventions.
-Being a question of classification, we can express a causal identity in relation to a $\mathfrak{v}$-task, for which the causal identity is a policy (expanding upon [35, def.
-10]), giving us definition 11.
-Example 1.
-Suppose we have organisms $\mathfrak{a}$ (Alice) and $\mathfrak{b}$ (Bob), and that the inputs Alice has experienced so far are $I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}$.
-These can be divided into those in which Bob affected Alice $S^{\mathfrak{b}}_{\mathfrak{a}}$ and those in which Bob did not $S^{\lnot\mathfrak{b}}_{\mathfrak{a}}=I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}-S^{\mathfrak{b}}_{\mathfrak{a}}$.
-By affecting Alice, Bob has intervened in Alice’s experience.
-Alice can construct a causal identity $b$ for Bob corresponding to interventions $INT=S^{\mathfrak{b}}_{\mathfrak{a}}$ and observations $OBS=S^{\lnot\mathfrak{b}}_{\mathfrak{a}}$.
-5.2 Ascribing Intent To Other Objects
-The “do” operator is necessary to discern the difference between an event one has caused, and that same event passively observed.
-However, what is passive observation if not the result of an intervention by something other than oneself?
-The distinction between “intervention” and not is misleading.
-Everything is an intervention.
-The question is not “is this an intervention” but “by whom was this intervention made?”.
-To illustrate this point we return to Alice, Bob and her raincoat. Earlier, we arrived at the following graph in which Bob’s intervention was given by $int$.
-$int$$obs$$rain$
-However, what if a third person Larry puts the coat on Alice?
-Surely Bob can observe this, and so Bob’s observation of Larry’s intervention is $v\in L_{\mathfrak{v}_{\mathfrak{a}}}$ such that $obs\subset v$.
-To account for this, Bob can construct a causal graph as below (with $b.$ representing Bob and $l.$ representing Larry).
-$b.\ int$$l.\ int$$obs$$rain$
-Bob’s causal identity for himself $c_{b}\subset b.\ int-obs$ only represents the intervention by himself.
-However, now we can see that Bob must also construct a causal identity $c_{l}$ for Larry, where the $c_{l}\subset l.\ int-obs$.
-More generally, for an organism $\mathfrak{a}$ with sensorimotor language $L_{\mathfrak{v}_{\mathfrak{a}}}$ to construct a causal identity for an object $\mathfrak{b}$, it must first be the case that $\mathfrak{a}$ is affected by $\mathfrak{b}$ [42], to satisfy the incentive precondition for causal identity.
-It depends on motive, or valence.
-Recall that to be affected is formally defined in definition 9.
-Assume an organism $\mathfrak{a}$ is affected by $\mathfrak{b}$ given inputs $INT$, and not affected given inputs $OBS$.
-To then attribute the contents of $INT$ to one specific entity, there must be something in common between the members of $INT$ caused by $\mathfrak{b}$ that is not shared by any member of $OBS$ caused by something else (in other words it must be at least possible for $\mathfrak{a}$ to discern the existence of $\mathfrak{b}$).
-The contents of $INT$ are “interventions” by $\mathfrak{b}$ and by learning $c$, a corresponding causal identity, $\mathfrak{a}$ can discern the existence of $\mathfrak{b}$.
-This is not to say that $\mathfrak{b}$ decides (chooses an output) or is even alive.
-What is important is that $\mathfrak{b}$ affects $\mathfrak{a}$, making it possible to discern when these interventions are a consequence of $\mathfrak{b}$’s existence.
-There are certain preconditions for the existence of a causal identity corresponding to $INT$ and $OBS$.
-5.3 Preconditions
-First, the vocabulary $\mathfrak{v}_{\mathfrak{a}}$ of an organism $\mathfrak{a}$ must be large and expressive enough to ensure that observations are distinguishable from interventions (in other words, the causal identity must be a subset of the vocabulary).
-We must have sufficient scale.
-Second, it must be in the organism’s interest to make this distinction.
-Inference is only possible if some states are preferable to others.
-It is a value judgement.
-As Hume pointed out, one cannot derive what “ought” to be from a statement of what “is”.
-Natural selection provides a notion of what ought to be, by eliminating anything which ought not.
-  1. 1.
-    The scale precondition requires $\mathfrak{v}$ contain the causal identity.
-  2. 2.
-    The incentive precondition is that fitness demands the causal identity.
-5.4 Realising Lower Order States And Higher Order Meta Representations
-Causal identities are not constrained to other organisms, but we use other organisms for intuition.
-This is because the causal identity for an object is a policy predicting its behaviour.
-This makes intuitive sense when that object is a self-organising system with goals.
-However, it also applies to inanimate objects like rocks.
-A rock still has behaviour, but it does not have intent.
-Yet by constructing a policy representing the rock’s intent, one may predict what the rock will do (e.g.
-damage a tree when thrown).
-That ability to predict is what matters.
-The weaker the causal identity, the more pervasive the “identity” in the sense of being part of more interventions.
-Learning is not just about constructing policies but joining them together into weaker, more abstract policies.
-By learning policies that correctly identify the causes of valence at different levels of abstraction, the organism engages in relevance realisation.
-Importantly, embodiment imposes severe limitations in the form of a finite vocabulary.
-In definition 6 we formalise this for the sake of explanation, formalising an organism $\mathfrak{o}$ with a set of policies the organism has realised.
-Each and every policy implies a $\mathfrak{v}$-task.
-A $\mathfrak{v}$-task is a triadic relation between inputs, outputs and policies which resembles Peircean semiosis [58, 42] of sign, referent and interpretant.
-Hence, in definition 7 we formalise this as a “protosymbol” [^13] Proto because it is something more primitive than a symbol as conceived of by Peirce.
-system $\mathfrak{s}_{\mathfrak{o}}$ for the organism $\mathfrak{o}$.
-Interaction is defined in terms of choosing an output, which means there is inevitably a policy and a policy implies a $\mathfrak{v}$-task, so definition 6 includes a preference order over tasks, which is used in definition 8 to formalise the interpretation of inputs (in terms of constraints, rather than an algorithm).
-In particular, the organism must always act according to an interpretation, and some interpretations imply others.
-Related as they are in a lattice, protosymbols are analogous lower order states and higher order meta representations.
-Tasks exist in a “generational hierarchy”.
-They are not mutually exclusive.
-Higher level tasks are more general, and have fewer policies because only very weak policies could complete them.
-Some have framed consciousness as a problem of moving from unary, to dyadic, to triadic relations [59].
-Similarly, we have gone from unary states, to dyadic declarative programs, to triadic tasks and protosymbols.
-The question now is when does an organism realise a causal identity for when it causes valence? When does it construct a causal identity for itself?
-6 Multi-Layered Self-Organisation
-In an interactive setting, it is necessary to correctly predict the effect of one’s actions [45] to satisfy even very basic goals.
-As capacity for weak policy optimisation (WPO) scales, a greater variety of concepts can be learned.
-Hence as WPO scales, progressively higher orders of ‘causal-identity’ for one’s self-related information processing are constructed [35].
-This lets us frame the construction of embodied selves in developmental [60] and evolutionary terms.
-We turn to this discussion now.
-6.1 The First Order Self
-A first order self (1ST henceforth, see figure 1) is functionally equivalent to reafference, which has been observed in the mammalian mid-brain and insect central cortex [37, 38].
-Reafference lets an organism discern the consequences of its actions, and so a 1ST order self is necessary for accurate inference in an interactive setting in the same sense as Pearl’s ‘do’ operator [45].
-A 1ST order self formally defined in definition 13.
-It serves as the locus of self-related information processing and experience [35], allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction.
-But also, to anticipate future planned interactions (for example, an insect navigating its environment [38]), and recollections of past interactions (it is a subset of all those relevant “statements” in the formal language).
-Figure 1: Visual intuition for a 1ST order self. The organism constructs a causal identity for itself and can relate that to observed events.
-Natural selection prefers efficiency.
-The absence of a single, centralised causal identity can create inefficiencies [^14] The effect of insufficiently weak representations can be observed in large language models like GPT-4 [42, 61], where they fail to make consistent statements about an object because represent it as multiple objects..
-Decentralised or asynchronous control might be advantageous in some circumstances (e.g.
-an ant colony), but not others (e.g.
-in an individual human body).
-An organism that has a distributed control system might have a “self” for each part of that system, and perhaps the co-ordination of parts might seem to suggest the existence of a centralised “self”.
-Redundancy might be useful in some circumstances, but there is also a cost - more data may be required to learn and adapt, because the optimal choice of policy for accurate generalisation is the “weakest” [34].
-If the scale and incentive preconditions are met for a causal identity for one’s self, then the organism must realise one.
-This might be partly hard-coded, but our focus is on inference by the organism itself.
-Preconditions are formally defined in 14.
-6.2 The Second Order Selves
-Survival may demand organism $\mathfrak{a}$ infer $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s interventions (to see one’s self as if through another’s eyes [42]).
-This is called a second order self (2ND henceforth).
-We argue that if access conscious contents are available for communication in the human sense, then they must be communicable in the Gricean sense [62, 63].
-Grice argued that communication is about the inference of intent.
-If person $\mathfrak{a}$ and $\mathfrak{b}$ are talking, then the meaning $m_{\mathfrak{a}}$ of what $\mathfrak{a}$ says is whatever $\mathfrak{a}$ intends $\mathfrak{b}$ understand.
-The meaning $m_{\mathfrak{b}}$ that $\mathfrak{b}$ understands is whatever $\mathfrak{b}$ thinks $\mathfrak{a}$ wants $\mathfrak{b}$ to think.
-$\mathfrak{b}$ has understood what $\mathfrak{a}$ means if $m_{\mathfrak{b}}$ approximates $m_{\mathfrak{a}}$.
-This can happen only if $\mathfrak{a}$ can predict with reasonable accuracy what $\mathfrak{b}$ thinks $\mathfrak{a}$ thinks, and $\mathfrak{b}$ can predict what $\mathfrak{a}$ thinks $\mathfrak{b}$ will think upon hearing an utterance.
-In other words, both $\mathfrak{a}$ and $\mathfrak{b}$ must have 2ND order selves that are good approximations.
-Yes, there are other aspects to communication.
-However, here we are talking about consciousness.
-Access conscious contents are those available for reasoning and report.
-It follows [^15] If the definition of access consciousness is to be consistent with reasoning and report as exhibited by conscious humans.
-that access conscious contents must in principle be communicable in the sense Grice described.
-As such, we argue contents available for communication can only be the contents of 2ND order selves, which means only an organism with 2ND order selves can be considered to have access consciousness.
-2ND order selves also explain attention and self-awareness.
-An organism can have many 2ND order selves because they depend upon who or what the organism is interacting with, just as the availability of information depends on context.
-Intuitively, where a 1ST order self might allow one to observe a cat and form plans regarding causal interactions with the cat, a 2ND order self would allow one to be consciously aware of the cat for the purpose of reasoning and report.
-One can know of the cat, and that another organism knows of the cat, but a 2ND order self is insufficient to be aware that one is aware of the cat [^16] This is ‘meta-self-reflexive consciousness’ as some have described it [39]..
-Figure 2: Visual intuition for a 2ND order self.
-The organism constructs a causal identity for itself, and for another object, and the causal identity for the other object includes a prediction of the organism itself from that object’s perspective.
-This would occur, for example, if one was a predator trying to predict the movements of prey in response to one’s own actions.
-It is an extension of the 1ST order self.
-More formally using the notation given in definition 15, assume $\mathfrak{a}$ and $\mathfrak{b}$ are organisms that evolved to accurately predict one another’s behaviour.
-Assume $\mathfrak{a}$ constructs a causal identity $c^{\mathfrak{b}}_{\mathfrak{a}}$ to predict $\mathfrak{b}$ given input $i_{\mathfrak{a}}\in I_{\mu_{\mathfrak{a}}}$, of which a second order self $c^{\mathfrak{ba}}_{\mathfrak{a}}$ is part.
-Likewise, $\mathfrak{b}$ constructs $c^{\mathfrak{a}}_{\mathfrak{b}}$ to predict $\mathfrak{a}$ given input $i_{\mathfrak{b}}\in I_{\mu_{\mathfrak{b}}}$, of which $c^{\mathfrak{ab}}_{\mathfrak{b}}$ is part.
-What is important here is that each organism’s intent is to some extent inferred by the other, and that fact inevitably changes the sorts of policies that will be “fit”.
-For example, second order self means each knows the other can anticipate manipulation, which means the optimal policy will often be to have rather than feign intent that aligns to some extent with the other party’s desires, to co-operate [64] [^17] Depending upon circumstances, for example organisms may co-operate in some circumstances but not others [42], and transient relations, information asymmetry and other factors can make deceit a more attractive option..
-Repeated interaction creates an iterated prisoner’s dilemma, incentivising co-operation and signals that both parties interpret similarly (the beginnings of language) [42].
-To communicate in Gricean terms, $\mathfrak{a}$ must intend to convey meaning $m_{\mathfrak{a}}$, and $\mathfrak{b}$ must recognise this intent.
-The incentive precondition explains why $\mathfrak{a}$ would form such intent (co-operation is often advantageous), while how may be understood as follows:
-  • $c^{\mathfrak{ba}}_{\mathfrak{a}}$ lets $\mathfrak{a}$ predict what $\mathfrak{b}$ will come to believe when it observes $\mathfrak{a}$’s behaviour.
-  • $c^{\mathfrak{ab}}_{\mathfrak{b}}$ then lets $\mathfrak{b}$ predict what $\mathfrak{a}$ intends that $\mathfrak{b}$ believe.
-$\mathfrak{a}$ can use $c^{\mathfrak{ba}}_{\mathfrak{a}}$ to infer behaviour to which $\mathfrak{b}$ will ascribe the intent to communicate $m_{\mathfrak{a}}$, and $c^{\mathfrak{ab}}_{\mathfrak{b}}$ lets $\mathfrak{b}$ infer that this is what $\mathfrak{a}$ intends.
-The “utterance” Grice refers to is how $\mathfrak{a}$ affects $\mathfrak{b}$ in accord with earlier definitions.
-Put another way, $\mathfrak{a}$ encodes $m_{\mathfrak{a}}$ into its behaviour in a manner that $\mathfrak{b}$ can decode (their respective second order selves act as encoders and decoders).
-By encode and decode, we mean a loose approximation of $m_{\mathfrak{a}}$ is communicated.
-There are of course shortcuts, for example of $\mathfrak{a}$ and $\mathfrak{b}$ are of the same species then they likely have similar motives and experiences, and so the efficient thing for each to do would be to use its own intent as an approximation of what the other might think.
-However, that does not obviate the need for second order selves, it just makes such things easier to realise.
-6.2.1 The Third Order Selves
-We can continue scaling WPO indefinitely.
-3RD and higher order selves can explain function.
-For example, meta self-awareness [39] is the awareness that one is self-aware.
-If self-awareness stems from 2ND order selves, then it follows that meta self-awareness requires 3RD order selves.
-Formally a third order self for $\mathfrak{a}$ reflecting off $c^{\mathfrak{ba}}_{\mathfrak{a}}$ lets $\mathfrak{b}$ is $c^{\mathfrak{baba}}_{\mathfrak{a}}$.
-It is $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$.
-Figure 3: Visual intuition for a 3RD order self. The organism constructs a 2ND order self for its 2ND order self, and so becomes aware that it is aware.
-7 The What and Why of Consciousness
-Up to now we have developed the conceptual toolbox that one can use to dissolve the hard problem.
-We started with the basic observation that self-organizing systems such as the human bodies constantly process information to maintain oneself in the face of constant change both inside and outside one’s body.
-An organism does not spring into existence understanding number systems or objects, but constructs them through constant interactions with the environment.
-One can regard them as policies according to which the environment is interpreted.
-To learn how to interpret the environment an organism must differentiate between states, must react to change, and learn policies according to the valence associated with that change.
-As we saw earlier, learned policy has valence, and it is a classifier of inputs.
-Those inputs are “information” in the mechanistic sense of the environment being in one state and not another, in the sense of axioms 1 and 2 (see section 3).
-Hence, a policy is a classifier of information, but that information is not in a language, and it is not yet something labelled or quantified (until there is a policy that labels or quantifies).
-Clearly, prelinguistic self-organising system must classify and attach value and disvalue to states and anticipated states to prioritise and make decisions.
-To learn to label and classify information an organism must sense that “something” has changed, and motivated by valence construct a policy classifying that “something”.
-In the absence of language, the difference between two states can only be qualitative.
-The bold claim here is that information processing at 1ST level biological self-organising systems such as the human body is necessarily qualitative by the very virtue of existing and experiencing (i.e.
-exploring one’s body and environment).
-This fundamental, basic way of engaging with the environment is what we call experience (see Ciaunica in prep.).
-Note that this is different from experience in the sense of experiential (i.e.
-phenomenal) as usually defined in the literature, which we take to be the second-order self-perception.
-Importantly, experience in the sense of experiential content cannot exist without experience in the sense of embodied exploration of the body and world as we defined here.
-To put it provocatively quality precedes quantity, and quantity is nothing more than the interpretation of quality.
-Quality comes first and experiences should be regarded on a continuum rather than a switch on/ switch off phenomena.
-All living systems experience the world through their bodies and as such, there is something what it is like to experience the world in that basic way (even when one is asleep or a baby).
-One can access those phenomenal, experiential aspects at a higher level, true, but by accessing them, it doesn’t mean that one ‘constructs’ consciousness or one becomes a conscious being.
-One is already consciously experiencing the world before one can explicitly access one’s own experiences as a 2ND order selves.
-It follows that every policy learned in this way must classify a quality.
-Hence every such policy is a local state.
-A causal identity imbued with meaning by valence.
-There would be a policy for one’s act of smelling coffee.
-For perceiving one’s friend.
-There would be something different it is like to interact with a hostile version of that very same friend.
-The 1ST order self-accompanies everything an organism does.
-It has a quality, so the 1ST order self is “what it is like” to be that organism.
-Put another way, Nagel’s question of “what it is like to be” a particular organism could be answered if one could somehow have that organism’s first order self [1].
-One’s 2ND order selves would also have a certain qualitative character, and one’s 3RD order too.
-There would, however, be a very clear delineation between conscious and not.
-The absence of a 1ST order self would mean there would be no policy linking all other policies.
-There could be no “self” to experience all of this.
-Hence a 1ST order self must precede all others.
-Furthermore, there could be no 2ND order self without a 1ST.
-Our framework thus makes a zombie impossible (there can be no perfect unconscious replica of a conscious being).
-The hard problem has it backwards.
-The question is not why qualia exist, but why anyone thinks representational contents [^18] Policies for interpreting the environment.
-can exist without first being learned through qualitative experience and discrimination.
-This ‘dissolves’ the hard problem by going a level down to the fundamental drive: stay alive!
-The imperative ‘stay alive’ involves sensing, classifying, evaluating, prioritising and acting.
-Our arguments align with those of Merker, Barron and Klein, who have linked subjective experience to reafference [36, 37, 38].
-We agree that reafference is key, but provide a very different explanation of why and how.
-Their work presents biological evidence for subjective experience in organisms with reafference.
-In contrast, we derive the 1ST order self from first principals and explain why and how it is a classifies “what it is like” to be a particular organism.
-The 1ST order self also happens to be equivalent to reafference, so we arrive at the same conclusion as Merker from very different, mathematical premises.
-Hence these are very complementary positions.
-We speculate an organism approaches the full richness of human subjective experience when different orders of self-interact [^19] As argued by Boltuc the full richness of human consciousness, or “hard” consciousness as he calls it, might be distinguished from phenomenal and access consciousness.
-It is this to which we refer now..
-Different orders of learned self might interact through hierarchical planning in a distributed system, perhaps something like a connective core [65].
-There is no possible way phenomenal consciousness can exist as we experience it without these different orders of self.
-There is also no way the adaptability we see in living organisms can be possible without the construction of such selves.
-We turn to this discuss now.
-8 From Rocks to Einstein: The Hierarchy of Being
-To illustrate how our argument applies in the real world we describe stages of conscious organism.
-Each stage follows from scaling up supply and demand for WPO, through natural selection (see figure 4):
-$0:$ Unconscious (e.g. a rock)
-$1:$ Hard Coded (e.g. protozoan)
-$2:$ Learning (e.g. nematode)
-$3:$ 1ST Order Self (e.g. housefly)
-$4:$ 2ND Order Selves (e.g. raven)
-$5:$ 3RD Order Selves (e.g. human)
-Stage 0: Unconsciousness
-Stage zero may be understood as the consciousness we are willing to assign to a rock: more exactly, the lack of consciousness so assigned.
-While some assert a thesis of panpsychism and claim everything, even a solitary hydrogen atom is conscious, we take rocks as expressing a baseline example of things in the universe that are not conscious at all.
-They do not sense.
-They do not think.
-They do not act.
-They have no 1ST order self.
-  • Example: A rock.
-Stage 1: Hard Coded
-Stage one refers to adaptations “hard-coded” or hard-wired by natural selection. This is the sort of preset behaviour that allows complexity to persist [66] in a reasonably consistent environment.
-  • What: Hard-coded adaptations. Habituation and sensitization.
-  • How: The extension of fit behaviour is learned by natural selection and hard-coded into the organism as a policy (in DNA, form, the local environment etc).
-  • Why: If the environment never changes, it makes more sense to just hard-code survivable behaviour.
-  • Example: Single-celled protozoan.
-Stage 2: Learning
-Stage two introduces learning.
-To learn an organism must store, classify and order historical examples by valence.
-There is not something it is like to be stage two, because there is no locus of “self”.
-A biological example of such a decentralised nervous system is the cubozoan box jellyfish Tripedalia cystophora.
-Even Tripedalia cystophora was recently shown to be capable of associative learning [40].
-An entirely distributed control system can “learn”.
-Likewise, stage two is exemplified by the nematode C.
-elegans [67, 68], which has a centralised nervous system and exhibits some ability to adapt with experience.
-However, the absence of a “self” prohibits cause and effect reasoning, which as others have already pointed out must limit spatial, navigational abilities [38].
-When starved C.
-elegans exhibit “increased locomotion and dispersal in a random, rather than directed, search” [38, 69, 70], whereas something like a bee or an ant can recall and navigate to previously discovered food [71, 72, 73].
-A nematode can learn, but the absence of a self for the purpose of causal reasoning limits adaptability.
-  • What: Mindless learning, with no centralised intent or locus “self”, to feel.
-  • How: Any system which learns will do. Search, approximation or biology. Learning is impossible without affect, reward or some other notion of value.
-  • Why: An organism that can learn can survive more than one that cannot.
-  • Examples: Jellyfish, nematode.
-Stage 3: 1ST Order Self
-This is where phenomenal (i.e.
-experiential) consciousness begins, with a 1ST order self.
-In biological terms this also implies reafference, which others have argued is the key to subjective experience, albeit for different reasons than what we have argued here [36, 37, 38].
-They identified a housefly as a good example of where subjective experience may begin, and we concur.
-We also hold this is where an organism might be said to have intent.
-Intuitively, the policy that motivated behaviour is the intent of that behaviour [^20] In the same way declarative and imperative programs are equivalent [74]..
-The “weaker” a policy is, the more behaviours it motivates.
-Conversely, if one’s actions share anything in common it is the intent that motivated them.
-The more diverse the actions, the more “general” or “high level” the intent they share.
-For example, the action of eating tends to involve the intent of satisfying hunger.
-If a policy is implied by all of an organism’s behaviour, then it might have motivated that behaviour.
-While this is an unusual form of words, we contend that an organism is such a policy.
-A stage three there is a self to feel simple hunger and pain.
-However, communication as described by Grice would be impossible [63], because a stage three organism has no self-awareness, to represent another’s perception of their intent [42].
-Nor would a stage three organism be able to conceive of its own death, or shame, because it cannot conceive of itself.
-  • What: A 1ST order self. Reafference. Phenomenal consciousness.
-  • How: Embodiment in which intervention is not identical to observation.
-  • Why: Accurate prediction of consequences of interventions. For example, a fly must distinguish between having moved, and the environment having moved, to navigate.
-  • Example: Housefly.
-Stage 4: 2ND Order Selves
-Stage four is the 2ND order self, and importantly this is where we hold access consciousness begins because it is where information is available for report in the Gricean sense.
-In other words, access consciousness follows phenomenal consciousness.
-The ability of ravens to intentionally deceive [75] suggests they are at least stage four.
-Raven $\mathfrak{a}$, aware that it is being observed by raven $\mathfrak{b}$, will act as if it is hiding food in one location to mislead $\mathfrak{b}$, but will then move the food in another location unobserved by $\mathfrak{b}$.
-$\mathfrak{a}$ seems to predict not just the intent of $\mathfrak{b}$ (to steal the food), but $\mathfrak{b}$’s perception of $\mathfrak{a}$.
-It seems intuitively likely that dogs and cats have second order selves, as they must hunt reasonably intelligent animals and must anticipate how their actions are perceived.
-For example, a cat anticipates its prey will flee when it is observed, and hides.
-  • What: Access consciousness. Theory of mind. Self-awareness. Inner narrative.
-  • How: Selection pressures that demand theory of mind.
-  • Why: A 2ND order self is necessary to anticipate, manipulate and communicate intent.
-It allows one to anticipate how others will perceive one’s behaviour, which may be necessary to survive in a social hierarchy.
-  • Example: Raven.
-Stage 5: 3RD Order Selves
-A 2ND order self for one’s 2ND order self. Humans appear to possess this, because we are aware that we are aware.
-  • What: Meta self-awareness. Inner narrative in which actors have inner narratives.
-  • How: More accurate prediction and planning.
-  • Why: Because a social organism must predict complex social dynamics.
-  • Example: Human.
-Figure 4: Overview of stages and orders of self.
-Here we suggested that one single formalism can explain so many different theories of consciousness.
-This unifies lower and higher order theories of consciousness, by scaling simple axioms from first principles.
-Importantly, this formalism is compatible both with the idea the environment is software running on a single Turing machine (a strong interpretation of pancomputationalism), and the idea of an environment as co-created by the interaction of independent entities (enactivism or a distributed computing system).
-This ensures there is simultaneous compatibility with enactivism and pancomputationalism.
-9 Unifying Lower and Higher Order Theories of Consciousness
-Our earlier research argues that assuming lower order states is like assuming an abstraction layer in a computer [35, 44].
-The behaviour of software is determined by the abstraction layer that “interprets” it.
-At the most foundational level software is machine code [^21] A “word” of machine code triggers a mechanistic process hardwired in the CPU by the human design.
-For example a word may copy the 32 bit value stored at memory address $X$ into the 32 bit register $Y$, which the next line of machine code “adds” to register $Z$ by looping over each bit in an “adder” circuit..
-Machine code is interpreted by hardware, and that hardware determines every aspect of what that machine code does.
-A word of machine code is a mechanical trigger that only “means” whatever we have designed the hardware to do when we input that word.
-That hardware is an abstraction layer in which the software exists (including “higher level” abstraction layers like the Python interpreter).
-In other words, what we call software is nothing more than the state of hardware [32].
-This has undermined all but the most subjective of claims regarding the behaviour of theorised, software superintelligence [76, 46, 77].
-It is a flaw in the very idea of intelligent software.
-The distinction between software mind and hardware embodiment is subsequently called computational dualism [32], because it is reminiscent of how Cartesian dualism conceives of a mental substance distinct from physical substance.
-If lower order mental states function as an abstraction layer, then HOT is a form of computational dualism.
-If we are to avoid computational dualism, then we should formalise mental activity as embedded, extending into the environment and enacted by interactions between bodies [29].
-From the enactive perspective we cannot assume lower orders as pregiven and fixed.
-Instead, contents are co-created through interaction between an organism and its environment [22, 78].
-The organism learns a world, or an “abstraction layer” in computing terms, that is relevant to its motivations.
-This process is sometimes called “relevance realisation” [79, 80, 81, 82].
-Some have argued enactivism is incompatible with computationalism because the set of possible interpretations that must be searched for relevance realisation is intractable [82].
-A computationalist who thinks of the mind as software must assume there is a hardware abstraction layer, and that abstraction layer hinders relevance realisation because it constrains the set of concepts that can be entertained [35].
-This objection to cognitive computationalism [47] may be understood as an alternative framing of the objection to computational dualism raised the context of artificial intelligence [32].
-However, just because relevance realisation is at odds with that sort of computational cognitivism does not mean enactivism is incompatible with the idea of computation [^22] There have also been efforts to combine enactivism with the Free Energy Principle (FEP) [83, 84].
-While our explicandum and formalization are substantially different the overall approach, explaining behaviour in terms of optimisation and embodiment, is similar.
-Hence, our approach and the FEP should be understood as complementary.
-in general [85].
-Importantly, our claim is not that everything is computation understood as minimal information processing (Williams & Ciaunica in prep.).
-Rather our claim is weaker, namely that one can provide one single unifying formalism in terms of computations linking lower to higher level information processing.
-We develop a formalisation [32] of representationless pancomputationalism.
-Computation as mechanics.
-Not representational, but physical in the sense described by Piccinini [30].
-Intuitively, if this is information processing, then it is in the sense that an apple falling to the ground is information interpreted by physics.
-Whatever causes the environment to act as it does might then be thought of as an interpreter in this mechanistic sense.
-Our radical and provocative claim is that phenomenal consciousness without access consciousness is likely very common, but the reverse is implausible.
-A zombie impossible.
-A data warehouse might have “access” to information, but that is not the same thing as access consciousness.
-We have classified this sort of mindless “access” to information as functional but not access consciousness.
-For example, Block holds that phenomenally conscious content is phenomenal, whereas access conscious content is representational.
-In pancomputational enactivism there are only states and the programs they form.
-This suggests the abstract “representations” we construct are just organised phenomenal content, clustered according to what causes valence.
-We don’t discard these two sorts of consciousness, but unify them by showing how phenomenal consciousness gives rise to access.
-Far from suggesting there is no such thing as qualia, this suggests instead that there is no such thing as purely representational content in anything but the tools we construct and our interpretations of their behaviour.
-Cells are a material with agency [86].
-A human is a “multiscale competency architecture” [87, 88, 24].
-Our intelligence is a swarm intelligence; the high level goal directed behaviour of a swarm of cells.
-When we embody human abstractions (e.g.
-arithmetic) in silico (e.g.
-the x86 instruction set), we disconnect the high level goal directed behaviour from the low level behaviour (and the motivating affect) that gave rise to it [44].
-We call the information embodied in the computer “representational” because it means something to us.
-However, if we view the computer as an organism, then it amounts a set of stage one adaptations none of which are the sort of general purpose solution we observe in the “agentic material” we call cells.
-When we embody our abstractions in silico we disconnect them from the affect that motivated their construction [43].
-This suggests there is no such thing as representational contents.
-They are a fiction we have invented because we struggle to reduce our own abstractions to their basic nature; the causes of valence.
-10 Conclusions and Outlook: Why Nature Does Not Like Zombies
-In this paper we provided a mathematical formalism uniting lower order states and higher order meta-representations, dissolving the hard problem of consciousness.
-Specifically, we described a multilayered formalism illustrating how biological self-organising systems become phenomenally conscious when they construct a 1ST order self.
-A human lacking a 1ST order self could not perform causal reasoning needed to adapt as humans evidently can [45].
-As previous research pointed out [32], the computer metaphor with the seminal distinction between hardware and software is a simplification.
-Rather, software is a state of hardware.
-Nature privileges efficiency over abstract simplification.
-Compare the vast quantities of both training data and energy required by a language model, to the small quantities humans need to solve a problem.
-Biological systems are more efficient, because they are adaptive at a every level [24, 44].
-Hence, rather than trying to explain the mind in the abstract like software, we started at the level of the embodied organism.
-One direct consequence of our approach is that it places the phenomenal quality of conscious experience before access consciousness.
-We have shown a consistent definition of access consciousness implies 2ND order selves, which imply a 1ST order self.
-Phenomenal consciousness arises first, access comes later.
-Unlike panpsychism, we don’t believe rocks are conscious but solely those self-organising systems that need to adapt, motivated by valence, while keeping track of the self.
-Consciousness is a necessary adaptation, and a zombie is impossible because some behaviour cannot be achieved without consciousness.
-There remain unanswered questions regarding whether the mere presence of 1ST, 2ND and 3RD order selves is sufficient for consciousness, but we hold they are at least necessary.
-Their absence should guarantee a system is not conscious.
-This serves to resolve questions about the consciousness of artificially intelligent systems, such as large language models.
-Such models are neither optimised to construct weak representations, nor have any incentive to construct selves, being passive mimics [42] of human behaviour.
-Future research should attempt to identify 2ND and 3RD order selves in biological systems.
-This will help establish levels of consciousness in different organisms.
-Merker, Barron and Klein have already nicely demonstrated the existence of a 1ST order self in humans and insects.
-Future research may also consider training organoids and artificial systems in a manner that should cause them to construct 1ST, 2ND and 3RD order selves.
-This may help whether what we have described is sufficient to achieve the outward behaviour of consciousness and what else, if anything, may be required to engineer it.
-Any difference between equivalent synthetic biological organoid and non-biological artificial intelligence will help resolve questions [44] about the difference between biological and non-biological substrates.
-Another important future avenue for research in this area revolves around synchronicity and centralisation.
-For example, our formalism suggests a moment in which an output is chosen.
-This means that there exists a state of the environment in which all the declarative programs that make up a policy are facts.
-These declarative programs are physically manifest in the part of the environment at that moment, and a “decision” made by an agent is either determined by or determines the state at that moment.
-This centralises control, in that parts synchronise to affect a decision and a coherent whole (relative to a state).
-Importantly, this approach is analogous to how cells form an organ in a multiscale competency architecture [88].
-When a cell is isolated from the informational structure of which it is part, it becomes cancer.
-Now, one important question for future research is what happens if some part is isolated from the informational structure of the whole in the case of consciousness as we have portrayed it?
-Is this informational structure confined to a point in time?
-For example, can a policy be spread over time as a sequence much like how a CPU in a computer might process information over time, as sequence?
-What exactly is a moment?
-It is arguable that everything is centralised, and nothing is, depending upon the perspective we adopt with respect to space and time.
-There does not seem to be anything like an “event horizon” for consciousness; no cut-off point where something cannot affect something else and so prevent the formation of selves.
-This raises interesting questions about consciousness at different scales of space-time.
-Given our formalism, it is conceivable that what we perceive as continuous time is in fact disjointed, and there are large gaps in our apparently continuous perception we are unable to perceive.
-Perhaps we are “prompted” to process information by some states of the environment and not others, because some states are indistinguishable from one another given the abstraction layer in which we exist.
-Future research might explore such questions.
-Looking further at the social aspect: if consciousness does not require some specific form of synchronicity, then what does that say about the action of populations?
-Where is the line between feeling and being?
-Feeling is being.
-But feeling comes first.
-We hold that valence informs learning at Stage Two.
-Being, at the human level, the Being of which the phenomenologists speak, does not emerge until Stage Three.
-Arguably, a group of humans is stage two, and yet an individual human is stage three.
-A group of humans cannot have a second order or first order self as we perceive it in time, but what of the effect of their actions in the environment across time?
-From the multiscale competency architecture perspective, the human mind is the mind of a collective of cells.
-It is at least conceivable, however implausible, that collectives of collectives of cells might be conscious, albeit at a very different timescale.
-Another area future research might explore is how subjective experience depends on the ability to simultaneously learn and decide (simultaneous “learning and inference”, or “induction and abduction”).
-If an organism ceases to learn, it has only the memory of affect.
-Information retrieval must trigger affect to some extent in order to retrieve the quality of a protosymbol (otherwise, the quality would not be remembered).
-In other words, the memory of a painful experience must invoke pain, or it is not really remembered.
-Hence, an organism that has learned must still have subjective experience even if all it can do is remember (the absence of any subjective experience would mean it could not remember subjective experience).
-Recall of valence still compels one to act, meaning one must be subject to it.
-A human does not operate in two distinct phases of learning and inference as a machine learning system does.
-We appear to learn and infer at all times.
-Were we to have separate phases we would be unable to remember anything that happens in the inference phase, or else we would be motivated by the new memories and so would still be learning.
-To put it provocatively, this suggests a “pure thinker” [89] cannot be conscious.
-Can we really consider the memory of consciousness to be the same thing as consciousness, if there is no awareness of the present which those memories are used for inference?
-If there are no new subjective experiences?
-Likewise, does “dropping Hume’s Guillotine” [44] to embody purely representative content in-silico eliminate the possibility of consciousness, by separating abstractions from the valence that motivated their construction?
-Our paper sparks perhaps more questions that it hasn’t answered.
-But our proposal lays the foundations of a formal science of consciousness, deeply connected with natural selection rather than abstract thinking, closer to human fact than zombie fiction.

### v3

@@
-hard problem of consciousness
+xxxxx, xxxxx, xxxx
 Abstract
-We tackle the hard problem of consciousness taking the naturally selected, self-organising, embodied organism as our starting point.
-We provide a mathematical formalism describing how biological systems self-organise to hierarchically interpret unlabelled sensory information according to valence and specific needs.
-Such interpretations imply behavioural policies which can only be differentiated from each other by the qualitative aspect of information processing.
-Selection pressures favour systems that can intervene in the world to achieve homeostatic and reproductive goals.
-Quality is a property arising in such systems to link cause to affect to motivate real world interventions.
-This produces a range of qualitative classifiers (interoceptive and exteroceptive) that motivate specific actions and determine priorities and preferences.
-Building upon the seminal distinction between access and phenomenal consciousness, our radical claim here is that phenomenal consciousness without access consciousness is likely very common, but the reverse is implausible.
+We tackle the hard problem of consciousness taking the naturally selected, embodied organism as our starting point.
+We provide a formalism describing how biological systems self-organise to hierarchically interpret unlabelled sensory information according to valence.
+Such interpretations imply behavioural policies which are differentiated from each other only by the qualitative aspect of information processing.
+Natural selection favours systems that intervene in the world to achieve homeostatic and reproductive goals.
+Quality is a property arising in such systems to link cause to affect to motivate interventions.
+This produces interoceptive and exteroceptive classifiers and determines priorities.
+In formalising the seminal distinction between access and phenomenal consciousness, we claim that access consciousness at the human level requires the ability to hierarchically model i) the self, ii) the world/others and iii) the self as modelled by others, and that this requires phenomenal consciousness.
+Phenomenal without access consciousness is likely common, but the reverse is implausible.
 To put it provocatively: death grounds meaning, and Nature does not like zombies.
-We formally describe the multilayered architecture of self-organisation from rocks to Einstein, illustrating how our argument applies in the real world.
-We claim that access consciousness at the human level is impossible without the ability to hierarchically model i) the self, ii) the world/others and iii) the self as modelled by others.
-Phenomenal consciousness is therefore required for human-level functionality.
-Our proposal lays the foundations of a formal science of consciousness, deeply connected with natural selection rather than abstract thinking, closer to human fact than zombie fiction.
+We then describe the multilayered architecture of self-organisation from rocks to Einstein, illustrating how our argument applies.
+Our proposal lays the foundation of a formal science of consciousness, closer to human fact than zombie fiction.
 keywords:
-hard problem of consciousness
+hard problem, consciousness
+journal: rsif
+{fmtext}
 1 Introduction
@@
 We are not suggesting biological systems are non-physical.
-systems process information, yet it seems that humans consciously experience in addition to merely processing information.
+systems process information, yet humans consciously experience as well as process information.
 Why?
 Living organisms are constantly processing self and world related information to survive in an ever-changing world.
-Human bodies share with all other physical systems the property of being instantiated in time and space, (e.g., our body occupies a given position and volume in space at a given time).
-Yet, unlike physical systems, biological, living systems are dissipative systems using energy to self-organise in the face of entropic decay and environmental perturbation [nagel1974, chalmers1995].
+Human bodies share with all other physical systems the property of being instantiated in time and space (e.g.
+occupy space).
+Yet, unlike physical systems, biological systems are dissipative systems using energy to self-organise in the face of entropic decay and environmental perturbation [nagel1974, chalmers1995].
 Initially formally introduced in the field of cybernetics [ashby1947, foerster1960] the notion of self-organisation has been applied to various disciplines including physics [haken1983], biology [camazine2003, bell2007] and neuroscience [kelso1997, Friston2010, tognoli2014].
-Self-organization is typically defined as the spontaneous emergence of spatiotemporal order or pattern-formation processes in physical and biological systems resulting from interactions of its components with the environment [camazine2001, seeley2002, rosas2018].
+Self-organisation is typically defined as the spontaneous emergence of spatiotemporal order or pattern-formation processes in physical and biological systems resulting from interactions of its components with the environment [camazine2001, seeley2002, rosas2018].
 Interestingly, much self and world related information processing goes on behind the scenes, or “in the dark” so to speak, that is, without being constantly present to our conscious minds.
 But why doesn’t all information processing go in the dark?
-This question is the subject of long-standing debates across disciplines [seth2022].
+This question is the subject of long-standing debate [seth2022].
 One highly influential view is that consciousness has two aspects [seth2022].
 The first is functional, by which we mean the ability to access and communicate information [block1995].
-How information processing is linked to consciousness is the “easy problem” of consciousness [chalmers1995].
+How function relates to consciousness is considered the “easy problem” of consciousness [chalmers1995].
 The second aspect is “what it is like” to consciously experience information processing, or phenomenal consciousness [chalmers1995, block1995, nagel1974, gallagher2021, fuchs2017].
-This doesn’t mean just global states like being awake, but more specific local states like smelling a cup of coffee.
+This doesn’t mean just global states like being awake, but more local states like smelling a cup of coffee.
 These local contents or “qualia” are characterised by what it is like to be in them [seth2022].
@@
 The hard problem has sparked a substantial body of work, detailed discussion of which [seth2022, northoff2014] lies beyond the scope of our paper.
-Rather, in what follows we build upon the useful distinction between lower and higher order theories of information processing in relation to conscious experiences.
-For example, higher order thought theory (HOT) [rosenthal2005, brown2019] holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states.
+In what follows we build upon the distinction between lower and higher order theories of information processing in relation to conscious experiences.
+Higher order thought theory (HOT) [rosenthal2005, brown2019] holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states.
 Lower order states may include emotions and perceptions, while higher order meta-representations reflect upon those.
@@
 Sense data is processed by the body resulting in lower order mental states, and then meta representations of those is where we might find more abstract conceptual or thought-like contents of consciousness.
-The distinction between lower and higher orders is a good starting point to understand why most biological information processing goes on “in the dark”.
-Yet, the interesting question in our view is why do these lower order states arise?
+HOTs are a good starting point to understand why most biological information processing goes on “in the dark”, but why do these lower order states arise?
 Can they occur in the absence of subjective, qualitative experience?
-It is widely agreed that higher order mental states such as desire to drink a coffee is accompanied by a qualitative aspect.
-But is there something it is like to “be in the dark”?
-That is: to process information at the lower bodily levels?
+Is there something it is like to “be in the dark” processing information at the lower bodily levels?
 In this paper we suggest a way to dissolve the “hard problem” of consciousness is to reverse the order and start with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states.
-If an explanation of how consciousness functions is to reveal why information processing sometimes has a qualitative aspect, then it must explain how we get local states [thompson2016], not begin by assuming local states.
-We need to go a level down and begin at the level of the embodied organism.
 Longtime considered a fringe approach, the embodied cognition paradigm [varela2016] has recently gained substantial influence in cognitive science and philosophy [ciaunica2023, levin2024, seth2018].
@@
 In other words, conscious experiences do not merely depend on bodily experiences as an external factor that can be replaced with a vat or an artificial system.
-One promising approach to this question is to try to rigorously define consciousness from first principles and show that some aspects of functional consciousness depend on phenomenal consciousness in a manner that makes zombies impossible, “dissolving” the hard problem by showing the phenomenal to be intrinsically functional [^2] Note that we are far from the first to claim that the phenomenal is functional.
+Our approach is to try to rigorously define consciousness from first principles and show that some aspects of functional consciousness depends on phenomenal consciousness in a manner that makes zombies impossible, “dissolving” the hard problem by showing the phenomenal to be intrinsically functional [^2] Note that we are far from the first to claim that the phenomenal is functional.
 For example, the proposed Conscious Turing Machine [blum2020] based on Global Workspace Theory [baars1997], and constructivist approaches to artificial intelligence [wang2020] take similar position.
 It is our explanation of how and why that is novel..
-Unifying them.
-We would need to establish axioms that hold in every conceivable environment and show that it is impossible to have the function of consciousness without the subjective experience of it.
-Such an answer must explain how consciousness functions, and why most information processing goes on “in the dark”.
-This paper lays the grounds of a computational model formalizing the link between lower and higher levels of information processing in relation to conscious experiences in a way that is compatible with the basic principles of embodied cognition and enactivism.
-Enactivism is roughly the view that information processing arises through a dynamic interaction between an acting organism and its environment [thompson2007].
-Now, the notion of computation is widely debated, and a detailed review of these discussions would lead us to a major digression [piccinini2015, piccinini2021].
-Here we define computation not in terms of abstract symbol shuffling and representation, but in concrete mechanistic terms where ‘information processing’ refers to the actual physical transition of a system from one state to another; a system such as a human body.
-Pancomputationalism is the idea that all dynamic systems are constituted by computation, which for our purposes is defined as above.
-This paper builds upon the formalism of pancomputational enactivism [bennett2024a].
-To answer “why is anything conscious” we proceed in two steps.
-First, we develop a formalism in which lower and higher order theories of consciousness, and phenomenal and access consciousness, are all derived from first principles.
-Each follows because of scaling natural selection pressures and the ability to adapt [^3] This is loosely inspired by a scale-based framing of machine learning [sutton2019] - we are applying Sutton’s “bitter lesson” to biology..
-Second, we provide an argument building on that to dissolve the hard problem of consciousness, answering why there must a qualitative aspect to information processing and in what circumstances.
-This paper is organised as follows.
-It should be noted first that the appendix at the end of this paper is a list of formal mathematical definitions, which the reader may refer to as needed.
-For readers without formal background, the appendix can be skipped.
-In Section 2 we extend separately published arguments, experiments and proof results [bennett2024a, bennett2023b, bennett2023c], developing a model of self-organising systems that holds across all conceivable environments.
+We establish axioms that hold in every conceivable environment and show that it is impossible to have the function of consciousness without the subjective experience of it, and why most information processing goes on “in the dark”.
+Enactivism is roughly the view that information processing arises through interaction between an organism and its environment [thompson2007], and it is considered by some to be incompatible with narrow definitions of computation.
+The notion of computation is widely debated, and a detailed review of these discussions would lead us to a major digression [piccinini2015, piccinini2021].
+Here we define computation not in terms of abstract symbol shuffling and representation, but in concrete mechanistic terms where ‘information processing’ refers to the actual physical transition of a system from one state to another, so it is compatible with enactivism.
+Pancomputationalism is the idea that all systems are computational.
+We build upon a formalism called pancomputational enactivism [bennett2024a].
+To answer “why is anything conscious”, we use this formalism to show how lower and higher order theories, phenomenal and access consciousness all follow from first principles, scaling natural selection pressures and the ability to adapt [^3] This is loosely inspired by a scale-based framing of machine learning [sutton2019] - we are applying Sutton’s “bitter lesson” to biology..
+First, in Section 2 we integrate and extend separately published works of narrower scope [bennett2024a, bennett2023b, bennett2023c] to justify our model of self-organising systems.
 In Sections 3 and 4 we explain how this model formalises relevance realisation and unifies lower and higher order theories.
 Rather than assuming abstract objects are primary and trying to learn causal relations between them, this assumes valence is primary and that abstract objects are constructed to classify cause and anticipate valence.
-This takes us from attraction and repulsion to contentless states, through lower order thoughts to higher order meta representations.
+This takes us from attraction and repulsion to physical states, through lower order thoughts to higher order meta representations.
 We call this the psychophysical principle of causality.
-In Section 5 we extend previous work [bennett2023c] on causal learning and the development of first (1ST), second (2ND) and third (3RD) order selves, and explain how they are constructed as a direct consequence of scaling the ability of a self-organising system to adapt with natural selection pressures.
+In Section 5 we extend previous work [bennett2023c] on causal learning and the development of first (1ST), second (2ND) and third (3RD) order selves, and explain how they are a consequence of scaling the ability of a self-organising system to adapt with natural selection pressures.
 In Sections 6 and 7 we explain how subjective experience requires a 1ST order self, and conversely why a 1ST order self implies there is something it is like to be an organism that has a 1ST order self.
 We then argue that access consciousness requires the ability to communicate meaning, which requires both 1ST and 2ND order selves, and show how a philosophical zombie is impossible.
-In Section 8 we describe the development of consciousness as we scale up the capacity to adapt with natural selection pressures.
-We discuss examples of each stage from nematodes to humans.
-Section 9 is concluding remarks.
+In Section 8 we describe the development of consciousness as we scale up the capacity to adapt with natural selection pressures, with examples from nematodes to humans.
 2 Back to Foundations
-Rather than presupposing local states or an abstraction layer, we start from first principles to formalise all conceivable environments [bennett2024a, bennett2024b], as in definition 1.
-This is based on two assumptions:
+To meet length restrictions, the full mathematical definitions and proofs have been placed in the appendices [bennett2024appendices].
+The core arguments should be understandable without the math, but for convenience here is a quick reference guide:
+  1. 1.
+    A set $\Phi$ is assumed, whose elements are called states.
+A declarative program is set $f$ of states, and $f$ is true relative to $\phi\in\Phi$ iff $\phi\in f$.
+$P$ is the set of all programs.
+Embodiment is formalised as an abstraction layer, for which we single out finite set $\mathfrak{v}\in 2^{\Phi}$ called a vocabulary.
+A choice of vocabulary $\mathfrak{v}$ implies a formal language $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap l\neq\emptyset\}$.
+  2. 2.
+    Members of $L_{\mathfrak{v}}$ are called statements, and a statement $l$ is considered true relative to a state $\phi$ iff $\phi\in\bigcup l$.
+A completion of a statement $l_{1}$ is any statement $l_{2}$ s.t.
+$l_{1}\subseteq l_{2}$.
+The extension of a statement $l$ is the set $E_{l}$ of all its completions, s.t.
+$x\in{L_{\mathfrak{v}}}$ is $E_{x}=\{y\in{L_{\mathfrak{v}}}:x\subseteq y\}$.
+The extension of a set $X\subseteq{L_{\mathfrak{v}}}$ of statements is $E_{X}=\bigcup\limits_{x\in X}E_{x}$.
+  3. 3.
+    A $\mathfrak{v}$-task $\alpha$ is a pair $\langle I_{\alpha},O_{\alpha}\rangle$ where $I_{\alpha}\subset L_{\mathfrak{v}}$ are the inputs, $E_{I_{\alpha}}$ is the outputs available given ${I_{\alpha}}$, and $O_{\alpha}\subset E_{I_{\alpha}}$ are the correct outputs.
+All the tasks an abstraction layer tasks form a hierarchy.
+If $\alpha\sqsubset\omega$ then $\omega$ is a parent of $\alpha$, meaning $I_{\alpha}\subset I_{\omega}$ and $O_{\alpha}\subseteq O_{\omega}$.
+A parent is higher level than its children.
+  4. 4.
+    A policy is a statement $\pi$ and a policy is correct for $\alpha$ iff $\pi\in{\Pi_{\alpha}}=\{\pi\in L_{\mathfrak{v}}:{E}_{I_{\alpha}}\cap E_{\pi}={O_{\alpha}}\}$.
+Given input $i\in I_{\alpha}$ and a policy $\pi$, to infer is to choose $o\in E_{i}$.
+A task is complete if $o\in O_{\alpha}$.
+Given a task definition $\alpha$ s.t.
+$\alpha\sqsubset\omega$, to learn is to choose a policy $\pi\in\Pi_{\alpha}$, and $\omega$ has been learned from $\alpha$ if one chooses $\pi\in\Pi_{\omega}$.
+  5. 5.
+    A causal identity for $\mathfrak{b}$ is a policy classifying the causal interventions undertaken by $\mathfrak{b}$.
+Subscript denotes the system that learned the causal identity, while superscript denotes the object it classifies.
+For example, $c^{\mathfrak{b}}_{\mathfrak{a}}$ denotes a causal identity learned by $\mathfrak{a}$ to classify $\mathfrak{b}$’s interventions, and $c^{\mathfrak{ba}}_{\mathfrak{a}}\subset c^{\mathfrak{b}}_{\mathfrak{a}}$ denotes $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$.
+We begin by formalising all conceivable environments [bennett2024a, bennett2024b] from first principles:
 Axiom 1: When there are things, we call these things the environment.
-Axiom 2: Where things differ, we have different states of the environment.
-Hence, the environment is a set $\Phi$ of contentless global states $\Phi$.
-We define a declarative “program” as a set of states and treat the states themselves as irreducible [^4] We do not concern ourselves with imperative programs, because there is an equivalence between declarative and imperative [howard1980]..
-The set $P$ contains all such programs.
-The powerset of $P$ is every aspect of the environment, because every aspect of any conceivable environment must be set of such programs, which is a subset of $P$.
-Programs return true or false, so a program is a “fact” if it is true, and the current state of the environment is the set of all “facts”.
-Truth is determined with respect to states, and states are assumed to be mutually exclusive.
-For example, if time is one way in which things differ, then there is one state at a time.
-It follows that every aspect of the environment is a set of programs, and those programs are facts relative to any physical state they contain.
-An aspect of the environment is realized by a state if the programs it contains are facts.
-When we say a state is contentless, we mean we assume nothing about what states contain or are.
-Their only significance is in how they relate to one another, and each such relation is a program.
-This is intentionally like to post-structuralism [eacho2020, hayles2005], in which the only meaning words have is given via their relations to each other.
-This is why previous work claims it applies to all conceivable environments [bennett2024a].
-Universality Claim: Axioms 1 and 2 hold for every conceivable environment.
-Intuitively, one might understand the set of all programs (the powerset of all states) as the set of all relations between these otherwise meaningless states.
+Axiom 2: Where things change or differ, we have different states of the environment.
+Hence, the environment is a set $\Phi$ of mutually exclusive states.
+We call a set of states a declarative “program” [^4] We do not concern ourselves with imperative programs, because there is an equivalence between declarative and imperative [howard1980]..
+A program is “true” relative to the states it contains.
+For example, if time is one dimension along which things change, then there is one state at a time.
+A “fact” is a program which contains the state at the present time.
+It follows that every aspect of the environment is a set of programs.
+An aspect of the environment is realised by (exists in) a state if the programs it contains are facts.
+Previous work argues this formalises every conceivable environment [bennett2024b].
+To give some intuition as to why, note that we assume nothing of what states might be or contain.
+We assume no differences within a state, only between states, and those differences are the programs of which aspects are formed.
+After all, one can only interact with aspects of one’s environment.
+If one were to try and pinpoint what an aspect was made of, the answer would be deferred to other aspects.
+Any description that sought to fully encapsulate the semantics, truth or unmediated pure experience of a state would always be delayed or incomplete.
+This does not render such attempts vacuous, but speaks to their inherent contingency and conditionality, which is what this formalism attempts to capture.
 2.1 Natural Selection and Embodiment
-We then assume a process of natural selection. This produces embodied organisms.
-A set of facts can represent anything [^5] Some may object, pointing out that this ignores composition.
-However, the application of a function is a fact regardless of whether it takes another function as input., hence the body of an organism must be a set of programs.
-We call this set a vocabulary, and it is formalised in definition 2 (see Appendix).
-The vocabulary can be seen as hardware and software operating together.
-Piccinini [piccinini2015, piccinini2021] identifies two sorts of computation: abstract and concrete.
+We assume a naturally selected, embodied organism.
+Every aspect of the environment is a set of facts [^5] Some may object, pointing out that this ignores composition.
+However, the application of a function is a fact regardless of whether it takes another function as input., hence the body of an organism that persists across multiple states must be a set of programs.
+We call this set a vocabulary, and it is formalised in definition 2 (see appendix [bennett2024appendices]).
+Piccinini [piccinini2015, piccinini2021] divides computation into abstract and concrete sorts.
 Abstract computations mean or represent whatever we say they do.
-In contrast concrete computations are physical.
-They are the action of physical systems over time.
-The programs we talk about here are concrete, physical computations in that they are the relations between physical states.
-It follows that the computations a body performs are also concrete, and so no two organisms have the same vocabulary, because that would mean they are the same body.
+In contrast concrete computations are the actions of physical systems.
+The declarative programs we describe are concrete.
+Each captures a single point of difference between physical states.
+As such no two organisms have the same vocabulary because that would mean they are the same body.
 This might seem inconvenient, as intuitively we might prefer to think about bodies as an abstract ‘type’.
 However, doing so would undermine any claim we might make, as earlier results in artificial intelligence (AI) show [bennett2023c, bennett2024c].
-The meaning of abstractions depends upon interpretation, much like the behavior of software is determined by the abstraction layer that “interprets” it.
+The meaning of abstractions depends upon interpretation, much like the behaviour of software is determined by the abstraction layer that “interprets” it.
 At its least abstract software is machine code [^6] A “word” of machine code triggers a mechanistic process hardwired in the CPU by the human design.
 For example a word may copy the 32 bit value stored at memory address $X$ into the 32 bit register $Y$, which the next line of machine code “adds” to register $Z$ by looping over each bit in an “adder” circuit..
-Machine code is interpreted by hardware, and that hardware determines every aspect of what that machine code does.
+Machine code is interpreted by hardware which determines every aspect of what that machine code does.
 A word of machine code is a mechanical trigger that only “means” whatever we have designed the hardware to do when we input that word.
@@
 In other words, what we call software is nothing more than the state of hardware [bennett2024a].
-This has undermined all but the most subjective of claims regarding the behavior of theorised, software superintelligence [hutter2010, leike2015, leike2018].
+This has undermined all but the most subjective of claims regarding the behaviour of theorised, software superintelligence [hutter2010, leike2015, leike2018].
 It is a flaw in the very idea of intelligent software.
 The distinction between software mind and hardware embodiment is subsequently called computational dualism [bennett2024a], because it is reminiscent of how traditional Cartesianism conceived of a mental substance distinct from physical.
-Here we suggest that must avoid computational dualism.
 If we are to dissolve the hard problem by starting with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states, then we cannot presuppose organisms use particular abstractions.
@@
 Axiom 3: All aspects of the environment are spatially extended [bennett2024b].
-This means that a body can occupy only a finite number of states, which is formalised in definition 2.
+This means that a body is realised by only a finite number of states (see def.
+2).
 A body is a vocabulary $\mathfrak{v}\subset P$ with finitely many elements.
 $\mathfrak{v}$ implies a formal language $L_{\mathfrak{v}}$ of interaction between body and environment.
-A statement in this formal language is an aspect of the environment.
+A statement in this language is an aspect of the environment.
 A set of “programs” from our vocabulary.
@@
 2.2 Self-Organising Systems as Self and World Constraints
-Both snowflakes and human bodies are self-organising systems, yet we regard only the latter as conscious.
-What separates the two?
-To tackle this question, we need to introduce the notion of self and world constraints.
-Given that we have an embodied formal language, we can talk about computation with inputs and outputs by just treating everything as embodied statements embedded and enacted within the environment.
-Given an input $i$, the set of all possible outputs is the extension $E_{i}$ of that input.
-This is because if $i$ is realised by the environment, then the environment is constrained to only those states that realise $i$, which constrains what other statements can be realised.
-We can use this fact to talk about “policies” as embodied [^7] As a state of hardware rather than software, which is how the term policy is normally used in reinforcement learning.
-constraints on behaviour.
-A policy is a statement whose extension constrains outputs, like a causal intermediary in functionalism [putnam1967].
+Both snowflakes and human bodies self-organise, yet we regard only the latter as conscious. What separates the two? To tackle this question we need self and world constraints.
+Using our embodied formal language $L_{\mathfrak{v}}$, we can talk about computation with inputs and outputs by treating everything as embodied statements embedded and enacted within the environment.
+Given an input $i\in L_{\mathfrak{v}}$, the set of all possible outputs is the extension $E_{i}$ of that input.
+This is because if $i$ is realised by the environment, then the environment is constrained to those states that realise $i$, which constrains what other statements are realised.
 If behaviour is “motivated”, then statements have valence determined by natural selection and only a subset of the statements a body could make are “fit”.
-As a result, an organism will self-organise to express some statements, but not others.
-Though our environmental states are contentless and their only significance is in how they relate to one another, natural selection must make them attractive or repulsive to a self-organising system.
+Hence an organism will self-organise to express some statements, but not others.
+Though environmental states are contentless and their only significance is in how they relate to one another, natural selection must make them attractive or repulsive to a self-organising system.
 This is formalised by the $\mathfrak{v}$-task as in definition 3777A $\mathfrak{v}$-task is called because a $\mathfrak{v}$-task because it exists in the context of a vocabulary $\mathfrak{v}$..
@@
 The extension $E_{I_{\mu}}$ of $I_{\mu}$ would be every output that it is possible to choose given the inputs $I_{\mu}$, but only a subset of those $O_{\mu}\subset E_{I_{\mu}}$ are correct outputs.
-In other words, this defines a self-organising behaviour as a constraint on outputs, given inputs [^9] Conversely, a system which considered all outputs to be correct would not be self-organising..
 2.3 Inference
@@
 We call a constraining statement a policy.
-A policy constrains outputs given inputs.
-A correct policy is one that constrains outputs to only correct outputs.
+A policy constrains outputs given inputs, and a correct policy is one that constrains outputs to only correct outputs.
 For the sake of intuition, think of “correct” as “fit” according to natural selection [^8] This need not always be the case..
-This is formally expressed in definition 4.
+This is formalised in definition 4.
+2.4 Learning
+If an organism remains alive, then its history is an ostensive definition of “fit” self-organising behaviour.
+That history is expressed as a $\mathfrak{v}$-task $\mathfrak{h}$, where each input and output in $\mathfrak{h}$ is an interaction between organism and environment.
+If the set of all fit behaviour is a task $\mu$, then the organism remains fit then $\mathfrak{h}\sqsubset\mu$.
 “I survive therefore my model is viable.” - Mark Solms [solms2021]
-If an organism remains alive, then its history is an ostensive definition of “fit” self-organising behaviour.
-That history of interactions is just a set of statements in the organism’s embodied language, which can be expressed as a $\mathfrak{v}$-task $\mathfrak{h}$.
-Each input and output in $\mathfrak{h}$ is behaviour the organism has exhibited and everything that was involved in that behaviour (the interaction between organism and environment).
-The set of all fit behaviour would then be another task $\mu$, and if the organism remains fit then $\mathfrak{h}\sqsubset\mu$ [^11] One task is a child of the other, meaning the inputs and outputs are subsets of the parents’ inputs and outputs..
-2.4 Learning
-Learning can be understood as the process of approximating the program which caused data.
-Constructing a “world” model.
-Here, we just need to explain how an organism gets $\mu$ from $\mathfrak{h}$.
-We don’t need a world model, but a way to constrain outputs to just those that are fit.
-Like everything else, that constraint is once again a statement in the formal language $L_{\mathfrak{v}}$.
+$\mathfrak{h}$ implies a set $\Pi_{\mathfrak{h}}$ of policies, and not all of those policies will constrain outputs the same way given new inputs.
+Some policies will “generalise” to imply fit behaviour in unfamiliar circumstances, meaning correct outputs given a new set of inputs.
+If those policies imply fit behaviour given new inputs, then the organism will remain fit.
+An organism learns fit behaviour by learning $\pi\in\Pi_{\mathfrak{h}}$ such that $\pi\in\Pi_{\mu}$.
 “The best model of the world is the world itself” - Rodney Brooks [dreyfus2007]
-Those examples in $\mathfrak{h}$ could have been generated by any one of a set of policies $\Pi_{\mathfrak{h}}$.
-These are not representations of possible causal intermediaries in the machine functionalist sense, but just parts of the outputs from which the outputs could be derived given the inputs.
-Every $\mathfrak{v}$-task (and thus every self-organising system) implies such a set of policies, and not all of those policies will constrain outputs the same way given new inputs.
-Some policies will “generalise” to imply fit behaviour in unfamiliar circumstances, meaning correct outputs given a new set of inputs (a parent task of the organism’s history $\mathfrak{h}$).
-If those policies imply fit behaviour given new inputs, then the organism will remain fit.
-If an organism learns, then it is adapting to embody a policy that it hopes will generalise to fit behaviour $\mu$.
-Earlier work [bennett2023b] showed that the optimal strategy to learn and adapt as fast as possible [^12] Meaning to converge on a fit policy given the smallest history possible, to realise fit behaviour.
+Earlier work [bennett2023b] showed that the optimal strategy to adapt as fast as possible [^9] Meaning to converge on a fit policy given the smallest history possible, to realise fit behaviour.
 is to prefer “weaker” policies, meaning those with larger extensions.
 The policy which generated outputs $O$ given inputs $I$ is most efficiently identified by constructing a policy $\pi$ such that $\pi$ generates $O$ from $I$, and $\pi$ implies the weakest constraint that can be implied while still generating $O$ from $I$.
-This maximises adaptability, because it allows a self-organising system to construct fit policies from a shorter history [bennett2023c].
-This formalised in definition 5 and illustrated by figure 1.
-Figure 1: Illustration of how child tasks are examples of parent tasks.
+This means a system that does so will construct fit policies from a shorter history [bennett2023c] than one that does not (see def.
+5 and proofs in appendix [bennett2024appendices]) [^10] The effect of insufficiently weak representations can be observed in large language models like GPT-4 [bennett2023d, gpt], where they fail to make consistent statements about an object because represent it as multiple objects..
 3 Relevance Realisation Through Causal Learning
@@
 To efficiently adapt, an organism must correctly anticipate valence, which means correctly identifying what causes valence.
-For convenience, we will now expand upon the concept of organism given in definition 6.
-All organisms must have preferences, because they will make some decisions and not others.
-Behavior implies policies, and every policy implies tasks, so we can define preferences as an binary relation over tasks.
-Natural selection favours adaptability, hence we assume it prefers organisims that prefer weaker policies, generalising to remain fit.
-When a system optimises for weak policies, we call it weak policy optimisation (WPO).
-If a policy is fit, then it must correctly predict causes of valence.
-Hence by constructing fit policies, an organism divides the world into relevant objects and properties.
+All organisms (see definition 6) must have preferences, because they will make some decisions and not others.
+Behaviour implies policies, and every policy implies tasks, so we can define preferences as an binary relation over tasks.
+Natural selection favours adaptability, hence it prefers organisms that optimise for weak policies, which we call weak policy optimisation (WPO).
+By constructing fit policies, an organism divides the world into relevant objects and properties.
 A weaker policy implies all more specific versions of itself, meaning those that more tightly constrain outputs by having a smaller extension.
@@
 Every fit policy would be a “causal identity” for something relevant, for example a specific object like “this coffee” or a weaker and more general concept like “all coffee”.
-We also need to introduce the notion of interaction between organisms.
-Organisms interact when they “affect” one another (in the physical rather than psychological sense of the word).
-This is formally defined in 9.
+Organisms interact when they “affect” one another (see definition 9) in the causal or physical rather than psychological sense of the word.
 Now, Artificial Intelligence (AI) and machine learning [russell2005ai, bishop2006, sutton2018, goertzel2014] are concerned with engineering adaptive agents.
@@
 Causal learning is demonstrably necessary to thrive in an interactive setting [pearl2018, richens2024robust].
-Where the causal graph is known in advance (for example if we were to be measuring the efficacy of medical interventions), this issue could be resolved using causal language (such as “Do Calculus” [pearl2009]) to represent an intervention from outside the system the graph describes.
-Such an intervention represent agency, in that an organism observing the environment draws conclusion about that environment, and then intervenes to change the environment.
-To illustrate this point, consider the following example from [bennett2023c].
-Suppose an organism named Bob is attempting to learn and predict the environment.
-Now assume Bob has observed Alice wearing a raincoat only when it rains, and that Bob has observed rain only at those times when Alice has been observed wearing a raincoat.
+Where the causal graph is known in advance (for example if measuring the efficacy of medical interventions), this issue could be resolved using causal language [pearl2009] to represent an intervention from outside the system the graph describes.
+To illustrate [bennett2023c], suppose Bob is attempting to learn and predict the environment.
+Bob has observed Alice wearing a raincoat only when it rains, and Bob has seen rain only when he observed Alice wearing a raincoat.
 If we represent the raincoat observation with a binary variable $O\in\{true,false\}$ and the advent of rain with a variable $R\in\{true,false\}$, and if Bob is a Bayesian, then Bob’s observations will lead to the conclusion that $p(R=true\mid O=true)=1$.
@@
 Assume Bob wants it to rain.
-Based on the belief $p(R=true\mid O=true)=1$, Bob may conclude that forcing $O=true$ by holding a gun to Alice’s head and demanding Alice wear a raincoat will cause it to rain.
+As $p(R=true\mid O=true)=1$, Bob may conclude that forcing $O=true$ by holding a gun to Alice’s head and demanding Alice wear a raincoat will cause it to rain.
 This is absurd.
@@
 To accurately represent the environment, we need to represent that $p(R=true\mid q)=p(R=true)\neq p(R=true\mid v)=1$, meaning both $q$ and $v$.
-To illustrate the problem visually, we started with the acyclic graph
+To illustrate visually, we started with the acyclic graph
 $O$$R$
@@
 Thus, the aforementioned $q$ is equivalent to $do[O=true]$, while $v$ is equivalent to $O=true$.
-The Psychophysical Principle of Causality
-We have explained why we need to account for causality.
-However, we have not explained how one might come to know all the objects involved (where did the variables come from?), or how they relate to one another causally (to form a graph).
+3.0.1 The Psychophysical Principle of Causality
+We have established the need for causal reasoning, but not how one might come to know the objects involved (where did the variables come from?), or how they relate to one another causally (to form a graph).
 There are two ways to learn causal graphs.
 One can either assume a set of variables and learn relations between them, or assume relations and learn the objects that fit them [bennett2023c].
-Beginning as we are with contentless states, we do the latter.
+We do the latter.
 Rather than assuming objects are primary and trying to learn causal relations between them, this assumes valence is primary and that abstract objects are constructed according to what causes valence.
@@
 Now assume Bob observes Alice taking an action that changes an aspect of the environment, represented by the variable $O$ (for example, Bob observes Alice putting on the raincoat).
-From the perspective of Bob, Alice’s action is the assignation of a value to a variable $A$, and to the variable $O$.
+From Bob’s perspective, Alice’s action assigns values to variables $A$ and $O$.
 There is no need to involve a do operator in this scenario because we can already represent that $p(R=true\mid A=x,O=true)=p(R=true)\neq p(R=true\mid A=y,O=true)=1$ (because Alice is part of the causal graph).
-This raises the question; if we do not require a $do$ operator to represent the actions of Alice, then why would we need it to represent an intervention by Bob?
-Couldn’t we represent the same information by introducing a new variable?
-The answer is yes [bennett2023c, dawid2002].
+Likewise, we can add variables for interventions of Bob [bennett2023c].
 $O$$B$$A$$R$
-Now we will show how it is possible to learn the relevant nodes to add, in the form of statements.
-The matter of which cause-and-effect relations are learned is determined by valence, and so the objects learned are statements classifying causes and valence.
+Now we will show how it is possible to learn these nodes.
+The matter of which cause-and-effect relations are learned is determined by valence, and so the nodes learned are statements classifying causes and valence.
 4 Relevant Causal Identities
-The formalism we employ began as an attempt to explain why some abstractions are formed, and not others [bennettmaruyama2022a].
-The optimal choice of policy for efficient adaptation is the weakest because it is implied by my statements in an embodied language.
-That language is dispositional, shaped by natural selection, and so a weaker policy is likely to be a fitter policy.
-As we have already shown, correct identification of cause and effect is often necessary to make accurate predictions and survive.
-WPO will thus correctly identify cause and effect where relevant.
-This is not just a case of substituting the “do” operator, but of simplifying the entire environment into objects and properties relevant to the organism’s motivations.
-We will now explain how relevant objects are classified, as “causal identities” [bennett2023c] and protosymbols [bennett2023d].
-4.1 Causal Learning
-In accord with [bennett2023d, def.
-6] we assume a vocabulary $\mathfrak{v}_{\mathfrak{b}}$ belonging to an organism $\mathfrak{b}$ (Bob).
-A “cause” in the context of this formalisation is not a variable but a statement $l\in L_{\mathfrak{v}_{\mathfrak{b}}}$ in that formal language.
+The optimal choice of policy for adaptability is the weakest because it is implied by more statements in an embodied language.
+That language is dispositional [bennettmaruyama2022a], shaped by natural selection, and so a weaker policy is likely to be a fit policy.
+WPO will correctly identify cause and effect where relevant, simplifying the entire environment into objects and properties relevant to the organism’s motivations.
+Assume a vocabulary $\mathfrak{v}_{\mathfrak{b}}$ belonging to an organism $\mathfrak{b}$ (Bob).
+A “cause” in the context of this formalisation is not a variable but a statement $l\in L_{\mathfrak{v}_{\mathfrak{b}}}$.
 The raincoat example would involve $obs,rain\in L_{\mathfrak{v}_{\mathfrak{b}}}$ such that:
@@
 $int$$obs$$rain$
-In other words, so long as $obs\neq int$, the intervention can be differentiated from the passive observation. We formally define this in definition 10.
+So long as $obs\neq int$, the intervention can be differentiated from the passive observation (see definition 10).
 This being the case, any set $c\subseteq int-obs$ could be used to identify the party undertaking the intervention, which is why $c$ is referred to as a “causal identity”.
@@
 A weaker or more general causal identity would be one that is shared by more interventions.
-For example, we might have two different interventions $a_{1}$ and $a_{2}$, with the observed effects $c_{1}\subset a_{1}$ and $c_{2}\subset a_{2}$, and causal identities $a_{1}-c_{1}=i_{1}$ and $a_{2}-c_{2}=i_{2}$.
-If $i_{3}=i_{1}\cap i_{2}\neq\emptyset$, then $i_{3}$ is a causal identity that is present in two interventions.
-Being a question of classification, we can express a causal identity in relation to a $\mathfrak{v}$-task, for which the causal identity is a policy (expanding upon [bennett2023c, def.
-10]), giving us definition 11.
-Example 1.
-Suppose we have organisms $\mathfrak{a}$ (Alice) and $\mathfrak{b}$ (Bob), and that the inputs Alice has experienced so far are $I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}$.
-These can be divided into those in which Bob affected Alice $S^{\mathfrak{b}}_{\mathfrak{a}}$ and those in which Bob did not $S^{\lnot\mathfrak{b}}_{\mathfrak{a}}=I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}-S^{\mathfrak{b}}_{\mathfrak{a}}$.
-By affecting Alice, Bob has intervened in Alice’s experience.
-Alice can construct a causal identity $b$ for Bob corresponding to interventions $INT=S^{\mathfrak{b}}_{\mathfrak{a}}$ and observations $OBS=S^{\lnot\mathfrak{b}}_{\mathfrak{a}}$.
+For example, suppose the inputs Alice has been subject to are $I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}$.
+These can be divided into those in which Bob affected Alice $I^{\mathfrak{b}}_{\mathfrak{a}}$ and those in which Bob did not $I^{\lnot\mathfrak{b}}_{\mathfrak{a}}=I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}-I^{\mathfrak{b}}_{\mathfrak{a}}$.
+Alice can construct a causal identity $b$ for Bob corresponding to interventions $INT=I^{\mathfrak{b}}_{\mathfrak{a}}$ and observations $OBS=I^{\lnot\mathfrak{b}}_{\mathfrak{a}}$ (see definition 11).
 4.1 Ascribing Intent to Other Objects
-The “do” operator is used to discern the difference between an event one has caused, and that same event passively observed.
-However, what is passive observation if not the result of an intervention by something other than one’s self?
 The distinction between “intervention” and not is misleading.
-Everything is an intervention.
-The question is not “is this an intervention” but “by whom was this intervention made?”.
+Passive observation is just bearing witness to an intervention by something other than one’s self.
+The question is not “is this an intervention" but “by whom was this intervention made?".
 Earlier, we arrived at the following graph in which Bob’s intervention was given by $int$.
 $int$$obs$$rain$
-However, what if a third person Larry puts the coat on Alice?
-Surely Bob can observe this, and so Bob’s observation of Larry’s intervention is $v\in L_{\mathfrak{v}_{\mathfrak{a}}}$ such that $obs\subset v$.
+What if a third person Larry puts the coat on Alice?
+Bob may observe this, and so Bob’s observation of Larry’s intervention is $v\in L_{\mathfrak{v}_{\mathfrak{a}}}$ such that $obs\subset v$.
 To account for this, Bob can construct a causal graph as below (with $b.$ representing Bob and $l.$ representing Larry).
@@
 Bob’s causal identity for himself $c_{b}\subset b.\ int-obs$ only represents the intervention by himself.
-However, now we can see that Bob must also construct a causal identity $c_{l}$ for Larry, where the $c_{l}\subset l.\ int-obs$.
-More generally, for an organism $\mathfrak{a}$ with an embodied language $L_{\mathfrak{v}_{\mathfrak{a}}}$ to construct a causal identity for an object $\mathfrak{b}$, it must first be the case that $\mathfrak{a}$ is affected by $\mathfrak{b}$ [bennett2023d], to satisfy the incentive precondition for causal identity.
-It depends on motive, or valence.
-Recall that to be affected is formally defined in definition 9.
+However, now we can see that Bob must also construct an identity $c_{l}$ for Larry, where the $c_{l}\subset l.\ int-obs$.
+For an organism $\mathfrak{a}$ with an embodied language $L_{\mathfrak{v}_{\mathfrak{a}}}$ to construct a causal identity for an object $\mathfrak{b}$, it must first be the case that $\mathfrak{a}$ is affected by $\mathfrak{b}$ [bennett2023d], to satisfy the incentive precondition for causal identity.
 Assume an organism $\mathfrak{a}$ is affected by $\mathfrak{b}$ given inputs $INT$, and not affected given inputs $OBS$.
@@
 The contents of $INT$ are “interventions” by $\mathfrak{b}$ and by learning $c$, a corresponding causal identity, $\mathfrak{a}$ can discern the existence of $\mathfrak{b}$.
-This is not to say that $\mathfrak{b}$ decides (chooses an output) or is even alive.
-What is important is that $\mathfrak{b}$ affects $\mathfrak{a}$, making it possible to discern when these interventions are a consequence of $\mathfrak{b}$’s existence.
+This is not to say that $\mathfrak{b}$ has intent, but intent can be ascribed to $\mathfrak{b}$ because $\mathfrak{b}$ affects $\mathfrak{a}$, who can then discern when interventions are a consequence of $\mathfrak{b}$.
 4.2 Preconditions
 There are preconditions for the existence of a causal identity.
-First, the vocabulary $\mathfrak{v}_{\mathfrak{a}}$ of an organism $\mathfrak{a}$ must be large and expressive enough to ensure that observations are distinguishable from interventions, meaning we have a vocabulary of sufficient scale.
+First, the vocabulary $\mathfrak{v}_{\mathfrak{a}}$ of an organism $\mathfrak{a}$ must be of sufficient scale to ensure that observations are distinguishable from interventions.
 Second, there must be an incentive to construct it.
 Inference is only possible if some states are preferable to others.
-It is a value judgement.
-As Hume pointed out, one cannot derive what “ought” to be from a statement of what “is”.
+One cannot derive what “ought” to be from a statement of what “is”.
 Natural selection provides a notion of what ought to be, by eliminating anything which ought not.
@@
 4.3 Realising Lower Order States And Higher Order Meta Representations
-The weaker the causal identity, the more pervasive the “identity” in the sense of being part of more interventions.
-Learning is not just about constructing policies but joining them together into weaker, more abstract policies.
-By learning policies that correctly identify the causes of valence at different levels of abstraction, the organism engages in relevance realisation.
-This is formalised in definition 6.
-An organism $\mathfrak{o}$ has a set of policies it has learned.
-Each and every policy implies a $\mathfrak{v}$-task.
+Each policy an organism $\mathfrak{o}$ learns implies $\mathfrak{v}$-tasks.
 A $\mathfrak{v}$-task is a triadic relation between inputs, outputs and policies which resembles Peircean semiosis [sep-peirce-semiotics, bennett2023d] of sign, referent and interpretant.
-Hence, in definition 7 we formalise this as a “protosymbol” [^13] Proto because it is something more primitive than a symbol as conceived of by Peirce.
-system $\mathfrak{s}_{\mathfrak{o}}$ for the organism $\mathfrak{o}$.
-Interaction is defined in terms of choosing an output, which means there is inevitably a policy and a policy implies a $\mathfrak{v}$-task, so definition 6 includes a preference order over tasks, which is used in definition 8 to formalise the interpretation of inputs (in terms of constraints, rather than an algorithm).
-In particular, the organism must always act according to an interpretation, and some interpretations imply others.
-Related as they are in a lattice, protosymbols are analogous lower order states and higher order meta representations.
+We formalise this as a “protosymbol” [^11] Proto because it is something more primitive than a symbol as conceived of by Peirce.
+system $\mathfrak{s}_{\mathfrak{o}}$ for the organism $\mathfrak{o}$ (see definitions 6 and 7).
 Tasks exist in a “generational hierarchy”.
@@
 Higher level tasks are more general, and have fewer policies because only very weak policies could complete them.
+Related as they are in a lattice, protosymbols are analogous lower order states and higher order meta representations.
 Some have framed consciousness as a problem of moving from unary, to dyadic, to triadic relations [goertzel2006].
 Similarly, we have gone from unary states to dyadic programs, to triadic tasks and protosymbols.
-Figure 2: Illustration of how the generational hierarchy of tasks is a lattice of protosymbols.
-The question now is when does an organism realise a causal identity for when it causes valence? When does it construct a causal identity for itself?
 5 Multi-Layered Self-Organisation
@@
 5.1 The First Order Self
-A first order self (1ST henceforth, see figure 3) is functionally equivalent to reafference, which has been observed in the mammalian mid-brain and insect central cortex [merker2007, barron2016].
-Reafference lets an organism discern the consequences of its actions, and so a 1ST order self is necessary for accurate inference in an interactive setting in the same sense as Pearl’s ‘do’ operator [pearl2018].
-A 1ST order self formally defined in definition 13.
-It serves as the locus of self related information processing and experience [bennett2023c], allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction.
-But also, to anticipate future planned interactions (for example, an insect navigating its environment [barron2016]), and recollect past interactions (it is a subset of all those relevant “statements” in the formal language).
+A first order self (1ST henceforth, formalised in definition 13) allows an organism discern the consequences of its actions.
+This serves as the locus of self related information processing and experience [bennett2023c], allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction.
+1ST order selves amount to reafference [merker2007, barron2016], observed in mammals and insects.
 Figure 1: Visual intuition for a 1ST order self. The organism constructs a causal identity for itself and can relate that to observed events.
-Natural selection prefers efficiency.
-The absence of a single, centralised causal identity can create inefficiencies [^14] The effect of insufficiently weak representations can be observed in large language models like GPT-4 [bennett2023d, gpt], where they fail to make consistent statements about an object because represent it as multiple objects..
-Decentralised or asynchronous control might be advantageous in some circumstances (e.g.
-an ant colony), but not others (e.g.
-in an individual human body).
-An organism that has a distributed control system might have a “self” for each part of that system, and perhaps the co-ordination of parts might seem to suggest the existence of a centralised “self”.
-Redundancy might be useful in some circumstances, but there is also a cost - more data may be required to learn and adapt, because the optimal choice of policy for accurate generalisation is the “weakest” [bennett2023b].
-If the scale and incentive preconditions are met for a causal identity for one’s self, then the organism must learn one.
-This might be partly hard-coded, but our focus is on inference by the organism itself.
+A 1ST order self may require centralisation and a “solid brain” with a persistent structure [sole2019] (an individual human), as opposed to a “liquid brain” (e.g.
+an ant colony or a population of humans).
 5.2 The Second Order Selves
@@
 It is an extension of the 1ST order self.
-More formally using the notation given in definition 15, assume $\mathfrak{a}$ and $\mathfrak{b}$ are organisms that evolved to accurately predict one another’s behaviour.
+More formally using the notation given in the appendix [bennett2024appendices] (see quick reference guide in section 2), assume $\mathfrak{a}$ and $\mathfrak{b}$ are organisms that evolved to accurately predict one another’s behaviour.
 Assume $\mathfrak{a}$ constructs a causal identity $c^{\mathfrak{b}}_{\mathfrak{a}}$ to predict $\mathfrak{b}$ given input $i_{\mathfrak{a}}\in I_{\mu_{\mathfrak{a}}}$, of which a second order self $c^{\mathfrak{ba}}_{\mathfrak{a}}$ is part.
 Likewise, $\mathfrak{b}$ constructs $c^{\mathfrak{a}}_{\mathfrak{b}}$ to predict $\mathfrak{a}$ given input $i_{\mathfrak{b}}\in I_{\mu_{\mathfrak{b}}}$, of which $c^{\mathfrak{ab}}_{\mathfrak{b}}$ is part.
-What is important here is that each organism’s intent is to some extent inferred by the other, and that fact inevitably changes the sorts of policies that will be “fit”.
+What is important here is that each organism’s intent is to some extent inferred by the other, changing the sorts of policies that are fit.
 For example, second order self means each knows the other can anticipate manipulation, which means the optimal policy will often be to have rather than feign intent that aligns to some extent with the other party’s desires, to co-operate [alexander2022] [^14] Depending upon circumstances, for example organisms may co-operate in some circumstances but not others [bennett2023d], and transient relations, information asymmetry and other factors can make deceit a more attractive option..
@@
 5.3 The Third Order Selves
-We can continue scaling the vocabulary and capacity for WPO indefinitely, however for the purposes of this paper we will stop at 3RD order selves, because this appears to be the highest level of human consciousness, known as Meta self-awareness [morin2006].
+We can scale preconditions indefinitely, however for the purposes of this paper we will stop at 3RD order selves, because this appears to be the highest level of human consciousness, known as Meta self-awareness [morin2006].
 This is the awareness that one is self-aware.
@@
 Up to now we have developed the conceptual toolbox that we use to dissolve the hard problem.
-We started with the observation that self-organizing systems such as a human body processes information to maintain itself in the face of change both within and without.
+We started with the observation that human body self organises to maintain itself in the face of change both within and without.
 An organism does not spring into existence understanding number systems or objects, but constructs them through interaction with the environment.
@@
 Quality comes first and experiences should be regarded on a continuum rather than a switch on/ switch off phenomena.
-All living systems experience the world through their bodies and as such, there is something what it is like to experience the world in that basic way (even when one is asleep or a baby).
+All living systems experience the world through their bodies and as such, there is something what it is like to experience the world in that basic way (even when one is asleep).
 One can access those phenomenal, experiential aspects at a higher level, true, but by accessing them, it doesn’t mean that one ‘constructs’ consciousness or one becomes a conscious being.
@@
 There would, however, be a very clear delineation between conscious and not.
-The absence of a 1ST order self would mean there would be no one policy we might point at to say “this is what it is like” to be an organism.
-No policy linking all interventions together, and so no “self” that experiences making them.
-Hence a 1ST order self must precede all others.
+The absence of a 1ST order self would mean there would be no policy linking all interventions together, and so no “self” that experiences making them.
+Hence a 1ST order self must precede higher orders.
 Our arguments align with those of Merker, Barron and Klein, who have linked subjective experience to reafference [merker2005, merker2007, barron2016].
@@
 7 Unifying Lower and Higher Order Theories of Consciousness
-The psychophysical principle of causality takes us from attraction to and repulsion from contentless physical states, up to higher order meta representations.
+The psychophysical principle of causality takes us from attraction to and repulsion from physical states, to higher order meta representations.
 This produces different orders of self.
-We suggest that the full richness of human subjective experience depends upon these different orders of self, as does human behavior [^18] We additionally speculate that different orders of learned self may interact through hierarchical planning in a distributed system, perhaps something like a connective core [shanahan2012].
-Arguably [boltuc2012], the full richness of human consciousness may be distinguished from phenomenal and access consciousness alone..
-Conscious “access” as a human has it involves meaningful report in the sense described by Grice, which as we’ve established requires 2ND order selves.
+We suggest that the full richness of human subjective experience depends upon these different orders of self [^15] Perhaps the full richness of “hard” consciousness [boltuc2012] depends upon the interaction of selves, perhaps in planning [shanahan2012]..
+Conscious “access” as a human has it involves meaningful report, which as we’ve established requires 2ND order selves.
 This is very different from the mere “access” to information.
-Hence, we hold that contents of which one is “access conscious”, in the sense that they are available for reasoning and report, can only be the contents of 2ND order selves.
 However, supposing we could somehow contrive a system with 2ND order selves but no 1ST, then 2ND order selves alone would still not allow one to reason about how one’s interventions might affect the contents of 2ND order selves to communicate meaning.
@@
 Hence access consciousness requires both 1ST and 2ND order selves.
-Hence our radical and provocative claim is that phenomenal consciousness without access consciousness is likely very common, but the reverse is impossible.
+Our radical and provocative claim is that phenomenal consciousness without access consciousness is likely very common, but the reverse is implausible.
 For example, Block holds that phenomenally conscious content is phenomenal, whereas access conscious content is representational.
@@
 Far from suggesting there is no such thing as qualia, this suggests instead that there is no such thing as purely representational content in anything but the tools we construct and our interpretations of their behaviour.
-Cells are a material with agency [ball2023].
-A human is a “multiscale competency architecture” [fields2020, levin2021, levin2024].
-Our intelligence is a swarm intelligence; the high level goal directed behaviour of a swarm of cells.
+Cells are a material with agency [ball2023], and human intelligence is the high level goal directed behaviour of a swarm of cells [fields2020, levin2021, levin2024].
 When we embody human abstractions (e.g.
@@
 To illustrate how our argument applies in the real world we describe stages of conscious organism.
-Each stage follows from scaling up supply and demand for WPO, through natural selection (see figure 6).
+Each stage follows from scaling up supply and demand for WPO, through natural selection.
 For each stage we point out animals which are likely to be at least so conscious:
-$0:$ Unconscious (e.g. a rock)
+$0:$ Inert (e.g. a rock)
 $1:$ Hard Coded (e.g. protozoan)
@@
 $5:$ 3RD Order Selves (e.g. human)
-Stage 0: Unconsciousness
-Stage zero may be understood as the consciousness we are willing to assign to a rock: more exactly, the lack of consciousness so assigned.
-  • Example: A rock.
 Stage 1: Hard Coded
-Stage one refers to adaptations hard-wired by natural selection. This is the sort of preset behaviour that allows complexity to persist [heylighen2023b] in a reasonably consistent environment.
+Stage one refers to adaptations hard-wired by natural selection, allowing complexity to persist [heylighen2023b] in a stable environment.
   • What: Hard-coded adaptations. Habituation and sensitization.
@@
 An entirely distributed control system can “learn”.
-Likewise, stage two is exemplified by the nematode C.
-elegans [rankin2022, willett2018], which has a centralised nervous system and exhibits some ability to adapt with experience.
-However, the absence of a “self” prohibits cause and effect reasoning, which as others have already pointed out must limit spatial, navigational abilities [barron2016].
+Likewise, stage two is exemplified by nematodes [rankin2022, willett2018] with a centralised nervous system and some ability to adapt with experience.
+However, the absence of a “self” limits causal reasoning, which as others have already pointed out must limit spatial, navigational abilities [barron2016].
 When starved C.
 elegans exhibit “increased locomotion and dispersal in a random, rather than directed, search” [barron2016, luersen2014, artyukhin2015], whereas something like a bee or an ant can recall and navigate to previously discovered food [wehner2013, seeley1995page295, oades1978].
-A nematode can learn, but the absence of a self for the purpose of causal reasoning limits adaptability.
   • What: Learning.
@@
 This is where phenomenal consciousness begins, with a 1ST order self.
-In biological terms this also implies reafference, which others have argued is the key to subjective experience, albeit for different reasons than what we have argued here [merker2005, merker2007, barron2016].
+In biological terms this implies reafference, which others have argued is the key to subjective experience, albeit for different reasons than what we have [merker2005, merker2007, barron2016].
 They identified a housefly as a good example of where subjective experience may begin, and we concur.
@@
 Intuitively, the policy that motivated behaviour is the intent of that behaviour [^16] In the same way declarative and imperative programs are equivalent [howard1980]..
-The “weaker” a policy is, the more behaviours it motivates.
-Conversely, if one’s actions share anything in common it is the intent that motivated them.
-The more diverse the actions, the more “general” or “high level” the intent they share.
-For example, the action of eating tends to involve the intent of satisfying hunger.
-If a policy is implied by all of an organism’s behaviour, then it might have motivated that behaviour.
-While this is an unusual form of words, we contend that an organism is such a policy.
-A stage three there is a self to feel simple hunger and pain.
-However, communication as described by Grice would be impossible [grice1969], because a stage three organism has no self-awareness, to represent another’s perception of their intent [bennett2023d].
-Nor would a stage three organism be able to conceive of its own death, or shame, because it cannot conceive of itself.
+For example, eating tends to involve the intent of satisfying hunger.
+A stage three organism can feel simple things like hunger, but cannot cannot conceive of itself from another’s perspective.
+Subsequently it cannot communicate in the Gricean sense [grice1969, bennett2023d], or conceive of its own death, or experience shame.
   • What: A 1ST order self. Reafference. Phenomenal consciousness.
@@
 Stage 4: 2ND Order Selves
-Stage four is the 2ND order self, and importantly this is where we hold access consciousness begins because it is where information is available for report in the Gricean sense.
-In other words, access consciousness follows phenomenal consciousness.
+Stage four is the 2ND order self, and this is where we hold access consciousness begins because it is where information is available for report in the Gricean sense.
 The ability of ravens to intentionally deceive [bugnyar2002] suggests they are at least stage four.
@@
 $\mathfrak{a}$ seems to predict not just the intent of $\mathfrak{b}$ (to steal the food), but $\mathfrak{b}$’s perception of $\mathfrak{a}$.
-It seems intuitively likely that dogs and cats have second order selves, as they must hunt reasonably intelligent animals and must anticipate how their actions are perceived.
+It seems likely that dogs and cats have second order selves, as they must hunt reasonably intelligent animals and must anticipate how their actions are perceived.
 For example, a cat anticipates its prey will flee when it is observed, and hides.
@@
   • How: Selection pressures that demand theory of mind.
-  • Why: A 2ND order self is necessary to anticipate, manipulate and communicate intent.
-It allows one to anticipate how others will perceive one’s behaviour.
-This could be anticipating the reactions of predator or prey, or navigating a social hierarchy.
+  • Why: A 2ND order self is necessary to anticipate, manipulate and communicate intent. This could be anticipating the reactions of predator or prey, or navigating a social hierarchy.
   • Example: Cats, dogs, ravens.
@@
 Humans appear to possess this, because we are aware that we are aware.
-The seemingly altruistic behavior observed in Australian magpies [potvin2022] suggests they may also have 3RD order selves.
+Altruistic behaviour observed in Australian magpies [potvin2022] suggests they may also have 3RD order selves.
   • What: Meta self-awareness. Inner narrative in which actors have inner narratives.
@@
 9 Conclusions: Why Nature Does Not Like Zombies
-In this paper we formally unified lower order states and higher order meta-representations, dissolving the hard problem of consciousness.
-We described a multilayered formalism illustrating how biological self-organising systems become phenomenally conscious when they construct a 1ST order self.
+We have described a multilayered formalism illustrating how biological self-organising systems become phenomenally conscious when they construct a 1ST order self.
 A human lacking a 1ST order self could not perform causal reasoning needed to adapt as humans evidently can [pearl2018].
-As previous research pointed out [bennett2024a], the computer metaphor with the seminal distinction between hardware and software is a simplification.
+As previous research pointed out [bennett2024a], the computer metaphor with hardware and software is a simplification.
 Rather, software is a state of hardware.
@@
 Only self-organising systems that need to adapt, motivated by valence, while keeping track of the self are conscious.
-Consciousness is a necessary adaptation, and a philosophical zombie is impossible.
+Consciousness is an adaptation, and a philosophical zombie is impossible.
 Furthermore, the conceivability of zombies depends on the ability to duplicate disposition but not phenomenal experience.
@@
 Our proposal lays the foundation of a formal science of consciousness, deeply connected with natural selection rather than abstract thinking, closer to the organic facts of humanity than the conceptual fiction of zombies.
-Compliance with Ethical Standards
 Funding
-This work was supported by a Fundaçao para a Ciencia e a Tecnologia (FCT) grant PTDC/FER-FIL/4802/2020 ; and 2020-02773 CEECIND FCT to A.C.
-Ethics Approval
-This article does not contain any studies with human participants or animals performed by any of the authors.
-Appendix
+Supported by a Fundaçao para a Ciencia e a Tecnologia (FCT) grant PTDC/FER-FIL/4802/2020; and 2020-02773 CEECIND FCT to A.C.
+Appendix A Appendix
 For convenience of reference, we have placed all definitions here.
@@
   • We say $x$ and $y$ are equivalent iff $E_{x}=E_{y}$.
-(notation) $E$ with a subscript is the extension of the subscript212121e.g. $E_{l}$ is the extension of $l$..
+(notation) $E$ with a subscript is the extension of the subscript181818e.g. $E_{l}$ is the extension of $l$..
 (intuitive summary) $L_{\mathfrak{v}}$ is everything which can be realised in this abstraction layer.
@@
 Definition 5 (learning).
-.
 A proxy $<$ is a binary relation on statements.
@@
     3RD and higher order selves: meta self-reflexive consciousness (human level hard consciousness) is achieved when $\mathfrak{o^{3}}\in\mathfrak{p}_{\mathfrak{o}}$.
-Proposition
+A.1 Proofs
+Versions of or alternatives to proofs 1-3 were published in [bennett2023b].
+Further details on these updated versions are available in on GitHub, along with the supporting experiments [bennett2024appendices].
+Theorem 1 (sufficiency).
+Assume $\alpha\sqsubset\omega$.
+The weakness proxy sufficient to maximise the probability that a parent $\omega$ is learned from a child $\alpha$ [^20] Assume there exist correct policies for $\omega$, because otherwise there would be no point in trying to learn it..
+Proof.
+You’re given the definition of $\mathfrak{v}$-task $\alpha$ from which you infer a hypothesis $\mathbf{h}\in{\Pi}_{\alpha}$. To learn $\omega$, you need $\mathbf{h}\in\Pi_{\omega}$:
+  1. 1.
+    For every $\mathbf{h}\in\Pi_{\alpha}$ there exists a $\mathfrak{v}$-task $\gamma_{\mathbf{h}}\in\Gamma_{\mathfrak{v}}$ s.t.
+$O_{\gamma_{\mathbf{h}}}=E_{\mathbf{h}}$, meaning $\mathbf{h}$ permits only correct outputs for that task regardless of input.
+We’ll call the highest level task $\gamma_{\mathbf{h}}$ s.t.
+$O_{\gamma_{\mathbf{h}}}=E_{\mathbf{h}}$ the policy task of $\mathbf{h}$ [^21] I’d like to give credit here to Nora Belrose for pointing out an error.
+Nora pointed out I was miscounting the number of tasks.
+As a result I realised I was not counting tasks, I was in fact counting policy tasks and had entirely neglected to mention this fact.
+This was a significant error which has now been corrected, with several additional steps added to account for equivalence..
+  2. 2.
+    $\omega$ is either the policy task of a policy in $\Pi_{\alpha}$, or a child thereof.
+  3. 3.
+    If a policy $\mathbf{h}$ is correct for a parent of $\omega$, then it is also correct for $\omega$.
+Hence we should choose $\mathbf{h}$ that has a policy task with the largest number of children.
+As tasks are uniformly distributed, that will maximise the probability that $\omega$ is $\gamma_{\mathbf{h}}$ or a child thereof.
+  4. 4.
+    For the purpose of this proof, we say one task is equivalent [^22] This is because switching from $\beta$ to $\zeta$ s.t.
+$I_{\beta}\neq I_{\zeta}$ and $O_{\beta}=O_{\zeta}$ would be to pursue the same goal in different circumstances.
+This is because inputs are subsets of outputs, so both sets of inputs are implied by the outputs.
+$O_{\zeta}$ implies $I_{\beta}$ and $O_{\beta}$ implies $I_{\zeta}$ to another if it has the same correct outputs.
+  5. 5.
+    No two policies in $\Pi_{\alpha}$ have the same policy task [^23] Every policy task for policies of $\alpha$ is non-equivalent from the others..
+This is because all the policies in $\Pi_{\alpha}$ are derived from the same set inputs, $I_{\alpha}$.
+  6. 6.
+    The set of statements which might be outputs addressing inputs in ${I}_{\omega}$ and not ${I}_{\alpha}$, is $\overline{E_{{I}_{\alpha}}}=\{l\in L_{\mathfrak{v}}:l\notin E_{{I}_{\alpha}}\}$ [^24] This is because $E_{{I}_{\alpha}}$ contains every statement which is a correct output or an incorrect output, and $\overline{E_{{I}_{\alpha}}}$ contains every statement which could possibly be in $I_{\omega}$, $E_{I_{\omega}}$ and thus $O_{\omega}$..
+  7. 7.
+    For any given $\mathbf{h}\in{\Pi}_{\alpha}$, the extension $E_{\mathbf{h}}$ of $\mathbf{h}$ is the set of outputs $\mathbf{h}$ implies.
+The subset of $E_{\mathbf{h}}$ which fall outside the scope of what is required for the known task $\alpha$ is $\overline{E_{{I}_{\alpha}}}\cap E_{\mathbf{h}}$ [^25] This is because $E_{{I}_{\alpha}}$ is the set of all conceivable outputs by which one might attempt to complete $\alpha$, and so the set of all outputs that can’t be made when undertaking $\alpha$ is $\overline{E_{{I}_{\alpha}}}$ because those outputs occur given inputs that aren’t part of ${I}_{\alpha}$..
+  8. 8.
+    $\left|\overline{E_{{I}_{\alpha}}}\cap E_{\mathbf{h}}\right|$ increases monotonically with $\left|E_{\mathbf{h}}\right|$, because for all $e\in E_{\mathbf{h}}$ is $e\notin\overline{E_{{I}_{\alpha}}}$ iff $e\in E_{{I}_{\alpha}}$.
+  9. 9.
+    $2^{\left|\overline{E_{{I}_{\alpha}}}\cap E_{\mathbf{h}}\right|}$ is the number of non-equivalent parents of $\alpha$ to which $\mathbf{h}$ generalises.
+It increases monotonically with the weakness of $\mathbf{h}$.
+  10. 10.
+    Given $\mathfrak{v}$-tasks are uniformly distributed and $\Pi_{\alpha}\cap\Pi_{\omega}\neq\emptyset$, the probability that $\mathbf{h}\in{\Pi}_{\alpha}$ generalises to $\omega$ is
+    $p(\mathbf{h}\in{\Pi}_{\omega}\mid\mathbf{h}\in{\Pi}_{\alpha},\alpha\sqsubset\omega)=\frac{2^{\left|\overline{E_{{I}_{\alpha}}}\cap E_{\mathbf{h}}\right|}}{2^{\left|\overline{E_{{I}_{\alpha}}}\right|}}$
+$p(\mathbf{h}\in{\Pi}_{\omega}\mid\mathbf{h}\in{\Pi}_{\alpha},\alpha\sqsubset\omega)$ is maximised when $\left|E_{\mathbf{h}}\right|$ is maximised.
+Recall from definition 4 that $<_{w}$ is the weakness proxy.
+For statements $l_{1},l_{2}$ we have $l_{1}<_{w}l_{2}$ iff $\lvert E_{l_{1}}\rvert<\lvert E_{l_{2}}\rvert$.
+$\mathbf{h}$ that maximises $<_{w}$ will also maximise $p(\mathbf{h}\in{\Pi}_{\omega}\mid\mathbf{h}\in{\Pi}_{\alpha},\alpha\sqsubset\omega)$.
+Hence the weakness proxy maximises the probability that [^26] Subsequently it also maximises the sample efficiency with which a parent $\omega$ is learned from a child $\alpha$.
+a parent $\omega$ is learned from a child $\alpha$.
+$\square$
+Theorem 2 (necessity).
+To maximise the probability of learning $\omega$ from $\alpha$, it is necessary to use weakness as a proxy.
+Proof.
+Let $\alpha$ and $\omega$ be defined exactly as they were in the proof of prop. 1.
+  1. 1.
+    If $\mathbf{h}\in{\Pi}_{\alpha}$ and $E_{{I}_{\omega}}\cap E_{\mathbf{h}}={O}_{\omega}$, then it must be he case that ${O}_{\omega}\subseteq E_{\mathbf{h}}$.
+  2. 2.
+    If $\left|E_{\mathbf{h}}\right|<\left|{O}_{\omega}\right|$ then generalisation cannot occur, because that would mean that ${O}_{\omega}\not\subseteq E_{\mathbf{h}}$.
+  3. 3.
+    Therefore generalisation is only possible if $\left|E_{\mathbf{h}}\right|\geq\left|{O}_{\omega}\right|$, meaning a sufficiently weak hypothesis is necessary to generalise from child to parent.
+  4. 4.
+    For any two hypotheses $\mathbf{h}_{1}$ and $\mathbf{h}_{2}$, if $\mid E_{\mathbf{h}_{1}}\mid<\mid E_{\mathbf{h}_{2}}\mid$ then the probability $p(\left|E_{\mathbf{h}_{1}}\right|\geq\left|{O}_{\omega}\right|)<p(\left|E_{\mathbf{h}_{2}}\right|\geq\left|{O}_{\omega}\right|)$ because tasks are uniformly distributed.
+  5. 5.
+    Hence the probability that $\left|E_{{m}}\right|\geq\left|{O}_{\omega}\right|$ is maximised when $\left|E_{{m}}\right|$ is maximised.
+To maximise the probability of learning $\omega$ from $\alpha$, it is necessary to select the weakest hypothesis.
+To select the weakest hypothesis, it is necessary to use the weakness proxy. $\square$
+Theorem 3 (simplicity sub-optimality).
+Description length is neither a necessary nor sufficient proxy for the purposes of maximising the probability that induction generalises.
+Proof.
+In propositions 1 and 2 we proved that weakness is a necessary and sufficient choice of proxy to maximise the probability of generalisation.
+It follows that either maximising $\frac{1}{\left|m\right|}$ (minimising description length) maximises $\left|E_{m}\right|$ (weakness), or minimisation of description length is unnecessary to maximise the probability of generalisation.
+Assume the former, and we’ll construct a counterexample with $\mathfrak{v}=\{a,b,c,d,e,f,g,h,j,k,z\}$ s.t.
+$L_{\mathfrak{v}}=\{\{a,b,c,d,j,k,z\},\{e,b,c,d,k\},\{a,f,c,d,j\},\{e,b,g,d,j,k,z\},\newline \{a,f,c,h,j,k\},\{e,f,g,h,j,k\}\}$ and a task $\alpha$ where
+  • ${I}_{\alpha}=\{\{a,b\},\{e,b\}\}$
+  • ${O}_{\alpha}=\{\{a,b,c,d,j,k,z\},\{e,b,g,d,j,k,z\}\}$
+  • ${\Pi}_{\alpha}=\{\{z\},\{j,k\}\}$
+Weakness as a proxy selects $\{j,k\}$, while description length as a proxy selects $\{z\}$.
+This demonstrates the minimising description length does not necessarily maximise weakness, and maximising weakness does not minimise description length.
+As weakness is necessary and sufficient to maximise the probability of generalisation, it follows that minimising description length is neither.
+$\square$
+The following sketch is more for intuition and explanation than anything formal.
+Theorem 4.
 An organism that uses weakness as its proxy will learn an $n^{th}$ order self if the incentive and scale preconditions are met for that order of self.
@@
 $\square$
+[…]

### v4

@@
-xxxxx, xxxxx, xxxx
+
 Abstract
@@
 keywords:
-hard problem, consciousness
-journal: rsif
-{fmtext}
+hard problem of consciousness
 1 Introduction
@@
 occupy space).
-Yet, unlike physical systems, biological systems are dissipative systems using energy to self-organise in the face of entropic decay and environmental perturbation [nagel1974, chalmers1995].
-Initially formally introduced in the field of cybernetics [ashby1947, foerster1960] the notion of self-organisation has been applied to various disciplines including physics [haken1983], biology [camazine2003, bell2007] and neuroscience [kelso1997, Friston2010, tognoli2014].
-Self-organisation is typically defined as the spontaneous emergence of spatiotemporal order or pattern-formation processes in physical and biological systems resulting from interactions of its components with the environment [camazine2001, seeley2002, rosas2018].
+Yet, unlike physical systems, biological systems are dissipative systems using energy to self-organise in the face of entropic decay and environmental perturbation [1, 2].
+Initially formally introduced in the field of cybernetics [3, 4] the notion of self-organisation has been applied to various disciplines including physics [5], biology [6, 7] and neuroscience [8, 9, 10].
+Self-organisation is typically defined as the spontaneous emergence of spatiotemporal order or pattern-formation processes in physical and biological systems resulting from interactions of its components with the environment [11, 12, 13].
 Interestingly, much self and world related information processing goes on behind the scenes, or “in the dark” so to speak, that is, without being constantly present to our conscious minds.
 But why doesn’t all information processing go in the dark?
-This question is the subject of long-standing debate [seth2022].
-One highly influential view is that consciousness has two aspects [seth2022].
-The first is functional, by which we mean the ability to access and communicate information [block1995].
-How function relates to consciousness is considered the “easy problem” of consciousness [chalmers1995].
-The second aspect is “what it is like” to consciously experience information processing, or phenomenal consciousness [chalmers1995, block1995, nagel1974, gallagher2021, fuchs2017].
+This question is the subject of long-standing debate [14].
+One highly influential view is that consciousness has two aspects [14].
+The first is functional, by which we mean the ability to access and communicate information [15].
+How function relates to consciousness is considered the “easy problem” of consciousness [2].
+The second aspect is “what it is like” to consciously experience information processing, or phenomenal consciousness [2, 15, 1, 16, 17].
 This doesn’t mean just global states like being awake, but more local states like smelling a cup of coffee.
-These local contents or “qualia” are characterised by what it is like to be in them [seth2022].
+These local contents or “qualia” are characterised by what it is like to be in them [14].
 It is unclear the extent to which functional and phenomenal aspects are independent.
-David Chalmers has influentially suggested that it may be possible to construct a “philosophical zombie” which acts in every way like a person but has no qualia [block1995].
+David Chalmers has influentially suggested that it may be possible to construct a “philosophical zombie” which acts in every way like a person but has no qualia [15].
 For example, a thermostat certainly detects heat and so processes information, but there presumably is not anything it is like to be a thermostat.
@@
 This is the “hard problem” of consciousness.
-The hard problem has sparked a substantial body of work, detailed discussion of which [seth2022, northoff2014] lies beyond the scope of our paper.
+The hard problem has sparked a substantial body of work, detailed discussion of which [14, 18] lies beyond the scope of our paper.
 In what follows we build upon the distinction between lower and higher order theories of information processing in relation to conscious experiences.
-Higher order thought theory (HOT) [rosenthal2005, brown2019] holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states.
+Higher order thought theory (HOT) [19, 20] holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states.
 Lower order states may include emotions and perceptions, while higher order meta-representations reflect upon those.
@@
 In this paper we suggest a way to dissolve the “hard problem” of consciousness is to reverse the order and start with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states.
-Longtime considered a fringe approach, the embodied cognition paradigm [varela2016] has recently gained substantial influence in cognitive science and philosophy [ciaunica2023, levin2024, seth2018].
+Longtime considered a fringe approach, the embodied cognition paradigm [21] has recently gained substantial influence in cognitive science and philosophy [22, 23, 24].
 The idea is that instead of considering the body as a mere device designed to fuel and contain the mind (a device that can be replaced with a vat or a robot, for example), one must consider the mind as serving the self-sustaining needs of a surviving body.
@@
 Our approach is to try to rigorously define consciousness from first principles and show that some aspects of functional consciousness depends on phenomenal consciousness in a manner that makes zombies impossible, “dissolving” the hard problem by showing the phenomenal to be intrinsically functional [^2] Note that we are far from the first to claim that the phenomenal is functional.
-For example, the proposed Conscious Turing Machine [blum2020] based on Global Workspace Theory [baars1997], and constructivist approaches to artificial intelligence [wang2020] take similar position.
+For example, the proposed Conscious Turing Machine [25] based on Global Workspace Theory [26], and constructivist approaches to artificial intelligence [27] take similar position.
 It is our explanation of how and why that is novel..
 We establish axioms that hold in every conceivable environment and show that it is impossible to have the function of consciousness without the subjective experience of it, and why most information processing goes on “in the dark”.
-Enactivism is roughly the view that information processing arises through interaction between an organism and its environment [thompson2007], and it is considered by some to be incompatible with narrow definitions of computation.
-The notion of computation is widely debated, and a detailed review of these discussions would lead us to a major digression [piccinini2015, piccinini2021].
+Enactivism is roughly the view that information processing arises through interaction between an organism and its environment [28], and it is considered by some to be incompatible with narrow definitions of computation.
+The notion of computation is widely debated, and a detailed review of these discussions would lead us to a major digression [29, 30].
 Here we define computation not in terms of abstract symbol shuffling and representation, but in concrete mechanistic terms where ‘information processing’ refers to the actual physical transition of a system from one state to another, so it is compatible with enactivism.
 Pancomputationalism is the idea that all systems are computational.
-We build upon a formalism called pancomputational enactivism [bennett2024a].
-To answer “why is anything conscious”, we use this formalism to show how lower and higher order theories, phenomenal and access consciousness all follow from first principles, scaling natural selection pressures and the ability to adapt [^3] This is loosely inspired by a scale-based framing of machine learning [sutton2019] - we are applying Sutton’s “bitter lesson” to biology..
-First, in Section 2 we integrate and extend separately published works of narrower scope [bennett2024a, bennett2023b, bennett2023c] to justify our model of self-organising systems.
+We build upon a formalism called pancomputational enactivism [31].
+To answer “why is anything conscious”, we use this formalism to show how lower and higher order theories, phenomenal and access consciousness all follow from first principles, scaling natural selection pressures and the ability to adapt [^3] This is loosely inspired by a scale-based framing of machine learning [32] - we are applying Sutton’s “bitter lesson” to biology..
+First, in Section 2 we integrate and extend separately published works of narrower scope [31, 33, 34] to justify our model of self-organising systems.
 In Sections 3 and 4 we explain how this model formalises relevance realisation and unifies lower and higher order theories.
@@
 We call this the psychophysical principle of causality.
-In Section 5 we extend previous work [bennett2023c] on causal learning and the development of first (1ST), second (2ND) and third (3RD) order selves, and explain how they are a consequence of scaling the ability of a self-organising system to adapt with natural selection pressures.
+In Section 5 we extend previous work [34] on causal learning and the development of first (1ST), second (2ND) and third (3RD) order selves, and explain how they are a consequence of scaling the ability of a self-organising system to adapt with natural selection pressures.
 In Sections 6 and 7 we explain how subjective experience requires a 1ST order self, and conversely why a 1ST order self implies there is something it is like to be an organism that has a 1ST order self.
@@
 2 Back to Foundations
-To meet length restrictions, the full mathematical definitions and proofs have been placed in the appendices [bennett2024appendices].
+To meet length restrictions, the full mathematical definitions have been placed in the appendices [35].
 The core arguments should be understandable without the math, but for convenience here is a quick reference guide:
@@
 For example, $c^{\mathfrak{b}}_{\mathfrak{a}}$ denotes a causal identity learned by $\mathfrak{a}$ to classify $\mathfrak{b}$’s interventions, and $c^{\mathfrak{ba}}_{\mathfrak{a}}\subset c^{\mathfrak{b}}_{\mathfrak{a}}$ denotes $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$.
-We begin by formalising all conceivable environments [bennett2024a, bennett2024b] from first principles:
+We begin by formalising all conceivable environments [31, 36] from first principles:
 Axiom 1: When there are things, we call these things the environment.
@@
 Hence, the environment is a set $\Phi$ of mutually exclusive states.
-We call a set of states a declarative “program” [^4] We do not concern ourselves with imperative programs, because there is an equivalence between declarative and imperative [howard1980]..
+We call a set of states a declarative “program” [^4] We do not concern ourselves with imperative programs, because there is an equivalence between declarative and imperative [37]..
 A program is “true” relative to the states it contains.
@@
 An aspect of the environment is realised by (exists in) a state if the programs it contains are facts.
-Previous work argues this formalises every conceivable environment [bennett2024b].
+Previous work argues this formalises every conceivable environment [36].
 To give some intuition as to why, note that we assume nothing of what states might be or contain.
@@
 However, the application of a function is a fact regardless of whether it takes another function as input., hence the body of an organism that persists across multiple states must be a set of programs.
-We call this set a vocabulary, and it is formalised in definition 2 (see appendix [bennett2024appendices]).
-Piccinini [piccinini2015, piccinini2021] divides computation into abstract and concrete sorts.
+We call this set a vocabulary, and it is formalised in definition 2 (see appendix [35]).
+Piccinini [29, 30] divides computation into abstract and concrete sorts.
 Abstract computations mean or represent whatever we say they do.
@@
 This might seem inconvenient, as intuitively we might prefer to think about bodies as an abstract ‘type’.
-However, doing so would undermine any claim we might make, as earlier results in artificial intelligence (AI) show [bennett2023c, bennett2024c].
+However, doing so would undermine any claim we might make, as earlier results in artificial intelligence (AI) show [34, 38].
 The meaning of abstractions depends upon interpretation, much like the behaviour of software is determined by the abstraction layer that “interprets” it.
@@
 That hardware is an abstraction layer in which the software exists (including “higher level” abstraction layers like the Python interpreter).
-In other words, what we call software is nothing more than the state of hardware [bennett2024a].
-This has undermined all but the most subjective of claims regarding the behaviour of theorised, software superintelligence [hutter2010, leike2015, leike2018].
+In other words, what we call software is nothing more than the state of hardware [31].
+This has undermined all but the most subjective of claims regarding the behaviour of theorised, software superintelligence [39, 40, 41].
 It is a flaw in the very idea of intelligent software.
-The distinction between software mind and hardware embodiment is subsequently called computational dualism [bennett2024a], because it is reminiscent of how traditional Cartesianism conceived of a mental substance distinct from physical.
+The distinction between software mind and hardware embodiment is subsequently called computational dualism [31], because it is reminiscent of how traditional Cartesianism conceived of a mental substance distinct from physical.
 If we are to dissolve the hard problem by starting with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states, then we cannot presuppose organisms use particular abstractions.
-Instead we must explain why particular abstractions are formed, in terms of contentless states, as the product of interaction between an organism and its environment [thompson2007, varela2016, rolla2021].
-This process is sometimes called “relevance realisation” [vervaeke2012, vervaeke2013a, vervaeke2013b].
-Some have argued enactivism is incompatible with computation because the set of possible abstractions that must be searched for relevance realisation is intractable [jaeger2024].
+Instead we must explain why particular abstractions are formed, in terms of contentless states, as the product of interaction between an organism and its environment [28, 21, 42].
+This process is sometimes called “relevance realisation” [43, 44, 45].
+Some have argued enactivism is incompatible with computation because the set of possible abstractions that must be searched for relevance realisation is intractable [46].
 For tractability, one must first isolate a “small world” of relevant information from the “big world” of all information.
 Part of the reason we can address this is because we formalise embodiment, in concrete terms from the level of irreducible physical states.
-Embodiment implies finite resource constraints, so we adopt one more axiom from [bennett2024b].
-Axiom 3: All aspects of the environment are spatially extended [bennett2024b].
+Embodiment implies finite resource constraints, so we adopt one more axiom from [36].
+Axiom 3: All aspects of the environment are spatially extended [36].
 This means that a body is realised by only a finite number of states (see def.
@@
 Most importantly, we avoid the distinction between software and hardware and so avoid computational dualism.
-Embodiment is dispositional.
-As natural selection evolves a vocabulary, it begins to isolate a “small world” relevant to survival [jaeger2024, bennett2024b, friston2023], from the intractable “big world” $P$ of all information.
+Embodiment is dispositional. As natural selection evolves a vocabulary, it begins to isolate a “small world” relevant to survival [46, 36, 47], from the intractable “big world” $P$ of all information.
 2.2 Self-Organising Systems as Self and World Constraints
@@
 If the set of all fit behaviour is a task $\mu$, then the organism remains fit then $\mathfrak{h}\sqsubset\mu$.
-“I survive therefore my model is viable.” - Mark Solms [solms2021]
+“I survive therefore my model is viable.” - Mark Solms [48]
 $\mathfrak{h}$ implies a set $\Pi_{\mathfrak{h}}$ of policies, and not all of those policies will constrain outputs the same way given new inputs.
@@
 An organism learns fit behaviour by learning $\pi\in\Pi_{\mathfrak{h}}$ such that $\pi\in\Pi_{\mu}$.
-“The best model of the world is the world itself” - Rodney Brooks [dreyfus2007]
-Earlier work [bennett2023b] showed that the optimal strategy to adapt as fast as possible [^9] Meaning to converge on a fit policy given the smallest history possible, to realise fit behaviour.
+“The best model of the world is the world itself” - Rodney Brooks [49]
+Earlier work [33] showed that the optimal strategy to adapt as fast as possible [^9] Meaning to converge on a fit policy given the smallest history possible, to realise fit behaviour.
 is to prefer “weaker” policies, meaning those with larger extensions.
 The policy which generated outputs $O$ given inputs $I$ is most efficiently identified by constructing a policy $\pi$ such that $\pi$ generates $O$ from $I$, and $\pi$ implies the weakest constraint that can be implied while still generating $O$ from $I$.
-This means a system that does so will construct fit policies from a shorter history [bennett2023c] than one that does not (see def.
-5 and proofs in appendix [bennett2024appendices]) [^10] The effect of insufficiently weak representations can be observed in large language models like GPT-4 [bennett2023d, gpt], where they fail to make consistent statements about an object because represent it as multiple objects..
+This means a system that does so will construct fit policies from a shorter history [34] than one that does not (see def.
+5 and proofs in appendix [35]) [^10] The effect of insufficiently weak representations can be observed in large language models like GPT-4 [50, 51], where they fail to make consistent statements about an object because represent it as multiple objects..
 3 Relevance Realisation Through Causal Learning
@@
 Organisms interact when they “affect” one another (see definition 9) in the causal or physical rather than psychological sense of the word.
-Now, Artificial Intelligence (AI) and machine learning [russell2005ai, bishop2006, sutton2018, goertzel2014] are concerned with engineering adaptive agents.
+Now, Artificial Intelligence (AI) and machine learning [52, 53, 54, 55] are concerned with engineering adaptive agents.
 In that context, causality has become a mainstream topic of research.
-Causal learning is demonstrably necessary to thrive in an interactive setting [pearl2018, richens2024robust].
-Where the causal graph is known in advance (for example if measuring the efficacy of medical interventions), this issue could be resolved using causal language [pearl2009] to represent an intervention from outside the system the graph describes.
-To illustrate [bennett2023c], suppose Bob is attempting to learn and predict the environment.
+Causal learning is demonstrably necessary to thrive in an interactive setting [56, 57].
+Where the causal graph is known in advance (for example if measuring the efficacy of medical interventions), this issue could be resolved using causal language [58] to represent an intervention from outside the system the graph describes.
+To illustrate [34], suppose Bob is attempting to learn and predict the environment.
 Bob has observed Alice wearing a raincoat only when it rains, and Bob has seen rain only when he observed Alice wearing a raincoat.
@@
 There are two ways to learn causal graphs.
-One can either assume a set of variables and learn relations between them, or assume relations and learn the objects that fit them [bennett2023c].
+One can either assume a set of variables and learn relations between them, or assume relations and learn the objects that fit them [34].
 We do the latter.
@@
 This is what will take us from attraction to and repulsion from contentless physical states, through lower order thoughts to higher order meta representations.
-We start with the complete absence of content [thompson2016], and once a system learns policies there are distinct “contents”.
+We start with the complete absence of content [59], and once a system learns policies there are distinct “contents”.
 Policies serve to classify objects and properties.
 Death grounds meaning.
-To develop this idea, we first show adding additional nodes to a causal graph is an adequate substitute for the “do” operator [dawid2002].
+To develop this idea, we first show adding additional nodes to a causal graph is an adequate substitute for the “do” operator [60].
 Assume we again have Bob who constructs a causal graph of the environment.
@@
 There is no need to involve a do operator in this scenario because we can already represent that $p(R=true\mid A=x,O=true)=p(R=true)\neq p(R=true\mid A=y,O=true)=1$ (because Alice is part of the causal graph).
-Likewise, we can add variables for interventions of Bob [bennett2023c].
+Likewise, we can add variables for interventions of Bob [34].
 $O$$B$$A$$R$
@@
 The optimal choice of policy for adaptability is the weakest because it is implied by more statements in an embodied language.
-That language is dispositional [bennettmaruyama2022a], shaped by natural selection, and so a weaker policy is likely to be a fit policy.
+That language is dispositional [61], shaped by natural selection, and so a weaker policy is likely to be a fit policy.
 WPO will correctly identify cause and effect where relevant, simplifying the entire environment into objects and properties relevant to the organism’s motivations.
@@
 Passive observation is just bearing witness to an intervention by something other than one’s self.
-The question is not “is this an intervention" but “by whom was this intervention made?".
+The question is not “is this an intervention” but “by whom was this intervention made?”.
 Earlier, we arrived at the following graph in which Bob’s intervention was given by $int$.
@@
 However, now we can see that Bob must also construct an identity $c_{l}$ for Larry, where the $c_{l}\subset l.\ int-obs$.
-For an organism $\mathfrak{a}$ with an embodied language $L_{\mathfrak{v}_{\mathfrak{a}}}$ to construct a causal identity for an object $\mathfrak{b}$, it must first be the case that $\mathfrak{a}$ is affected by $\mathfrak{b}$ [bennett2023d], to satisfy the incentive precondition for causal identity.
+For an organism $\mathfrak{a}$ with an embodied language $L_{\mathfrak{v}_{\mathfrak{a}}}$ to construct a causal identity for an object $\mathfrak{b}$, it must first be the case that $\mathfrak{a}$ is affected by $\mathfrak{b}$ [50], to satisfy the incentive precondition for causal identity.
 Assume an organism $\mathfrak{a}$ is affected by $\mathfrak{b}$ given inputs $INT$, and not affected given inputs $OBS$.
@@
 Each policy an organism $\mathfrak{o}$ learns implies $\mathfrak{v}$-tasks.
-A $\mathfrak{v}$-task is a triadic relation between inputs, outputs and policies which resembles Peircean semiosis [sep-peirce-semiotics, bennett2023d] of sign, referent and interpretant.
+A $\mathfrak{v}$-task is a triadic relation between inputs, outputs and policies which resembles Peircean semiosis [62, 50] of sign, referent and interpretant.
 We formalise this as a “protosymbol” [^11] Proto because it is something more primitive than a symbol as conceived of by Peirce.
@@
 Related as they are in a lattice, protosymbols are analogous lower order states and higher order meta representations.
-Some have framed consciousness as a problem of moving from unary, to dyadic, to triadic relations [goertzel2006].
+Some have framed consciousness as a problem of moving from unary, to dyadic, to triadic relations [63].
 Similarly, we have gone from unary states to dyadic programs, to triadic tasks and protosymbols.
@@
 As the vocabulary and capacity for WPO scale, a greater variety of concepts can be learned.
-Progressively higher orders of ‘causal-identity’ for one’s self related information processing are constructed [bennett2023c].
-This lets us frame the construction of embodied selves in developmental [ciaunica2019] and evolutionary terms.
+Progressively higher orders of ‘causal-identity’ for one’s self related information processing are constructed [34].
+This lets us frame the construction of embodied selves in developmental [64] and evolutionary terms.
 5.1 The First Order Self
 A first order self (1ST henceforth, formalised in definition 13) allows an organism discern the consequences of its actions.
-This serves as the locus of self related information processing and experience [bennett2023c], allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction.
-1ST order selves amount to reafference [merker2007, barron2016], observed in mammals and insects.
+This serves as the locus of self related information processing and experience [34], allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction.
+1ST order selves amount to reafference [65, 66], observed in mammals and insects.
 Figure 1: Visual intuition for a 1ST order self. The organism constructs a causal identity for itself and can relate that to observed events.
-A 1ST order self may require centralisation and a “solid brain” with a persistent structure [sole2019] (an individual human), as opposed to a “liquid brain” (e.g.
-an ant colony or a population of humans).
+A 1ST order self may require centralisation and a “solid brain” with a persistent structure [67] (an individual human), as opposed to a “liquid brain” (e.g. an ant colony or a population of humans).
 5.2 The Second Order Selves
-Survival may demand organism $\mathfrak{a}$ infer $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s interventions (to see one’s self as if through another’s eyes [bennett2023d]).
+Survival may demand organism $\mathfrak{a}$ infer $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s interventions (to see one’s self as if through another’s eyes [50]).
 This is called a second order self (2ND henceforth).
-We argue that if access conscious contents are available for communication in the human sense, then they must be communicable in the Gricean sense [grice1957, grice1969].
+We argue that if access conscious contents are available for communication in the human sense, then they must be communicable in the Gricean sense [68, 69].
 Grice argued that communication is about the inference of intent.
@@
 Where a 1ST order self might allow one to observe a cat and form plans regarding causal interactions with the cat, a 2ND order self would allow one to be consciously aware of the cat for the purpose of reasoning and report.
-One can know of the cat, and that another organism knows of the cat, but a 2ND order self is insufficient to be aware that one is aware of the cat [^13] This is ‘meta-self-reflexive consciousness’ as some have described it [morin2006]..
+One can know of the cat, and that another organism knows of the cat, but a 2ND order self is insufficient to be aware that one is aware of the cat [^13] This is ‘meta-self-reflexive consciousness’ as some have described it [70]..
 Figure 2: Visual intuition for a 2ND order self.
@@
 It is an extension of the 1ST order self.
-More formally using the notation given in the appendix [bennett2024appendices] (see quick reference guide in section 2), assume $\mathfrak{a}$ and $\mathfrak{b}$ are organisms that evolved to accurately predict one another’s behaviour.
+More formally using the notation given in the appendix [35] (see quick reference guide in section 2), assume $\mathfrak{a}$ and $\mathfrak{b}$ are organisms that evolved to accurately predict one another’s behaviour.
 Assume $\mathfrak{a}$ constructs a causal identity $c^{\mathfrak{b}}_{\mathfrak{a}}$ to predict $\mathfrak{b}$ given input $i_{\mathfrak{a}}\in I_{\mu_{\mathfrak{a}}}$, of which a second order self $c^{\mathfrak{ba}}_{\mathfrak{a}}$ is part.
@@
 What is important here is that each organism’s intent is to some extent inferred by the other, changing the sorts of policies that are fit.
-For example, second order self means each knows the other can anticipate manipulation, which means the optimal policy will often be to have rather than feign intent that aligns to some extent with the other party’s desires, to co-operate [alexander2022] [^14] Depending upon circumstances, for example organisms may co-operate in some circumstances but not others [bennett2023d], and transient relations, information asymmetry and other factors can make deceit a more attractive option..
-Repeated interaction creates an iterated prisoner’s dilemma, incentivising co-operation and signals that both parties interpret similarly (the beginnings of language) [bennett2023d].
+For example, second order self means each knows the other can anticipate manipulation, which means the optimal policy will often be to have rather than feign intent that aligns to some extent with the other party’s desires, to co-operate [71] [^14] Depending upon circumstances, for example organisms may co-operate in some circumstances but not others [50], and transient relations, information asymmetry and other factors can make deceit a more attractive option..
+Repeated interaction creates an iterated prisoner’s dilemma, incentivising co-operation and signals that both parties interpret similarly (the beginnings of language) [50].
 To communicate in Gricean terms, $\mathfrak{a}$ must intend to convey meaning $m_{\mathfrak{a}}$, and $\mathfrak{b}$ must recognise this intent.
@@
 5.3 The Third Order Selves
-We can scale preconditions indefinitely, however for the purposes of this paper we will stop at 3RD order selves, because this appears to be the highest level of human consciousness, known as Meta self-awareness [morin2006].
+We can scale preconditions indefinitely, however for the purposes of this paper we will stop at 3RD order selves, because this appears to be the highest level of human consciousness, known as Meta self-awareness [70].
 This is the awareness that one is self-aware.
@@
 It has a quality, so the 1ST order self is “what it is like” to be that organism.
-Put another way, Nagel’s question of “what it is like to be” a particular organism could be answered if one could somehow have that organism’s first order self [nagel1974].
+Put another way, Nagel’s question of “what it is like to be” a particular organism could be answered if one could somehow have that organism’s first order self [1].
 An organism does not make a decision to interpret information, it just reacts.
@@
 Hence a 1ST order self must precede higher orders.
-Our arguments align with those of Merker, Barron and Klein, who have linked subjective experience to reafference [merker2005, merker2007, barron2016].
+Our arguments align with those of Merker, Barron and Klein, who have linked subjective experience to reafference [72, 65, 66].
 We agree that reafference is key, but we provide a very different explanation of why and how.
@@
 This produces different orders of self.
-We suggest that the full richness of human subjective experience depends upon these different orders of self [^15] Perhaps the full richness of “hard” consciousness [boltuc2012] depends upon the interaction of selves, perhaps in planning [shanahan2012]..
+We suggest that the full richness of human subjective experience depends upon these different orders of self [^15] Perhaps the full richness of “hard” consciousness [73] depends upon the interaction of selves, perhaps in planning [74]..
 Conscious “access” as a human has it involves meaningful report, which as we’ve established requires 2ND order selves.
@@
 Far from suggesting there is no such thing as qualia, this suggests instead that there is no such thing as purely representational content in anything but the tools we construct and our interpretations of their behaviour.
-Cells are a material with agency [ball2023], and human intelligence is the high level goal directed behaviour of a swarm of cells [fields2020, levin2021, levin2024].
+Cells are a material with agency [75], and human intelligence is the high level goal directed behaviour of a swarm of cells [76, 77, 23].
 When we embody human abstractions (e.g.
 arithmetic) in silico (e.g.
-the x86 instruction set), we disconnect the high level goal directed behaviour from the low level behaviour (and the motivating valence) that gave rise to it [bennett2024c].
+the x86 instruction set), we disconnect the high level goal directed behaviour from the low level behaviour (and the motivating valence) that gave rise to it [38].
 We call the information embodied in the computer “representational” because it means something to us.
-However when we embody our abstractions in silico we disconnect them from the valence that motivated their construction [bennett2024b].
+However when we embody our abstractions in silico we disconnect them from the valence that motivated their construction [36].
 This suggests there is no such thing as representational contents.
@@
 Stage 1: Hard Coded
-Stage one refers to adaptations hard-wired by natural selection, allowing complexity to persist [heylighen2023b] in a stable environment.
+Stage one refers to adaptations hard-wired by natural selection, allowing complexity to persist [78] in a stable environment.
   • What: Hard-coded adaptations. Habituation and sensitization.
@@
 A biological example of such a decentralised nervous system is the cubozoan box jellyfish Tripedalia cystophora.
-Even Tripedalia cystophora was recently shown to be capable of associative learning [bielecki2023].
+Even Tripedalia cystophora was recently shown to be capable of associative learning [79].
 An entirely distributed control system can “learn”.
-Likewise, stage two is exemplified by nematodes [rankin2022, willett2018] with a centralised nervous system and some ability to adapt with experience.
-However, the absence of a “self” limits causal reasoning, which as others have already pointed out must limit spatial, navigational abilities [barron2016].
+Likewise, stage two is exemplified by nematodes [80, 81] with a centralised nervous system and some ability to adapt with experience.
+However, the absence of a “self” limits causal reasoning, which as others have already pointed out must limit spatial, navigational abilities [66].
 When starved C.
-elegans exhibit “increased locomotion and dispersal in a random, rather than directed, search” [barron2016, luersen2014, artyukhin2015], whereas something like a bee or an ant can recall and navigate to previously discovered food [wehner2013, seeley1995page295, oades1978].
+elegans exhibit “increased locomotion and dispersal in a random, rather than directed, search” [66, 82, 83], whereas something like a bee or an ant can recall and navigate to previously discovered food [84, 85, 86].
   • What: Learning.
@@
 This is where phenomenal consciousness begins, with a 1ST order self.
-In biological terms this implies reafference, which others have argued is the key to subjective experience, albeit for different reasons than what we have [merker2005, merker2007, barron2016].
+In biological terms this implies reafference, which others have argued is the key to subjective experience, albeit for different reasons than what we have [72, 65, 66].
 They identified a housefly as a good example of where subjective experience may begin, and we concur.
 We also hold this is where an organism might be said to have intent.
-Intuitively, the policy that motivated behaviour is the intent of that behaviour [^16] In the same way declarative and imperative programs are equivalent [howard1980]..
+Intuitively, the policy that motivated behaviour is the intent of that behaviour [^16] In the same way declarative and imperative programs are equivalent [37]..
 For example, eating tends to involve the intent of satisfying hunger.
 A stage three organism can feel simple things like hunger, but cannot cannot conceive of itself from another’s perspective.
-Subsequently it cannot communicate in the Gricean sense [grice1969, bennett2023d], or conceive of its own death, or experience shame.
+Subsequently it cannot communicate in the Gricean sense [69, 50], or conceive of its own death, or experience shame.
   • What: A 1ST order self. Reafference. Phenomenal consciousness.
@@
 Stage four is the 2ND order self, and this is where we hold access consciousness begins because it is where information is available for report in the Gricean sense.
-The ability of ravens to intentionally deceive [bugnyar2002] suggests they are at least stage four.
+The ability of ravens to intentionally deceive [87] suggests they are at least stage four.
 Raven $\mathfrak{a}$, aware that it is being observed by raven $\mathfrak{b}$, will act as if it is hiding food in one location to mislead $\mathfrak{b}$, but will then move the food in another location unobserved by $\mathfrak{b}$.
@@
 Humans appear to possess this, because we are aware that we are aware.
-Altruistic behaviour observed in Australian magpies [potvin2022] suggests they may also have 3RD order selves.
+Altruistic behaviour observed in Australian magpies [88] suggests they may also have 3RD order selves.
   • What: Meta self-awareness. Inner narrative in which actors have inner narratives.
@@
 We have described a multilayered formalism illustrating how biological self-organising systems become phenomenally conscious when they construct a 1ST order self.
-A human lacking a 1ST order self could not perform causal reasoning needed to adapt as humans evidently can [pearl2018].
-As previous research pointed out [bennett2024a], the computer metaphor with hardware and software is a simplification.
+A human lacking a 1ST order self could not perform causal reasoning needed to adapt as humans evidently can [56].
+As previous research pointed out [31], the computer metaphor with hardware and software is a simplification.
 Rather, software is a state of hardware.
@@
 Compare the vast quantities of both training data and energy required by a large language model, to the small quantities humans need to solve a problem.
-Biological systems are more efficient, because they are adaptive at a every level [levin2024, bennett2024c].
+Biological systems are more efficient, because they are adaptive at a every level [23, 38].
 Hence, rather than trying to explain the mind in the abstract like software, we started at the level of the embodied organism.
@@
 Furthermore, the conceivability of zombies depends on the ability to duplicate disposition but not phenomenal experience.
-Carruth argued that if qualia are dispositional, then that is impossible [carruth2016].
+Carruth argued that if qualia are dispositional, then that is impossible [89].
 Our arguments align with Carruth’s.
@@
 Supported by a Fundaçao para a Ciencia e a Tecnologia (FCT) grant PTDC/FER-FIL/4802/2020; and 2020-02773 CEECIND FCT to A.C.
-Appendix A Appendix
+10 Appendix
 For convenience of reference, we have placed all definitions here.
 We have aimed to make the gist of the paper understandable without the math, hence the reader can either skip this section or refer back to it if needed.
-Many of the definitions have been adapted from a variety of preceding work [bennett2024a, bennett2023b, bennett2023c, bennett2023d, bennett2024b, bennett2024c].
+Many of the definitions have been adapted from a variety of preceding work [31, 33, 34, 50, 36, 38].
 They are referred to in the body of the paper when they become relevant, in the order in which they appear here.
@@
 (optimal proxy) There is no proxy more sample efficient than weakness.
-The weakness proxy formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [bennett2023b]).
+The weakness proxy formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [33]).
 (intuitive summary) Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy.
@@
 The lower level the child from which an agent generalises to parent, the “faster” it learns, the more sample efficient the proxy.
-The most sample efficient proxy is weakness [bennett2023b, prop.
+The most sample efficient proxy is weakness [33, prop.
 1, 2], which is why we’re using it here.
@@
 If $obs\in L_{\mathfrak{v}}$ is sensorimotor activity we interpret as an “observed event”, and $int\in L_{\mathfrak{v}}$ is an intervention to cause that event, then $obs\subset int$ (because $int$ could not be said to cause $obs$ unless $obs\subset int$).
-(intuitive summary) An intervention is action undertake an organism or other agency, in the sense described by Pearl [pearl2018].
+(intuitive summary) An intervention is action undertake an organism or other agency, in the sense described by Pearl [56].
 Intuitively, if “$int$” and “$obs$” are events which have happened, then we say that $int$ has caused $obs$ if $obs$ would not have happened in the absence of $int$ (counterfactual).
@@
     3RD and higher order selves: meta self-reflexive consciousness (human level hard consciousness) is achieved when $\mathfrak{o^{3}}\in\mathfrak{p}_{\mathfrak{o}}$.
-A.1 Proofs
-Versions of or alternatives to proofs 1-3 were published in [bennett2023b].
-Further details on these updated versions are available in on GitHub, along with the supporting experiments [bennett2024appendices].
-Theorem 1 (sufficiency).
+10.1 Proofs
+Versions of proofs 1-3 were published in proceedings [33]. Further details are available in on GitHub, along with the supporting experiments [35].
+Proposition 1 (sufficiency).
 Assume $\alpha\sqsubset\omega$.
@@
 $\square$
-Theorem 2 (necessity).
+Proposition 2 (necessity).
 To maximise the probability of learning $\omega$ from $\alpha$, it is necessary to use weakness as a proxy.
@@
 To select the weakest hypothesis, it is necessary to use the weakness proxy. $\square$
-Theorem 3 (simplicity sub-optimality).
+Proposition 3 (simplicity sub-optimality).
 Description length is neither a necessary nor sufficient proxy for the purposes of maximising the probability that induction generalises.
@@
 The following sketch is more for intuition and explanation than anything formal.
-Theorem 4.
+Proposition 4.
 An organism that uses weakness as its proxy will learn an $n^{th}$ order self if the incentive and scale preconditions are met for that order of self.
@@
   3. 3.
-    From [bennett2023b, prop.
+    From [33, prop.
 3] we have proof that weakness is the optimal choice of proxy to maximise the probability of generalisation from child to parent is the weakest policy.
@@
 $\square$
-[…]

### v5

@@
 The hard problem has sparked a substantial body of work, detailed discussion of which [14, 18] lies beyond the scope of our paper.
-In what follows we build upon the distinction between lower and higher order theories of information processing in relation to conscious experiences.
-Higher order thought theory (HOT) [19, 20] holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states.
+In what follows we start with the observation that the mental/physical distinction that fuels the classical debates, ranging from panpsychism [19] to physicalism à la Dennett [20] may not be the only way to tackling the hard problem of consciousness.
+We remain agnostic to the question whether conceivability arguments such as the zombie thought experiment [21] or the knowledge argument [22, 23] are the best way to approach the problem of phenomenal consciousness.
+For example, one may argue that artificial systems such as Large Language Models (LLMs) conceive entities that humans never conceived or never will.
+It remains an open question on whether these entities exist just on the basis that a system can conceive them abstractedly or not.
+Our aim is more modest and starts with low level biological facts rather than high level philosophical abstractions to reformulate the problem of consciousness in a way that does not take for granted the mental/physical division in the first place.
+In this paper we suggest a way to dissolve the “hard problem” of consciousness is to reverse the order and start with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states.
+Take for example the higher order thought theory (HOT) [24, 25] holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states.
 Lower order states may include emotions and perceptions, while higher order meta-representations reflect upon those.
@@
 Is there something it is like to “be in the dark” processing information at the lower bodily levels?
-In this paper we suggest a way to dissolve the “hard problem” of consciousness is to reverse the order and start with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states.
 Longtime considered a fringe approach, the embodied cognition paradigm [26] has recently gained substantial influence in cognitive science and philosophy [27, 28, 29].
@@
 Some have argued enactivism is incompatible with computation because the set of possible abstractions that must be searched for relevance realisation is intractable [51].
-For tractability, one must first isolate a “small world” of relevant information from the “big world” of all information.
+For tractability, one must first isolate a “small world” of relevant information from the “big world” of all information [52].
 Part of the reason we can address this is because we formalise embodiment, in concrete terms from the level of irreducible physical states.
@@
 Behaviour implies policies, and every policy implies tasks, so we can define preferences as an binary relation over tasks.
-Natural selection favours adaptability, hence it prefers organisms that optimise for weak policies, which we call weak policy optimisation (WPO).
+Natural selection favours adaptability, hence it prefers organisms that optimise for weak policies, which we call weak policy optimisation (w-maxing).
 By constructing fit policies, an organism divides the world into relevant objects and properties.
@@
 That language is dispositional [67], shaped by natural selection, and so a weaker policy is likely to be a fit policy.
-WPO will correctly identify cause and effect where relevant, simplifying the entire environment into objects and properties relevant to the organism’s motivations.
+w-maxing will correctly identify cause and effect where relevant, simplifying the entire environment into objects and properties relevant to the organism’s motivations.
 Assume a vocabulary $\mathfrak{v}_{\mathfrak{b}}$ belonging to an organism $\mathfrak{b}$ (Bob).
@@
 5 Multi-Layered Self-Organisation
-As the vocabulary and capacity for WPO scale, a greater variety of concepts can be learned.
+As the vocabulary and capacity for w-maxing scale, a greater variety of concepts can be learned.
 Progressively higher orders of ‘causal-identity’ for one’s self related information processing are constructed [39].
@@
 5.1 The First Order Self
-A first order self (1ST henceforth, formalised in definition 13) allows an organism discern the consequences of its actions.
+A first order self (1ST henceforth, formalised in definition 13) allows an organism to discern the consequences of its actions.
 This serves as the locus of self related information processing and experience [39], allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction.
@@
 Up to now we have developed the conceptual toolbox that we use to dissolve the hard problem.
-We started with the observation that human body self organises to maintain itself in the face of change both within and without.
+We started with the observation that a human body self organises to maintain itself in the face of change both within and without.
 An organism does not spring into existence understanding number systems or objects, but constructs them through interaction with the environment.
@@
 Hence, a policy is a classifier of information, but that information is not in a language, and it is not yet something labelled or quantified (until there is a policy that labels or quantifies).
+For biological, living self-organising systems, being alive is intrinsically good, being dead or ill is bad.
+For example, in developing self-organising systems such as the human body, it is good to have a better developed sense of smell than vision at early stages of life (babies are more accurate smellers than adults).
+Different senses have different degrees of valences at different stages of the developmental ladder.
+Hence there is “something what it is like” to be a baby which is different from something what it is like to be an adult.
+However, the idea here is that basic experiences (i.e.
+the self-organising process of staying alive) have valence at their very core (good=stay alive; bad=dead).
+This culminates in experiential, phenomenal ‘quality’ of appearance, of what it is likeness and how thing appear to the system (i.e.
+phenomena).
+That is the aspect that philosophers typically tackle.
+Our aim is not to dismiss the latter; our claim is that one cannot have the latter without the former.
 Clearly, prelinguistic self-organising systems must classify and attach value and disvalue to states and anticipated states to prioritise and make decisions.
 To learn is to sense “something” and, motivated by valence, construct a policy classifying that “something”.
-Physical states are contentless.
-In the absence of language, the difference between any two states can only be qualitative.
-In other words we hold that quality is fundamentally valenced, i.e.
-the product of actively sensing and being physically attracted and repelled within a given environment.
+When we say physical states are contentless, we are not suggesting physical objects and properties don’t exist.
+Objectively, a physical object is an aspect of the environment and every aspect is realized by one state or another.
+However, previous work has shown that for every aspect of the environment there is an equivalent program, meaning from an objective point of view what is or is not a physical object or property is a matter of interpretation [41].
+Subjectively an object, property or quality only exists if an organism constructs a causal identity for it.
+At the most basic level, the foundation of all of this is physical attraction and repulsion.
+An organism hard-wired by natural selection can be physically attracted or repelled without any concept of what it is physically attracted or repelled from.
+It doesn’t know to what it is attracted or repelled from in the sense that it has not constructed a causal identity for that object.
+There is no named object which has the property ‘attractive’.
+The organism itself is simply attracted or repelled according to hard-wired policy.
+For example, a single-celled organism might be attracted to glucose, and its response to ‘tumble’ or ‘stop’ would be hard-wired.
+It does not satisfy the incentive or scale preconditions to construct a causal identity for glucose, or ‘sweet’, or hunger, or anything else.
+Intuitively, one might think of this like ‘one dimensional’ valence.
+To the bacteria, there is no subjective categorical variable or object to which it is attracted, though in the objective sense such a thing does exist.
+It is simply attracted or repelled.
+Over the course of this paper we have sought to explain how we get from simple ‘one dimensional’ hard-wired valence to a property or quality that has valence.
+Categorical variables for properties or ‘qualities’ like ‘the colour red’ or ‘the smell of coffee’ only exist when we have a causal identity for them, which requires scale and incentive.
+It can be hard-wired, or learned.
+When we have only one cell we do not have a vocabulary capable of satisfying the scale precondition.
+However self-organising biological systems are collective [78].
+Control is delegated and distributed among smaller components [43] which operate concurrently.
+An individual cell might only have ‘one dimensional’ valence, but as soon as we have two cells we have something analogous to a second ‘dimension’.
+Cellular collectives are polycomputational systems, meaning one cell can play a part in more than one computation simultanously, at different scales [79] [^15] This is what the weakness of a policy captures; the number of interactions of which it can be a part..
+As we scale up the collective, its is impelled by the cells of which it is formed.
+Instead of ‘one dimensional’ valence, we now have a rich tapestry of competing drives.
+It is this rich tapestry that is a property or ‘quality’ of a state of the environment interpreted by the organism.
+It is this rich tapestry that can be reduced to causal identities for properties like ‘the color red’, ‘hunger’ or ‘thirst’.
+A higher level of abstraction is formed from this rich tapestry of valence and how it changes over many interactions (the different inputs and outputs in a task that imply generalizable causal identities).
+Hunger and thirst might have the same overall intensity and so one could say they have the same overall valence, but they are qualitatively different because they are different ‘tapestries’ of valence at a lower level of abstraction.
+This is what is meant by The Psychophysical Principle of Causality in section 3.
+In a self-organising system of sufficient scale with sufficient incentive, what starts out as simple ‘one dimensional’ valence culminates in categorical variables that serve different organismic needs.
+This is why the sound of middle C is different from the color blue.
+They have different causal identities.
+Middle C results from sound waves hitting the ear, blue results from light waves hitting the eye.
+These variations produce different qualia.
+Qualia do not need to be perfect.
+They simply need to be “good enough” to motivate actions necessary for survival in the environment of the organism.
+Qualia are produced by the sensory and nervous systems of the body.
+The valence of red light waves misses the eyes of dogs entirely.
+Dogs rely more on smell than sight.
+This suits their environmental niche.
+At a sufficient scale causal identities can be constructed, stored and recalled by their quality and valence.
 The bold claim here is that information processing (i.e.
@@
 Hence a 1ST order self must precede higher orders.
-Our arguments align with those of Merker, Barron and Klein, who have linked subjective experience to reafference [72, 65, 66].
+Our arguments align with those of Merker, who has linked subjective experience to reafference [80, 71, 72].
 We agree that reafference is key, but we provide a very different explanation of why and how.
@@
 To illustrate how our argument applies in the real world we describe stages of conscious organism.
-Each stage follows from scaling up supply and demand for WPO, through natural selection.
+Each stage follows from scaling up supply and demand for w-maxing, through natural selection.
 For each stage we point out animals which are likely to be at least so conscious:
@@
   • We say $x$ and $y$ are equivalent iff $E_{x}=E_{y}$.
-(notation) $E$ with a subscript is the extension of the subscript181818e.g. $E_{l}$ is the extension of $l$..
+(notation) $E$ with a subscript is the extension of the subscript191919e.g. $E_{l}$ is the extension of $l$..
 (intuitive summary) $L_{\mathfrak{v}}$ is everything which can be realised in this abstraction layer.
@@
 The lower level the child from which an agent generalises to parent, the “faster” it learns, the more sample efficient the proxy.
-The most sample efficient proxy is weakness [33, prop.
+The most sample efficient proxy is weakness [38, prop.
 1, 2], which is why we’re using it here.
@@
 10.1 Proofs
-Versions of proofs 1-3 were published in proceedings [33]. Further details are available in on GitHub, along with the supporting experiments [35].
+Versions of proofs 1-3 were published in [38]. Further details are available in on GitHub, along with the supporting experiments [40].
 Proposition 1 (sufficiency).
@@
   3. 3.
-    From [33, prop.
-3] we have proof that weakness is the optimal choice of proxy to maximise the probability of generalisation from child to parent is the weakest policy.
+    From [38] we have proof that weakness is the optimal choice of proxy to maximise the probability of generalisation from child to parent is the weakest policy.
 It follows that $\mathfrak{a}$ will generalise from $\mathfrak{h}_{<t_{\mathfrak{a}}}$ to $\mu_{\mathfrak{a}}$ given the smallest history of interventions with which it is possible to do so (meaning the smallest possible ostensive definition, or cardinality $\lvert D_{\alpha}\rvert$).

### v6

@@
-
+Why is anything conscious?
 Abstract
-We tackle the hard problem of consciousness taking the naturally selected, embodied organism as our starting point.
-We provide a formalism describing how biological systems self-organise to hierarchically interpret unlabelled sensory information according to valence.
-Such interpretations imply behavioural policies which are differentiated from each other only by the qualitative aspect of information processing.
-Natural selection favours systems that intervene in the world to achieve homeostatic and reproductive goals.
-Quality is a property arising in such systems to link cause to affect to motivate interventions.
-This produces interoceptive and exteroceptive classifiers and determines priorities.
-In formalising the seminal distinction between access and phenomenal consciousness, we claim that access consciousness at the human level requires the ability to hierarchically model i) the self, ii) the world/others and iii) the self as modelled by others, and that this requires phenomenal consciousness.
-Phenomenal without access consciousness is likely common, but the reverse is implausible.
-To put it provocatively: death grounds meaning, and Nature does not like zombies.
-We then describe the multilayered architecture of self-organisation from rocks to Einstein, illustrating how our argument applies.
+We tackle the problem of consciousness by taking the naturally selected, embodied organism as our starting point.
+We provide a formalism describing how biological systems such as human bodies self-organize to hierarchically interpret unlabelled sensory information according to valence.
+The system is attracted and repelled at different spatial and temporal scales.
+This is a qualitative interpretation of an unlabelled physical state.
+We show how such interpretations imply behavioural policies which are differentiated from each other only by this qualitative aspect of information processing.
+Natural selection favours systems that actively intervene in the world to achieve homeostatic and reproductive goals.
+Put provocatively, death grounds meaning.
+This means that in living systems information processing is necessarily subjective, that is, it has quality embedded into its very core.
+Qualitative information processing involves interoceptive and exteroceptive classifiers, and determines priorities for self-survival.
+We formulate The Psychophysical Principle of Causality as a theorem, and prove generalisation optimal learning forces this valence first ontology.
+Qualitative good or bad processing necessarily comes before quality neutral representations of properties (i.e.
+“red” is constructed from valence).
+Under selection pressures like sophisticated predation this produces a hierarchy of selves, of which reafference and reflective self awareness are a consequence.
+We discuss this in light of the seminal distinction between phenomenal and access consciousness.
+We claim that phenomenal consciousness without access is likely common, but the reverse is implausible.
 Our proposal lays the foundation of a formal science of consciousness, closer to human fact than zombie fiction.
-keywords:
-hard problem of consciousness
 1 Introduction
 Why is anything conscious?
-Both biological and other physical [^1] Physical here just means non-biological.
-We are not suggesting biological systems are non-physical.
-systems process information, yet humans consciously experience as well as process information.
+Both biological and other physical systems process information, yet humans consciously experience as well as process the information.
 Why?
@@
 occupy space).
-Yet, unlike physical systems, biological systems are dissipative systems using energy to self-organise in the face of entropic decay and environmental perturbation [1, 2].
-Initially formally introduced in the field of cybernetics [3, 4] the notion of self-organisation has been applied to various disciplines including physics [5], biology [6, 7] and neuroscience [8, 9, 10].
-Self-organisation is typically defined as the spontaneous emergence of spatiotemporal order or pattern-formation processes in physical and biological systems resulting from interactions of its components with the environment [11, 12, 13].
+Yet, unlike physical systems, biological systems are dissipative systems using energy to self-organise in the face of entropic decay and environmental perturbation Nagel (1974); Chalmers (1995).
+Initially formally introduced in the field of cybernetics Ashby (1947); von Foerster (1960) the notion of self-organisation has been applied to various disciplines including physics Haken (1983), biology Camazine (2003); Bell and Deater-Deckard (2007) and neuroscience Kelso (1995); Friston (2010); Tognoli and Kelso (2014).
+Self-organisation is typically defined as the spontaneous emergence of spatiotemporal order or pattern-formation processes in physical and biological systems resulting from interactions of its components with the environment Camazine et al.
+(2001); Seeley (2002); Rosas et al.
+(2018).
 Interestingly, much self and world related information processing goes on behind the scenes, or “in the dark” so to speak, that is, without being constantly present to our conscious minds.
 But why doesn’t all information processing go in the dark?
-This question is the subject of long-standing debate [14].
-One highly influential view is that consciousness has two aspects [14].
-The first is functional, by which we mean the ability to access and communicate information [15].
-How function relates to consciousness is considered the “easy problem” of consciousness [2].
-The second aspect is “what it is like” to consciously experience information processing, or phenomenal consciousness [2, 15, 1, 16, 17].
+This question is the subject of long-standing debate Seth and Bayne (2022).
+One highly influential view is that consciousness has two aspects Seth and Bayne (2022).
+The first is functional, by which we mean the ability to access and communicate information Block (1995).
+How function relates to consciousness is considered the “easy problem” of consciousness Chalmers (1995).
+The second aspect is “what it is like” to consciously experience information processing, or phenomenal consciousness Chalmers (1995); Block (1995); Nagel (1974); Gallagher and Zahavi (2021); Fuchs (2017).
 This doesn’t mean just global states like being awake, but more local states like smelling a cup of coffee.
-These local contents or “qualia” are characterised by what it is like to be in them [14].
+These local contents or “qualia” are characterised by what it is like to be in them Seth and Bayne (2022).
 It is unclear the extent to which functional and phenomenal aspects are independent.
-David Chalmers has influentially suggested that it may be possible to construct a “philosophical zombie” which acts in every way like a person but has no qualia [15].
+David Chalmers has influentially suggested that it may be possible to construct a “philosophical zombie” which acts in every way like a person but has no qualia Block (1995).
 For example, a thermostat certainly detects heat and so processes information, but there presumably is not anything it is like to be a thermostat.
 Hence the question “why is anything conscious” may be understood as “why is there sometimes a qualitative aspect to information processing?”.
-This is the “hard problem” of consciousness.
-The hard problem has sparked a substantial body of work, detailed discussion of which [14, 18] lies beyond the scope of our paper.
-In what follows we start with the observation that the mental/physical distinction that fuels the classical debates, ranging from panpsychism [19] to physicalism à la Dennett [20] may not be the only way to tackling the hard problem of consciousness.
-We remain agnostic to the question whether conceivability arguments such as the zombie thought experiment [21] or the knowledge argument [22, 23] are the best way to approach the problem of phenomenal consciousness.
-For example, one may argue that artificial systems such as Large Language Models (LLMs) conceive entities that humans never conceived or never will.
-It remains an open question on whether these entities exist just on the basis that a system can conceive them abstractedly or not.
-Our aim is more modest and starts with low level biological facts rather than high level philosophical abstractions to reformulate the problem of consciousness in a way that does not take for granted the mental/physical division in the first place.
-In this paper we suggest a way to dissolve the “hard problem” of consciousness is to reverse the order and start with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states.
-Take for example the higher order thought theory (HOT) [24, 25] holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states.
+This is somewhat related to the “hard problem” of consciousness.
+However the hard problem has sparked a substantial body of work, detailed discussion of which Seth and Bayne (2022); Northoff (2014) lies beyond the scope of our paper.
+In what follows we largely avoid the mental/physical distinction that fuels the classical debates, ranging from panpsychism Goff (2019) to physicalism à la Dennett Dennett (1991).
+Likewise we remain agnostic to the question whether conceivability arguments such as the zombie thought experiment Chalmers (1996) or the knowledge argument Jackson (1982, 1986) are the best way to approach the problem of phenomenal consciousness.
+How we do so can be understood in terms of the hardware and software metaphor.
+The distinction between mental and physical that fuels classical debates can be understood in computational terms.
+In this view mental activity is like software, which interacts with the physical world through hardware Hutter (2010); Bennett (2025c); Orseau and Ring (2012).
+Hardware acts as an interpreter between software and the physical environment.
+This is somewhat like the Pineal gland in Cartesian dualism, and so this view is called “computational dualism” Bennett (2024b).
+While software is a useful frame for building standardised products, formalisms of intelligence based on computational dualism are inherently flawed Bennett (2024b); Leike and Hutter (2015).
+The problem is that software is not in fact a separate substance.
+In actuality, software is a state of the hardware on which it runs.
+To explain, computers can be understood as “stacks” of abstraction layers, where each layer is simplified subsets of behaviour of the layer below.
+For example Python code relies on interpretation by a program written in C, which is built on Assembly, which in turn relies on machine code and thus hardware.
+In this sense, software is an “abstraction layer” over the hardware on which it runs.
+Nature privileges adaptation, and working only at high levels of abstraction can limit adaptability, decreasing efficiency Bennett (2025a).
+Compare the vast quantity of data and energy required to train a large language model to solve a problem, to the small quantities humans need to solve a problem.
+Biological systems are more efficient, because they delegate adaptation to lower levels of abstraction Bennett (2024a).
+For example, an organism can be understood as an abstraction layer over organs, which are a layer over cells.
+Adaptation takes place at every level of this biological “stack”, whereas a software “intelligence” in a computer is adaptive only at a a high level of abstraction, in software.
+Our ability to model and understand problems is limited by the level of abstraction at which we view them Bennett (2025a).
+Efficiency problems in Artificial Intelligence (AI) can be addressed by accounting for abstraction layers Bennett (2024b).
+Mental activity can be framed as a layer over the body, so rather than trying to explain the mind in the abstract like software, we start at the level of the embodied organism.
+In short, our explanation of consciousness asserts that world has valence.
+Physical states are attractive or repulsive to a self-organising adaptive system.
+Organisms are such self-organising systems which learn to classify causes of valence that affect their decision procedures, to have a better chance of surviving.
+Qualia and affects, the atomic building blocks of phenomenal consciousness, are qualitative interpretations rather than abstract logical or mathematical representations.
+Qualitative processing conveys an advantage in evolutionary terms.
+At high levels of abstraction these appear to be irreducible qualities to which valence seems attached.
+However at lower levels of abstraction these qualities are the various states of attraction and repulsion happening in a body at a given point in time.
+Put another way, sensory stimuli are attractive or repulsive at multiple spatial and temporal scales within a self-organising system Bennett (2025a).
+Qualia and affect are easily and obviously linked to affordances in the world by organisms.
+Phenomenology has a ‘logic’ – not the symbols and rules of the supersets of first order logic that make up programming languages but a presymbolic ‘logic’ of cause and effect.
+These are modelled in the decision procedures of prelinguistic organisms as cause and affect.
+At a high level, organic decision procedures boil down to affect-infused qualia, the key components of phenomenology.
+Natural selection favours organisms with phenomenology.
+Not always but often.
+Our position is that phenomenology is functional and decisive, not epiphenomenal or illusory, but this is not to deny that conscious decisions have subconscious components.
+We see consciousness as the “executive” top level of a “stack” of biological components.
+We bypass the “computational dualism” of hardware and software Bennett (2024b) by starting with low level biological facts rather than high level philosophical abstractions to reformulate the problem of consciousness in a way that does not take for granted the mental/physical division in the first place.
+Take for example the higher order thought theory (HOT) Rosenthal (2005); Brown et al.
+(2019) holds that the information of which a conscious being is aware are higher order “meta-representations” of lower order “local” mental states.
 Lower order states may include emotions and perceptions, while higher order meta-representations reflect upon those.
@@
 Sense data is processed by the body resulting in lower order mental states, and then meta representations of those is where we might find more abstract conceptual or thought-like contents of consciousness.
-HOTs are a good starting point to understand why most biological information processing goes on “in the dark”, but why do these lower order states arise?
+HOTs might be a good starting point to understand why most biological information processing goes on “in the dark”, but why do these lower order states arise?
 Can they occur in the absence of subjective, qualitative experience?
 Is there something it is like to “be in the dark” processing information at the lower bodily levels?
-Longtime considered a fringe approach, the embodied cognition paradigm [26] has recently gained substantial influence in cognitive science and philosophy [27, 28, 29].
+Long considered a fringe approach, the embodied cognition paradigm Varela et al.
+(2016) has recently gained substantial influence in cognitive science and philosophy Ciaunica and others (2023); McMillen and Levin (2024); Seth and Tsakiris (2018).
 The idea is that instead of considering the body as a mere device designed to fuel and contain the mind (a device that can be replaced with a vat or a robot, for example), one must consider the mind as serving the self-sustaining needs of a surviving body.
-If this is so then understanding consciousness must start with understanding the ‘humble’ lower bodily levels of information processing, not the higher order levels of information processing only.
+If this is so then unravelling consciousness must start with understanding the ‘humble’ lower bodily levels of information processing, not the higher order, cognitive levels of information processing.
 In other words, conscious experiences do not merely depend on bodily experiences as an external factor that can be replaced with a vat or an artificial system.
-Our approach is to try to rigorously define consciousness from first principles and show that some aspects of functional consciousness depends on phenomenal consciousness in a manner that makes zombies impossible, “dissolving” the hard problem by showing the phenomenal to be intrinsically functional [^2] Note that we are far from the first to claim that the phenomenal is functional.
-For example, the proposed Conscious Turing Machine [30] based on Global Workspace Theory [31], and constructivist approaches to artificial intelligence [32] take similar position.
+Rather they fundamentally constitute what a conscious experience is.
+How can we prove that?
+Our aim here is to try to rigorously define consciousness from first principles and show that some aspects of functional consciousness depends on phenomenal consciousness in a manner that makes zombies impossible, “dissolving” the problem by showing the phenomenal to be intrinsically functional [^1] Note that we aren’t the first to claim that the phenomenal is functional.
+For example, the Conscious Turing Machine Blum and Blum (2020) based on Global Workspace Theory Baars (1997), and constructivist approaches to AI Wang (2020) take similar positions.
 It is our explanation of how and why that is novel..
-We establish axioms that hold in every conceivable environment and show that it is impossible to have the function of consciousness without the subjective experience of it, and why most information processing goes on “in the dark”.
-Enactivism is roughly the view that information processing arises through interaction between an organism and its environment [33], and it is considered by some to be incompatible with narrow definitions of computation.
-The notion of computation is widely debated, and a detailed review of these discussions would lead us to a major digression [34, 35].
-Here we define computation not in terms of abstract symbol shuffling and representation, but in concrete mechanistic terms where ‘information processing’ refers to the actual physical transition of a system from one state to another, so it is compatible with enactivism.
-Pancomputationalism is the idea that all systems are computational.
-We build upon a formalism called pancomputational enactivism [36].
-To answer “why is anything conscious”, we use this formalism to show how lower and higher order theories, phenomenal and access consciousness all follow from first principles, scaling natural selection pressures and the ability to adapt [^3] This is loosely inspired by a scale-based framing of machine learning [37] - we are applying Sutton’s “bitter lesson” to biology..
-First, in Section 2 we integrate and extend separately published works of narrower scope [36, 38, 39] to justify our model of self-organising systems.
+We establish axioms that hold in every conceivable environment.
+We then argue that if those axioms hold there is no conceivable environment where the function of consciousness without the subjective experience of it, and why most information processing goes on “in the dark” [^2] That is, without explicit conscious awareness of it..
+Box 1: Supporting formal results and provenance This paper is primarily philosophical.
+However, the key claims are also supported by formal theorems, proofs and experiments provided in the Supplementary Information (ESM).
+Many of these experiments and proofs were peer reviewed and published in the PhD thesis “How To Build Conscious Machines” Bennett (2025a), and in the proceedings of the Spring 2026 AAAI Workshop on Machine Consciousness Bennett (2026b).
+The main manuscript also introduces new formal results.
+These include the proof of The Psychophysical Principle of Causality and substantially improved propositions showing how the 1ST, 2ND and 3RD order selves follow from generalisation optimal learning under the scale and incentive preconditions.
+Full proofs are included in the ESM.
+The reformulation of informal The Psychophysical Principle of Causality as a formal theorem, and the resulting proof, are dependent upon the already published theorem The Law of the Stack Bennett (2024a), so The Law of the Stack is also detailed in the ESM.
+Boxes 2 to 4 summarise the new theorem, the other proof results, and the experiments.
+Enactivism is roughly the view that information processing arises through interaction between an organism and its environment Thompson (2007), and it is considered by some to be incompatible with narrow definitions of computation.
+The notion of computation is widely debated, and a detailed review of these discussions would lead us to a major digression Piccinini (2015); Piccinini and Maley (2021).
+Here we define computation not in terms of abstract symbol shuffling and representation, but in concrete mechanistic terms where ‘information processing’ refers to the actual physical transition of a system from one state to another.
+To formalise enactivism [^3] By formalising something akin to “intelligence without representation” as described by Brooks Brooks (1991), we end up with physical interpretations rather than representations., we build upon Pancomputational Enactivism Bennett (2024b, a) and Stack Theory Bennett (2025a), which frame intelligence in terms of embodied abstraction layers.
+To answer “why is anything conscious”, we use this formalism to show how lower and higher order theories, phenomenal and access consciousness all follow from first principles, scaling natural selection pressures and the ability to adapt [^4] This is loosely inspired by a scale-based framing of machine learning Sutton (2019) - we are applying Sutton’s “bitter lesson” to biology..
+First, in Section 2 we integrate and extend separately published works of narrower scope Bennett (2024b, 2023c, 2023a) to justify our model of self-organising systems.
 In Sections 3 and 4 we explain how this model formalises relevance realisation and unifies lower and higher order theories.
@@
 This takes us from attraction and repulsion to physical states, through lower order thoughts to higher order meta representations.
-We call this the psychophysical principle of causality.
-In Section 5 we extend previous work [39] on causal learning and the development of first (1ST), second (2ND) and third (3RD) order selves, and explain how they are a consequence of scaling the ability of a self-organising system to adapt with natural selection pressures.
+In the associated thesis Bennett (2025a) we call this the psychophysical principle of causality.
+Here we formalise this as a theorem of the same name (theorem 3 in the ESM).
+In Section 5 we extend previous work Bennett (2023a) on causal learning and the development of first (1ST), second (2ND) and third (3RD) order selves, and explain how they are a consequence of scaling the ability of a self-organising system to adapt with natural selection pressures.
 In Sections 6 and 7 we explain how subjective experience requires a 1ST order self, and conversely why a 1ST order self implies there is something it is like to be an organism that has a 1ST order self.
@@
 2 Back to Foundations
-To meet length restrictions, the full mathematical definitions have been placed in the appendices [40].
-The core arguments should be understandable without the math, but for convenience here is a quick reference guide:
+This paper can be read in purely narrative form, meaning the core argument should be understandable without the math.
+However for readers who do wish to delve into the mathematical details, a short summary of definitions is given below.
+The full set of definitions used in this paper, along with full proofs and the full experimental protocols and plots, is in the Electronic Supplementary Material (ESM).
+The longer Stack Theory appendix Bennett (2025a) is the master reference.
   1. 1.
-    A set $\Phi$ is assumed, whose elements are called states.
-A declarative program is set $f$ of states, and $f$ is true relative to $\phi\in\Phi$ iff $\phi\in f$.
-$P$ is the set of all programs.
-Embodiment is formalised as an abstraction layer, for which we single out finite set $\mathfrak{v}\in 2^{\Phi}$ called a vocabulary.
-A choice of vocabulary $\mathfrak{v}$ implies a formal language $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap l\neq\emptyset\}$.
+    An environment is a nonempty set $\Phi$ of mutually exclusive states.
+A (declarative) program is any set of states $p\subseteq\Phi$.
+Let $\mathcal{P}:=2^{\Phi}$ be the set of all programs.
+An abstraction layer is a finite vocabulary $\mathfrak{v}\subseteq\mathcal{P}$, meaning the finite set of programs a body can stably implement.
+A vocabulary induces an embodied formal language
+    $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap_{p\in l}p\neq\emptyset\}.$
   2. 2.
-    Members of $L_{\mathfrak{v}}$ are called statements, and a statement $l$ is considered true relative to a state $\phi$ iff $\phi\in\bigcup l$.
-A completion of a statement $l_{1}$ is any statement $l_{2}$ s.t.
-$l_{1}\subseteq l_{2}$.
-The extension of a statement $l$ is the set $E_{l}$ of all its completions, s.t.
-$x\in{L_{\mathfrak{v}}}$ is $E_{x}=\{y\in{L_{\mathfrak{v}}}:x\subseteq y\}$.
-The extension of a set $X\subseteq{L_{\mathfrak{v}}}$ of statements is $E_{X}=\bigcup\limits_{x\in X}E_{x}$.
+    Members of $L_{\mathfrak{v}}$ are called statements. The truth set of a statement $l$ is
+    $T(l):=\bigcap_{p\in l}p.$
+    We take $T(\emptyset)=\Phi$ by convention.
+The statement $l$ is true relative to a state $\phi\in\Phi$ iff $\phi\in T(l)$.
+A completion of $x$ is any statement $y$ such that $x\subseteq y$.
+The extension of a statement $x$ is the set of all its completions
+    $E_{x}:=\{y\in L_{\mathfrak{v}}:x\subseteq y\}.$
+    For a set $X\subseteq L_{\mathfrak{v}}$ we write $E_{X}:=\bigcup_{x\in X}E_{x}$. Two statements are equivalent when they have the same extension.
   3. 3.
-    A $\mathfrak{v}$-task $\alpha$ is a pair $\langle I_{\alpha},O_{\alpha}\rangle$ where $I_{\alpha}\subset L_{\mathfrak{v}}$ are the inputs, $E_{I_{\alpha}}$ is the outputs available given ${I_{\alpha}}$, and $O_{\alpha}\subset E_{I_{\alpha}}$ are the correct outputs.
-All the tasks an abstraction layer tasks form a hierarchy.
+    A $\mathfrak{v}$-task $\alpha$ is a pair $\langle I_{\alpha},O_{\alpha}\rangle$ where $I_{\alpha}\subseteq L_{\mathfrak{v}}$ are the inputs and $O_{\alpha}\subseteq E_{I_{\alpha}}$ are the correct outputs.
+Tasks form a hierarchy.
 If $\alpha\sqsubset\omega$ then $\omega$ is a parent of $\alpha$, meaning $I_{\alpha}\subset I_{\omega}$ and $O_{\alpha}\subseteq O_{\omega}$.
-A parent is higher level than its children.
   4. 4.
-    A policy is a statement $\pi$ and a policy is correct for $\alpha$ iff $\pi\in{\Pi_{\alpha}}=\{\pi\in L_{\mathfrak{v}}:{E}_{I_{\alpha}}\cap E_{\pi}={O_{\alpha}}\}$.
-Given input $i\in I_{\alpha}$ and a policy $\pi$, to infer is to choose $o\in E_{i}$.
-A task is complete if $o\in O_{\alpha}$.
-Given a task definition $\alpha$ s.t.
-$\alpha\sqsubset\omega$, to learn is to choose a policy $\pi\in\Pi_{\alpha}$, and $\omega$ has been learned from $\alpha$ if one chooses $\pi\in\Pi_{\omega}$.
+    A policy is a statement $\pi\in L_{\mathfrak{v}}$. A policy is correct for $\alpha$ iff
+    $\pi\in{\Pi_{\alpha}}=\{\pi\in L_{\mathfrak{v}}:{E}_{I_{\alpha}}\cap E_{\pi}={O_{\alpha}}\}.$
+    Given input $i\in I_{\alpha}$ and a policy $\pi$, to infer is to choose an output $o\in E_{i}\cap E_{\pi}$. A task is complete if $o\in O_{\alpha}$.
   5. 5.
-    A causal identity for $\mathfrak{b}$ is a policy classifying the causal interventions undertaken by $\mathfrak{b}$.
-Subscript denotes the system that learned the causal identity, while superscript denotes the object it classifies.
-For example, $c^{\mathfrak{b}}_{\mathfrak{a}}$ denotes a causal identity learned by $\mathfrak{a}$ to classify $\mathfrak{b}$’s interventions, and $c^{\mathfrak{ba}}_{\mathfrak{a}}\subset c^{\mathfrak{b}}_{\mathfrak{a}}$ denotes $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$.
-We begin by formalising all conceivable environments [36, 41] from first principles:
+    Learning must choose among many correct policies.
+A standard proxy is weakness $w(\pi):=|E_{\pi}|$, which counts how many further commitments remain compatible with $\pi$ while staying correct.
+Weakness maximisation (w-maxing) chooses a correct policy that maximises $w(\pi)$.
+A causal identity is a policy $c^{\mathfrak{b}}_{\mathfrak{a}}\in L_{\mathfrak{v}}$ that classifies interventions by some source $\mathfrak{b}$ as distinct from matched observations, relative to a learner $\mathfrak{a}$.
+Nested causal identities represent higher orders of modelling.
+For example $c^{\mathfrak{ba}}_{\mathfrak{a}}\subseteq c^{\mathfrak{b}}_{\mathfrak{a}}$ is a refinement of $\mathfrak{a}$’s model of $\mathfrak{b}$ that depends on $\mathfrak{a}$’s prediction of how $\mathfrak{b}$ models $\mathfrak{a}$.
+We begin by formalising all conceivable environments Bennett (2024b, c) from first principles:
 Axiom 1: When there are things, we call these things the environment.
@@
 Hence, the environment is a set $\Phi$ of mutually exclusive states.
-We call a set of states a declarative “program” [^4] We do not concern ourselves with imperative programs, because there is an equivalence between declarative and imperative [42]..
+We call a set of states a declarative “program” [^5] We do not concern ourselves with imperative programs, because there is an equivalence between declarative and imperative Howard (1980)..
 A program is “true” relative to the states it contains.
@@
 An aspect of the environment is realised by (exists in) a state if the programs it contains are facts.
-Previous work argues this formalises every conceivable environment [41].
-To give some intuition as to why, note that we assume nothing of what states might be or contain.
+Given each state is a difference, it follows that states are mutually exclusive, which elsewhere has been used to portray time as difference and each state as a point in time [^6] Each possible sequence of states might be thought of as a timeline, and the set of all timelines can be thought of as the set of all conceivable universes (infinite if the set of states is infinite)..
+Such foundational arguments are largely beyond the scope of this paper and are discussed at length elsewhere Bennett (2025a).
+However, this frame allows us to talk about “computation” in a much more general sense than Turing computation.
+By computation, we mean physical cause and effect, dictated by the state transitions of whatever environment we might conceive of.
+In the frame of Stack Theory Bennett (2025a, 2024a), this is like saying we don’t know where the base abstraction layer is, whether there even is such a layer, or what it might entail.
+We are merely saying that no matter what abstraction layer is considered, there must be change or difference for there to be anything at all.
+That entails there must be states, which entails there are ontological “facts” or “true programs” we can formalise as sets of states, every aspect of the environment that exists is a set of such facts, and finally this lets us talk about computation in a much more general sense than Turing computation [^7] Related work holds each timeline is a “cosmic ought” from which goals and value judgements follow Bennett (2025a).
+From a subjective perspective of a body (an aspect of the environment), “truth” is unknown (there can be many possible timelines in which a given body exists, seeming non-deterministic), but from the objective perspective of a given timeline truth is deterministic.
+Deterministic and non-deterministic views can be reconciled, just as the set of all languages recognisable as DFAs (deterministic finite automata) is the same as that recognisable as NFAs (non-deterministic) Rabin and Scott (1959).
+In this view, the universe has a sort of “opinion” in the sense that a given timeline preserves some aspects of the environment and destroys others as it progresses.
+The set of all possible timelines would be the set of all such “opinions”..
+We can talk about any given environment, or any given physical body within that environment, as an abstraction layer that performs a unique set of computations [^8] Put another way, no two bodies perform exactly the same set of computations..
+To give some more intuition as to why we can talk about any body, or any conceivable environment in this frame, note that we assume nothing of what states might ontologically be or epistemologically contain.
 We assume no differences within a state, only between states, and those differences are the programs of which aspects are formed.
@@
 However, the application of a function is a fact regardless of whether it takes another function as input., hence the body of an organism that persists across multiple states must be a set of programs.
-We call this set a vocabulary, and it is formalised in definition 2 (see appendix [40]).
-Piccinini [34, 35] divides computation into abstract and concrete sorts.
-Abstract computations mean or represent whatever we say they do.
-In contrast concrete computations are the actions of physical systems.
-The declarative programs we describe are concrete.
-Each captures a single point of difference between physical states.
-As such no two organisms have the same vocabulary because that would mean they are the same body.
-This might seem inconvenient, as intuitively we might prefer to think about bodies as an abstract ‘type’.
-However, doing so would undermine any claim we might make, as earlier results in artificial intelligence (AI) show [39, 43].
+We call this set a vocabulary (formally defined in the appendix Bennett (2025a) as well as the summary at the beginning of this chapter).
 The meaning of abstractions depends upon interpretation, much like the behaviour of software is determined by the abstraction layer that “interprets” it.
@@
 That hardware is an abstraction layer in which the software exists (including “higher level” abstraction layers like the Python interpreter).
-In other words, what we call software is nothing more than the state of hardware [36].
-This has undermined all but the most subjective of claims regarding the behaviour of theorised, software superintelligence [44, 45, 46].
+As stated earlier, what we call software is nothing more than the state of hardware Bennett (2024b).
+This has undermined all but the most subjective of claims regarding the behaviour of theorised, software superintelligence Hutter (2010); Leike and Hutter (2015, 2018).
 It is a flaw in the very idea of intelligent software.
-The distinction between software mind and hardware embodiment is subsequently called computational dualism [36], because it is reminiscent of how traditional Cartesianism conceived of a mental substance distinct from physical.
-If we are to dissolve the hard problem by starting with the ‘impure’ embodied biological organism instead of the ‘pure’ abstract mental states, then we cannot presuppose organisms use particular abstractions.
-Instead we must explain why particular abstractions are formed, in terms of contentless states, as the product of interaction between an organism and its environment [33, 26, 47].
-This process is sometimes called “relevance realisation” [48, 49, 50].
-Some have argued enactivism is incompatible with computation because the set of possible abstractions that must be searched for relevance realisation is intractable [51].
-For tractability, one must first isolate a “small world” of relevant information from the “big world” of all information [52].
-Part of the reason we can address this is because we formalise embodiment, in concrete terms from the level of irreducible physical states.
-Embodiment implies finite resource constraints, so we adopt one more axiom from [41].
-Axiom 3: All aspects of the environment are spatially extended [41].
-This means that a body is realised by only a finite number of states (see def.
-2).
-A body is a vocabulary $\mathfrak{v}\subset P$ with finitely many elements.
+The distinction between software mind and hardware embodiment is subsequently called computational dualism Bennett (2024b), because it is reminiscent of how traditional Cartesianism conceived of a mental substance distinct from physical.
+If we start with the embodied biological organism instead of the abstract mental states, then we cannot presuppose organisms use particular abstractions.
+Instead we must explain why particular abstractions are formed, in terms of contentless states, as the product of interaction between an organism and its environment Thompson (2007); Varela et al.
+(2016); Rolla and Figueiredo (2021).
+This process is sometimes called “relevance realisation” Vervaeke et al.
+(2012); Vervaeke and Ferraro (2013b, a), and this shift in perspective is key for our argument here.
+Indeed, the idea that conscious experiences depend upon interactions with the environment has been defended by a longstanding tradition in embodied cognition and enactivism Thompson (2007).
+For example, some have argued enactivism is incompatible with computation at a fundamental physical level, because the set of possible abstractions that must be searched for relevance realisation is intractable Jaeger et al.
+(2024).
+For tractability, one must first isolate a “small world” of relevant information from the “big world” of all information Savage (1954).
+To address this we formalise embodiment, in concrete terms from the level of irreducible physical states.
+Embodiment implies finite resource constraints Bennett (2022b, 2025a); Ciaunica (2025), hence we adopt one more axiom.
+Axiom 3: All aspects of the environment are spatially extended Bennett (2024c).
+This means that a body is realised by only a finite number of states.
+See the ESM for the self contained definition, and the Stack Theory appendix Bennett (2025a) for the full development.
+A body is a vocabulary $\mathfrak{v}\subset\mathcal{P}$ with finitely many elements.
 $\mathfrak{v}$ implies a formal language $L_{\mathfrak{v}}$ of interaction between body and environment.
@@
 Most importantly, we avoid the distinction between software and hardware and so avoid computational dualism.
-Embodiment is dispositional. As natural selection evolves a vocabulary, it begins to isolate a “small world” relevant to survival [51, 41, 53], from the intractable “big world” $P$ of all information.
+To summarise.
+Embodiment is dispositional Bennett (2024b, 2025b).
+As natural selection evolves a vocabulary, it begins to isolate a “small world” relevant to survival Jaeger et al.
+(2024); Bennett (2024c); Friston et al.
+(2023), from the intractable “big world” $\mathcal{P}$ of all information.
 2.2 Self-Organising Systems as Self and World Constraints
@@
 Hence an organism will self-organise to express some statements, but not others.
-Though environmental states are contentless and their only significance is in how they relate to one another, natural selection must make them attractive or repulsive to a self-organising system.
-This is formalised by the $\mathfrak{v}$-task as in definition 3777A $\mathfrak{v}$-task is called because a $\mathfrak{v}$-task because it exists in the context of a vocabulary $\mathfrak{v}$..
+Crucially, though environmental states are contentless they imply content by how they relate to one another.
+As states transition from one to another this preserves some aspects of the environment, and deletes others (which would happen when an aspect is made of facts that don’t hold at a particular moment in time, given the state at that time).
+As argued elsewhere, the environment preserves that which preserves itself Bennett (2025a).
+It follows that natural selection must make states attractive or repulsive to a self-organising system if that self-organising system is to be deemed ‘fit’.
+This is formalised by the $\mathfrak{v}$-task as in definition 3111111A $\mathfrak{v}$-task is called a $\mathfrak{v}$-task because it is defined relative to a vocabulary $\mathfrak{v}$..
 If $\mathfrak{v}$ is the vocabulary of a body, and $\mathfrak{v}$-task $\mu=\langle I_{\mu},O_{\mu}\rangle$ is fit behaviour, then $I_{\mu}$ is all the statements that body can express in which it is possible to remain fit, and $O_{\mu}$ is all the statements that body can express in which it remains fit.
@@
 That history is expressed as a $\mathfrak{v}$-task $\mathfrak{h}$, where each input and output in $\mathfrak{h}$ is an interaction between organism and environment.
-If the set of all fit behaviour is a task $\mu$, then the organism remains fit then $\mathfrak{h}\sqsubset\mu$.
-“I survive therefore my model is viable.” - Mark Solms [54]
+If the set of all fit behaviour is a task $\mu$, and the organism remains fit throughout its history, then $\mathfrak{h}\sqsubset\mu$.
+“I survive therefore my model is viable.” - Mark Solms Solms (2021)
 $\mathfrak{h}$ implies a set $\Pi_{\mathfrak{h}}$ of policies, and not all of those policies will constrain outputs the same way given new inputs.
@@
 An organism learns fit behaviour by learning $\pi\in\Pi_{\mathfrak{h}}$ such that $\pi\in\Pi_{\mu}$.
-“The best model of the world is the world itself” - Rodney Brooks [55]
-Earlier work [38] showed that the optimal strategy to adapt as fast as possible [^9] Meaning to converge on a fit policy given the smallest history possible, to realise fit behaviour.
+“The best model of the world is the world itself” - Rodney Brooks Dreyfus (2007)
+Earlier work Bennett (2023c) showed that the optimal strategy to adapt as fast as possible [^13] Meaning to converge on a fit policy given the smallest history possible, to realise fit behaviour.
 is to prefer “weaker” policies, meaning those with larger extensions.
 The policy which generated outputs $O$ given inputs $I$ is most efficiently identified by constructing a policy $\pi$ such that $\pi$ generates $O$ from $I$, and $\pi$ implies the weakest constraint that can be implied while still generating $O$ from $I$.
-This means a system that does so will construct fit policies from a shorter history [39] than one that does not (see def.
-5 and proofs in appendix [40]) [^10] The effect of insufficiently weak representations can be observed in large language models like GPT-4 [56, 57], where they fail to make consistent statements about an object because represent it as multiple objects..
+This means a system that does so will construct fit policies from a shorter history Bennett (2023a) than one that does not.
+Full proofs are in the ESM.
+[^14] The effect of insufficiently weak interpretations can be observed in large language models like GPT-3 and GPT-4 Floridi and Chiriatti (2020); Bennett (2023b), where they can fail to make consistent statements about an object because they represent it as multiple objects.
 3 Relevance Realisation Through Causal Learning
@@
 To efficiently adapt, an organism must correctly anticipate valence, which means correctly identifying what causes valence.
-All organisms (see definition 6) must have preferences, because they will make some decisions and not others.
-Behaviour implies policies, and every policy implies tasks, so we can define preferences as an binary relation over tasks.
-Natural selection favours adaptability, hence it prefers organisms that optimise for weak policies, which we call weak policy optimisation (w-maxing).
+All organisms must have preferences, meaning some futures are better for them than others.
+Formally, we model an organism as an embodied learner defined by an embodied vocabulary, a viability task, and a preference ordering over tasks.
+The full tuple definition is in the ESM.
+Behaviour implies policies, and every policy implies tasks, so preferences can be formalised as a binary relation over tasks.
+Natural selection favours adaptability, hence it prefers organisms that optimise for weak correct policies, which is referred to elsewhere as weakness maximisation (w-maxing) Bennett (2025c).
 By constructing fit policies, an organism divides the world into relevant objects and properties.
@@
 Every fit policy would be a “causal identity” for something relevant, for example a specific object like “this coffee” or a weaker and more general concept like “all coffee”.
-Organisms interact when they “affect” one another (see definition 9) in the causal or physical rather than psychological sense of the word.
-Now, Artificial Intelligence (AI) and machine learning [58, 59, 60, 61] are concerned with engineering adaptive agents.
+Organisms interact when they can affect one another in the causal or physical rather than psychological sense of the word.
+Here affect means one organism’s interventions can change the other organism’s inputs, hence changing which outputs are viable.
+The formal definition we use is in the ESM.
+Now, AI and machine learning Russell and Norvig (2020); Bishop (2006); Sutton and Barto (2018); Goertzel (2014) are concerned with engineering adaptive agents.
 In that context, causality has become a mainstream topic of research.
-Causal learning is demonstrably necessary to thrive in an interactive setting [62, 63].
-Where the causal graph is known in advance (for example if measuring the efficacy of medical interventions), this issue could be resolved using causal language [64] to represent an intervention from outside the system the graph describes.
-To illustrate [39], suppose Bob is attempting to learn and predict the environment.
+Causal learning is demonstrably necessary to thrive in an interactive setting Pearl and Mackenzie (2018); Richens and Everitt (2024).
+Where the causal graph is known in advance (for example if measuring the efficacy of medical interventions), this issue could be resolved using causal language Pearl (2009) to represent an intervention from outside the system the graph describes.
+To illustrate Bennett (2023a), suppose Bob is attempting to learn and predict the environment.
 Bob has observed Alice wearing a raincoat only when it rains, and Bob has seen rain only when he observed Alice wearing a raincoat.
@@
 This means Bob believes that if Alice wears a raincoat, then it must be raining.
-Now lets permit Bob to interact with its environment.
+Now let us permit Bob to interact with his environment.
 Assume Bob wants it to rain.
@@
 This is absurd.
-$O=true$ does not represent the event $q=$“I coerced Alice into wearing a raincoat”, but an entirely different event $v=$“Alice decided to put on a raincoat for the same reason I have observed Alice wearing a raincoat in the past”.
+$O=true$ does not represent the event q=q=“I coerced Alice into wearing a raincoat”, but an entirely different event v=v=“Alice decided to put on a raincoat for the same reason I have observed Alice wearing a raincoat in the past”.
 To accurately represent the environment, we need to represent that $p(R=true\mid q)=p(R=true)\neq p(R=true\mid v)=1$, meaning both $q$ and $v$.
-To illustrate visually, we started with the acyclic graph
-$O$$R$
-and our intervention disconnected rain from the choice of clothing:
-$O$$R$
-This can be resolved by introducing a “do” operator that we apply to a variable $O$ to obtain $do[O=o]$, to represent the fact that an agency from outside the system has intervened to assign a value to $O$, so that we can represent $p(R=true\mid do[O=true])=p(R=true)\neq p(R=true\mid O=true)=1$.
+To illustrate visually, we started with the acyclic graph OORR and our intervention disconnected rain from the choice of clothing: OORR This can be resolved by introducing a “do” operator that we apply to a variable $O$ to obtain $do[O=o]$, to represent the fact that an agency from outside the system has intervened to assign a value to $O$, so that we can represent $p(R=true\mid do[O=true])=p(R=true)\neq p(R=true\mid O=true)=1$.
 Thus, the aforementioned $q$ is equivalent to $do[O=true]$, while $v$ is equivalent to $O=true$.
-3.0.1 The Psychophysical Principle of Causality
+3.1 The Psychophysical Principle of Causality
 We have established the need for causal reasoning, but not how one might come to know the objects involved (where did the variables come from?), or how they relate to one another causally (to form a graph).
 There are two ways to learn causal graphs.
-One can either assume a set of variables and learn relations between them, or assume relations and learn the objects that fit them [39].
+One can either assume a set of variables and learn relations between them, or assume relations and learn the objects that fit them Bennett (2023a).
 We do the latter.
@@
 This is what will take us from attraction to and repulsion from contentless physical states, through lower order thoughts to higher order meta representations.
-We start with the complete absence of content [65], and once a system learns policies there are distinct “contents”.
+We start with the complete absence of content Thompson (2016), and once a system learns policies there are distinct “contents”.
 Policies serve to classify objects and properties.
 Death grounds meaning.
-To develop this idea, we first show adding additional nodes to a causal graph is an adequate substitute for the “do” operator [66].
+To develop this idea, we first show adding additional nodes to a causal graph is an adequate substitute for the “do” operator Dawid (2002).
 Assume we again have Bob who constructs a causal graph of the environment.
@@
 From Bob’s perspective, Alice is just a part of the environment represented by a variable $A$ in Bob’s causal graph.
-$O$$A$$R$
-Now assume Bob observes Alice taking an action that changes an aspect of the environment, represented by the variable $O$ (for example, Bob observes Alice putting on the raincoat).
+OOAARR Now assume Bob observes Alice taking an action that changes an aspect of the environment, represented by the variable $O$ (for example, Bob observes Alice putting on the raincoat).
 From Bob’s perspective, Alice’s action assigns values to variables $A$ and $O$.
 There is no need to involve a do operator in this scenario because we can already represent that $p(R=true\mid A=x,O=true)=p(R=true)\neq p(R=true\mid A=y,O=true)=1$ (because Alice is part of the causal graph).
-Likewise, we can add variables for interventions of Bob [39].
-$O$$B$$A$$R$
+Likewise, we can add variables for interventions of Bob Bennett (2023a).
+OOBBAARR
 Now we will show how it is possible to learn these nodes.
@@
 The optimal choice of policy for adaptability is the weakest because it is implied by more statements in an embodied language.
-That language is dispositional [67], shaped by natural selection, and so a weaker policy is likely to be a fit policy.
+That language is dispositional Bennett and Maruyama (2022), shaped by natural selection, and so a weaker policy is likely to be a fit policy.
 w-maxing will correctly identify cause and effect where relevant, simplifying the entire environment into objects and properties relevant to the organism’s motivations.
@@
 A “cause” in the context of this formalisation is not a variable but a statement $l\in L_{\mathfrak{v}_{\mathfrak{b}}}$.
-The raincoat example would involve $obs,rain\in L_{\mathfrak{v}_{\mathfrak{b}}}$ such that:
-[…]
-[…]
-[…]
-$int$$obs$$rain$
-So long as $obs\neq int$, the intervention can be differentiated from the passive observation (see definition 10).
+The raincoat example would involve $obs,rain\in L_{\mathfrak{v}_{\mathfrak{b}}}$ such that: obs ↔“Alice put on a raincoat” and rain ↔“It rained” obsobs and rainrain have truth values in accord with the definition.
+As in the earlier example Bob’s passive observation implies $p(rain\mid obs)=1$.
+The statement $obs=\textit{``Alice put on a raincoat''}$ can be made true by either passive observation or intervention.
+However, the statement which is true in the case of intervention not only obsobs, but $int\in L$ such that $obs\subseteq int$ and: int ↔“Alice is wearing a raincoat because of Bob’s actions” $\mathrm{int}$$\mathrm{obs}$$\mathrm{rain}$ So long as $obs\neq int$, the intervention can be differentiated from the passive observation at the more abstract representational scale of the vocabulary Bennett (2024a).
 This being the case, any set $c\subseteq int-obs$ could be used to identify the party undertaking the intervention, which is why $c$ is referred to as a “causal identity”.
-It distinguishes the intervention $int$ from the passively observed effect $obs$, like reafference in living organisms.
+It distinguishes the intervention intint from the passively observed effect obsobs, like reafference in living organisms.
 However, the above only considers one intervention.
 A weaker or more general causal identity would be one that is shared by more interventions.
+A formal proposition shows that, when intervention episodes share at least one distinguishing signature, a w-max learner selects an inclusion minimal causal identity.
+In plain English, it learns the least committal intervention tag that still separates intervention from matched observation.
+See Box 3 for the sketch and see the ESM propositions on causal identity minimality for the full proof.
+If a statement $l$ holds in both observation and intervention episodes, then $l$ alone cannot distinguish them.
+If $c\in\mathcal{C}(INT,OBS)$, then $l\cup c$ is true on every $int\in INT$ and false on every $obs\in OBS$.
+So $l\cup c$ plays the role of a do tag at the abstraction scale of $L_{\mathfrak{v}_{\mathfrak{a}}}$.
 For example, suppose the inputs Alice has been subject to are $I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}$.
 These can be divided into those in which Bob affected Alice $I^{\mathfrak{b}}_{\mathfrak{a}}$ and those in which Bob did not $I^{\lnot\mathfrak{b}}_{\mathfrak{a}}=I_{\mathfrak{h}_{<t_{\mathfrak{a}}}}-I^{\mathfrak{b}}_{\mathfrak{a}}$.
-Alice can construct a causal identity $b$ for Bob corresponding to interventions $INT=I^{\mathfrak{b}}_{\mathfrak{a}}$ and observations $OBS=I^{\lnot\mathfrak{b}}_{\mathfrak{a}}$ (see definition 11).
+Alice can construct a causal identity $b$ for Bob corresponding to interventions $INT=I^{\mathfrak{b}}_{\mathfrak{a}}$ and observations $OBS=I^{\lnot\mathfrak{b}}_{\mathfrak{a}}$.
+The full formalism for building these sets from a history is in the ESM.
 4.1 Ascribing Intent to Other Objects
@@
 The question is not “is this an intervention” but “by whom was this intervention made?”.
-Earlier, we arrived at the following graph in which Bob’s intervention was given by $int$.
-$int$$obs$$rain$
-What if a third person Larry puts the coat on Alice?
+Earlier, we arrived at the following graph in which Bob’s intervention was given by intint.
+$\mathrm{int}$$\mathrm{obs}$$\mathrm{rain}$ What if a third person Larry puts the coat on Alice?
 Bob may observe this, and so Bob’s observation of Larry’s intervention is $v\in L_{\mathfrak{v}_{\mathfrak{a}}}$ such that $obs\subset v$.
 To account for this, Bob can construct a causal graph as below (with $b.$ representing Bob and $l.$ representing Larry).
-$b.\ int$$l.\ int$$obs$$rain$
-Bob’s causal identity for himself $c_{b}\subset b.\ int-obs$ only represents the intervention by himself.
+$b.\,\mathrm{int}$$l.\,\mathrm{int}$$\mathrm{obs}$$\mathrm{rain}$ Bob’s causal identity for himself $c_{b}\subset b.\ int-obs$ only represents the intervention by himself.
 However, now we can see that Bob must also construct an identity $c_{l}$ for Larry, where the $c_{l}\subset l.\ int-obs$.
-For an organism $\mathfrak{a}$ with an embodied language $L_{\mathfrak{v}_{\mathfrak{a}}}$ to construct a causal identity for an object $\mathfrak{b}$, it must first be the case that $\mathfrak{a}$ is affected by $\mathfrak{b}$ [56], to satisfy the incentive precondition for causal identity.
+For an organism $\mathfrak{a}$ with an embodied language $L_{\mathfrak{v}_{\mathfrak{a}}}$ to construct a causal identity for an object $\mathfrak{b}$, it must first be the case that $\mathfrak{a}$ is affected by $\mathfrak{b}$ Bennett (2023b), to satisfy the incentive precondition for causal identity.
 Assume an organism $\mathfrak{a}$ is affected by $\mathfrak{b}$ given inputs $INT$, and not affected given inputs $OBS$.
@@
 This is not to say that $\mathfrak{b}$ has intent, but intent can be ascribed to $\mathfrak{b}$ because $\mathfrak{b}$ affects $\mathfrak{a}$, who can then discern when interventions are a consequence of $\mathfrak{b}$.
+A likelihood ratio bound shows that decoder mismatch can hide minds.
+If every mind-like hypothesis in an evaluator’s model fits an observed trace worse than a noise hypothesis by a large factor, Bayes rule forces the posterior away from mind (see Box 3 and ESM for technical details).
 4.2 Preconditions
@@
 Natural selection provides a notion of what ought to be, by eliminating anything which ought not.
+In a creature that can die, death grounds meaning Bennett (2025a).
   1. 1.
@@
 Each policy an organism $\mathfrak{o}$ learns implies $\mathfrak{v}$-tasks.
-A $\mathfrak{v}$-task is a triadic relation between inputs, outputs and policies which resembles Peircean semiosis [68, 56] of sign, referent and interpretant.
-We formalise this as a “protosymbol” [^11] Proto because it is something more primitive than a symbol as conceived of by Peirce.
-system $\mathfrak{s}_{\mathfrak{o}}$ for the organism $\mathfrak{o}$ (see definitions 6 and 7).
+A $\mathfrak{v}$-task is a triadic relation between inputs, outputs and policies.
+This resembles Peircean semiosis Atkin (2024); Bennett (2023b) of sign, referent and interpretant.
+We call the resulting family of reusable tasks a “protosymbol” [^15] Proto because it is more primitive than a Peircean symbol.
+system $\mathfrak{s}_{\mathfrak{o}}$ for the organism $\mathfrak{o}$.
+Intuitively, a protosymbol is a behavioural template that can be reused across contexts (full definition in ESM).
 Tasks exist in a “generational hierarchy”.
@@
 Related as they are in a lattice, protosymbols are analogous lower order states and higher order meta representations.
-Some have framed consciousness as a problem of moving from unary, to dyadic, to triadic relations [69].
+Some have framed consciousness as a problem of moving from unary, to dyadic, to triadic relations Goertzel (2006).
 Similarly, we have gone from unary states to dyadic programs, to triadic tasks and protosymbols.
+Box 2: Theorem 3.
+The Psychophysical Principle of Causality.
+One can learn causality by learning the objects and properties that fit a causal relation, or one can learn causal relations that fit a set of presupposed objects and properties.
+However if you want generalisation optimal learning, you cannot hard code a quality neutral ontology (no presupposed objects and properties).
+Only the good or bad signal can be a fixed representational primitive from which more abstract classifications are constructed as needed.
+This is because valence is the signal that determines how representations are judged, and so any fixed abstraction which is not a direct consequence of that signal only serves to obfuscate the signal and impose a floor on variational free energy Bennett (2024a).
+Continued existence and its lack determine what is attractive to whatever remains.
+Hence causal relation of valence is given.
+In a living system a causal identity is constructed of valence.
+It is a pattern in valence.
+A sufficiently complex mixture of repulsive and attractive signals, can together amount to a qualitatively neutral classification like “the colour red”.
+Attraction and repulsion is how an organism carves up its world into objects and properties.
+Objects and properties are made of valence, and so are qualitative rather than quantitative information.
+This is the informal version of The Psychophysical Principle of Causality Bennett (2025a).
+Here we reformulate this idea as an equivalent mathematical theorem, and prove (rather than just argue) that optimality demands valence first learning (at least, within the formal confines of the Stack Theory).
+We number this as Theorem 3 to match the numbering in the Supplementary Information.
+Theorem 3 is a corollary to The Law of the Stack (see ESM), but is itself a novel contribution of this paper.
+Definition (hard primitive commitments).
+Let $\alpha$ be a task and let $\Pi_{\alpha}$ be its set of correct policies.
+Let $C\subseteq L_{\mathfrak{v}}$ be a nonempty set of commitments.
+The restricted policy class that bakes in $C$ is $\Pi_{\alpha}^{C}:=\{\pi\in\Pi_{\alpha}:\forall c\in C,c\subseteq\pi\}.$ We call $C$ a set of hard primitives when we enforce this restriction regardless of the evidence.
+Setup.
+Let $\mu$ be a viability statement encoding the most basic valence partition.
+Viable versus not viable.
+For any policy $\pi$, define the viable continuation set $\Omega_{\pi}:=E_{\pi}\cap E_{\mu}.$ This is the set of viable future completions still compatible with $\pi$.
+Define the free energy proxy in bits $F_{2}(\pi):=\log_{2}|E_{\mu}|-\log_{2}|\Omega_{\pi}|.$ This equals $D_{\mathrm{KL},2}(q_{\pi}\|p_{\mu})$.
+Here $q_{\pi}$ is uniform on $\Omega_{\pi}$ and $p_{\mu}$ is uniform on $E_{\mu}$.
+Theorem 3 (The Psychophysical Principle of Causality).
+Assume $C\neq\emptyset$ contains at least one commitment that is not forced by viability.
+This means there exists $o\in E_{\mu}$ such that $c\not\subseteq o$ for some $c\in C$.
+Then any policy that is forced to include $C$ has strictly higher $F_{2}$ whenever $C$ excludes at least one viable continuation that would otherwise remain possible.
+So a generalisation optimal learner cannot keep such non valence primitives fixed.
+Proof sketch.
+Adding a commitment intersects the viable continuation set with a smaller extension.
+That can only shrink $\Omega_{\pi}$.
+Since $F_{2}(\pi)$ is monotone in $|\Omega_{\pi}|$, the free energy increases strictly if any viable continuation is removed.
+See ESM Theorem 3 for the full proof.
+Interpretation.
+The only safe primitive is the qualitative viability signal.
+That is what we call valence.
+Quality neutral representations must therefore be constructed later, as learned causes of valence.
 5 Multi-Layered Self-Organisation
 As the vocabulary and capacity for w-maxing scale, a greater variety of concepts can be learned.
-Progressively higher orders of ‘causal-identity’ for one’s self related information processing are constructed [39].
-This lets us frame the construction of embodied selves in developmental [70] and evolutionary terms.
+Progressively higher orders of ‘causal-identity’ for one’s self related information processing are constructed Bennett (2023a).
+This lets us frame the construction of embodied selves in developmental Ciaunica and Crucianelli (2019) and evolutionary terms.
 5.1 The First Order Self
-A first order self (1ST henceforth, formalised in definition 13) allows an organism to discern the consequences of its actions.
-This serves as the locus of self related information processing and experience [39], allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction.
-1ST order selves amount to reafference [71, 72], observed in mammals and insects.
-Figure 1: Visual intuition for a 1ST order self. The organism constructs a causal identity for itself and can relate that to observed events.
-A 1ST order self may require centralisation and a “solid brain” with a persistent structure [73] (an individual human), as opposed to a “liquid brain” (e.g. an ant colony or a population of humans).
+A first order self (1ST henceforth) allows an organism to discern the consequences of its actions.
+In our formalism it is the weakest causal identity that distinguishes self generated interventions from matched observations.
+The full definition and proof are in the ESM.
+This serves as the locus of self related information processing and experience Bennett (2023a), allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction.
+1ST order selves amount to reafference Merker (2007); Barron and Klein (2016), observed in mammals and insects.
+A formal role for the 1ST order self.
+If an organism can improve expected utility by treating intervention episodes differently from observation episodes, then some internal variable must encode that difference.
+Otherwise the organism would be forced to map the same embodied statement to the same internal state, and so could not act differently in the two cases.
+Under the scale and incentive preconditions already described, the candidate set of self intervention signatures is nonempty.
+A w-max learning rule then selects a maximally weak signature from that set.
+Any w-maximiser is inclusion minimal among candidates.
+If a strictly smaller signature existed, it would have a strictly larger extension and would have been preferred.
+This is the operational content of reafference in our framework.
+Why self is a fixed point of w-maxing.
+A short proposition shows that if there exists a minimal self intervention tag that is implied by every viable policy, and that tag is itself viable, then it is the weakest viable policy and w-maxing converges to it.
+See Box 3 for the sketch and see the ESM proposition titled “Self fixed point” for the full proof.
+A 1ST order self may require centralisation and a “solid brain” with a persistent structure Solé et al.
+(2019) (an individual human), as opposed to a “liquid brain” (e.g.
+an ant colony or a population of humans).
+A solid brain would be required for the signals that contribute to a first order self to be co-instantiated in proper time Bennett (2026a), and causally affect one another Bennett (2026c).
 5.2 The Second Order Selves
-Survival may demand organism $\mathfrak{a}$ infer $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s interventions (to see one’s self as if through another’s eyes [56]).
+Survival may demand organism $\mathfrak{a}$ infer $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s interventions (to see one’s self as if through another’s eyes Bennett (2023b)).
 This is called a second order self (2ND henceforth).
-We argue that if access conscious contents are available for communication in the human sense, then they must be communicable in the Gricean sense [74, 75].
+We argue that if access conscious contents are available for communication in the human sense, then they must be communicable in the Gricean sense Grice (1957, 1969).
 Grice argued that communication is about the inference of intent.
@@
 Where a 1ST order self might allow one to observe a cat and form plans regarding causal interactions with the cat, a 2ND order self would allow one to be consciously aware of the cat for the purpose of reasoning and report.
-One can know of the cat, and that another organism knows of the cat, but a 2ND order self is insufficient to be aware that one is aware of the cat [^13] This is ‘meta-self-reflexive consciousness’ as some have described it [76]..
-Figure 2: Visual intuition for a 2ND order self.
-The organism constructs a causal identity for itself, and for another object, and the causal identity for the other object includes a prediction of the organism itself from that object’s perspective.
-This would occur, for example, if one was a predator trying to predict the movements of prey in response to one’s own actions.
-It is an extension of the 1ST order self.
-More formally using the notation given in the appendix [40] (see quick reference guide in section 2), assume $\mathfrak{a}$ and $\mathfrak{b}$ are organisms that evolved to accurately predict one another’s behaviour.
+One can know of the cat, and that another organism knows of the cat, but a 2ND order self is insufficient to be aware that one is aware of the cat [^17] This is ‘meta-self-reflexive consciousness’ as some have described it Morin (2006)..
+More formally using the notation given in the appendix Bennett (2025a) (see quick reference guide in section 2), assume $\mathfrak{a}$ and $\mathfrak{b}$ are organisms that evolved to accurately predict one another’s behaviour.
 Assume $\mathfrak{a}$ constructs a causal identity $c^{\mathfrak{b}}_{\mathfrak{a}}$ to predict $\mathfrak{b}$ given input $i_{\mathfrak{a}}\in I_{\mu_{\mathfrak{a}}}$, of which a second order self $c^{\mathfrak{ba}}_{\mathfrak{a}}$ is part.
@@
 What is important here is that each organism’s intent is to some extent inferred by the other, changing the sorts of policies that are fit.
-For example, second order self means each knows the other can anticipate manipulation, which means the optimal policy will often be to have rather than feign intent that aligns to some extent with the other party’s desires, to co-operate [77] [^14] Depending upon circumstances, for example organisms may co-operate in some circumstances but not others [56], and transient relations, information asymmetry and other factors can make deceit a more attractive option..
-Repeated interaction creates an iterated prisoner’s dilemma, incentivising co-operation and signals that both parties interpret similarly (the beginnings of language) [56].
+For example, second order self means each knows the other can anticipate manipulation, which means the optimal policy will often be to have rather than feign intent that aligns to some extent with the other party’s desires, to co-operate Alexander et al.
+(2022) [^18] Depending upon circumstances, for example organisms may co-operate in some circumstances but not others Bennett (2023b), and transient relations, information asymmetry and other factors can make deceit a more attractive option..
+Repeated interaction creates an iterated prisoner’s dilemma, incentivising co-operation and signals that both parties interpret similarly (the beginnings of language) Bennett (2023b).
 To communicate in Gricean terms, $\mathfrak{a}$ must intend to convey meaning $m_{\mathfrak{a}}$, and $\mathfrak{b}$ must recognise this intent.
@@
 However, that does not obviate the need for second order selves, it just makes them easier to realise.
+Formal and experimental results supporting our claims regarding 2ND order selves are given in the ESM.
+A minimal separation task.
+Consider a speaker who must transmit a private bit to a listener.
+The listener can be in one of two decoder modes.
+One decodes literally and the other flips the bit.
+If the speaker cannot condition on any evidence about which decoder is active, its success is at most one half.
+If the speaker can send one probe signal and observe the listener’s response, it can infer the decoder and succeed with probability one.
+That conditioning variable is the smallest possible second order self.
+It is an internal handle for how the audience is interpreting the speaker.
+Decoder mismatch in Monte Carlo benchmarks.
+Monte Carlo benchmarks in ESM Experiment S2 show that a couple of short probes largely eliminate the cost of decoder mismatch.
+Passive signalling stays near chance, while probing approaches oracle performance.
+See Box 4 for the headline results, and see ESM Experiment S2 for the full protocol and plots.
 5.3 The Third Order Selves
-We can scale preconditions indefinitely, however for the purposes of this paper we will stop at 3RD order selves, because this appears to be the highest level of human consciousness, known as Meta self-awareness [76].
-This is the awareness that one is self-aware.
-If self-awareness stems from 2ND order selves, then it follows that meta self-awareness requires 3RD.
-Formally a 3RD order self for $\mathfrak{a}$ reflecting off $c^{\mathfrak{ba}}_{\mathfrak{a}}$ lets $\mathfrak{b}$ is $c^{\mathfrak{baba}}_{\mathfrak{a}}$.
+We can scale preconditions indefinitely.
+For the purposes of this paper we stop at 3RD order because this is where a self can extend across time as a stable narrative.
+It is where an agent can treat its present actions as evidence about its future behaviour.
+It is also where the agent can deliberately bind itself to reduce what it could do later.
+This is what makes long horizon cooperation and trust possible.
+Formally, a 3RD order self for $\mathfrak{a}$ relative to an audience $\mathfrak{b}$ is the iterated causal identity c^baba_a.
 It is $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$.
-Figure 3: Visual intuition for a 3RD order self. The organism constructs a 2ND order self for its 2ND order self, and so becomes aware that it is aware.
+$\mathfrak{a}$’s action now changes what $\mathfrak{b}$ expects of $\mathfrak{a}$ later, and $\mathfrak{a}$ knows that $\mathfrak{b}$ knows that $\mathfrak{a}$ knows.
+A formal separation in the ESM shows why narrative, 3RD order selves matter for trust.
+Even with perfect audience modelling, nonbinding promises do not support trust in equilibrium when exploitation remains available.
+If an agent can publicly bind itself by paying a one time cost that removes an exploit option, a trust equilibrium exists.
+Monte Carlo benchmarks in ESM Experiment S3 show the same pattern across randomised trust games.
+See Box 3 for the formal sketch and Box 4 for the experimental headline results.
+See ESM Experiment S3 for the full protocol and plots.
+Box 3: Other formal proofs in support of the main text claims The main text uses proof results besides Theorem 3.
+They constrain and support the narrative arguments of this paper.
+Here we give only the sketch level version.
+Complete proofs are in the ESM.
+Reminder.
+Weakness maximisation, called w-maxing, means selecting a correct policy with maximal weakness $w(\pi)=|E_{\pi}|$.
+In plain English, it keeps the largest set of future refinements open while staying correct.
+To tie this to related biological frameworks on cognition spaces Solé et al.
+(2026), weakness corresponds to the size of the “cognitive light cone” of an embodied system.
+Result 1.
+w-maxing learns causal intervention tags.
+Given matched intervention and observation episodes, a w-maximised causal identity exists in any finite embodied language and is inclusion minimal among candidates.
+In plain English, it is the least committal signature that still separates self caused change from external change.
+This proves an earlier but less formal claim Bennett (2023a).
+Result 2.
+Decoder mismatch can hide minds.
+A Bayesian likelihood ratio bound shows that an evaluator can be forced toward a noise hypothesis when every mind-like hypothesis in its model class fits the observed trace worse than noise by a constant factor.
+In other words, if you decode the signal with the wrong key, Bayes rule will tell you there is no sender [^19] This creates mind-blindness and relates to The Fermi Paradox Bennett (2022a)..
+Result 3.
+The 1ST order self is a w-max fixed point.
+If there exists a minimal self intervention tag that is implied by every viable policy and is itself viable, then it is the weakest viable policy and a w-max learner converges to it.
+In plain English, reafference becomes a stable self tag when survival depends on it.
+It is a mathematical result that predicts the reafference we observe in living organisms.
+Result 4.
+Gricean report and trust force higher order selves.
+In Gricean communication, what counts as a correct report depends on the listener’s inference, not a fixed signal to world mapping.
+We show that any signalling policy that is robust across variation in an audience’s epistemic state must condition on an audience model.
+In our formalism, that audience model is a 2ND order self.
+We also give a minimal commitment game in which even perfect audience modelling cannot sustain trust when exploitation remains available.
+If an agent can publicly bind itself by paying a one time cost that removes an exploit option, a trust equilibrium exists.
+In other words, reportable access is not just the ability to say a thing.
+It is saying the thing knowing a listener will understand, by predicting what they predict you intend by what you say, and sometimes even binding yourself so they can believe you (because they will predict your self bind and observe your behaviour is consistent with it).
+Taken together, Results 3 and 4 make the self hierarchy a derived necessity.
+It is the minimal structure that survives under generalisation optimal adaptation when the niche includes self intervention, report, and trust.
+See the ESM propositions on causal identity minimality, decoder mismatch, self fixed points, Gricean report, and credible commitment.
 6 The What and Why of Consciousness
-Up to now we have developed the conceptual toolbox that we use to dissolve the hard problem.
-We started with the observation that a human body self organises to maintain itself in the face of change both within and without.
-An organism does not spring into existence understanding number systems or objects, but constructs them through interaction with the environment.
-One can regard them as policies according to which the environment is interpreted.
+Up to now we have developed the conceptual toolbox that we use to cross the “explanatory gap” and explain consciousness from an embodied perspective.
+We started with the observation that a human body is a system that self-organises to maintain itself in the face of change both within and without its finite boundaries.
+To achieve this vital goal, the embodied system needs to process information qualitatively, meaning it is intrinsically valanced in service of survival.
+An organism does not spring into existence “objectively” by understanding objects and properties, but rather constructs them through subject-oriented interaction with the environment.
+One can regard objects and properties as policies according to which the environment is subjectively interpreted.
 To learn how to interpret the environment an organism must differentiate between states, must react to change, and learn policies according to the valence associated with that change.
-We will now argue that: i) there is something it is like to be an organism with a 1ST order self, ii) that subjective experience is the process of learning and enacting a hierarchy or causal identities, and iii) that access consciousness requires both 1ST and 2ND order selves.
-Because we do not presuppose representations, we explain how they emerge from attraction to and repulsion from contentless physical states.
-This means the foundation of information processing is valence.
-As we saw earlier, learned policy has valence, and it is a classifier of inputs.
-Those inputs are “information” in the mechanistic sense of the environment being in one state and not another.
-Hence, a policy is a classifier of information, but that information is not in a language, and it is not yet something labelled or quantified (until there is a policy that labels or quantifies).
+But how do we get from these embodied subjective experiences to the subjectivity that is usually linked to phenomenal consciousness in awake adult minds?
+We will now argue the following.
+  1. 1.
+    There is something it is like to be an organism with a 1ST order self.
+  2. 2.
+    Subjective experience is the process of learning and enacting a hierarchy of causal identities.
+  3. 3.
+    Access consciousness requires both 1ST and 2ND order selves.
+Because we do not presuppose representations, we explain how qualitative interpretations emerge from attraction to and repulsion from contentless physical states.
+This means the foundation of information processing in biological self-organising systems is valence.
 For biological, living self-organising systems, being alive is intrinsically good, being dead or ill is bad.
-For example, in developing self-organising systems such as the human body, it is good to have a better developed sense of smell than vision at early stages of life (babies are more accurate smellers than adults).
+For example, in developing self-organising systems such as the human body, it is good to have a better developed sense of smell than vision at early stages of life (babies are more accurate smellers than adults Schaal et al.
+(2020)).
 Different senses have different degrees of valences at different stages of the developmental ladder.
@@
 However, the idea here is that basic experiences (i.e.
-the self-organising process of staying alive) have valence at their very core (good=stay alive; bad=dead).
+the bodily information processing subserving the ultimate goal of staying alive) have valence at their very core (good=stay alive; bad=dead).
 This culminates in experiential, phenomenal ‘quality’ of appearance, of what it is likeness and how thing appear to the system (i.e.
@@
 Our aim is not to dismiss the latter; our claim is that one cannot have the latter without the former.
+As we saw earlier, learned policy has valence, and it is a classifier of inputs.
+Those inputs are what one may call “information” in the mechanistic sense of the environment being in one state and not another.
+Hence, a policy is a classifier of information, but that information is not in a language, and it is not yet something labelled or quantified (until there is a policy that labels or quantifies).
 Clearly, prelinguistic self-organising systems must classify and attach value and disvalue to states and anticipated states to prioritise and make decisions.
@@
 Objectively, a physical object is an aspect of the environment and every aspect is realized by one state or another.
-However, previous work has shown that for every aspect of the environment there is an equivalent program, meaning from an objective point of view what is or is not a physical object or property is a matter of interpretation [41].
-Subjectively an object, property or quality only exists if an organism constructs a causal identity for it.
+However, previous work has shown that for every aspect of the environment there is an equivalent program, meaning from an objective point of view what is or is not a physical object or property is a matter of interpretation Bennett (2024c).
+The radical and novel claim here is that, subjectively, an object or property only exists to the extent that that organism constructs a causal identity for it.
 At the most basic level, the foundation of all of this is physical attraction and repulsion.
 An organism hard-wired by natural selection can be physically attracted or repelled without any concept of what it is physically attracted or repelled from.
-It doesn’t know to what it is attracted or repelled from in the sense that it has not constructed a causal identity for that object.
+The system doesn’t know to what it is attracted or repelled from in the sense that it has not constructed a causal identity for that object.
 There is no named object which has the property ‘attractive’.
@@
 Over the course of this paper we have sought to explain how we get from simple ‘one dimensional’ hard-wired valence to a property or quality that has valence.
-Categorical variables for properties or ‘qualities’ like ‘the colour red’ or ‘the smell of coffee’ only exist when we have a causal identity for them, which requires scale and incentive.
-It can be hard-wired, or learned.
+We suggest that categorical variables for properties or ‘qualities’ like ‘the colour red’ or ‘the smell of coffee’ only exist when we have a causal identity for them, which requires scale and incentive.
+It can be hard-wired or learned.
 When we have only one cell we do not have a vocabulary capable of satisfying the scale precondition.
-However self-organising biological systems are collective [78].
-Control is delegated and distributed among smaller components [43] which operate concurrently.
-An individual cell might only have ‘one dimensional’ valence, but as soon as we have two cells we have something analogous to a second ‘dimension’.
-Cellular collectives are polycomputational systems, meaning one cell can play a part in more than one computation simultanously, at different scales [79] [^15] This is what the weakness of a policy captures; the number of interactions of which it can be a part..
-As we scale up the collective, its is impelled by the cells of which it is formed.
-Instead of ‘one dimensional’ valence, we now have a rich tapestry of competing drives.
-It is this rich tapestry that is a property or ‘quality’ of a state of the environment interpreted by the organism.
-It is this rich tapestry that can be reduced to causal identities for properties like ‘the color red’, ‘hunger’ or ‘thirst’.
-A higher level of abstraction is formed from this rich tapestry of valence and how it changes over many interactions (the different inputs and outputs in a task that imply generalizable causal identities).
-Hunger and thirst might have the same overall intensity and so one could say they have the same overall valence, but they are qualitatively different because they are different ‘tapestries’ of valence at a lower level of abstraction.
+However self-organising biological systems are collective Fields and Levin (2020).
+Control is delegated and distributed among smaller components Bennett (2024a) which operate concurrently.
+An individual cell might only have ‘one dimensional’ valence [^20] Debatably, as even single cells have complex sub-components that interact with each other., but as soon as we have two cells we have something analogous to a second ‘dimension’.
+Cellular collectives are polycomputational systems, meaning one cell can play a part in more than one computation simultaneously, at different scales Bongard and Levin (2023).
+As we scale up the collective, it is impelled by the cells of which it is formed.
+Instead of ‘one dimensional’ valence, we now have a rich tapestry of competing drives Bennett (2025a) [^21] The contributing parts of this tapestry are co-instantiated in proper time Bennett (2026a) and causally influence one another in space Bennett (2026c).
+This is important to understand how it relates to notions of conscious “unity” and “integration” in other theories Dennett and Kinsbourne (1992); Varela et al.
+(2001); Singer and Gray (1995); Fries (2005); Baars (1988); Dehaene and Naccache (2001); Tononi (2004); Lamme (2006) and psychophysics Eagleman and Sejnowski (2000); Pöppel (1997); Vroomen and Keetels (2010); Wallace and Stevenson (2014)..
+Our position is that it is this rich tapestry that is a property or ‘quality’ of a state of the environment interpreted by the organism.
+It is this rich tapestry that can be explained through causal identities for properties like ‘the color red’, ‘hunger’ or ‘thirst’.
+A higher level of abstraction may be formed from this rich tapestry of valence and how this changes over many interactions (the different inputs and outputs in a task that imply generalizable causal identities).
+For example, hunger and thirst might have the same overall intensity and so one could say they have the same overall valence, but they are qualitatively different because they are different ‘tapestries’ of valence at a lower level of abstraction.
 This is what is meant by The Psychophysical Principle of Causality in Section 3.
-In a self-organising system of sufficient scale with sufficient incentive, what starts out as simple ‘one dimensional’ valence culminates in categorical variables that serve different organismic needs.
-This is why the sound of middle C is different from the color blue.
+In a self-organising system of sufficiently complex scale with sufficient incentive [^22] Here we refer to the scale and incentive preconditions formally defined earlier., what starts out as simple ‘one dimensional’ valence culminates in categorical variables that serve different organismic needs.
+This is why the sound of middle C is phenomenologically different from the color blue.
 They have different causal identities.
 Middle C results from sound waves hitting the ear, blue results from light waves hitting the eye.
-These variations produce different qualia.
-Qualia do not need to be perfect.
-They simply need to be “good enough” to motivate actions necessary for survival in the environment of the organism.
-Qualia are produced by the sensory and nervous systems of the body.
-The valence of red light waves misses the eyes of dogs entirely.
-Dogs rely more on smell than sight.
-This suits their environmental niche.
-At a sufficient scale causal identities can be constructed, stored and recalled by their quality and valence.
+These variations produce different subjective embodied experiences.
 The bold claim here is that information processing (i.e.
@@
 Note that this is different from experience in the sense usually defined in the literature.
-Experience in the sense of experiential phenomenological content cannot exist without experience in the sense of embodied exploration of the body and world as we defined here (see also Ciaunica in prep).
 To put it provocatively quality precedes quantity, and quantity is nothing more than the interpretation of quality.
 Quality comes first and experiences should be regarded on a continuum rather than a switch on / switch off phenomena.
-All living systems experience the world through their bodies and as such, there is something what it is like to experience the world in that basic way (even when one is asleep).
+All living systems experience the world through their bodies and as such, there is something it is like to experience the world in that basic way (even when one is asleep).
 One can access those phenomenal, experiential aspects at a higher level, true, but by accessing them, it doesn’t mean that one ‘constructs’ consciousness or one becomes a conscious being.
@@
 Hence every such policy is a local state.
-A causal identity imbued with disposition by valence.
+A causal identity is inherently dispositional.
+It is a product of valence in self-organising systems such as human bodies.
 There would be a policy for one’s act of smelling coffee.
@@
 It has a quality, so the 1ST order self is “what it is like” to be that organism.
-Put another way, Nagel’s question of “what it is like to be” a particular organism could be answered if one could somehow have that organism’s first order self [1].
-An organism does not make a decision to interpret information, it just reacts.
+Put another way, Nagel’s question of “what it is like to be” a particular organism could be answered if one could somehow have that organism’s first order self Nagel (1974).
+An organism does not make a decision to interpret information, it just does it.
 What we call subjective experience is the process of learning and enacting a lattice of causal identities, and there is something that “has” all of those subjective experiences once there is a 1ST order self that is part of them all.
-This is where phenomenal consciousness begins, with experience as learning and exploration.
-One’s 2ND order selves would also have a certain qualitative character, and one’s 3RD order too.
-There would, however, be a very clear delineation between conscious and not.
+This is where phenomenal consciousness begins, with subjective embodied experience as learning and exploration.
+Likewise one’s 2ND order selves would have a certain qualitative character, and one’s 3RD order too.
+There would, however, be a very clear delineation between the presence and absence of conscious subjective experience.
 The absence of a 1ST order self would mean there would be no policy linking all interventions together, and so no “self” that experiences making them.
-Hence a 1ST order self must precede higher orders.
-Our arguments align with those of Merker, who has linked subjective experience to reafference [80, 71, 72].
+Hence a 1ST order self must precede higher order selves and not the other way around.
+Our approach thus stays in stark contrast with Higher Order Thought Theories of Consciousness (HOT) Fleming (2019).
+Rather our arguments align with those of Merker, who has linked subjective experience to reafference Merker (2005, 2007); Barron and Klein (2016).
 We agree that reafference is key, but we provide a very different explanation of why and how.
@@
 Hence these are complementary positions.
-7 Unifying Lower and Higher Order Theories of Consciousness
-The psychophysical principle of causality takes us from attraction to and repulsion from physical states, to higher order meta representations.
+7 Collapsing The Abstraction Layers
+Our theory of embodied consciousness places the subjective experience of embodied organisms in the physical world and not in some parallel, dualist world.
+This means roughly that the psychophysical principle of causality takes us from attraction to and repulsion from physical states, to higher order meta representations.
 This produces different orders of self.
-We suggest that the full richness of human subjective experience depends upon these different orders of self [^16] Perhaps the full richness of “hard” consciousness [81] depends upon the interaction of selves, perhaps in planning [82]..
+We suggest that the full richness of human subjective experience depends upon these different orders of self [^23] Perhaps the full richness of “hard” consciousness Boltuc (2012) depends upon the interaction of selves, perhaps in planning Shanahan (2012)..
 Conscious “access” as a human has it involves meaningful report, which as we’ve established requires 2ND order selves.
-This is very different from the mere “access” to information.
+This is very different from the mere “access” to information, in the sense of impinging signals upon the system or stored in memory.
 However, supposing we could somehow contrive a system with 2ND order selves but no 1ST, then 2ND order selves alone would still not allow one to reason about how one’s interventions might affect the contents of 2ND order selves to communicate meaning.
@@
 Our radical and provocative claim is that phenomenal consciousness without access consciousness is likely very common, but the reverse is implausible.
-For example, Block holds that phenomenally conscious content is phenomenal, whereas access conscious content is representational.
-In pancomputational enactivism there are only contentless states and the programs they imply.
-Any abstract “representational” content is just organised phenomenal content, clustered according to what causes valence.
-We don’t discard these two sorts of consciousness, but unify them by showing how phenomenal consciousness gives rise to access.
-Our framework thus makes a philosophical zombie impossible because there can be no perfect unconscious replica of a conscious human.
-The hard problem has it backwards.
-The question is not why qualia exist, but why anyone thinks representational contents can exist without first being learned through qualitative experience and discrimination.
-Far from suggesting there is no such thing as qualia, this suggests instead that there is no such thing as purely representational content in anything but the tools we construct and our interpretations of their behaviour.
-Cells are a material with agency [83], and human intelligence is the high level goal directed behaviour of a swarm of cells [78, 84, 28].
-When we embody human abstractions (e.g.
-arithmetic) in silico (e.g.
-the x86 instruction set), we disconnect the high level goal directed behaviour from the low level behaviour (and the motivating valence) that gave rise to it [43].
-We call the information embodied in the computer “representational” because it means something to us.
-However when we embody our abstractions in silico we disconnect them from the valence that motivated their construction [41].
-This suggests there is no such thing as representational contents.
-They are a fiction we have invented because we struggle to reduce our own abstractions to their basic nature: the causes of valence.
-This ‘dissolves’ the hard problem by going a level down to the fundamental drive: stay alive!
+Our framework thus makes a philosophical zombie impossible because there can be no perfect unconscious replica of a conscious embodied organism but without the qualitative aspect of information processing attached to it [^24] Removing the qualitative aspects would introduce unnecessary extra steps and thus inefficiencyBennett (2025a)..
+In a self-organising system is causality all the way down.
+Layer upon layer of abstraction classifies the causes of valence across spatial and temporal scales, from simple one-dimensional signals up to complex tapestries of valence that are the phenomena that make up a self-organising system’s experience.
+Related work argues qualia are nothing more than the tapestries of valence we describe Bennett (2025a).
+Put another way, it claims qualia are reducible to valence.
+If one were to subscribe to this view, then one might conclude it dissolves hard problem of consciousness, because the formalism [^25] Stack Theory, and Pancomputational Enactivism within it.
+describes every conceivable world and implies a philosophical zombie with access but not phenomenal consciousness is impossible in all of them.
+A philosophical zombie acts exactly as a human but without qualia.
+If one takes qualia to be tapestries of valence, then a system with a 1ST-order-self and the efficient, qualitative information processing we have described would have qualia.
+Thus a human without qualia would be less efficient than its conscious counterpart and thus not strictly speaking a philosophical zombie.
 8 From Rocks to Einstein: The Hierarchy of Being
-To illustrate how our argument applies in the real world we describe stages of conscious organism.
+In the last section of the paper we suggest the problem of phenomenal consciousness has it backwards.
+The question is not why qualia exist, but why anyone thinks representational contents can exist in a prelinguistic organism without direct modelling of qualities (i.e.
+qualia) and discrimination via information processing of cause and affect.
+To illustrate how our argument applies in the real world we need to go back to the simplest basic level and describe stages of conscious organism.
 Each stage follows from scaling up supply and demand for w-maxing, through natural selection.
@@
 $5:$ 3RD Order Selves (e.g. human)
-Stage 1: Hard Coded
-Stage one refers to adaptations hard-wired by natural selection, allowing complexity to persist [85] in a stable environment.
+Stage 1: Hard coded.
+Stage one refers to adaptations hard-wired by natural selection, allowing complexity to persist Heylighen (2022) in a stable environment.
   • What: Hard-coded adaptations. Habituation and sensitization.
@@
   • Example: Single-celled protozoan.
-Stage 2: Learning
+Stage 2: Learning.
 Stage two introduces learning.
 To learn an organism must store, classify and order historical examples by valence.
-However there is not something it is like to be stage two, because there is no locus of “self”.
+However there is not “something it is like” to be stage two, because there is no locus of “self”.
 A biological example of such a decentralised nervous system is the cubozoan box jellyfish Tripedalia cystophora.
-Even Tripedalia cystophora was recently shown to be capable of associative learning [86].
+Even Tripedalia cystophora was recently shown to be capable of associative learning Bielecki and others (2023).
 An entirely distributed control system can “learn”.
-Likewise, stage two is exemplified by nematodes [87, 88] with a centralised nervous system and some ability to adapt with experience.
-However, the absence of a “self” limits causal reasoning, which as others have already pointed out must limit spatial, navigational abilities [72].
+Likewise, stage two is exemplified by nematodes Yu and Rankin (2022); Willett et al.
+(2018) with a centralised nervous system and some ability to adapt with experience.
+However, the absence of a “self” limits causal reasoning, which as others have already pointed out must limit spatial, navigational abilities Barron and Klein (2016).
 When starved C.
-elegans exhibit “increased locomotion and dispersal in a random, rather than directed, search” [72, 89, 90], whereas something like a bee or an ant can recall and navigate to previously discovered food [91, 92, 93].
+elegans exhibit “increased locomotion and dispersal in a random, rather than directed, search” Barron and Klein (2016); Lüersen et al.
+(2014); Artyukhin et al.
+(2015), whereas something like a bee or an ant can recall and navigate to previously discovered food Wehner (2013); Seeley (1995); Oades and Isaacson (1978).
   • What: Learning.
@@
   • Examples: Jellyfish, nematode.
-Stage 3: 1ST Order Self
+Stage 3: 1ST order self.
 This is where phenomenal consciousness begins, with a 1ST order self.
-In biological terms this implies reafference, which others have argued is the key to subjective experience, albeit for different reasons than what we have [80, 71, 72].
+In biological terms this implies reafference, which others have argued is the key to subjective experience, albeit for different reasons than what we have Merker (2005, 2007); Barron and Klein (2016).
 They identified a housefly as a good example of where subjective experience may begin, and we concur.
 We also hold this is where an organism might be said to have intent.
-Intuitively, the policy that motivated behaviour is the intent of that behaviour [^17] In the same way declarative and imperative programs are equivalent [42]..
-For example, eating tends to involve the intent of satisfying hunger.
-A stage three organism can feel simple things like hunger, but cannot cannot conceive of itself from another’s perspective.
-Subsequently it cannot communicate in the Gricean sense [75, 56], or conceive of its own death, or experience shame.
-  • What: A 1ST order self. Reafference. Phenomenal consciousness.
+Intuitively, the policy that motivated behaviour is the intent of that behaviour [^26] In the same way declarative and imperative programs are equivalent Howard (1980)..
+For example, eating tends to involve the intent of satisfying hunger, which in turn satisfy the basic policy ‘stay alive’.
+A stage three organism can feel simple things like hunger, but cannot conceive of itself from another’s perspective.
+Subsequently it cannot communicate in the Gricean sense Grice (1969); Bennett (2023b), or conceive of its own death, or experience shame.
+  • What: 1ST order self. Reafference. Phenomenal consciousness.
   • How: Embodiment in which intervention is not identical to observation.
@@
   • Example: Housefly.
-Stage 4: 2ND Order Selves
-Stage four is the 2ND order self, and this is where we hold access consciousness begins because it is where information is available for report in the Gricean sense.
-The ability of ravens to intentionally deceive [94] suggests they are at least stage four.
+Stage 4: 2ND order selves.
+Stage four is the 2ND order self, and this is where Block’s reportable access consciousness begins in our formal account, because information is now available for report in the Gricean sense.
+In the ESM we prove that robust Gricean report requires an audience model, which is exactly a 2ND order self.
+The ability of ravens to intentionally deceive Bugnyar and Kotrschal (2002) suggests they are at least stage four.
 Raven $\mathfrak{a}$, aware that it is being observed by raven $\mathfrak{b}$, will act as if it is hiding food in one location to mislead $\mathfrak{b}$, but will then move the food in another location unobserved by $\mathfrak{b}$.
@@
 For example, a cat anticipates its prey will flee when it is observed, and hides.
-  • What: Access consciousness. Theory of mind. Self-awareness. Inner narrative.
+  • What: Reportable access consciousness. Theory of mind. Self-awareness.
   • How: Selection pressures that demand theory of mind.
-  • Why: A 2ND order self is necessary to anticipate, manipulate and communicate intent. This could be anticipating the reactions of predator or prey, or navigating a social hierarchy.
+  • Why: A 2ND order self is necessary to infer, communicate and manipulate intent.
   • Example: Cats, dogs, ravens.
-Stage 5: 3RD Order Selves
-A 3RD order self facilitates complex planning, particularly of a social nature.
-A 3RD order self is a 2ND order self for one’s 2ND order self.
-Humans appear to possess this, because we are aware that we are aware.
-Altruistic behaviour observed in Australian magpies [95] suggests they may also have 3RD order selves.
-  • What: Meta self-awareness. Inner narrative in which actors have inner narratives.
+Stage 5: 3RD order selves (narrative selves).
+A 3RD order self is a narrative self.
+It lets an organism treat itself as persisting across time, with a story that links past interventions to future commitments.
+Formally, a 3RD order self is a 2ND order self for one’s 2ND order self.
+In the ESM we give a minimal trust and commitment game that separates 2ND from 3RD order selves.
+It shows that even perfect audience modelling cannot sustain trust unless a publicly legible self binding move exists.
+Humans appear to possess this.
+Altruistic behaviour observed in Australian magpies Crampton et al.
+(2022) suggests they may also have narrative selves.
+  • What: Impelling narrative self. Meta self-awareness. Trust.
   • How: More accurate prediction and planning.
@@
   • Example: Human. Highly intelligent animals such as Australian magpies may also be this conscious.
-Figure 4: Overview of stages and orders of self.
+Hierarchy of Selves 0TH. No self
+No stable self tag. Behaviour is reactive and not model based.
+1ST. Self tag
+Separates self generated interventions from matched observations. This is reafference. Grounds phenomenal consciousness.
+2ND. Audience model
+Models how an audience will decode the agent. Enables Gricean report. Grounds reportable access consciousness.
+3RD. Narrative self
+Stabilises an identity across time. Enables long horizon planning, self binding, trust, and an impelling narrative.
+NTH. Order selves
+Higher order selves may facilitate as yet unexplored capabilities.
+Figure 1: Overview of orders of self as nested causal identities.
+0TH is an embodied agent without an explicit self tag.
+1ST adds a reafference like separator between self generated interventions and matched observations.
+On our account this grounds minimal phenomenal consciousness.
+2ND adds a model of how an audience will decode the agent.
+This enables Gricean report and grounds reportable access consciousness.
+3RD is a narrative self that stabilises an identity across time.
+This enables long horizon planning, self binding, trust, and an impelling narrative.
+Formal definitions and proofs are in the ESM, where we derive these orders from generalisation optimal learning under the scale and incentive preconditions.
+Picture kindly provided by Ricard Solé of The Santa Fe Institute.
 9 Conclusions: Why Nature Does Not Like Zombies
-We have described a multilayered formalism illustrating how biological self-organising systems become phenomenally conscious when they construct a 1ST order self.
-A human lacking a 1ST order self could not perform causal reasoning needed to adapt as humans evidently can [62].
-As previous research pointed out [36], the computer metaphor with hardware and software is a simplification.
-Rather, software is a state of hardware.
-Nature privileges efficiency over abstract simplification.
-Compare the vast quantities of both training data and energy required by a large language model, to the small quantities humans need to solve a problem.
-Biological systems are more efficient, because they are adaptive at a every level [28, 43].
-Hence, rather than trying to explain the mind in the abstract like software, we started at the level of the embodied organism.
-One consequence of our approach is that it places the phenomenal quality of conscious experience before access consciousness.
+In this paper we have proposed a novel, causal identity theory of consciousness, taking as a starting point not the conscious mind but the embodied experiencing organism.
+Building on Stack Theory Bennett (2025a), we have illustrated in formal terms how biological self-organising systems become phenomenally conscious when they construct a 1ST order self.
+A human lacking a 1ST order self could not perform causal reasoning needed to adapt as healthy humans can Pearl and Mackenzie (2018).
+Box 4: Experimental results at a glance We report three simple Monte Carlo benchmarks from the associated workshop paper Bennett (2026b).
+They are minimal tests of the formal claims in noisy randomised settings.
+Experiment 1.
+Weakness beats simplicity.
+Across randomised abstraction layers, w-maxing has zero regret under a natural ignorance prior.
+A shortest description length selector incurs large regret.
+In one benchmark, that regret is about $13.4$ bits.
+That is an odds penalty of about $2^{13.4}\approx 10{,}800$.
+In other words description length depends on the encoding, while weakness is an invariant of the embodied language.
+Experiment 2.
+Probing defeats decoder mismatch.
+When a listener’s decoding rule is hidden, passive signalling stays near chance.
+With two short probes, a probing policy reaches about $0.93$ success across both stationary and nonstationary decoder settings.
+An oracle decoder reaches about $0.96$ in the stationary setting.
+In other words, causal intervention helps ascertain the listener’s decoder for communication.
+Experiment 3.
+Binding creates trust.
+In randomised trust games, nonbinding promises do not generate trust in equilibrium.
+Binding moves do.
+In one benchmark sample, binding yields trust about $0.95$ by horizon $H=5$.
+Third order selves thus help navigate complex social environments.
+Protocols and plots.
+See ESM Experiments S1 to S3, with Tables S1 to S4 and Figures S1 to S4.
+One consequence of our approach is that it places the qualitative aspect of conscious embodied experience [^27] Inherently valenced information processing through the lens of self-preservation.
+before access consciousness.
+Ontologically, this means that before one can access information, one needs to experience information through bodily and causal interaction with the world.
 We have shown a consistent definition of access consciousness requires 1ST and 2ND order selves.
@@
 Unlike panpsychism, we don’t believe rocks are conscious.
-Only self-organising systems that need to adapt, motivated by valence, while keeping track of the self are conscious.
-Consciousness is an adaptation, and a philosophical zombie is impossible.
-Furthermore, the conceivability of zombies depends on the ability to duplicate disposition but not phenomenal experience.
-Carruth argued that if qualia are dispositional, then that is impossible [96].
-Our arguments align with Carruth’s.
-A causal identity is inherently dispositional, as is embodiment.
-In summary, we have argued that access consciousness at the human level is impossible without the ability to hierarchically model i) the self, ii) the world and others and iii) the self as modelled by others.
+Only self-organising systems that need to adapt, motivated by valence, while keeping track of self-relevant information are conscious.
+Consciousness is a key adaptation of a vulnerable organism in a physical world.
+In summary, we have argued that access consciousness at the human level is impossible without the ability to hierarchically model the following.
+  1. 1.
+    The self.
+  2. 2.
+    The world and others.
+  3. 3.
+    The self as modelled by others.
 These follow from learning and enacting a lattice of causal identities, if the scale and incentive preconditions are met.
-There is something it is like to be an organism with a self, and that is where phenomenal consciousness begins.
-Our proposal lays the foundation of a formal science of consciousness, deeply connected with natural selection rather than abstract thinking, closer to the organic facts of humanity than the conceptual fiction of zombies.
-Funding
+There is something it is like to be an organism with a self.
+Paradoxically, phenomenal consciousness begins in organisms that have an end and a primary goal to avoid that end.
+Our proposal lays the foundation of a formal science of consciousness, deeply connected with natural selection rather than abstract thinking, closer to the facts of embodiment than the fiction of zombies.
+Acknowledgements
 Supported by a Fundaçao para a Ciencia e a Tecnologia (FCT) grant PTDC/FER-FIL/4802/2020; and 2020-02773 CEECIND FCT to A.C.
-10 Appendix
-For convenience of reference, we have placed all definitions here.
-We have aimed to make the gist of the paper understandable without the math, hence the reader can either skip this section or refer back to it if needed.
-Many of the definitions have been adapted from a variety of preceding work [36, 38, 39, 56, 41, 43].
-They are referred to in the body of the paper when they become relevant, in the order in which they appear here.
-Definition 1 (environment).
-  • We assume a set $\Phi$ whose elements we call states.
-  • A declarative program is $f\subseteq\Phi$, and we write $P$ for the set of all programs (the powerset of $\Phi$).
-  • By a truth or fact about a state $\phi$, we mean $f\in P$ such that $\phi\in f$.
-  • By an aspect of a state $\phi$ we mean a set $l$ of facts about $\phi$ s.t.
-$\phi\in\bigcap l$.
-By an aspect of the environment we mean an aspect $l$ of any state, s.t.
-$\bigcap l\neq\emptyset$.
-We say an aspect of the environment is realised [^18] Realised meaning it is made real, or brought into existence.
-by state $\phi$ if it is an aspect of $\phi$.
-Definition 2 (abstraction layer).
-  • We single out a subset $\mathfrak{v}\subseteq P$ which we call the vocabulary of an abstraction layer. The vocabulary is finite.
-  • $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap l\neq\emptyset\}$ is a set of aspects in $\mathfrak{v}$. We call $L_{\mathfrak{v}}$ a formal language, and $l\in L_{\mathfrak{v}}$ a statement.
-  • We say a statement is true given a state iff it is an aspect realised by that state.
-  • A completion of a statement $x$ is a statement $y$ which is a superset of $x$. If $y$ is true, then $x$ is true.
-  • The extension of a statement $x\in{L_{\mathfrak{v}}}$ is $E_{x}=\{y\in{L_{\mathfrak{v}}}:x\subseteq y\}$. $E_{x}$ is the set of all completions of $x$.
-  • The extension of a set of statements $X\subseteq{L_{\mathfrak{v}}}$ is $E_{X}=\bigcup\limits_{x\in X}E_{x}$.
-  • We say $x$ and $y$ are equivalent iff $E_{x}=E_{y}$.
-(notation) $E$ with a subscript is the extension of the subscript191919e.g. $E_{l}$ is the extension of $l$..
-(intuitive summary) $L_{\mathfrak{v}}$ is everything which can be realised in this abstraction layer.
-The extension $E_{x}$ of a statement $x$ is the set of all statements whose existence implies $x$, and so it is like a truth table.
-Intuitively a sensorimotor system is an abstraction layer.
-Likewise, a computer (each statement asserting a state of the computer, or a part thereof).
-Definition 3 ($\mathfrak{v}$-task).
-For a chosen $\mathfrak{v}$, a task $\alpha$ is a pair $\langle{I}_{\alpha},{O}_{\alpha}\rangle$ where:
-  • ${I}_{\alpha}\subset L_{\mathfrak{v}}$ is a set whose elements we call inputs of $\alpha$.
-  • ${O_{\alpha}}\subset E_{I_{\alpha}}$ is a set whose elements we call correct outputs of $\alpha$.
-${I_{\alpha}}$ has the extension $E_{I_{\alpha}}$ we call outputs, and ${O_{\alpha}}$ are outputs deemed correct. $\Gamma_{\mathfrak{v}}$ is the set of all tasks given $\mathfrak{v}$.
-[…]
-(notation) If $\omega\in\Gamma_{\mathfrak{v}}$, then we will use subscript $\omega$ to signify parts of $\omega$, meaning one should assume $\omega=\langle{I}_{\omega},{O}_{\omega}\rangle$ even if that isn’t written.
-(intuitive summary) To reiterate and summarise the above:
-  • An input is a possibly incomplete description of a world.
-  • An output is a completion of an input [def. 2].
-  • A correct output is a correct completion of an input.
-A $\mathfrak{v}$-task is a formal, behavioural description of goal directed behaviour.
-For example, an organism could be described by all behaviour in which it remains alive.
-Likewise, a $\mathfrak{v}$-task could describe a Turing machine.
-Definition 4 (inference).
-  • A $\mathfrak{v}$-task policy is a statement $\pi\in L_{\mathfrak{v}}$. It constrains how we complete inputs.
-  • $\pi$ is a correct policy iff the correct outputs $O_{\alpha}$ of $\alpha$ are exactly the completions $\pi^{\prime}$ of $\pi$ such that $\pi^{\prime}$ is also a completion of an input.
-  • The set of all correct policies for a task $\alpha$ is denoted $\Pi_{\alpha}$.
-[^20] To repeat the above definition in set builder notation: ${\Pi_{\alpha}}=\{\pi\in L_{\mathfrak{v}}:{E}_{I_{\alpha}}\cap E_{\pi}={O_{\alpha}}\}$
-Assume $\mathfrak{v}$-task $\omega$ and a policy $\pi\in L_{\mathfrak{v}}$. Inference proceeds as follows:
-  1. 1.
-    we are presented with an input ${i}\in{I}_{\omega}$, and
-  2. 2.
-    we must select an output $e\in E_{i}\cap E_{\pi}$.
-  3. 3.
-    If $e\in{O}_{\omega}$, then $e$ is correct and the task “complete”.
-$\pi\in{\Pi}_{\omega}$ implies $e\in{O}_{\omega}$, but $e\in{O}_{\omega}$ doesn’t imply $\pi\in{\Pi}_{\omega}$ (an incorrect policy can imply a correct output).
-(intuitive summary) To reiterate and summarise the above:
-  • A policy constrains how we complete inputs.
-  • A correct policy is one that constrains us to correct outputs.
-In functionalist terms, a policy is a “causal intermediary” between inputs and outputs.
-Definition 5 (learning).
-A proxy $<$ is a binary relation on statements.
-In this paper we use only one proxy, called the weakness proxy, which compares the cardinality of a statement’s extension.
-For statements $l_{1},l_{2}$ we have $l_{1}<l_{2}$ iff $\lvert Z_{l_{1}}\rvert<\lvert Z_{l_{2}}\rvert$.
-Whenever we use $<$ to compare statements, we are referring to the aforementioned weakness proxy.
-(generalisation) A statement $l$ generalises to a $\mathfrak{v}$-task $\alpha$ iff $l\in\Pi_{\alpha}$.
-We speak of learning $\omega$ from $\alpha$ iff, given a proxy $<$, $\pi\in{\Pi}_{\alpha}$ maximises $<$ relative to all other policies in ${\Pi}_{\alpha}$, and $\pi\in{\Pi}_{\omega}$.
-(probability of generalisation) We assume a uniform distribution over $\Gamma_{\mathfrak{v}}$.
-If $l_{1}$ and $l_{2}$ are policies, we say it is less probable that $l_{1}$ generalizes than that $l_{2}$ generalizes, written $l_{1}<_{g}l_{2}$, iff, when a task $\alpha$ is chosen at random from $\Gamma_{\mathfrak{v}}$ (using a uniform distribution) then the probability that $l_{1}$ generalizes to $\alpha$ is less than the probability that $l_{2}$ generalizes to $\alpha$.
-(sample efficiency) Suppose $\mathfrak{app}$ is the set of all pairs of policies. Assume a proxy $<$ returns $1$ iff true, else 00. Proxy $<_{a}$ is more sample efficient than $<_{b}$ iff
-$\left(\sum_{(l_{1},l_{2})\in\mathfrak{app}}\lvert(l_{1}<_{g}l_{2})-(l_{1}<_{a}l_{2})\rvert-\lvert(l_{1}<_{g}l_{2})-(l_{1}<_{b}l_{2})\rvert\right)<0$
-(optimal proxy) There is no proxy more sample efficient than weakness.
-The weakness proxy formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [38]).
-(intuitive summary) Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy.
-Humans typically learn from “examples”.
-An example of a task is a correct output and input.
-A collection of examples is a child task, so “learning” is an attempt to generalise from a child to one of its parents.
-The lower level the child from which an agent generalises to parent, the “faster” it learns, the more sample efficient the proxy.
-The most sample efficient proxy is weakness [38, prop.
-1, 2], which is why we’re using it here.
-Definition 6 (organism).
-We can describe the circumstances of an organism $\mathfrak{o}$ as $\langle\mathfrak{v}_{\mathfrak{o}},\mu_{\mathfrak{o}},\mathfrak{p}_{\mathfrak{o}},<_{\mathfrak{o}}\rangle$ where:
-  • ${O}_{\mu_{\mathfrak{o}}}$ contains every output which qualifies as “fit” according to natural selection.
-  • $\mathfrak{p}_{\mathfrak{o}}$ is the set of policies an organism knows, s.t. $\mathfrak{p}_{\mathfrak{o}}\subset\mathfrak{p}_{n.s.}\cup\mathfrak{p}_{\mathfrak{h}_{{<t}_{\mathfrak{o}}}}$ and:
-    • $\mathfrak{p}_{n.s.}\subset L_{\mathfrak{v}_{\mathfrak{o}}}$ is reflexes hard coded from birth by natural selection.
-    • $\mathfrak{p}_{\mathfrak{h}_{{<t}_{\mathfrak{o}}}}=\underset{\zeta\in\mathfrak{h}_{{<t}_{\mathfrak{o}}}}{\bigcup}\Pi_{\zeta}$ is the set of policies it is possible to learn from a history of past interactions represented by a task $\mathfrak{h}_{{<t}_{\mathfrak{o}}}$.
-    • If $\Pi_{{\mathfrak{h}_{<t_{\mathfrak{o}}}}}\not\subset(\mathfrak{p}_{\mathfrak{o}}-\mathfrak{p}_{n.s.})$ then the organism has selective memory.
-It can “forget” outputs, possibly to productive ends if they contradict otherwise good policies.
-  • $<_{\mathfrak{o}}$ is a binary relation over $\Gamma_{\mathfrak{v}_{\mathfrak{o}}}$ we call preferences.
-(intuitive summary) Strictly speaking an organism $\mathfrak{o}$ would be a policy, but we can describe the circumstances of its existence as a task $\mu$ that describes all “fit” behaviour for that organism.
-We can also identify policies the organism “knows”, because these are implied by the policy that is the organism.
-Likewise, we can represent lossy memory by having the organism “know” fewer policies than are implied by its history of interactions.
-Finally, preferences are the particular “protosymbol” the organism will use to “interpret” an input in later definitions.
-Definition 7 (protosymbol system).
-Assume an organism $\mathfrak{o}$.
-For each policy $p\in\mathfrak{p}_{\mathfrak{o}}$ there exists a set $\mathfrak{s}_{p}=\{\alpha\in\Gamma_{\mathfrak{v}_{\mathfrak{o}}}:p\in\Pi_{\alpha}\}$ of all tasks for which $p$ is a correct policy.
-The union of all such sets is
-$\mathfrak{s}_{\mathfrak{o}}=\underset{p\in\mathfrak{p}_{\mathfrak{o}}}{\bigcup}\{\alpha\in\Gamma_{\mathfrak{v}_{\mathfrak{o}}}:p\in\Pi_{\alpha}\}$
-We call $\mathfrak{s}_{\mathfrak{o}}$ a “protosymbol system”.
-A $\mathfrak{v}$-task $\alpha\in\mathfrak{s}_{\mathfrak{o}}$ is called a “protosymbol”, and is “more abstract” if it is higher in the generational hierarchy.
-Definition 8 (interpretation).
-Interpretation is an activity undertaken by an organism $\mathfrak{o}=\langle\mathfrak{v}_{\mathfrak{o}},\mu_{\mathfrak{o}},\mathfrak{p}_{\mathfrak{o}},<_{\mathfrak{o}}\rangle$, as follows:
-  1. 1.
-    Assume an input $i\in L_{\mathfrak{v}_{\mathfrak{o}}}$.
-  2. 2.
-    We say that $i$ signifies a protosymbol $\alpha\in\mathfrak{s}_{\mathfrak{o}}$ if $i\in I_{\alpha}$.
-  3. 3.
-    $\mathfrak{s}_{\mathfrak{o}}^{i}=\{\alpha\in\mathfrak{s}_{\mathfrak{o}}:i\in I_{\alpha}\}$ is the set of all protosymbols which $i$ signifies.
-  4. 4.
-    If $\mathfrak{s}_{\mathfrak{o}}^{i}\neq\emptyset$ then $i$ means something to the organism in the intuitive sense that there is “affect” or “value” compelling the organism to act.
-  5. 5.
-    If $i$ means something, then $\mathfrak{o}$ chooses $\alpha\in\mathfrak{s}_{\mathfrak{o}}^{i}$ that maximises its preferences $<_{\mathfrak{o}}$.
-  6. 6.
-    The organism then infers an output $o\in E_{i}\cap E_{\Pi_{\alpha}}$.
-(intuitive summary) Interpretation is inference, with the additional step of choosing policies according to preference.
-This allows for irrational and instinctive choices, as well as rational ones.
-Intuitively, $i$ is every aspect of the context in which the organism finds itself; everything that can influence its interpretation.
-Definition 9 (to affect).
-Suppose we have two organisms, $\mathfrak{a}$ (Alice) and $\mathfrak{b}$ (Bob). Suppose $\mathfrak{a}$ interprets $i\in L_{\mathfrak{v}_{\mathfrak{o}}}$ as an output $o$, then:
-  • a statement $v\subset i$ affects $\mathfrak{a}$ if $\mathfrak{a}$ would have interpreted $e=i-v$ as a different output $g\neq o$.
-  • an organism $\mathfrak{b}$ has affected $\mathfrak{a}$ by making an output $k$ if, because of $k$, there exists $v\subset s$ which affects $\mathfrak{a}$.
-Definition 10 (intervention).
-By event we mean a statement in $L_{\mathfrak{v}}$, and an event happens or is observed iff it is a true statement given a state $\phi$.
-If $obs\in L_{\mathfrak{v}}$ is sensorimotor activity we interpret as an “observed event”, and $int\in L_{\mathfrak{v}}$ is an intervention to cause that event, then $obs\subset int$ (because $int$ could not be said to cause $obs$ unless $obs\subset int$).
-(intuitive summary) An intervention is action undertake an organism or other agency, in the sense described by Pearl [62].
-Intuitively, if “$int$” and “$obs$” are events which have happened, then we say that $int$ has caused $obs$ if $obs$ would not have happened in the absence of $int$ (counterfactual).
-Definition 11 (causal identity).
-If $obs\in L_{\mathfrak{v}}$ is an observed event, and $int\in L_{\mathfrak{v}}$ is in intervention causing $obs$, then $c\subseteq int-obs$ “identifies” or “names” the intervening agency.
-If $c=\emptyset$ then we have no way of knowing the intervening agency, if there is one.
-We call $c$ a causal identity corresponding to $int$ and $obs$.
-Suppose $INT$ and $OBS$ are sets of statements, and we assume $OBS$ contains observed events and $INT$ interventions, then a causal identity corresponding to $INT$ and $OBS$ is $c\neq\emptyset$ s.t.
-$\forall i\in INT(c\subset int)$ and $\forall obs\in OBS(c\cap obs=\emptyset)$ (we can attempt to construct a causal identity for any $INT$ and $OBS$).
-If a policy is a causal identity, then the associated task is to classify interventions.
-Definition 12 (purpose, goal or intent).
-We consider a policy $c$ which is a causal identity corresponding to $INT$ and $OBS$ to be the intent, purpose or goal ascribed to the interventions.
-$c$ is what the interventions share in common, meaning the “name” or “identity” of behaviour is the “intent”, “goal” or “purpose” of behaviour.
-Just as an intervention caused an observation, the intent which motivated the agency undertaking the intervention is what caused it (to correctly infer intent, one must infer a causal identity that implies subsequent interventions).
-Definition 13 (1ST order self).
-If $c$ is the lowest level causal identity corresponding to $INT$ and $OBS$, and $INT$ is every intervention an organism could make (not just past interventions, but all potential future interventions), then we consider $c$ to be the system’s 1ST order self.
-If $c\in\mathfrak{p}_{\mathfrak{o}}$ then an organism has constructed a 1ST order self.
-A 1ST order self for an organism $\mathfrak{o}$ is denoted $\mathfrak{o^{1}}$.
-An organism has at most one 1ST order self.
-(intuitive summary) Intuitively, $\mathfrak{o^{1}}$ is where we draw the line between what the organism can intend and what it cannot.
-It is conceivable we might have two “organisms” in the same body by this definition, each with its own 1ST order causal identity.
-Ultimately, where an organism begins or ends remains malleable.
-Definition 14 (preconditions).
-If $\mathfrak{o}$ is an organism, and $c$ is a causal identity, the $\mathfrak{o}$ will construct $c$ only if the representation and incentive preconditions below are met:
-  • the scale precondition is met iff $c\in L_{\mathfrak{v_{o}}}$, and
-  • the incentive precondition is met if $\mathfrak{o}$ must learn $c$ to remain “fit”.
-(intuitive summary) If $c$ is a 1ST order self, then these are the preconditions that must be met for an organism to construct $c$. Likewise, any other sort of causal identity.
-Definition 15 (chain notation).
-Suppose we have two organisms, $\mathfrak{a}$ (Alice) and $\mathfrak{b}$ (Bob).
-$c^{\mathfrak{b}}_{\mathfrak{a}}$ denotes a causal identity for $\mathfrak{b}$ constructed by $\mathfrak{a}$ (what Alice thinks Bob intends).
-Subscript denotes the organism who constructs the causal identity, while superscript denotes the object.
-The superscript can be extended to denote chains of predicted causal identity.
-For example, $c^{\mathfrak{ba}}_{\mathfrak{a}}\subset c^{\mathfrak{b}}_{\mathfrak{a}}$ denotes $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a^{1}}$ (what Alice thinks Bob thinks Alice intends).
-The superscript of $c^{\mathfrak{*}}_{\mathfrak{a}}$ can be extended indefinitely to indicate recursive predictions; however the extent recursion is possible is determined by $\mathfrak{a}$’s vocabulary $\mathfrak{v}_{\mathfrak{a}}$.
-Finally, Bob need not be an organism.
-Bob can be anything for which Alice constructs a causal identity.
-[…]
-An $n^{th}$ order self for $\mathfrak{a}$ is $\mathfrak{a^{n}}=c^{\mathfrak{*a}}_{\mathfrak{a}}$ where * is replaced by a chain, and $n$ denotes the number of reflections.
-For example, a 2ND order self $\mathfrak{a^{2}}=c^{\mathfrak{ba}}_{\mathfrak{a}}$, and a 3RD order self $\mathfrak{a^{3}}=c^{\mathfrak{baba}}_{\mathfrak{a}}$.
-We use $\mathfrak{a^{2}}$ to refer to any 2ND order self, and chain notation to refer to a specific 2ND order self, for example $c^{\mathfrak{ba}}_{\mathfrak{a}}$.
-The union of two $n^{th}$ order selves is also considered to be an $n^{th}$ order self, for example $\mathfrak{a^{3}}=c^{\mathfrak{baba}}_{\mathfrak{a}}\cup c^{\mathfrak{dada}}_{\mathfrak{a}}$, and the weaker or higher level a self is in the generational hierarchy, the more selves there are of which it is part.
-Definition 17 (stages of consciousness).
-We argue the following stages by scaling the ability to learn weak policies:
-  1. 1.
-    Hard Coded: organism that acts but does not learn, meaning $\mathfrak{p}_{\mathfrak{o}}$ is fixed from birth.
-  2. 2.
-    Learning: an organism that learns, but $\mathfrak{o^{1}}\not\in\mathfrak{p}_{\mathfrak{o}}$ either because $\mathfrak{o^{1}}\not\in L_{\mathfrak{v_{o}}}$ (failing the “scale precondition”) or because the organism is not incentivised to construct $\mathfrak{o^{1}}$ (failing the “incentive precondition”).
-  3. 3.
-    1ST order self: reafference and phenomenal or core consciousness are achieved when $\mathfrak{o^{1}}\in\mathfrak{p}_{\mathfrak{o}}$ is learned by an organism because of attraction to and repulsion from statements in $L_{\mathfrak{v}_{\mathfrak{o}}}$.
-  4. 4.
-    2ND order selves:
-    1. (a)
-      access or self-reflexive consciousness is achieved when $\mathfrak{o^{2}}\in\mathfrak{p}_{\mathfrak{o}}$.
-    2. (b)
-      hard consciousness is achieved when a phenomenally conscious organism learns a 2ND order self (an organism is consciously aware of the contents of 2ND order selves, which must have quality if learned through phenomenal conscious).
-  5. 5.
-    3RD and higher order selves: meta self-reflexive consciousness (human level hard consciousness) is achieved when $\mathfrak{o^{3}}\in\mathfrak{p}_{\mathfrak{o}}$.
-10.1 Proofs
-Versions of proofs 1-3 were published in [38]. Further details are available in on GitHub, along with the supporting experiments [40].
-Proposition 1 (sufficiency).
-Assume $\alpha\sqsubset\omega$.
-The weakness proxy sufficient to maximise the probability that a parent $\omega$ is learned from a child $\alpha$ [^21] Assume there exist correct policies for $\omega$, because otherwise there would be no point in trying to learn it..
-Proof.
-You’re given the definition of $\mathfrak{v}$-task $\alpha$ from which you infer a hypothesis $\mathbf{h}\in{\Pi}_{\alpha}$. To learn $\omega$, you need $\mathbf{h}\in\Pi_{\omega}$:
-  1. 1.
-    For every $\mathbf{h}\in\Pi_{\alpha}$ there exists a $\mathfrak{v}$-task $\gamma_{\mathbf{h}}\in\Gamma_{\mathfrak{v}}$ s.t.
-$O_{\gamma_{\mathbf{h}}}=E_{\mathbf{h}}$, meaning $\mathbf{h}$ permits only correct outputs for that task regardless of input.
-We’ll call the highest level task $\gamma_{\mathbf{h}}$ s.t.
-$O_{\gamma_{\mathbf{h}}}=E_{\mathbf{h}}$ the policy task of $\mathbf{h}$ [^22] I’d like to give credit here to Nora Belrose for pointing out an error.
-Nora pointed out I was miscounting the number of tasks.
-As a result I realised I was not counting tasks, I was in fact counting policy tasks and had entirely neglected to mention this fact.
-This was a significant error which has now been corrected, with several additional steps added to account for equivalence..
-  2. 2.
-    $\omega$ is either the policy task of a policy in $\Pi_{\alpha}$, or a child thereof.
-  3. 3.
-    If a policy $\mathbf{h}$ is correct for a parent of $\omega$, then it is also correct for $\omega$.
-Hence we should choose $\mathbf{h}$ that has a policy task with the largest number of children.
-As tasks are uniformly distributed, that will maximise the probability that $\omega$ is $\gamma_{\mathbf{h}}$ or a child thereof.
-  4. 4.
-    For the purpose of this proof, we say one task is equivalent [^23] This is because switching from $\beta$ to $\zeta$ s.t.
-$I_{\beta}\neq I_{\zeta}$ and $O_{\beta}=O_{\zeta}$ would be to pursue the same goal in different circumstances.
-This is because inputs are subsets of outputs, so both sets of inputs are implied by the outputs.
-$O_{\zeta}$ implies $I_{\beta}$ and $O_{\beta}$ implies $I_{\zeta}$ to another if it has the same correct outputs.
-  5. 5.
-    No two policies in $\Pi_{\alpha}$ have the same policy task [^24] Every policy task for policies of $\alpha$ is non-equivalent from the others..
-This is because all the policies in $\Pi_{\alpha}$ are derived from the same set inputs, $I_{\alpha}$.
-  6. 6.
-    The set of statements which might be outputs addressing inputs in ${I}_{\omega}$ and not ${I}_{\alpha}$, is $\overline{E_{{I}_{\alpha}}}=\{l\in L_{\mathfrak{v}}:l\notin E_{{I}_{\alpha}}\}$ [^25] This is because $E_{{I}_{\alpha}}$ contains every statement which is a correct output or an incorrect output, and $\overline{E_{{I}_{\alpha}}}$ contains every statement which could possibly be in $I_{\omega}$, $E_{I_{\omega}}$ and thus $O_{\omega}$..
-  7. 7.
-    For any given $\mathbf{h}\in{\Pi}_{\alpha}$, the extension $E_{\mathbf{h}}$ of $\mathbf{h}$ is the set of outputs $\mathbf{h}$ implies.
-The subset of $E_{\mathbf{h}}$ which fall outside the scope of what is required for the known task $\alpha$ is $\overline{E_{{I}_{\alpha}}}\cap E_{\mathbf{h}}$ [^26] This is because $E_{{I}_{\alpha}}$ is the set of all conceivable outputs by which one might attempt to complete $\alpha$, and so the set of all outputs that can’t be made when undertaking $\alpha$ is $\overline{E_{{I}_{\alpha}}}$ because those outputs occur given inputs that aren’t part of ${I}_{\alpha}$..
-  8. 8.
-    $\left|\overline{E_{{I}_{\alpha}}}\cap E_{\mathbf{h}}\right|$ increases monotonically with $\left|E_{\mathbf{h}}\right|$, because for all $e\in E_{\mathbf{h}}$ is $e\notin\overline{E_{{I}_{\alpha}}}$ iff $e\in E_{{I}_{\alpha}}$.
-  9. 9.
-    $2^{\left|\overline{E_{{I}_{\alpha}}}\cap E_{\mathbf{h}}\right|}$ is the number of non-equivalent parents of $\alpha$ to which $\mathbf{h}$ generalises.
-It increases monotonically with the weakness of $\mathbf{h}$.
-  10. 10.
-    Given $\mathfrak{v}$-tasks are uniformly distributed and $\Pi_{\alpha}\cap\Pi_{\omega}\neq\emptyset$, the probability that $\mathbf{h}\in{\Pi}_{\alpha}$ generalises to $\omega$ is
-    $p(\mathbf{h}\in{\Pi}_{\omega}\mid\mathbf{h}\in{\Pi}_{\alpha},\alpha\sqsubset\omega)=\frac{2^{\left|\overline{E_{{I}_{\alpha}}}\cap E_{\mathbf{h}}\right|}}{2^{\left|\overline{E_{{I}_{\alpha}}}\right|}}$
-$p(\mathbf{h}\in{\Pi}_{\omega}\mid\mathbf{h}\in{\Pi}_{\alpha},\alpha\sqsubset\omega)$ is maximised when $\left|E_{\mathbf{h}}\right|$ is maximised.
-Recall from definition 4 that $<_{w}$ is the weakness proxy.
-For statements $l_{1},l_{2}$ we have $l_{1}<_{w}l_{2}$ iff $\lvert E_{l_{1}}\rvert<\lvert E_{l_{2}}\rvert$.
-$\mathbf{h}$ that maximises $<_{w}$ will also maximise $p(\mathbf{h}\in{\Pi}_{\omega}\mid\mathbf{h}\in{\Pi}_{\alpha},\alpha\sqsubset\omega)$.
-Hence the weakness proxy maximises the probability that [^27] Subsequently it also maximises the sample efficiency with which a parent $\omega$ is learned from a child $\alpha$.
-a parent $\omega$ is learned from a child $\alpha$.
-$\square$
-Proposition 2 (necessity).
-To maximise the probability of learning $\omega$ from $\alpha$, it is necessary to use weakness as a proxy.
-Proof.
-Let $\alpha$ and $\omega$ be defined exactly as they were in the proof of prop. 1.
-  1. 1.
-    If $\mathbf{h}\in{\Pi}_{\alpha}$ and $E_{{I}_{\omega}}\cap E_{\mathbf{h}}={O}_{\omega}$, then it must be he case that ${O}_{\omega}\subseteq E_{\mathbf{h}}$.
-  2. 2.
-    If $\left|E_{\mathbf{h}}\right|<\left|{O}_{\omega}\right|$ then generalisation cannot occur, because that would mean that ${O}_{\omega}\not\subseteq E_{\mathbf{h}}$.
-  3. 3.
-    Therefore generalisation is only possible if $\left|E_{\mathbf{h}}\right|\geq\left|{O}_{\omega}\right|$, meaning a sufficiently weak hypothesis is necessary to generalise from child to parent.
-  4. 4.
-    For any two hypotheses $\mathbf{h}_{1}$ and $\mathbf{h}_{2}$, if $\mid E_{\mathbf{h}_{1}}\mid<\mid E_{\mathbf{h}_{2}}\mid$ then the probability $p(\left|E_{\mathbf{h}_{1}}\right|\geq\left|{O}_{\omega}\right|)<p(\left|E_{\mathbf{h}_{2}}\right|\geq\left|{O}_{\omega}\right|)$ because tasks are uniformly distributed.
-  5. 5.
-    Hence the probability that $\left|E_{{m}}\right|\geq\left|{O}_{\omega}\right|$ is maximised when $\left|E_{{m}}\right|$ is maximised.
-To maximise the probability of learning $\omega$ from $\alpha$, it is necessary to select the weakest hypothesis.
-To select the weakest hypothesis, it is necessary to use the weakness proxy. $\square$
-Proposition 3 (simplicity sub-optimality).
-Description length is neither a necessary nor sufficient proxy for the purposes of maximising the probability that induction generalises.
-Proof.
-In propositions 1 and 2 we proved that weakness is a necessary and sufficient choice of proxy to maximise the probability of generalisation.
-It follows that either maximising $\frac{1}{\left|m\right|}$ (minimising description length) maximises $\left|E_{m}\right|$ (weakness), or minimisation of description length is unnecessary to maximise the probability of generalisation.
-Assume the former, and we’ll construct a counterexample with $\mathfrak{v}=\{a,b,c,d,e,f,g,h,j,k,z\}$ s.t.
-$L_{\mathfrak{v}}=\{\{a,b,c,d,j,k,z\},\{e,b,c,d,k\},\{a,f,c,d,j\},\{e,b,g,d,j,k,z\},\newline \{a,f,c,h,j,k\},\{e,f,g,h,j,k\}\}$ and a task $\alpha$ where
-  • ${I}_{\alpha}=\{\{a,b\},\{e,b\}\}$
-  • ${O}_{\alpha}=\{\{a,b,c,d,j,k,z\},\{e,b,g,d,j,k,z\}\}$
-  • ${\Pi}_{\alpha}=\{\{z\},\{j,k\}\}$
-Weakness as a proxy selects $\{j,k\}$, while description length as a proxy selects $\{z\}$.
-This demonstrates the minimising description length does not necessarily maximise weakness, and maximising weakness does not minimise description length.
-As weakness is necessary and sufficient to maximise the probability of generalisation, it follows that minimising description length is neither.
-$\square$
-The following sketch is more for intuition and explanation than anything formal.
-Proposition 4.
-An organism that uses weakness as its proxy will learn an $n^{th}$ order self if the incentive and scale preconditions are met for that order of self.
-Proof sketch.
-Assume we have an organism $\mathfrak{a}$ that learns using “weakness” as a proxy.
-A $\mathfrak{v}_{\mathfrak{a}}$-task $\mathfrak{h}_{<t_{\mathfrak{a}}}$ represents the history of $\mathfrak{a}$ (meaning $\mathfrak{h}_{<t_{\mathfrak{a}}}\sqsubset\mu_{\mathfrak{a}}$ and $\mathfrak{h}_{<t_{\mathfrak{a}}}$ is an ostensive definition of $\mu_{\mathfrak{a}}$, because $\mathfrak{a}$ remains alive).
-The organism explores the environment, intervening to maintain homeostasis.
-As it does so, more and more inputs and outputs are included in $\mathfrak{h}_{<t_{\mathfrak{a}}}$.
-It follows that:
-  1. 1.
-    From the scale precondition we have that there exists a $n^{th}$ order self $\mathfrak{a^{n}}\in L_{\mathfrak{v}_{\mathfrak{a}}}$.
-  2. 2.
-    To remain fit, $\mathfrak{a}$ must “generalise” to $\mu_{\mathfrak{a}}$ from $\mathfrak{h}_{<t_{\mathfrak{a}}}$.
-According to the incentive precondition, generalisation to $\mu_{\mathfrak{a}}$ requires $\mathfrak{a}$ learn the $n^{th}$ order self, which is when $\mathfrak{a^{n}}\in\mathfrak{p_{\mathfrak{a}}}$.
-  3. 3.
-    From [38] we have proof that weakness is the optimal choice of proxy to maximise the probability of generalisation from child to parent is the weakest policy.
-It follows that $\mathfrak{a}$ will generalise from $\mathfrak{h}_{<t_{\mathfrak{a}}}$ to $\mu_{\mathfrak{a}}$ given the smallest history of interventions with which it is possible to do so (meaning the smallest possible ostensive definition, or cardinality $\lvert D_{\alpha}\rvert$).
-Were we to assume learning under the above conditions does not construct an $n^{th}$ order self for $\mathfrak{a}$, then one of the three statements above would be false and we would have a contradiction.
-It follows that the proposition must be true.
-$\square$